### v1

# Pritam Goswami    Adri Bhattacharya    Raja Das    Partha Sarathi Mandal

Abstract
Perpetual exploration is a fundamental problem in the domain of mobile agents, where an agent needs to visit each node infinitely often. This issue has received lot of attention, mainly for ring topologies, presence of black holes adds more complexity. A black hole can destroy any incoming agent without any observable trace. In [2, 18], the authors considered this problem in the context of Periodic data retrieval. They introduced a variant of black hole called gray hole (where the adversary chooses whether to destroy an agent or let it pass) among others and showed that 4 asynchronous and co-located agents are essential to solve this problem (hence perpetual exploration) in presence of such a gray hole if each node of the ring has a whiteboard. This paper investigates the exploration of a ring in presence of a “byzantine black hole”. In addition to the capabilities of a gray hole, in this variant, the adversary chooses whether to erase any previously stored information on that node. Previously, one particular initial scenario (i.e., agents are co-located) and one particular communication model (i.e., whiteboard) are investigated. Now, there can be other initial scenarios where all agents may not be co-located. Also, there are many weaker models of communications (i.e., Face-to-Face, Pebble) where this problem is yet to be investigated. The agents are synchronous. The main results focus on minimizing the agent number while ensuring that perpetual exploration is achieved even in presence of such a node under various communication models and starting positions. Further, we achieved a better upper and lower bound result (i.e., 3 agents) for this problem (where the malicious node is a generalized version of a gray hole), by trading-off scheduler capability, for co-located and in presence of a whiteboard.
keywords:
Mobile Agents, Exploration, Ring, Black Hole, Byzantine Fault
1 Introduction
Exploring a set of nodes in a network is one of the fundamental tasks in the domain of distributed computing by mobile agents, formulated in the year of 1951 by Shannon [20]. Now the of security of these mobile agents while exploring these networks is one of the fundamental issues to address. Among all the possible security threats that is addressed yet in literature, two among them are the most prominent, i.e., the threats from a malicious agent [17] and the threats from a malicious host [13]. In this paper, we are interested in the latter case, where the threats are from a malicious host. This host is a stationary node in the network, which has the ability to destroy any incoming agents without leaving any trace of its existence. So, first task of the mobile agents operating in the network, must be to locate this malicious node. Note that, the most trivial optimisation parameter to ensure while locating the black hole, is that minimum of number of agents gets consumed or destroyed by this node. This problem of locating the black hole by mobile agents is termed as black hole search (also termed as BHS problem) problem. BHS problem is studied from the year of 2006, when Dobrev et al. [12] first introduced it. After which, till date there has been many variations to this problem, some of them are [7, 8, 11, 13, 15]. This problem has various real life implications, such as the black hole can be a virus in the network or it can be some kind of crash failure, such that this node resembles the characteristic of a black hole, after the failure.
Observe that, in order to detect the black hole, there needs to be some agent which has to visit that particular node. Further, since any agent visiting the node gets destroyed, so there must be some communication tool, which can render this information to other alive agents, such that at least one agent remains alive, knowing the location of the black hole. Three such communication tool has been predominantly used in literature: a) whiteboard model [14], in which there is a storage capacity at each node, which an agent can use to leave a message by reading its contents and writing some new information, b) pebble model [16], an agent can carry a movable token from one node to another, c) face-to-face model [10], in this case an agent can share and communicate with another agent when they are at the same node and at the same time. In addition to the communication tools, the initial locations of the agents (i.e., whether the agents are initially scattered [15] or they are co-located [10]) is also one of the important parameters, generally studied in literature.
Further, the most studied version of black hole has fairly a basic nature, i.e., only destroying any incoming agent. Note that, in reality black hole’s may not be such simple, it may have many ways to disrupt the movement or harm an agent. Considering this phenomenon, we in this paper have tried to consider a black hole which has more capabilities other than just destroying any incoming agent. In our case, black hole may or may not kill any incoming agent, it may do so based on an adversarial scheduler which decides when to destroy an incoming agent and when not to. Whenever it decides not to destroy an agent, it simply behaves as any other node in the network, disguising it from the rest of the nodes, creating no anomaly for the visiting agent. In addition to this, we have also considered that the black hole has further capabilities, it can also choose whether to destroy the message (i.e., stored data in case of whiteboard, and placed token in case of pebble) at that node along with the incoming agent or not. This choice is also maintained by an adversarial scheduler. We call this kind of black hole as a byzantine black hole.
Our aim in this paper is to solve the problem of perpetual exploration in a network, i.e., visiting every node in the network infinitely often by the mobile agents, while detecting the presence of this byzantine black hole. After detection, the task of the agents must be visit the remaining nodes of the network infinitely often, whereas avoiding this byzantine black hole node. Previously, in [2, 18] the authors introduced this model of black hole, which has more capabilities other than just destroying a black hole. They considered the following black hole characteristics: a black hole can fake agents, change the whiteboard contents, or change the ports different from the requested ones. In those papers, their aim was to achieve perpetual exploration in a ring (which they term as periodic data retrieval problem) by a team of asynchronous mobile agents, which are initially co-located, under these various black hole characteristics. In our paper, we chose one such black hole (which has more capability than a normal black hole) and we investigated this perpetual exploration problem in a ring, by a team of synchronous agents, under all possible initial configurations (i.e., the agents can be co-located or the agents can be scattered), as well as under all possible communication tools (i.e., the agents can either communicate among themselves via whiteboard, or using a pebble, or while they meet in a face-to-face manner). In each case, we are able to design algorithms, which serve our purpose of detection of this malicious node, as well as execute perpetual exploration.
1.1 Related Works
The black hole search (i.e., BHS) problem is a prominent sub-problem of exploration problem, studied in literature. A survey of which can be found in [19]. This problem is investigated under various topologies (such as trees [8], rings [13], tori [5] and in arbitrary and unknown networks [7, 11]). All these discussed networks yet are static in nature. Recently, there has been a lot of interest on dynamic networks. The following papers [3, 4, 10], studied the BHS problem on dynamic ring, dynamic torus and dynamic cactus graph, where the underlying condition is that, irrespective of how many edges are dynamic in nature, the network must remain connected at any time interval (which is also termed as 1-interval connected). In rings, the BHS problem has been studied for different variants, the most predominant among them are choice of schedulers (i.e., synchronous [6] and asynchronous [1]), communication tools (i.e., face-to-face [7], pebble [16] and whiteboard [1]) and initial position of the agents (i.e., colocated [1] and scattered [6]).
The most relevant papers, related to our work are the papers by Královič et al. [18] and by Bampas et al. [2]. The paper by Královič et al. [18] is the first to introduce a variant of this black hole, where the black hole has the ability to either choose to destroy an agent or let it pass (which they term as gray hole). Further they extended the notion of gray hole, where the gray hole has these following additional capabilities: it has the ability to alter the run time environment (i.e., changing the whiteboard information), or it has the ability to not to maintain communication protocol (i.e., do not maintain FIFO order). They solved this problem under asynchronous scheduler on a ring, only when the agents are initially co-located and each node in the network has a whiteboard. The following results are obtained by them, they gave an upper bound of 9 agents for performing periodic data retrieval (i.e., which is equivalent to perpetual exploration) in presence of a gray hole, further, in addition to gray hole, when the whiteboard is unreliable as well, they proposed an upper bound of 27 agents. Next, Bampas et al. [2] significantly improved the earlier results. They showed a non-trivial lower bound of 4 agents and 5 agents for gray hole case and for gray hole with unreliable whiteboard case, respectively. Further, with 4 agents as well, they obtained an optimal result for the gray hole case, whereas with 7 agents proposed a protocol for the case with gray hole and unreliable whiteboard. As far as we are aware, we are the first to investigate the perpetual exploration problem of a ring under different communication tools (i.e., face-to-face, pebble and whiteboard) as well as for different initial positions (i.e., co-located and scattered), for a variant of gray hole, where it can erase any previously stored information but can not alter it. We term this type of gray hole as a byzantine black hole. In the following part, we discuss the results we have obtained.
Our Contribution: We, in this paper, investigate the perpetual exploration problem, by a team of synchronous mobile agents, of a ring $R$ of size $n$, in presence of a byzantine black hole. First, we consider the case when the agents are initially co-located. We obtain the following results.
A: For Face-to-Face model of communication we obtain that $\lceil\log(n-1)\rceil+3$ agents are sufficient to perpetually explore $R$.
B: For Pebble model of communication, we obtain that 3 agents are necessary and sufficient to perpetually explore $R$.
C: For Whiteboard model as well, we achieve the same lower and upper bounds as mentioned in B. This result shows that, by weakening the scheduler from asynchronous (as assumed in [2]) to synchronous, the tight bound on number of agents in order to perpetually explore $R$, reduces from 4 to 3.
Next, we consider the case, when the agents are initially scattered, and in this context, we obtain the following results:
D: For Pebble model of communication, we show that 4 agents are necessary and sufficient to perpetually explore $R$.
E: For Whiteboard model of communication, we obtain an improved bound of 3 agents (in comparison to D), which is necessary and sufficient to perpetually explore the ring $R$.
In the following table, we have summarized the results.
Whiteboard Pebble Face-to-Face
Co-located Upper Bound 3 3 $\lceil\log(n-1)\rceil+3$
Lower Bound 3 3
Scattered Upper Bound 3 4
Lower Bound 3 4 Non-Constant [9]
Table 1: Table of results
2 Model and Preliminaries
In this paper , the considered underlying topology of the network is an oriented ring $R=\{v_{0},v_{1},\dots v_{n-1}\}$. Each node $v_{i}$ (where $i\in\{0,1,\dots,n-1$}) is unlabeled and has two ports connecting $v_{(i-1)\bmod{n}}$ and $v_{(i+1)\bmod{n}}$ labeled left and right consistently. A set $A=\{a_{0},a_{1},\dots a_{k-1}\}$ of $k$ agents operate in $R$. We consider two types initial position for the set $A$ of agents. In the first type we consider, each agent in $A$ is co-located at a node, which we term as $home$. In the second type we consider, the agents can start from several distinct nodes, which we term as scattered initial position. Each agent has the knowledge of the underlying topology $R$ and has some computational capabilities, thus it requires $\mathcal{O}(\log n)$ bits of internal memory. The agents have unique IDs of size $\lceil\log k\rceil$ bits taken from the set $[1,k]$, which are perceived by other agents when they are co-located. The agents are autonomous and they execute same set of rules (i.e., they execute the same algorithm).
We consider three types of communication that the agents have, in order to communicate with other agents. These models of communication are as follows:
  • Face-to-Face: In this model of communication, an agent can communicate with another agent when they are co-located.
  • Pebble: In this model, the agents are equipped with a movable token (also termed as pebble), which signifies a single bit of information, the agents can use this pebble in order to communicate with other agents.
  • Whiteboard: In this case, each node of $R$ contains $\mathcal{O}(\log k)$ bits of memory, which can be used to store and maintain information. Any agent can read the existing information or write any new information at the whiteboard of its current node. Note that, fair mutual exclusion is maintained, i.e., concurrent access of the whiteboard data is not permitted.
The agents operate in synchronous rounds and in each round every agent becomes active and takes a local snapshot of its surroundings. For an agent at a node $v$ in some round $r$, the snapshot contains two ports incident to $v$, already stored data’s on the memory of $v$ (if any, only in case of whiteboard model of communication), number of pebbles located at $v$ (if any, only in case of pebble model of communication), contents from its own local memory and IDs of other agents on $v$. Based on this snapshot an agent executes some action. This actions includes a communication step and a move step. In a communication step an agent can communicate implicitly or explicitly with other agents according to the communication models discussed above. In the move step the agent can move to a neighbour node by following a port incident to $v$. Thus if an agent at node $v$ in round $r$, decides to move during the move step, at round $r+1$ it resides on a neighbour node of $v$. All these actions are atomic, so an agent cannot distinguish another agent concurrently passing through the same edge, instead it can only interact with another agent (based on its communication model) only when it reaches another node.
A black hole is a stationary malicious node in an underlying graph, which has the ability to destroy any visiting agent, without leaving any trace of its existence. In this paper we have considered one unique node of $R$ to be a black hole. Further, the black hole nature of the node is controlled by an adversary. We term this kind of node as Byzantine Black Hole, which is defined as follows.
Definition 2.1 (Byzantine Black Hole).
Let $G$ be an underlying graph of a network with agents. A node $u$ of $G$ is called a Byzantine black hole if sometimes it behaves as a black hole, whereas, in rest of the time it behaves as a normal node. Moreover, this transition between the black hole nature and normal nature of the node is controlled by the adversary.
We, in this paper, have assumed that the underlying graph contains a single byzantine black hole, whereas the other nodes are normal nodes, and they are termed as safe nodes. Note that, the starting positions of each agent must be a safe node. Here, we assume that, if the adversary decides to activate the black hole nature of the byzantine black hole node, then it does so at the beginning of its corresponding round, and the node retains that nature until the end of this current round. Furthermore, we have considered that our byzantine black hole has the ability to always destroy any incoming agent, during its black hole nature. In addition it also has the ability to choose whether to destroy any information present on that node during the black hole nature of that node.
So, our aim in this paper, is to perpetually explore the ring $R$, with minimum number of agents. Next, we formally define our problem.
Definition 2.2 (PerpExploration-BBH).
Given a ring network $R$ with $n$ nodes, where one node ($v_{b}$) is a Byzantine black hole, and with a set of agents $A$ positioned on R, the PerpExploration-BBH asks the agents in $A$ to move such a way that, each node of $R$ except $v_{b}$ is visited by at least one agent infinitely often.
3 Impossibility Results
In the first result we prove that two agents can not solve PerpExploration-BBH in the strongest communication model of whiteboard.
Theorem 3.1.
A set of two synchronous agents in a ring $R$ of size $n$ cannot solve PerpExploration-BBH, even in presence of whiteboard if number of possible consecutive black hole positions is greater or equals to 3.
Proof 3.2.
Let $v_{1},v_{2}$ and $v_{3}$ be three possible consecutive black hole positions in a ring $R$. Let two agents $a_{1}$ and $a_{2}$ are sufficient to solve the PerpExploration-BBH problem on $R$. Thus there exists an algorithm $\mathcal{A}$ such that $a_{1}$ and $a_{2}$ can solve PerpExploration-BBH by executing $\mathcal{A}$. Without loss of generality let $a_{1}$ explores $v_{2}$ and it moves to $v_{2}$ for the first time at round $t$ from vertex $v_{1}$. Then $a_{1}$ must have been at the vertex $v_{1}$ at round $t-1$. Let us take two copies $R_{1}$ and $R_{2}$ of the same ring $R$. In $R_{1}$, $v_{1}$ is the black hole and in $R_{2}$, $v_{2}$ is the black hole. Let the adversary in $R_{1}$ destroys $a_{1}$ at round $t-1$ and in $R_{2}$ destroys $a_{1}$ in round $t$. We claim that at round $t$ and $t-1$, $a_{2}$ can not be on either of $v_{1}$ or $v_{2}$. Note that at $t$-th round $a_{2}$ can not be on $v_{2}$ otherwise, both $a_{1}$ and $a_{2}$ gets destroyed at the same round and no other agent is left to explore further. Similarly at round $t-1$ $a_{2}$ can not be on $v_{1}$. Now consider the case when at $t-$ th round $a_{2}$ is at $v_{1}$. Since at round $t-1$, $a_{2}$ can not be on $v_{2}$ there can be two cases either at round $t-1$, $a_{2}$ was in $u_{0}$ ($u_{0}$ is another adjacent node of $v_{1}$ except $v_{2}$) or was in $v_{1}$ itself. We first argue that at round $t-1$, $a_{2}$ can not be at $u_{0}$. Otherwise, since the whiteboard content at nodes except at $v_{1}$ and $v_{2}$ are same in $R_{1}$ and $R_{2}$ at round $t$ (due to same execution by $a_{1}$ and $a_{2}$ upto round $t-1$ and at round $t$ for $a_{2}$ in both $R_{1}$ and $R_{2}$). Now since at round $t$ all previous datas on whiteboard except $v_{1}$ and $v_{2}$ does not help $a_{2}$ to distinguish between $R_{1}$ and $R_{2}$, the problem now can be thought of solving PerpExploration-BBH with one agent where number of possible consecutive black hole position is atleast two. Now this is impossible to solve. So, at $t-1$-th round $a_{2}$ can not be on $u_{0}$ We now only have to prove that $a_{2}$ can not be at $v_{2}$ during round $t-1$. Let in $R_{2}$, adversary activates the black hole at $v_{2}$ at round $t-1$ which destroys $a_{2}$ at round $t-1$. Now since the effect of this destruction of agent $a_{2}$ does not effect the inputs of $a_{1}$ at round $t-1$ on $v_{1}$, it move to $v_{2}$ at round $t$ where the black hole is activated again by adversary destroying $a_{1}$ too. Thus $a_{2}$ must remain outside of $v_{1}$ and $v_{2}$ during round $t$ and $t-1$ while executing $\mathcal{A}$. So, at round $t$ since all whiteboard content are same for $R_{1}$ and $R_{2}$ (except for nodes $v_{1}$ and $v_{2}$ whose content only differs after round $t-2$ for $R_{1}$ and $R_{2}$ which can not be known by $a_{2}$ as it was not there during that rounds), they can not help $a_{2}$ to distinguish between $R_{1}$ and $R_{2}$. So the problem now can be thought of as solving PerpExploration-BBH with one agent where number of possible consecutive black hole position is at least 2. And this is impossible. So there doesn’t exist any algorithm that solves PerpExploration-BBH with two agents on a Ring $R$ where number of possible consecutive positions of black hole is at least 3.
Note that since the $n$ can be sufficiently large so that initially number of possible consecutive black hole positions is atleast three we have the following corollary.
Corollary 3.3.
It is impossible to solve PerpExploration-BBH on a ring $R$ with $n$ nodes using two synchronous agents even if the nodes are equipped with whiteboards.
Corollary 3.4.
A set of three synchronous agents are required to solve the PerpExploration-BBH on a ring $R$ with $n$ nodes, where each node has a whiteboard.
Corollary 3.5.
A set of two agents, each equipped with $O(\log n)$ pebbles can not solve the PerpExploration-BBH problem on a ring $R$ with $n$ nodes.
Now the following theorem improves the lower bound for the case when agents are scattered and each agent is equipped with a pebble, whereas there is no whiteboard present at the nodes of $R$. The theorem is as follows.
Theorem 3.6.
A set of 3 scattered agents, each equipped with a pebble can not solve the PerpExploration-BBH problem on a ring $R$ with $n$ nodes.
Proof 3.7.
Let $a_{1},a_{2}$ and $a_{3}$ be three agents equipped with a pebble each. The agents are placed on three nodes $h_{1},h_{2}$ and $h_{3}$ initially in such a way that distance between $h_{i}$ and $h_{j}$ is same for all $i,j\in\{1,2,3\}$, where we consider $h_{j}$ to be the nearest node of $h_{i}$ in the clockwise direction. Without loss of generality let this distances be sufficiently large. Further, let there exists an algorithm $\mathcal{A}$ that solves the PerpExploration-BBH problem in this setting. Let without loss of generality $a_{1}$ be the first agent to explore the third node from its corresponding starting position (i.e., $h_{1}$) in any of the clockwise or counter-clockwise direction, when each agent starts executing the algorithm $\mathcal{A}$. Suppose by following $\mathcal{A}$, $a_{1}$ can visit the third node in clockwise direction (without loss of generality) first at a round say $t>0$. Let $v_{1},v_{2}$ and $v_{3}$ be those set of three nodes from $h_{1}$ in the clockwise direction. Let $C_{1}$,$C_{2}$ and $C_{3}$ be three scenarios where in $C_{i}$, $v_{i}$ is the black hole. We claim that, $a_{1}$ can not carry its pebble during any execution of $\mathcal{A}$. Otherwise in scenario $C_{1}$, it would be consumed along with its pebble, and since the distances between two consecutive $h_{i}$ are sufficiently large, hence other agents would have no idea that an agent is already consumed. This is equivalent to solving the PerpExploration-BBH with two agents having a pebble each and as $n$ is sufficiently large this is impossible due to Corollary 3.5. Now suppose adversary chooses to activate the black hole whenever $a_{1}$ reaches there for the first time. In this case for all $C_{1}$, $C_{2}$ and $C_{3}$, at round $t$, agents $a_{2}$ and $a_{3}$ has no idea about where $a_{1}$ is consumed even if it knows that $a_{1}$ is consumed, as the distances between two consecutive $h_{i}$ are sufficiently large so their exploration region does not intersects till round $t$ and timeout (or,waiting for other agent) strategy do not work as for all $C_{i}$ an agent can get same timeout output. Also for all $C_{i}$s the position of the pebbles and alive agents will be same at round $t$. Now for the alive agents, the number of nodes for possible position of the black hole must be greater than or equal to 3 at round $t$. So the situation at round $t$ is similar to the problem of solving PerpExploration-BBH on a ring $R$ with two agents having a total of 3 pebbles where the number of possible consecutive positions of black hole is greater or equal to 3. Since we have assumed $\mathcal{A}$ solves the problem thus, $\mathcal{A}$ can also solve the problem of PerpExploration-BBH with two agents where number of possible positions of black hole is greater or equal to 3. But due to Theorem 3.1 it is impossible. Hence there can never exists any algorithm that solves PerpExploration-BBH with three scattered agents each of which are equipped with a pebble.
Corollary 3.8.
A set of four scattered agents each with a pebble, are required to solve PerpExploration-BBH on a ring $R$ with $n$ nodes.
4 Co-located Agents
4.1 Communication: Face-to-Face
In this section, we discuss the PerpExploration-BBH problem, under the conditions, where the agents are initially located at a safe node $home$, and they can only communicate with another agent whenever they are at the same node at the same round. Here considering this scenario we provide an algorithm PerpExplore-F2F (Algorithm 1) to solve the PerpExploration-BBH problem with $\lceil\log(n-1)\rceil+3$ agents.
Description of PerpExplore-F2F: Given a ring $R$ with a byzantine black hole node $v_{b}$, a set of agents $A=\{a_{1},a_{2},\dots,a_{k}\}$, where $k=\lceil\log(n-1)\rceil+3$, are initially co-located at a safe node, which we term as $home$. Note that, the agents have no knowledge about the position of $v_{b}$. The agents have the knowledge of $n$ (which is the total number of nodes in $R$), whereas they know that $home$ is safe, as it is the initial position of the set of agents, $A$. Moreover, they have no idea about the remaining $n-1$ nodes of $R$, as $v_{b}$ can be any node among these $n-1$ nodes. We call this arc containing these $n-1$ nodes as the suspicious region, and it is denoted by $S$ with cardinality $|S|$.
The main idea of the PerpExplore-F2F algorithm is from the $home$ only two agents, say $a_{1}$ and $a_{2}$ starts exploring the ring perpetually. The algorithm goes like this, one agent (preferably the agent with lowest Id, among the set of agents which are yet to be consumed by $v_{b}$) moves clockwise until it reaches the ’middle node’ of the suspicious region $S$, while the other agent (preferably the agent with second lowest Id, among the set of agents which have not been consumed by $v_{b}$) also performs the same movement, but along counter-clockwise direction, after which they return to the $home$ and this process perpetually continues. Now while executing this movement by these two agents, there can be two possible scenarios, the moment at least one agent gets destroyed by the byzantine black hole, $v_{b}$. Firstly, both the agents can get destroyed (this happens only when $v_{b}$ is the middle node of $S$), whereas secondly, only one agent is destroyed while the other agent following its usual course of algorithm, returns to $home$. For the initial case, the remaining agents at home can uniquely determines the byzantine black hole, $v_{b}$, as the middle node of $S$ and consecutively starts exploring the ring while avoiding $v_{b}$. For the latter case, the agents at the $home$ along with the agent that has returned to $home$, eventually gets to know that, the other agent has not returned and which in turn concludes that the arc, along which the missing agent was exploring contains the node $v_{b}$. This way our algorithm guarantees that, in the worst case, the size of the suspicious region decreases by a factor of at least half of the previous size, in the expense of one agent. So this implies, when $\lceil\log n\rceil$ agents are destroyed in $v_{b}$, the size of the suspicious region becomes exactly 1 (where this node in the suspicious region is nothing but $v_{b}$). In this case the remaining alive agent can determine the exact position of $v_{b}$ and hence, can perpetually explore the ring $R$ avoiding $v_{b}$.
Next we will give a brief description of our algorithm PerpExplore-F2F. The algorithm works as follows, initially each agent at $home$ starts from state Initial0. In this state, an agent initializes the following variables: $S_{lft}$ to 1, $S_{rgt}$ to $n-1$,where $S_{lft}$ and $S_{rgt}$ denotes the distance to the left and the right end nodes of the suspicious region $S$ from the $home$, in the clockwise direction. Also in this state an agent declares its current node as the $home$. Note that, the agents stay at this state only for one round at the beginning of the execution of our algorithm, after which they enter the state Initial1. Now, after an agent enters the state Initial1, it calculates $|S|$ (where $|S|=|S_{rgt}-S_{lft}+1|$), initializes $Tnodes$ to 0 (where $Tnodes$ is the distance an agent traverses in order to reach its destination node). Next, it checks the variables $|S|$ and $k$ (where $k$ is the number of agents present at the $Current-Node$, i.e., $home$). If it finds $|S|=2$ and $k=2$, i.e., the length of the suspicious region is 2 and the total number of agents present at $home$ is also 2, then both the agents execute the sequence $SEQ0$ of states, where $SEQ0$ is defined as follows:
[…]
According to this sequence, the agent moves in to state Detection (after satisfying both the conditions $|S|=2$ and $k=2$ in state Initial1). In this state, the lowest Id agent is instructed to move clockwise and reach its destination position $S_{lft}$, whereas, the other agent reaches its destination $S_{rgt}$ along a counter-clockwise direction, while after each movement updating $Tnodes=Tnodes+1$. Now, after reaching their respective positions, the agents return back to $home$ along the same path in which they reached their destination nodes, again incrementing $Tnodes$ by 1 after each move. So, the total number of rounds required by the agent, starting from Initial1 in order to reach $home$ is $Tnodes+1$. After reaching $home$, they are further instructed to wait for $22n-Tnodes+$ 2 rounds. In the following round it again checks whether, #$\#Agent$ at $home$ is still $k$ i.e., 2, if so then again it changes to state Initial1, and the sequence $SEQ0$ is executed repeatedly. Otherwise, if the current agent is of lowest Id then declares the byzantine black hole node is at the location of $S_{rgt}$, otherwise if the current agent is not of lowest Id, then it declares that this byzantine black hole node is at $S_{lft}$. Irrespective of which, they continue performing perpetual exploration of $R$, by avoiding this black hole node. Hence, the total number of rounds required to complete the execution of this sequence is $42n+$ 4.
On the contrary, if either $|S|\neq 2$ or $k\neq 2$. Note that, if $|S|=1$, then this node is nothing but the byzantine black hole node, and the remaining agents (such agent always exists, refer Lemma 4.11) detect this, and further continue perpetual exploration, avoiding this node. Otherwise, for $|S|\geq 2$ and $k\geq 3$ in state Initial1, the agents update the following variables: $k$ to $k-2$ (where $k-2$ is the number of agent which enters the next state Stagnant), locates the next destination node, where it calculates the variable $Dest=S_{lft}+\lceil\frac{|S|}{2}\rceil-1$, which is the clockwise distance of the destination node from $home$, $W_{time}$ to 00 (where $W_{time}$ signifies the waiting time at destination node). Agents in Initial1 state can change their state in to one of the following three states: Forward-Left, Forward-Right and Stagnant. An agent moves in to Forward-Left, if it is the agent with lowest Id among all the agents present at $home$, whereas an agent moves in to Forward-Right, if the agent is having second lowest Id among all the agents present at $home$. Lastly, an agent enters the state Stagnant, if it is neither with lowest Id nor with second lowest Id.
According to the state an agent enters after Initial1, we define three sequence of states $SEQ1$, $SEQ2$ and $SEQ3$, respectively, which an agent can follow in the exact order after Initial1 state, based on whether the agent enters Forward-Left or Forward-Right or Stagnant. If the agent enters the state Forward-Left then it starts following $SEQ1$, or if it enters the state Forward-Right then it follows $SEQ2$ or if it enters Stagnant state then it follows $SEQ3$, respectively. Note that, in each sequence of states the agent eventually reaches back to Initial1, if not it is destroyed at the node $v_{b}$, while performing any of the states: Forward-Left (resp, Forward-Right) or Wait-Left (resp, Wait-Right) or Backtrack-Left (resp, Backtrack-Right). Also note that if one agent stated executing sequence $SEQ1$, then there must be another agent that executes sequence $SEQ2$ and the other agents execute the sequence $SEQ3$ starting from the same round (refer Line 10-17 of Algorithm  1). We define $SEQ1$, $SEQ2$ and $SEQ3$ as follows:
[…]
Our algorithm guarantees that, each agent starting from Initial1 again enters the state Initial1 at the same round, after following any of these three sequences.
Definition 4.1 (Iteration).
We define an iteration of our algorithm as the number of rounds required to execute any one of these sequences (i.e., $SEQ0$, $SEQ1$, $SEQ2$ and $SEQ3$) by an agent.
Note that, each iteration has exactly $42n+$ 4 rounds. Each agent performs these iterations until this byzantine black hole is detected, after which, they explore the ring $R$ perpetually just by avoiding this black hole node.
Next, we describe how the agents execute the sequences in order to perform the PerpExplore-F2F algorithm in one iteration.
1 Input: $n,k$;
2 States: {Initial0, Initial1, Stagnant, Forward-Left, Forward-Right, Backtrack-Left, Backtrack-Right, Wait-Left, Wait-Right, Wait1-Left, Wait1-Right}
3
4In State Initial0:
5 $S_{lft}=$ 1, $S_{rgt}=$ $n-1$;
6
7Declare $Current-Node$ as $home$ and update to state Initial1.
8
9In State Initial1:
10 $|S|=|S_{rgt}-S_{lft}+1|$, $Tnodes=0$;
11 if $S_{lft}\neq S_{rgt}$ then
12       if $|S|=2\wedge k=2$ then
13             Move to state Detection.
14
15      else
16             $k=k-2$, $Dest=S_{lft}+\lceil\frac{|S|}{2}\rceil-1$, $W_{time}=0$;
17             if $LowestID$ then
18                   Update state to Forward-Left.;
19
20             else if $SecondLowestID$ then
21                   Update to state Forward-Right.;
22
23             else
24                   Move to state Stagnant.;
25
26
27
28else
29       Declare $S_{lft}$ as the byzantine black hole location and continue exploration avoiding that node.
30
31
32In State Stagnant:
33 Wait at the $Current-Node$ for $22n+$ 2 rounds.
34 if #$\#Agent=k$ then
35       Declare $Dest$ is the byzantine black hole location and continue exploration avoiding that node.
36
37 else if #$\#Agent=k+1$ then
38       Update $k=k+1$ and collect the updated $S_{lft}$, $S_{rgt}$ and update accordingly.
39       Move to state Initial1.
40
41 else
42       Update $k=k+2$ and move to state Initial1.
43
44
[…]
47
48In State Wait-Left:
49
50if $W_{time}<n-Tnodes$  then
51       $W_{time}++$
52else
53       Enter state Backtrack-Left and move $right$
54
55
56In State Wait-Right:
57
58if  $W_{time}<n-Tnodes$  then
59       $W_{time}++$
60else
61       Enter state Backtrack-Right and move $left$
62
63
Algorithm 1 PerpExplore-F2F
Execution of the sequence $SEQ1$: As discussed earlier, if the agent is of lowest Id in the in the state Initial1, then it is instructed to change their state to Forward-Left. In state Forward-Left, an agent reaches the destination node, which is $Dest$ distance away from $home$, along the clockwise direction, while after each move incrementing the variable $Tnodes$ by 1. Now, the moment it reaches the destination node, it is further instructed to change its state to Wait-Left. Note that, an agent stays in Forward-Left state for exactly $Tnodes+1$ rounds while following the sequence $SEQ1$ in an iteration. After changing its state to Wait-Left, the agent waits at the destination node for $n-Tnodes-1$ rounds, after which it is again instructed to move in to another state, which is Backtrack-Left while already starting to move one hop along counter-clockwise direction. Observe that an agent spends exactly $n-Tnodes$ round in state Wait-Left while executing $SEQ1$ in an iteration. Next in the state Backtrack-Left, an agent immediately starts moving along counter-clockwise direction, until the agent reaches the $home$ node, which is $Tnodes-1$ distance apart from its current node. While it reaches $home$, it again changes to state Wait1-Left. An agent stays at the Backtrack-Left state for exactly $Tnodes$ rounds. State Wait1-Left, instructs the agent to wait at the current node, i.e., $home$ for $n-Tnodes$ many rounds. This waiting time is enough for another agent that started executing sequence $SEQ2$ to return back to the $home$ if not destroyed by the byzantine black hole. After waiting, the agent checks the number of agent present at the current node. If #$\#Agent$ increases to $k+1$ (where $k$ agents are in state Stagnant), then in that case, it updates $k$ to $k+1$ and understands that the second lowest Id agent has not returned to $home$. In this situation, the only possibility of the other agent not returning back is that it has entered the black hole, along its traversal. In this situation it updates $S_{lft}$ to the destination node, which is $Dest$ distance apart from $home$ (i.e., basically the suspicious region becomes half, because initially $S_{lft}$ was 1, whereas now it is updated to $Dest$), and also communicates this information to the remaining $k-1$ agents at $home$, and directly moves into state Initial1.
42
[…]
45 In State Wait1-Left:
46 Update $d=0$ and wait at the $Current-Node$ for $n-Tnodes$ rounds.
47 if #$\#Agent=k+1$ then
48       Update $k=k+1$ and $S_{lft}=Dest$ and communicate $S_{lft}$ to remaining agents at $Current-Node$.
49       In the next round move to state Initial1.
50
51else
52       Update $k=k+2$.
53       In the next round move to state Initial1.
54
55
56In State Wait1-Right:
57 Update $d=0$ and wait at the $Current-Node$ for $n-Tnodes$ rounds.
58 if #$\#Agent=k+1$ then
59       Update $k=k+1$ and $S_{rgt}=Dest$ and communicate $S_{rgt}$ to remaining agents at $Current-Node$.
60       In the next round move to state Initial1.
61
62else
63       Update $k=k+2$.
64       In the next round move to state Initial1.
65
66
67In State Detection:
68 If lowest Id then move to $S_{lft}$ clockwise, otherwise move to $S_{rgt}$ counter-clockwise, and after each move, update $Tnodes=Tnodes+1$.
69 Return back to $home$ and again update $Tnodes=Tnodes+1$ after each move.
70 if $Current-Node$ is $home$ then
71       Wait at $home$ for $22n-Tnodes+$ 2 rounds.
72       if #$\#Agent=k$ then
73             Move to state Initial1.
74
75      else
76             If lowest Id then declare byzantine black hole is $S_{rgt}$, otherwise declare byzantine black hole is $S_{lft}$, and then continue exploration avoiding this node.
77
78
79
On the other hand, if #$\#Agent$ at $home$ increases from $k$ to $k+2$ (i.e., both lowest Id and second lowest Id has returned back to $home$) then the agent simply waits for one round, after which it directly moves to state Initial1. Here note that an agent stays in the state Wait1-Left for $n-Tnodes+2$ rounds. Hence for an agent executing sequence $SEQ1$, the length of an iteration is $42n+$ 4.
Execution of the sequence $SEQ2$. This execution if followed by the agent, which in state Initial1 has the second lowest Id. In this sequence, the agent from Initial1 moves to state Forward-Right. The state Forward-Right is exactly the same as the state Forward-Left which has been discussed above, the only difference is that the agent moves along a counter-clockwise direction, instead of clockwise direction. After this state, the agent moves to state Wait-Right, in this state again similar to the earlier state Wait-Left, the agent waits for $n-Tnodes-1$ many rounds at the destination node. After which it directly moves along clockwise direction for one hop and moves to state Backtrack-Right. This state resembles Backtrack-Left, instead the fact that in this state the agent moves along clockwise direction for $Tnodes-1$ distance, in which case it reaches $home$. After reaching $home$, it again waits for $n-Tnodes$ rounds while in state Wait1-Right, and after which performs similar checks as discussed earlier in Wait1-Left, but here, if after the waiting period, the total number of agents at $home$ is $k+1$ (where $k$ many agents are in state Stagnant), it similarly modifies the suspicious region by updating $S_{rgt}$ to $Dest$ (while earlier $S_{lft}$ was updated to $Dest$ in state Wait1-Left). Lastly, it ends this sequence $SEQ2$ by moving in to state Initial1 again. Similar to the length of an iteration for an agent executing $SEQ1$, the length of an iteration for any agent executing $SEQ2$ is again $42n+$ 4.
Execution of the sequence $SEQ3$. This execution is followed by the agent, which in state Initial1 finds that it is neither of lowest Id nor of second lowest Id. In this state, the agent is instructed to move to state Stagnant. While in state Stagnant, it waits for $22n+$ 2 rounds at $home$, after which in $(2n+3)$-th round it checks the number of agent present at $home$. If ##\##Agent is $k$ (i.e., only the agents which moved to state Stagnant), then the agent understands that the lowest Id agent (which started sequence $SEQ1$) and second lowest Id agent (which started sequence $SEQ2$), both has entered the black hole, and since $R$ has only one black hole node, so this scenario of both agents entering the black hole can only occur when the node is common to both of these agents. This common node is nothing but the destination node, hence if this scenario happens, then the agent in state Stagnant, concludes that the black hole is none other than the destination node, and continues perpetual exploration avoiding this node. On the contrary, if #$\#Agent$ is $k+1$ (i.e., one among the agents executing $SEQ1$ and $SEQ2$ has returned, whereas the other agent has entered the black hole). In this scenario, updates either $S_{lft}$ or $S_{rgt}$ (based on the fact, whether the agent executing $SEQ1$ or the agent executing $SEQ2$ has returned), and then moves to state Initial1 while updating $k$ to $k+1$, ending sequence $SEQ3$. On the other hand, if both the agents executing $SEQ1$ and $SEQ2$ return to $home$ (i.e., ##\##Agent is $k+2$), then the agents in state Stagnant simply moves to state Initial1 while updating $k$ to $k+2$, and ending the sequence $SEQ3$. Note that here also, the iteration is of length $42n+$ 4 rounds, which leads to the following observation.
{observation}
For any $i$ ($i>0$), the $i$-th iteration for each of the $k$ alive agents, starts at the same round and at the same node, i.e., $home$.
{observation}
For any $i$ ($i>0$), each agent moves to the state Initial1 at the end of this $i$-th iteration.
{observation}
If $|S|=2$ and $k=2$, then both these agents follow the sequence $SEQ0$, whereas, if $|S|\geq 2$ and $k\geq 3$, then one agent follows $SEQ1$, another agent follows $SEQ2$, whereas the remaining agents follow the sequence $SEQ3$. That is,during an iteration if an agent follows $SEQ0$ then no agent follows $SEQ1$, $SEQ2$ or $SEQ3$ in that iteration, and vice-versa.
4.1.1 Correctness and Complexity
In this section we give the correctness and complexity analysis of our algorithm PerpExplore-F2F.
Lemma 4.2.
During any iteration, if exactly one agent gets destroyed by the byzantine black hole, then at the end of this iteration, the length of the suspicious region becomes at most $\lceil\frac{|S|}{2}\rceil+1$, where $|S|\geq 2$ is the length of the suspicious region at the start of this iteration.
Proof 4.3.
Let us consider at the $i$-th iteration of our Algorithm 1, the suspicious region be $S$ (where, $|S|>1$), so the agent starts from the state Initial1, and can encounter one of the following scenarios: first, $|S|=2$ (i.e., $S=\{S_{lft},S_{rgt}\}$) and $k=2$, second, $|S|\geq 2$ and $k>2$ (refer line 8 of Algorithm 1).
If the first scenario is satisfied, then both the agents execute the sequence $SEQ0$ (refer description of PerpExplore-F2F), in which they perpetually visit the nodes $S_{lft}$ and $S_{rgt}$, respectively and return back to $home$, until one of these agent fails to return $home$, i.e., enters the byzantine black hole (refer state Detection in Algorithm 1). So, while the other agent returns $home$ (which was visiting $S_{rgt}$, without loss of generality), it is instructed to wait for $22n-Tnodes+$ 2 rounds (refer line 63 of Algorithm 1). After waiting, it finds that the agent visiting $S_{lft}$ is yet to return $home$, then this phenomenon helps the alive agent conclude that the agent visiting $S_{lft}$ must have entered the black hole (because $22n-Tnodes+$ 2 rounds is sufficient number of rounds for any agent visiting the node $S_{lft}$ to return back to $home$, if not consumed by the black hole), which is none other than the node $S_{lft}$ (as it is the only node in the suspicious region that this agent visits). Hence, in this case the black hole node is detected and also $|S|$ becomes 1 at the end of $i$-th iteration, which satisfies our claim.
On the otherhand, if $|S|\geq 2$ and $k>3$, the agents are either following $SEQ1$ or $SEQ2$ or $SEQ3$, among them exactly one agent follows $SEQ1$ (say, $a_{1}$), exactly one agent follows $SEQ2$ (say, $a_{2}$), whereas the rest of the agent follows $SEQ3$ (where these agents stay at $home$ during this iteration, refer the description of execution of $SEQ3$). This means that the agent which can be destroyed by the black hole can be either $a_{1}$ or $a_{2}$ (as $home$ is always a safe node). If $a_{1}$ is destroyed, then $a_{2}$ during this iteration, after returning back to $home$ (while in state Backtrack-Right), waits at $home$ for $n-Tnodes$ many rounds (while in state Wait1-Right) for $a_{1}$ to return. Since, $a_{1}$ has already been consumed by the black hole, so it fails to return to $home$ within the waiting time of $a_{2}$. This observation of $a_{2}$ at $home$ instigates it to modify the suspicious region by updating $S_{rgt}$ to $Dest$, and $a_{2}$ communicates this updated information to the rest of the agents executing $SEQ3$ at $home$. It is because, $a_{1}$ only explores the nodes of the suspicious region that lies between $S_{lft}$ and $Dest$. Let $S_{new}$ be the modified suspicious region, then $|S_{new}|=|Dest-S_{lft}+1|=|S_{lft}+\lceil\frac{|S|}{2}\rceil-1-S_{lft}+1|=\lceil\frac{|S|}{2}\rceil<\lceil\frac{|S|}{2}\rceil+1$.
Now consider the other case, where instead of $a_{1}$, $a_{2}$ be the agent which enters the black hole during an iteration. In this situation as well, similar to the argument explained above in case of $a_{2}$, $a_{1}$ will get to know $a_{2}$ has entered the black hole during the iteration, while $a_{1}$ waits for $n-Tnodes$ rounds at $home$. Observe, $a_{2}$ traverses the nodes of the suspicious region that lies between $Dest$ and $S_{rgt}$. So, $a_{1}$ updates $S_{lft}$ to $Dest$ in order to get the updated suspicious region $S_{new}$ and communicate with the rest of the agents executing $SEQ3$ at $home$. This implies $|S_{new}|=|S_{rgt}-Dest+1|=|S_{rgt}-S_{lft}-\lceil\frac{|S|}{2}\rceil+2|=||S|-\lceil\frac{|S|}{2}\rceil+1|\leq\lceil\frac{|S|}{2}\rceil+1$.
So, in each scenario, our claim holds.
{observation}
Note that, if $|S|$ is odd then $|S_{new}|=\lceil\frac{|S|}{2}\rceil$, otherwise, $\lceil\frac{|S|}{2}\rceil\leq|S_{new}|\leq\lceil\frac{|S|}{2}\rceil+1$.
Lemma 4.4.
If during an iteration, exactly two agents gets destroyed by the byzantine black hole, then in that iteration, the position of the black hole is detected by the remaining alive agents.
Proof 4.5.
This situation of both agents being consumed by the black hole cannot arise when $|S|=2$ and $k=2$, because in this case both agents explore disjoint nodes of the suspicious region, in which only one node can be the black hole.
On the other hand when $|S|\geq 2$ and $k>3$, this situation of two agents entering the black hole, can only arise for the agents $a_{1}$ and $a_{2}$, where $a_{1}$ and $a_{2}$ are following $SEQ1$ and $SEQ2$, respectively. Now, since the node (say $v_{b}$) at distance $Dest$ in the clockwise direction from $home$ is the only common node that both $a_{1}$ and $a_{2}$ visits while executing an iteration. So, if both these agents gets destroyed, then $v_{b}$ must be the node, in which the black hole is located. Now, as per our algorithm, the agents executing $SEQ3$ waits at $home$ for $22n+$ 2 rounds, after which they find that both $a_{1}$ and $a_{2}$ are yet to return $home$, whereas $22n+$ 2 rounds are sufficient for these two agents to return back to $home$ while executing their respective sequences, if they have not entered the black hole. This scenario helps the remaining agent conclude that both $a_{1}$ and $a_{2}$ has entered the black hole, and the node $v_{b}$ is the exact position of the byzantine black hole.
Lemma 4.6.
Let $|S_{0}|$ be the initial length of the suspicious region where $|S_{0}|=2^{p}+l$, $p\geq 1$ and $0\leq l<2^{p}$. Let $S_{i}$ be the suspicious region after it is modified for the $i-$th time. Then, $|S_{i}|\leq\lceil\frac{|S_{0}|}{2^{i}}\rceil+2$.
Proof 4.7.
Let $|S_{0}|$ be the initial length of the suspicious region where $|S_{0}|=2^{p}+l$, $p\geq 1$ and $0\leq l<2^{p}$. Let $S_{i}$ be the suspicious region after $i$ modifications. We now prove this lemma by induction. For the base case let $i=1$. Then from Lemma 4.2, $|S_{1}|\leq\lceil\frac{|S_{0}|}{2}\rceil+1<\lceil\frac{|S_{0}|}{2}\rceil+2$. So the lemma is true for $i=1$. Let the lemma be true for $i=j-1$. That is, $|S_{j-1}|\leq\lceil\frac{|S_{0}|}{2^{j-1}}\rceil+2$. Now to prove the lemma we have to show that $|S_{j}|\leq\lceil\frac{|S_{0}|}{2^{j}}\rceil+2$. Now by Lemma 4.2, $|S_{j}|\leq\lceil\frac{|S_{j-1}|}{2}\rceil+1\leq\lceil\frac{\lceil\frac{|S_{0}|}{2^{j-1}}\rceil+2}{2}\rceil+1\leq\lceil\frac{|S_{0}|}{2^{j}}\rceil+2$. Hence, by mathematical induction we can conclude that $|S_{i}|\leq\lceil\frac{|S_{0}|}{2^{i}}\rceil+2$, for any $i\geq 1$
Lemma 4.8.
If $|S_{t}|>2$ , then there exists at least one agent which is following the sequence $SEQ3$ in our Algorithm 1 in any iteration until the next modification of $S_{t}$, where $S_{t}$ is the suspicious region after $t$ modifications .
Proof 4.9.
Let $S_{0}$ be the initial suspicious region. Then, $|S_{0}|=2^{p}+l$ where $p\geq 1$ and $0\leq l<2^{p}$. Let $S_{i}$ be the suspicious region after $i$ modifications of it. Let for some $t>0$, $|S_{t}|>2$. This implies that during all the previous $t$ iterations where the suspicious region is modified, only one agent is consumed by the byzantine black hole. As otherwise $|S_{j}|=1$ for some $0<j<t$, and already the byzantine black hole is detected. This implies after the $t$-th modification, number of agents consumed is exactly $t$. So the remaining number of agent after $t$-th modification of suspicious region is exactly $\lceil\log|S_{0}|\rceil+3-t$ (since $\lceil\log|S_{0}|\rceil+3$, are the total number of agents executing PerpExplore-F2F along $R$). Now in any iteration until the next modification of $S_{t}$, there is an agent that follows $SEQ3$ if $\lceil\log|S_{0}|\rceil+3-t\geq 3\implies t\leq\lceil\log|S_{0}|\rceil$.
Now we suppose, let $t>\lceil\log|S_{0}|\rceil$. Now there are two cases.
Case-I : $|S_{0}|=2^{p}$ (i.e., $l=0$). Then, $t>p$. Now by Lemma 4.6, $|S_{p}|\leq 3$. Also, number of alive agent is exactly 3 after the $p-$th modification. Now if $|S_{p}|$ is not less than or equal to 2, then the agents will again follow the sequences $SEQ1,SEQ2$ and $SEQ3$ until the next modification of $S_{p}$. Now , $|S_{p+1}|\leq 2$ ( Observation 4.1.1). Now this is a contradiction as $2<|S_{t}|\leq|S_{p+1}|\leq 2$.
Case-II : Now let us consider the case where $|S_{0}|=2^{p}+l$ where $02^{p}>l>$ 0. Then $t>p+1$. Now by Lemma 4.6, $|S_{p}|\leq 4$. Also, number of alive agent is exactly 4 after the $p-$th modification. Now if $|S_{p}|$ is not less than or equal to 2, then the agents will again follow the sequences $SEQ1,SEQ2$ and $SEQ3$ until the next modification of $S_{p}$. Now, $|S_{p+1}|\leq 3$ (by Observation 4.1.1). Again if $|S_{P+1}|$ is $\geq 2$ then the agents will again follow the sequences $SEQ1,SEQ2$ and $SEQ3$ until the next modification of $S_{p+1}$. Now again by Observation 4.1.1, $|S_{P+2}|\leq 2$. Now this again is a contradiction as $2<|S_{t}|\leq|S_{p+2}|\leq 2$.
Hence $t\leq\lceil\log|S_{0}|\rceil$. Thus, there must exist an agent following $SEQ3$ in any iteration after the $t-$th modification of suspicious region until the next one where $|S_{t}|>2$
From the above proof we can have the following corollary
Corollary 4.10.
If $|S_{t}|>2$ where $S_{t}$ is the suspicious region after $t$ modifications, then there is atleast 3 alive agents for all iterations after the $t$-th modification of the suspicious region and until the next modification.
Lemma 4.11.
Let $S_{t}$ be the suspicious region after $t$-th modification of $S_{0}$, where $S_{0}$ is the initial suspicious region. If $|S_{t}|=2$, then Algorithm 1, ensures that the number of agent which has not been consumed by the byzantine black hole is at least 2, for all iterations after the $t$-th modification of the suspicious region and until the next modification.
Proof 4.12.
As per our claim, $|S_{t}|=2$ after the $t$-th modification of $S_{0}$. This implies that after the $(t-1)$-th modification, we have either $|S_{t-1}|=3$ or $|S_{t-1}|=4$. Now we have the following cases.
Case-I: If $|S_{0}|=2^{p}$ and $|S_{t-1}|=4$, then by Lemma 4.6, we have $|S_{t-1}|\leq\lceil\frac{|S_{0}|}{2^{t-1}}\rceil+2\implies 2\leq 2^{p-t+1}\implies 1\leq p-t+1$, i.e., $t-1\leq p-1$. Now since in each iteration where the suspicious region is modified exactly one agent gets destroyed, this means after $t-1$ modifications of suspicious region, at most $p-1$ agents gets destroyed by the black hole. This implies out of $p+3$ agents executing Algorithm 1 on $S_{0}$, now at least $4$ agents are alive to explore the suspicious region $S_{t-1}$. Hence, in the worst case, exactly one agent among them can enter the black hole at the $t$-th (note that 2 agents cannot enter the black hole, because in that case $|S_{t}|$ will be 1, by Lemma 4.4) modification whereas the updated length of suspicious region $S_{t}$, i.e., $|S_{t}|=2$. Now for the case when $|S_{t-1}|=3$, by similar argument $t-1\leq p$, so at least 3 agents among $p+3$ are now alive to explore $S_{t-1}$. Again, by consumption of one agent in the worst case, the updated suspicious region length $|S_{t}|$ will be 2. So, in both the scenarios, we show that when $|S_{t}|=2$ at least 2 agents are alive.
Case-II: If $|S_{0}|=2^{p}+l<2^{p+1}$, where $0<l<2^{p}$ and $|S_{t-1}|=4$, in this case we have $p+4$ agents operating along $S_{0}$ (since, $\lceil\log|S_{0}|\rceil+3$ are the total number of agents) then again by Lemma 4.6, we have $|S_{t-1}|\leq\lceil\frac{|S_{0}|}{2^{t-1}}\rceil+2\implies 2<2^{p-t+2}\implies t-1<p$, so by earlier argument discussed in case-I, at least 5 (=$p+4-(p-1)$) agents are alive. So, in the worst case, exactly one agent gets consumed among 5 agents for the modification of $S_{t-1}$, which leads to the $t$-th modification, where $|S_{t}|=2$. On the contrary, if $|S_{t-1}|=3$, then similarly we have $11<2^{p-t+2}\implies t-1<p+$ 1, i.e., at least 4 agents are alive ($=p+4-p$ agents), so again in the worst case exactly one agent gets consumed among 4 agents for the modification of $S_{t-1}$, which leads to the $t$-th modification, where $|S_{t}|=2$. So, in this case as well, in both the scenarios when $|S_{t}|=2$, at least 2 agents are alive. This holds our claim.
Theorem 4.13.
PerpExplore-F2F solves the PerpExplore-BBH problem with $\lceil\log(n-1)\rceil+3$ agents having face to face communication and starting from the same node on a ring $R$ with $n$ nodes.
Proof 4.14.
We prove this theorem, by first ensuring that in any iteration of our Algorithm 1, every node is visited by at least one agent. Second, we show that, in any iteration there exists at least one alive agent.
Let $|S|$ be the length of the suspicious region and $k$ be the number of alive agents at the beginning of iteration $i$. Let $|S|=2$ and $k=2$, in this case we have, $S=\{S_{lft},S_{rgt}\}$. So, the two agents $a_{1}$ and $a_{2}$ (say) follows the sequence $SEQ0$, in which $a_{1}$ is instructed to visit $S_{lft}$ along clockwise direction, whereas $a_{2}$ is instructed to visit $S_{rgt}$ along counter-clockwise direction, while performing an iteration. This implies that, if $S_{lft}$ is at distance $l$ ($l>0$) in a clockwise direction, $a_{1}$ after traversing a distance of $l$ reaches $S_{lft}$, whereas $a_{2}$ traverses a distance of $n-l-1$ in counter-clockwise direction to visit $S_{rgt}$, in an iteration. This in turn shows that, the agents starting from $home$, visits each node of $R$ while performing an iteration. Note that, $|S|>2$ and $k=2$ never occurs by Corollary 4.10. So, the only possibility remains is that $|S|>2$ and $k>2$, in this situation, one among them say, $a_{1}$ executes the sequence $SEQ1$, another say, $a_{2}$ executes the sequence $SEQ2$, whereas the remaining agents does not leave $home$ while executing $SEQ3$, in an iteration. Observe that, $a_{1}$ traverses in a clockwise direction for $Dest$ distance, and reaches the destination node. Similarly, $a_{2}$ also reaches this same node, while traversing along a counter-clockwise direction for $n-Dest$ distance. This shows that, in any iteration, there exists two such agents, which explore the whole ring, hence exploration is guaranteed in each such iteration.
Next, we show that in every iteration, number of alive agents is at least 1. Let us consider during the $i$-th iteration, all the agents gets consumed by the byzantine black hole.
Note that, at the beginning of the $i-$th iteration, number of alive agent can not be 1. By contradiction, let us suppose there is only one agent alive at the beginning of $i-$th iteration which gets consumed by the byzantine black hole during this iteration. This implies $|S|\geq 2$ at the beginning of this iteration. Now by Lemma 4.11 and Corollary 4.10, there must exists at least two alive agents at the beginning of the $i-$th iteration which contradicts our assumption of one alive agent. Also observe that number of alive agent at the start of iteration $i$, can not be strictly greater than 2. Otherwise, let number of alive agents $k>2$ at the beginning of iteration $i$. Since during iteration $i$ all the $k$ agents are destroyed, this implies $|S|\geq 2$ Now since $|S|\geq 2$ and $k>2$, our algorithm ensures that, in the worst case during this iteration at most two agents can be consumed by the byzantine black hole. So after the iteration number of alive agents is at least $k-2\geq 1$. This contradicts our assumption that all agents are destroyed at the end of this iteration. Now the only remaining case is that exactly two agents are alive at the beginning of $i-$th iteration. Similarly as both of the agents are destroyed during the $i-$th iteration $|S|\geq 2$ at the beginning of this iteration. Note that $|S|$ will be exactly 2, for $k=2$, as by Corollary 4.10, we have if $|S|>2$ then $k$ must be at least 3. Now according to the algorithm, both the agents follows $SEQ0$, which instructs them to follow disjoint paths. Hence, this concludes that during this iteration, only can be consumed by the black hole, while there exists another agent which is alive even after the iteration, contradicting our assumption. Hence for all iterations number of alive agents must be greater or equals to 1. Thus, algorithm PerpExplore-F2F solves the PerpExplore-BBH problem with $\lceil\log(n-1)\rceil+3$ agents having face to face communication and starting from the same node on a ring $R$ with $n$ nodes.
4.2 Communication: Pebble
In this section, we consider the model, where each agent has an access to a movable token (termed as pebble) which can be carried by an agent from one node to another. This pebble acts as a mode of communication for the agents, as the agents can perceive the presence of a pebble at the current node. Moreover, an agent can also perceive the presence of other agents which are co-located at the current round (i.e., gather the Ids of the other co-located agents). Next, we discuss the idea of our algorithm PerpExplore-Coloc-Pbl (Algorithm 2) which solves the PerpExploration-BBH problem with 3 agents, each accompanying a pebble.
Description of PerpExplore-Coloc-Pbl: Given a ring $R$ with a byzantine black hole node $v_{b}$, a set of agents $A=\{a_{1},a_{2},a_{3}\}$ (where w.l.o.g ID of $a_{1}\leavevmode\nobreak\ <$ ID of $a_{2}\leavevmode\nobreak\ <$ ID of $a_{3}$), are initially co-located at a safe node, which we term as $home$. In this case as well, the agents have no knowledge about the position of the node $v_{b}$, the only set of knowledge the agents have are: the total number of nodes in the underlying topology $R$ (i.e., $n$), and also the knowledge that the initial node, i.e., $home$ is safe. So, the remaining arc of $n-1$ nodes in $R$ is a suspicious region (which we again term as $S$, where the cardinality of this arc is defined by $|S|$) for each of the three agents.
The idea of this algorithm is as follows: if all the three agents are at $home$ along with their pebble, then in the $r-$th round ($r>0$) $a_{1}$ moves clockwise with its pebble. In the next round, i.e., $(r+1)-$th round, $a_{2}$, leaving its pebble at $home$ moves to the next node along clockwise direction to meet $a_{1}$. In the subsequent round, i.e., $(r+2)-$th round, $a_{1}$ after meeting $a_{2}$ moves again to the next node in the clockwise direction. Note that $a_{1}$ moves to next node even if $a_{2}$ meets $a_{1}$ without pebble if and only if $a_{1}$ is not at $home$, and waits for 3 rounds (i.e., till $(r+5)-$th round). In $(r+3)-$th round, $a_{2}$ leaves its current node, moves one step in counter-clockwise direction, collects the left behind pebble, and in $(r+4)-$th round moves again to the earlier node in clockwise direction. Subsequently, in $(r+5)-$th round, $a_{2}$ again leaves the pebble at its current node, and moves in a clockwise direction, to meet $a_{1}$. This process continues, until each of $a_{1}$ and $a_{2}$ along with the two pebbles they carried, reaches $home$. This movement of $a_{1}$ and $a_{2}$ guarantees exploration unless at least one agent is destroyed by the black hole, while they are exploring the ring by following the above mentioned rules. In the meantime the third agent $a_{3}$ along with a pebble waits at $home$ for these remaining agents to return. Note that, if all the agents along with the pebbles carried by them, return back to $home$, then again they start the above mentioned execution from $home$, this guarantees that until an agent enters the black hole, perpetual exploration is performed.
Now, if any agent ($a_{1}$ or $a_{2}$ or both $a_{1}$ and $a_{2}$), enters the black hole, then our algorithm guarantees that either perpetual exploration is performed by the remaining alive agents or, the byzantine black hole will be detected within finite rounds by at least one alive agent, and it will continue performing perpetual exploration, avoiding the byzantine black hole, and independent of the position of other alive agents or pebbles.
In the following part we give a detailed description of our algorithm PerpExplore-Coloc-Pbl. Initially all the agents are in state Initial. In this state, an agent first declares its $Current-Node$ as $home$, after which initializes the variable $T_{time}=0$ (where, $T_{time}$ is the number of rounds passed since the agent has moved from state Initial) and gathers the ID of the remaining agents currently at $home$. Next, the agent with lowest ID, i.e., $a_{1}$ moves to state Leader, the agent with second lowest ID, i.e., $a_{2}$ moves to state Follower-Find, whereas the remaining agent, i.e., $a_{3}$ moves to state Backup. We now define an iteration for this algorithm. An iteration is defined to be a collection of $14n+$ 1 consecutive rounds starting from the latest round where all 3 agents along with 3 pebbles are at $home$ in state Initial. Note that, a new iteration fails to execute, if either at least one pebble or an agent gets destroyed by the black hole in the current iteration. More precisely, the meaning of failing an iteration implies that, an agent after ending its current iteration does not again start a new iteration by moving in to state Initial.
Suppose, all the agents along with the pebbles successfully execute the $i$-th iteration and reaches $home$, i.e., neither any pebble nor any agent gets destroyed by the black hole. So, according to our algorithm, $a_{1}$ being the lowest ID, follows the following sequence $SEQ1$, whereas $a_{2}$ being the second lowest ID follows $SEQ2$, and the $a_{3}$ follows the sequence $SEQ3$. The above sequences are as follows:
  1. 1.
    $SEQ1$: Initial $\rightarrow$ Leader $\rightarrow$ Initial.
  2. 2.
    $SEQ2$: Initial $\rightarrow$ $({Follower-SEQ})^{n}$rightarrow Initial.
    1. (a)
      ${Follower-SEQ}$: Follower-Find $\rightarrow$ Follower-Collect $\rightarrow$ Follower-Find.
  3. 3.
    $SEQ3$: Initial $\rightarrow$ Backup $\rightarrow$ Initial.
Execution of sequence $SEQ1$: $a_{1}$ performs this sequence, in which it first changes to state Leader, after finding it as the lowest ID agent at $home$ in state Initial. In state Leader, the agent performs the following checks: if the current node is $home$, then it checks that if the current node has the agent with second lowest ID, whereas the number of pebble at $home$ is 3 and $T_{time}\geq 2$. If all these conditions are satisfied, then the agent waits till $T_{time}=4n$ after which it moves to state Initial during $T_{time}=4n+1$. Otherwise, if it finds that it is with the second lowest ID agent and number of pebble at $home$ is also 3, but $T_{time}<2$, in this case the agent initializes $W_{time}=0$ (where $W_{time}$ is the waiting time of the agent) and then moves one hop along with the pebble to the next clockwise node. Otherwise, if the agent finds that neither it is with the second lowest ID agent nor the number of pebbles at $home$ is 3, then it waits for 5 rounds (i.e., waits when $W_{time}<6$), after which if still the above condition persists, then concludes $S_{lft}$ as the node which is at a clockwise distance of $n-2$ from $home$ and $S_{rgt}$ the node at a clockwise distance of $n-1$. Further, it updates $W_{time}=0$ and moves to the next state Detection.
On the contrary, if the current node of $a_{1}$ is not $home$, then also it performs the following checks: first, it checks whether it is with the second lowest ID agent (i.e., $a_{2}$ in this case), if so, then initializes $W_{time}=0$ and then moves one hop clockwise to the next node along with the pebble it is accompanying. Otherwise, if it is not with the second lowest ID agent, then waits for 3 rounds (i.e., waits when $W_{time}<4$), after which if still it is without the second lowest ID agent, then the agent moves to state Report-Leader, while leaving the pebble at the current node and moving one hop to the clockwise node.
Note that until and unless there is no anomalies detected by the agent $a_{1}$, it ends an iteration, by changing to state Initial from Leader. Otherwise, if any anomaly is detected, then $a_{1}$ ends the current iteration by either changing to state Detection or Report-Leader from the state Leader, in which case, it never changes to state either Leader or Initial.
Execution of sequence $SEQ2$: $a_{2}$ being the second lowest Id agent in state Initial, performs this sequence in an iteration. On successful completion of this iteration, an agent starting from Initial, performs the sub-sequence ${Follower-SEQ}$, $n$ times and then again changes to state Initial. The sub-sequence ${Follower-SEQ}$ symbolizes the sequence Follower-Find $\rightarrow$ Follower-Collect $\rightarrow$ Follower-Find, which an agent performs $n$ times until it again changes to state Initial (i.e., ${Follower-SEQ}$ is denoted by ${Follower-SEQ}^{n}$). While executing ${Follower-SEQ}$, the agent in state Follower-Find checks, whether the current node does not contain the lowest ID agent and the $Move$ parameter is also set to 0 (the parameter $Move$ is either 0 or 1, when 1 it symbolizes that the current node must contain the lowest ID agent, when 0, it means the current node is the adjacent counter clockwise node with respect to the node which contains the lowest ID agent), if so then it moves one hop clockwise leaving the pebble at the current node, while updating $Move$ to 1. Otherwise, if it finds $Move$ to be 1 whereas the current node does not contain the lowest ID agent, then it stops the current iteration, detects that current node is black hole and starts performing perpetual exploration, avoiding the black hole node. On the other hand, if the current node contains the lowest ID agent, then $a_{2}$ directly changes to state Follower-Collect. In state Follower-Collect, $a_{2}$ first moves one hop counter-clockwise, then in the current node either finds a pebble or not. If it finds a pebble, then moves one hop clockwise along with the pebble. Now, if $T_{time}=4n+1$ and current node is $home$, then stops performing $Follower-SEQ$ and moves to state Initial. Otherwise, updates $Move$ to 0 and changes to state Follower-Find. If the pebble is not found, then it concludes current node to be the black hole and starts performing perpetual exploration avoiding this node.
Note that, the agent only changes to state Initial only when it reaches $home$ again, which is after it performs ${Follower-SEQ}$ for $n$ times starting from $SEQ2$. If at any point, if $a_{2}$ finds any irregularities, based on whether a pebble is present or not, then only it directly concludes the black hole position. In all other cases, it follows the ${Follower-SEQ}$ sequence and changes to state Follower-Collect.
Exploration of sequence $SEQ3$: This sequence is only performed by the highest ID agent, which is $a_{3}$ in this case. In this sequence, after being in state Initial, it changes to state Backup, in which the agent first waits at $home$ till $T_{time}=4n$, after which it checks the following details, and accordingly either moves to state Initial if no irregularities are detected, otherwise moves to state Find-Pebble or Find-BH, based on the irregularity it detects.
  • The current node, i.e., $home$ has only the lowest and highest ID agents, i.e., $a_{1}$ and $a_{3}$, whereas the number of pebble at $home$ is 1. If this condition is satisfied, then the agent changes its state to Find-Pebble. Note that, this condition is satisfied only when $a_{2}$ fails to reach $home$ while performing this iteration, i.e., enters the black hole, while $a_{1}$ has detected some anomalies, which instigates it to leave the pebble at the current position at which it detects the anomaly, and henceforth moves back to $home$ while in state Report-Leader.
  • $home$ has only the highest and lowest ID agents, i.e., $a_{1}$ and $a_{3}$, but number of pebble at $home$ is 2, which implies that $a_{2}$ has failed to return to $home$ even when $T_{time}=4n$. In this case, $a_{3}$ concludes $S_{lft}$ to be the node which is at a clockwise distance of $n-2$ from $home$, whereas $S_{rgt}$ to be the node which is at a clockwise distance of $n-1$ from $home$. Finally, it updates $W_{time}=0$ and then moves to state Detection.
  • $home$ has all the three agents and moreover, number of pebbles present at $home$ is also 3. In this case, $a_{3}$ changes to state Initial, completing the iteration without detecting any anomalies.
  • After $4n$ rounds, there is no other agent except $a_{3}$ present at $home$. In this case, the agent move one hop, clockwise without the pebble present at $home$, and then changes to state Find-BH. This case arises only when both the agents $a_{1}$ and $a_{2}$ fail to reach $home$, while performing this iteration.
Now, next we define the states Report-Leader, Find-Pebble, Find-BH, Detection. An agent moves to either of these states only when it detects some anomalies.
The state Report-Leader, is executed by only the lowest ID agent, i.e., $a_{1}$, while it is in state Leader. The anomaly detected is as follows, $a_{1}$, after reaching a new node that is not $home$, along with the pebble it is carrying, waits for at most 3 rounds, for second lowest ID agent, i.e., $a_{2}$ to arrive (which is in state Follower-Find). If $a_{2}$ does not arrive even after the waiting, $a_{1}$ concludes that the agent (i.e., $a_{2}$) has entered the black hole, this triggers $a_{1}$ to move to state Report-Leader. In this state, it moves clockwise until it finds the highest ID agent, i.e., $a_{3}$ at $home$. Then waits until $T_{time}=4n$, after which changes state to Find-Pebble.
1 Input: $n,\leavevmode\nobreak\ k=3$;
2 States:{Initial, Leader, Follower-Find, Follower-Collect, Backup, Report-Leader,Find-Pebble, Detection, Find-BH}
3 In State Initial:
$T_{time}=0$ // $T_{time}$ is the number of rounds elapsed since the agent has moved from state Initial
4 Declare $Current-Node$ as $home$.
5 Gather the Ids of the remaining agents at $home$.
6 if lowest ID then
7       Move to state Leader.
8
9 else if second lowest ID then
10       Set $Move=0$.
11       Move to state Follower-Find.
12
13 else
14       Move to state Backup.
15
16In State Leader:
17 if $Current-Node=home$ then
18       if $With\leavevmode\nobreak\ Second\leavevmode\nobreak\ Lowest\leavevmode\nobreak\ Id\wedge\#Pebble=3\wedge T_{time}\geq 2$  then
19             Wait at the current node till $T_{time}=4n$.
20             Move to state Initial.
21
22       else if $With\leavevmode\nobreak\ Second\leavevmode\nobreak\ Lowest\leavevmode\nobreak\ Id\wedge\#Pebble=3\wedge T_{time}<2$ then
23             $W_{time}=0$.
24             Move one hop along with the pebble to the next node along clockwise direction.
25
26       else if $Not\leavevmode\nobreak\ with\leavevmode\nobreak\ Second\leavevmode\nobreak\ Lowest\leavevmode\nobreak\ Id\lor\#Pebble\neq 3$ then
27             if $W_{time}<6$ then
28                   $W_{time}=W_{time}+1$
29            else
30                   $S_{lft}=$ node at a clockwise distance of $n-2$ from $home$
31                   $S_{rgt}=$ node at a clockwise distance of $n-1$ from $home$
32                   Update $W_{time}=0$ then move to state Detection.
33
34
35
36else
37       if with second lowest ID then
38             $W_{time}=0$.
39             Move one hop along with the pebble to the next node along clockwise direction.
40
41      else
42             if $W_{time}<4$ then
43                   Stay at the current node, $W_{time}=W_{time}+1$.
44
45            else
46                   Move to state Report-Leader while moving one hop in clockwise direction along without pebble.
47
48
49
50In State Follower-Find:
51 if not with lowest ID $\wedge$ $Move=0$ then
52       Move one hop clockwise, without pebble and update $Move=1$.
53 else if not with lowest ID $\wedge$ $Move=1$ then
54       Detect the $Current-Node$ as black hole and continue exploring the ring perpetually avoiding this node.
55
56 else
57       Move to state Follower-Collect.
58
59In State Follower-Collect:
60 Move one hop in counter-clockwise direction.
61 if pebble is found then
62       Collect the pebble, and move one hop clockwise along with the pebble.
63       if $T_{time}=4n+1\wedge Current-Node=home$ then
64             Change to state Initial.
65
66      Update $Move=0$ and change to state Follower-Find.
67
68else
69       Detect the $Current-Node$ as black hole and continue perpetually exploring the ring avoiding this node.
Algorithm 2 PerpExplore-Coloc-Pbl
54 In State Report-Leader:
55 if $T_{time}<4n+1$ then
56       Move clockwise until finds the agent with highest ID.
57       Wait until $T_{time}=4n$.
58
59Change to state Find-Pebble.
60
61In State Backup:
62 Wait until $T_{time}=4n$.
63 if $Current-Node$ has only the lowest and highest ID agents and #$\#Pebble=1$ then
64       Move to state Find-Pebble.
65
66 else if $Current-Node$ has only the lowest and highest ID agents and #$\#Pebble=2$ then
67       $S_{lft}=$ node at a clockwise distance of $n-2$ from $home$
68       $S_{rgt}=$ node at a clockwise distance of $n-1$ from $home$
69       Update $W_{time}=0$ $home\_away=1$ and move to state Detection.
70
71 else if $Current-Node$ has both lowest and second lowest Id agent and #$\#Pebble=3$ then
72       Change to state Initial.
73
74 else if $Current-Node$ has no other agent then
75       Move in a clockwise direction without pebble and change to state Find-BH.
76
77
78In State Find-BH:
79
80if a pebble is found then
81       Conclude the next node along clockwise direction is the black hole, and continue perpetual exploration avoiding the black hole node.
82
83else
84       Move along a clockwise direction.
85
86
87In State Find-Pebble:
88 Move counter-clockwise without the pebble.
89 if a pebble is found then
90       Declare $Current-Node$ as $home$.
91       $S_{lft}=$ node at a clockwise distance of $n-2$ from $home$
92       $S_{rgt}=$ node at a clockwise distance of $n-1$ from $home$
93       Update $W_{time}=0$ and move to state Detection.
94
95else
96       Move along a counter-clockwise direction.
97
98In State Detection:
99
100if lowest ID then
101       Moves clockwise till the node at a distance $n-2$ from $home$ then returns back to $home$ along the same path and checks for the other agent.
102       If other agent is there it repeats the same procedure as stated in line 86 of Algorithm 2. Otherwise detects $S_{rgt}$ as the black hole and start exploring avoiding that node.
103else
104       Moves one hop counter-clockwise from $home$ then returns back to $home$ along the same path then waits for $62n-$ 6 rounds and checks for the other agent.
105       If other agent is there it repeats the same procedure as stated in line 90 of Algorithm 2. Otherwise detects $S_{lft}$ as the black hole and start exploring avoiding that node.
The state Find-Pebble, is performed together by only $a_{1}$ and $a_{3}$, i.e., the lowest and highest ID agent. This state can be reached by $a_{1}$ only via Report-Leader state, whereas by $a_{3}$ from the state Backup. The main idea for the agents in this state is to move counter-clockwise till they find the pebble left by $a_{1}$ when it detects some anomaly (i.e., $a_{2}$ has entered the byzantine black hole) and moved into state Report-Leader. This pebble acts as a marker which indicates that black hole is either the counter-clockwise neighbour of this node or, it is at a counter clockwise distance of 2 hop from it. This node at which it finds the pebble is declared to be the new $home$, from which they conclude $S_{lft}$ and $S_{rgt}$ to be the nodes which are at a clockwise distance of $n-2$ and $n-1$, respectively from this new $home$. Finally, they updates $W_{time}=0$ and then changes to state Detection. Note that, this state is performed by an agent during the last round of the current iteration, (i.e., when $T_{times}=4n+1$).
The state Find-BH, is only executed by the highest ID agent, i.e., $a_{3}$. This state is executed by $a_{3}$, only when it finds no other agent at $home$, at the last round of the current iteration. This situation can only occur, when both the agents $a_{1}$ and $a_{2}$ has entered the black hole. So, this triggers $a_{3}$ to change to state Find-BH, in which it moves along a clockwise direction, until a pebble is found. Whenever, a pebble is found, it concludes the next node to be the black hole, and starts performing perpetual exploration avoiding this node.
Finally, the state Detection, is performed by an agent (i.e., $a_{1}$ or $a_{3}$) only if they know the position of $S_{lft}$ and $S_{rgt}$, which are not only two consecutive nodes in a suspicious region $S$. Moreover, these two nodes are exactly at a distance of clockwise $n-1$ and $n-2$ distance from $home$ (the $home$ can be the initial starting node, or it can be the updated $home$ as well). In this state, the agent with lowest ID, i.e., $a_{1}$, moves clockwise till it reaches $S_{rgt}$, and then again returns back to $home$. After returning back, if it finds $a_{3}$, then again it performs the same procedure. On the contrary, if it does not find $a_{3}$, then it understands that $S_{rgt}$ is the black hole node, and hence continues perpetual exploration avoiding this node.
On the other hand, $a_{3}$ in this state moves counter-clockwise direction from $home$, i.e., reaches $S_{lft}$, after which it returns back to $home$ again following the same path and then waits for $62n-$ 6 rounds at $home$ and then checks for the other agent, i.e., $a_{1}$. If it finds $a_{1}$, then again continues to perform the same movement, otherwise it concludes that $S_{lft}$ is the black hole node, and hence starts the perpetual exploration of the ring just by avoiding this black hole node.
4.2.1 Correctness and Complexity
In this section we discuss the correctness and complexity of our algorithm PerpExplore-Coloc-Pbl.
Lemma 4.15.
If no agent gets consumed by the byzantine black hole, while performing an iteration of Algorithm 2, then our algorithm ensures that in that iteration, either the ring $R$ is explored or there is one agent that knows exactly the location of the black hole.
Proof 4.16.
Suppose till the completion of $i$-th iteration $(i>0)$ no agent is consumed by the black hole, then in any iteration $0<j\leq i$, the agent $a_{1}$ follows the sequence $SEQ1$, whereas $a_{2}$ follows the sequence $SEQ2$, as instructed starting from the initial $home$ node. Considering this scenario we have two cases:
  • No agents and no pebbles are destroyed: In this case, while executing the sequence $SEQ1$, the agent after changing its state to Leader from Initial, visits a new node after every 4 rounds in clockwise direction. Since, no agent has been consumed in this iteration as well, so $a_{1}$ un-obstructively visits each node in a clockwise direction, and reaches $home$ to finally end this sequence $SEQ1$, while again changing its state to Initial. This shows that, at least one agent visits each node of $R$ in an iteration completing the exploration of $R$.
  • No agent gets destroyed but some pebbles are consumed by the black hole: Note that, this situation can only occur when the black hole consumes the pebble carried and released by $a_{2}$, while $a_{2}$ is not at the node containing the pebble, executing the sequence $SEQ2$. It is because, in an iteration, only $a_{1}$ and $a_{2}$ moves away from $home$ while executing their respective sequences $SEQ1$ and $SEQ2$. And the pebble carried by $a_{1}$ ( i.e., the agent with smallest ID) cannot be destroyed without destroying $a_{1}$, as $a_{1}$ always carries the pebble at each node while executing its respective sequence. So, in this situation, while $a_{2}$ ( i.e., the agent with second lowest ID) in state Follower-Collect moves back to the adjacent node in counter-clockwise direction, in order to collect the already released pebble, absence of which triggers $a_{2}$ to conclude the current node to be the black hole, and which leads the agent to immediately leave the black hole node. In this case, the consumption of a pebble leads to the black hole detection.
{observation}
If no agent gets consumed by the byzantine black hole while performing an iteration of Algorithm 2 but some pebbles are consumed by the black hole, then by Lemma 4.15, one agent (more precisely, the agent with second lowest ID, i.e., $a_{2}$) detects the exact location of the black hole. Further, this agent continues to perpetually explore the ring $R$ by avoiding the black hole node.
Lemma 4.17.
If exactly one agent enters the byzantine black hole while performing an iteration of Algorithm 2, then within finite additional rounds any of the two following conditions hold:
  1. 1.
    The exact location of the black hole is detected by at least one agent.
  2. 2.
    All alive agents become colocated and they agree about two consecutive nodes $S_{lft}$ and $S_{rgt}$, one among which is the black hole.
Proof 4.18.
Note that the agent which has entered the black hole is either $a_{1}$ or $a_{2}$, based on which we have the following conditions:
  • $a_{1}$ falls in to the black hole: This situation can only occur when $a_{1}$ following sequence $SEQ1$ is in state Leader, visits the black hole node along with its pebble. Now, based on the adversarial choice of byzantine black hole, the pebble may or may not be destroyed by the black hole. Irrespective of which, $a_{2}$ within 3 additional rounds reaches the black hole node while in state Follower-Find, and absence of $a_{1}$ triggers the agent $a_{2}$ to determine the current node as black hole, and immediately leave the current node.
  • $a_{2}$ falls in to the black hole: This can only happen when $a_{2}$ is not with $a_{1}$ at the black hole node. Otherwise, both agents will be destroyed, contradicting our claim. Note that, $a_{2}$ can either be with the pebble or alone, while it is consumed by the black hole. Also note that, since $a_{2}$ is executing the sequence $SEQ2$, hence the distance between $a_{1}$ and $a_{2}$ can be at most 2. So during an iteration, if $a_{2}$ is destroyed by a black hole, then the black hole must be any one of the two consecutive counter-clockwise nodes, from the current position of $a_{1}$. If the current node is not $home$, in this case, $a_{1}$ gets to know about this fact, only when it finds the absence of $a_{2}$ even after waiting for 3 rounds, at the current node. Now, according to our algorithm, it leaves the pebble accompanied by it, and moves from the current node in clockwise direction to $home$ with state Report-Leader. After, reaching $home$, it waits until the end of this iteration, i.e., until $4n$ rounds from the start of this iteration and moves to state Find-Pebble. During the $14n+$ 1-th round from the start of the current iteration, $a_{3}$ finds that it is only with $a_{1}$, whereas only one pebble exits, this leads the agent to change its state to Find-Pebble. In this state, both the agents move counter-clockwise from $home$ until they encounter a pebble. Note that, this pebble is the one left by $a_{1}$ after determining the fact that $a_{2}$ has entered the black hole. So, now both these agents $a_{1}$ and $a_{3}$, declare the node with pebble as $home$, whereas also denote the adjacent counter-clockwise node from the current node as $S_{lft}$, whereas the other node adjacent to $S_{lft}$ in counter-clockwise direction as $S_{rgt}$. On the other hand, if the current node of $a_{1}$ is $home$, while $a_{2}$ is destroyed by the black hole, then $a_{2}$ can never reach $home$ along with the pebble it was carrying by following the sequence $SEQ2$. So, at the end of this iteration, both $a_{1}$ and $a_{3}$ finds that there are two agents (i.e., $a_{1}$ and $a_{3}$) and two pebbles, respectively at $home$. This leads both of them to conclude that, the adjacent counter-clockwise node to be $S_{rgt}$, whereas the adjacent counter-clockwise node of $S_{rgt}$ as $S_{lft}$.
So, in each case we showed that either of our two conditions hold.
Remark 4.19.
Following from Lemma 4.17, if exactly one agent gets destroyed by the byzantine black hole and the exact location of the black hole is detected by at least one agent, then our algorithm ensures that the ring is perpetually explored by at least one alive agent while avoiding the black hole node. On the contrary, all alive agents become co-located and they agree about two consecutive nodes $S_{lft}$ and $S_{rgt}$, among which one is a black hole, then until and unless another agent falls in to the black hole, our algorithm ensures that the ring is perpetually explored, by the remaining alive agents. Note that, by Lemma 4.20 while any one among them gets destroyed, then the other agent detects the exact black hole node, and then further perpetually explores the ring $R$, avoiding the black hole node.
Lemma 4.20.
Algorithm PerpExplore-Coloc-Pbl ensures that if two among three agents gets consumed by the byzantine black hole, then the remaining agent knows the exact location of the black hole within some additional finite rounds without being destroyed.
Proof 4.21.
Let us suppose, by contradiction, there exists a round $r$ ($>0$), within which two agents gets destroyed by the black hole, whereas the alive agent is unable to detect the exact location. Now we have the following cases:
  • Both agents gets consumed by the black hole in the same iteration: Note that in this case, the consumed agents are $a_{1}$ and $a_{2}$, respectively, which while executing the sequences $SEQ1$ and $SEQ2$ enters the black hole. Also note that, $a_{1}$ and $a_{2}$ can only be together at a node, when $a_{2}$ has left the pebble it is carrying in the counter-clockwise node. So, now $a_{3}$ after the end of this iteration, finds that no agent among $a_{1}$ and $a_{3}$ has reached $home$, which triggers $a_{3}$ to change to state Find-BH. In this state, it moves clockwise, until it finds a pebble (this pebble is the one left by $a_{2}$ in the counter-clockwise node of the black hole). Whenever a pebble is found, $a_{3}$ concludes that the next node is the black hole, which contradicts our claim.
  • There exists an iteration where exactly one agent is destroyed and the other one is destroyed during round $r$: Now in this case let $r^{\prime}<r$ be a round when exactly an agent is destroyed during an iteration, say $i>0$. Then by Lemma 4.17, within finitely many additional rounds $r_{0}$, where $r_{0}+r^{\prime}<r$, both the alive agents becomes colocated and agree about two consecutive nodes $S_{lft}$ and $S_{rgt}$, one among which must be the black hole. Observe that according to our algorithm, the node which is clockwise adjacent to $S_{rgt}$ is determined to be the $home$ (if not already so). Note that the other case according to Lemma  4.17 where after exactly one agent is destroyed by the black hole during an iteration at a round $r^{\prime}<r$, there is an agent that knows the exact location of the black hole within a finite additional rounds and the agent starts exploring the ring avoiding the black hole and so even after round $r$ it stays alive knowing the exact location of the black hole contradicting our assumption. Now, when both the agents are colocated during a round say $r_{0}+r^{\prime}<r$ and knows that the black hole must be any of the two consecutive nodes $S_{lft}$ or, $S_{rgt}$. In this case, they move in to state Detection. After which one agent moves clockwise until it reaches $S_{lft}$ and the other one moves counter clockwise until it reaches $S_{right}$ and moves back to $home$. The agent that visits $S_{rgt}$ waits for the other agent at $home$ for $62n-$ 6 rounds which is sufficient for the other agent to return back at $home$. Now when during round $r$, another agent among these two gets destroyed, within at most $(2n-4)$ rounds they fails to meet at $home$. Which trigers the remaining alive agent to determine exactly which one among $S_{lft}$ or, $S_{rgt}$ is the black hole. This also contradicts the claim we made earlier. Thus, The algorithm PerpExplore-Coloc-Pbl ensures that, if two agent among three agents gets consumed by the byzantine black hole, the remaining agent must know the exact location of the black hole in some finite additional rounds without getting destroyed.
Corollary 4.22.
Number of alive agents must be at least 2, when the exact position of the byzantine black hole is not known to them.
From Lemma 4.20 and algorithm PerpExplore-Coloc-Pbl we have the following theorem
Theorem 4.23.
A set of 3 synchronous colocated agents equipped with one pebble each are necessary and sufficient to solve PerpExploration-BBH on a ring $R$.
Remark 4.24 (Algorithm on whiteboard model of communication).
The PerpExplore-Coloc-Pbl algorithm (i.e., Algorithm 2 ) also works for colocated agents with whiteboard model of communication. The agents can easily simulate a pebble on a node by a bit marked on that nodes memory. Also, carrying a pebble from one node to another can be simulated by removing the marking from the current node’s memory and move to a new node where the agent writes a new bit. Thus 3 colocated agents are also sufficient to solve Perpexplore-BBH when the nodes are equipped with whiteboards.
5 Scattered Agents
5.1 Communication: Pebble
In this section, we discuss the model where the agents are placed arbitrarily along the nodes of the ring $R$ (note that each such node must be a safe node), where each agent has a movable token, which it can carry along with it, and acts as a mode of inter agent communication. Moreover, the agent can gather the IDs of other agents which are currently at the same node at the same round. With this context, in the following part we show that 4 scattered agents with a pebble each is sufficient to solve PerpExploration-BBH on $R$, using our algorithm PerpExplore-Scat-Pbl.
If 4 agents are scattered in more than one nodes initially then there are 3 cases. Firstly, all four agents are at different nodes. Secondly, Four agents are scattered in three different nodes nodes initially. For this case there exists exactly one node with two agents and remaining two nodes with exactly one agents each. In the third case, four agents are scattered in two nodes initially. In this case either each of the two nodes contains exactly two agents or, one node contains three agents and the other one has exactly one agent. We here describe and present the pseudocode considering the first case only. The case with 3 agents in one node can be dealt by instructing the colocated agents to execute PerpExplore-Coloc-Pbl when there are three agents on the current node (i.e., from the beginning). The algorithm we discuss here can also be used for the remaining cases with slight modifications. These modifications are described in the Remark 5.7
The main idea of the PerpExplore-Scat-Pbl algorithm is that 4 agents perpetually explore the ring, when no agents are consumed by the black hole. Our algorithm ensures that during this exploration at most one agent can be destroyed. In this scenario, the remaining agents within further finite time, gather at a single node (which is the starting node of either of these 4 agents, hence it is safe), and after which they together start executing the algorithm PerpExplore-Coloc-Pbl.
At the beginning of the Algorithm 3 all agents are in state Initial1. In this state, the agents declare their current node as $home$, moreover they initialize the variables $size=0$ and $s=0$, where $size$ variable is used by the agents in order to store the length of the path between it’s own $home$ and the nearest $home$ of another agent in the clockwise direction. Further the variable $s$ indicates the state transition. If the agent is at state Forward and $s=$ 0 then the agent moved from state Initial1 to its current state (i.e., state Forward), otherwise if the agent is at state Forward and $s=1$ then this indicates it changes to state Forward from state Wait2. An agent $a_{1}$ in state Forward, initializes $W_{time}=0$ (which stores the waiting time) and checks whether $s=0$ or $s=1$. Note that, if $s=0$ and the current node is $home$, then the $a_{1}$ is first instructed to leave the pebble at current node and then move clockwise direction, after increasing $size$ by 1. The next node where $a_{1}$ has reached can either be $home$ of another agent, say $a_{2}$, or it is not $home$ for any other agents. For the first case $a_{1}$ would find a pebble on its $Current-Node$ which was left by $a_{2}$. In this case $a_{1}$ updates $s$ to 1 and moves into state Wait1. On the other hand, if the $Current-Node$ is not $home$ of any other agents then $a_{1}$ finds no pebble at the $Current-Node$ and thus moves clockwise after increasing the variable $size$ by 1. Note that, when $s=0$ and an agent is in state Forward it always finds another pebble on some node, if it is not destroyed (this pebble must be left by another agent which is simultaneously executing Algorithm 3 and currently in state Forward). This node must be the clockwise nearest $home$ (i.e., $home$ of another agent which is clockwise nearest to own $home$). Now since during each move at state Forward, an agent increases the variable $size$ by 1, the $size$ variable must store the length between an agents own $home$ and the clockwise nearest $home$ when the agent finds another pebble. Beyond this point (i.e., when $s=1$) the agents will use the $size$ variable as reference to calculate how much it should move, as after this, as there is a possibility that the agent may not find a pebble at the clockwise nearest $home$.
On the contrary, if $s=1$, this implies the agent already knows how much distance it should travel from its own $home$ in order to reach the clockwise nearest $home$. So, now if the current node is $home$ it just moves clockwise by initializing the variable $distance$ to 1 (where the variable $distance$ stores the amount of distance the agent has traversed from its own $home$). Otherwise, if current node is not $home$, then the agent continues to move clockwise while updating $distance$ by 1, until it reaches the clockwise nearest $home$ (i.e., when $distance=size$), after which it further changes to state Wait1.
In state Wait1, the agent is instructed to wait at the current node (more specifically, the current node is the clockwise nearest $home$ of the agent) for $n-size-1$ rounds. After which it changes to state Fetch. Note that, if each agent starts executing the state Forward at the same round, this implies that these agents change to state Fetch together at the same round (which is precisely at the $n+1$-th round since the start of state Forward). Thus we have the following observation. {observation} If all agents moved into the Forward state together at a round $t$ then, all of them move to state Fetch together at the round $t+n+1$.
In state Fetch, the agent first checks if the current node has a pebble, if so then it further checks whether current node is $home$ or not. If the current node is not $home$, then it moves counter-clockwise with pebble, until it reaches $home$. Otherwise, if current node is $home$ then it checks whether there are more than one pebble present at the current node or not. If so, then the agent changes its state to Gather1. Otherwise, if the current node does not have more than one pebble, then it initializes $W_{time}$ to 0 and changes to state Wait2.
In state Wait2, the agent first checks whether $W_{time}<2n-size$, if so then it further checks whether #$\#Agent$ at current node is exactly 1. If these two conditions are satisfied then the agent just increments $W_{time}$ by 1. On the other hand, if it finds that $W_{time}<2n-size$ but #$\#Agent$ at the current node is 2, then it changes its state to Gather2 and updates $W_{time}$ to 0. Otherwise, if it finds $W_{time}<2n-size$ but #$\#Agent$ at the current node is 3 then it updates $W_{time}$ to 0, whereas changes its state to Coloc. Lastly, whenever it finds $W_{time}=2n-size$, it changes its state to Forward.
Note that if no agent gets destroyed by the black hole, then our algorithm guarantees that an agent can neither changes its state to Gather1 or in state Gather2. It is because, an agent, say $a_{1}$ can only find more than one pebble at its $home$ node, only when another agent say $a_{2}$ (exploring the counter-clockwise nodes of $home$ of $a_{1}$), which was supposed to fetch the pebble left by $a_{1}$ at its $home$, is destroyed. In this case, whenever $a_{1}$ returns while in state fetch with another pebble, it finds more than one pebble. So the first agent, which first finds this anomaly changes its state to Gather1 from the state Fetch. Now this agent must initiate gathering with the remaining two agents (as it is the one which has first detected the anomaly of more than one pebble at its $home$). Note that, if while $a_{1}$ starts its gathering phase, moving in a clockwise direction, there is a possibility that other agents are still moving either in clockwise or counter-clockwise direction. So, to address this challenge, the algorithm instructs $a_{1}$ (i.e., the agent which is in state Gather1) to wait at the current node (i.e., its $home$) for precisely $n-size-1$ rounds, and then start moving clockwise direction. This waiting time of $a_{1}$ guarantees all the other alive agents to reach their respective $home$ while in state Fetch. Now, in order to avoid further movement of remaining agents (i.e., again change to state Forward, since these agents are yet to detect any anomaly, which makes gathering a bit challenging), our algorithm instructs the remaining agents to wait at the current node for a certain number of rounds (i.e., more precisely, $12n-size-$ 1). This waiting period is sufficient, for the agent $a_{1}$ to meet with the other alive agents while in state Gather1. Note that, if no agent detects any anomaly, then they directly change their state to Wait2 from the state Fetch. If no such agent exists which encounters any anomaly and changes its state to Gather1 or Gather2 then this waiting time of $12n-size-$ 1 rounds, also ensures that all of them must move to state Forward at the exact same round (i.e, ($23n+$ 2)-th round from the previous time it moved to state Forward) [^1] At $(n+1)$-th round an agent moves to state Fetch after that it moves for $size$ rounds to reach its $home$ in counter-clockwise direction and changes to state Wait2. So at round $n+3+size$ an agent changes to state Wait2. After that waits for further $12n-size-$ 1 rounds and changes to state forward, taking a total of $23n+$ 2 rounds . Now we have the following observation {observation} Let $t$ be a round when all alive agents moved into state Forward. If no agent detects any anomaly (i.e., no agent finds more than one pebble at home) then, at round $t+3n+2$ all the agents together move into state Forward again.
An agent in state Gather1, is the first agent to detect the anomaly, hence it is first instructed to wait at the current node, i.e., its $home$ for the first $n-size-1$ rounds, and then start moving in a clockwise direction, while accompanying all the pebbles at the current node  [^2] Note that we have considered that in this case an agent, while in state Gather1 or Gather2 can carry more than one pebble, which can be easily restricted to one agent can carry only one pebble, but in that case the agent needs to return and carry one pebble each time and reach back to its current node. This movement in the clockwise direction continues until the number of agent at the current node is exactly 3, after which it changes to state Coloc.
An agent in state Gather2 (i.e., the alive agents which are not the first to detect the anomaly), first checks whether the current node has exactly 3 agents on it, if not then it moves in a counter-clockwise direction, while accompanying the pebbles present at its $home$,
1 Input: $n$
2 State:{Initial1, Forward, Fetch, Wait1, Wait2, Coloc, Gather}
3 In State Initial1:
4 Declare $Current-Node$ as $home$.
5 Initialize $size=0$ and $s=0$, then change to state Forward.
6 In State Forward:
7 Initialize $W_{time}=0$.
8 if $s=0$ then
9
10      if $Current-Node$ is not $home$ and has a pebble then
11             update $s=1$ and Change state to Wait1
12
13       else if $Current-Node$ is $home$ then
14             Leave the pebble at $home$.
15             Move in a clockwise direction and update $size=size+1$.
16
17       else
18             Move in a clockwise direction and update $size=size+1$.
19
20
21else
22       if $Current-Node$ is $home$ then
23             Initialize $distance=1$ and move in a clockwise direction leaving the pebble.
24
25       else if $distance<size$ then
26             Move in a clockwise direction and update $distance=distance+1$.
27
28       else if $distance=size$ then
29             Change State to Wait1.
30
31
32In State Wait1:
33 if $W_{time}<n-size$ then
34       $W_{time}=W_{time}+1$.
35
36else
37       Change state to Fetch.
38
39In State Fetch:
40 if $Current-Node$ has a pebble then
41       if $Current-Node$ is not $home$ then
42             Move counter-clockwise with the pebble.
43
44      else
45             if $home$ has more than one pebble then
46                   Change to state Gather1.
47
48            else
49                   Initialize $W_{time}=0$ and change to state Wait2.
50
51
52
53else
54       if $Current-Node$ is not $home$ then
55             Move in a counter-clockwise direction.
56
57      else
58             Initialize $W_{time}=0$ and change to state Wait2.
59
60
61In State Wait2:
62 if $W_{time}<2n-size\wedge\#Agent$ at $Current-Node$ is 1 then
63       $W_{time}=W_{time}+1$.
64
65 else if $W_{time}<2n-size\wedge\#Agent$ at $Current-Node$ is 2 then
66       Initialize $W_{time}=0$ and change to state Gather2.
67
68 else if $W_{time}<2n-size\wedge\#Agent$ at $Current-Node$ is 3 then
69       Initialize $W_{time}=0$ and change to state Coloc.
70
71 else if $W_{time}=2n-size$ then
72       Change to state Forward.
73
74In State Gather1:
75 if $W_{time}<n-size$ then
76       $W_{time}=W_{time}+1$
77else
78       if #$\#Agent$ at $Current-Node<$3 then
79             Move in a clockwise direction with all the pebbles at the $Current-Node$
80
81      else
82             Change to state Coloc.
83
84In State Gather2:
85 if #$\#Agent$ at $Current-Node$ < 3 then
86       Move in a clockwise direction with all the pebbles at the $Current-Node$.
87
88else
89       Change to state Coloc.
90
91In State Coloc:
92 Declare $Current-Node$ as $home$.
Change state to Initial and execute PerpExplore-Coloc-Pbl.// State Initial is the state defined in PerpExplore-Coloc-Pbl algorithm
93
Algorithm 3 PerpExplore-Scat-Pbl
until it satisfies the condition, where it finds a node on which a total of 3 agents exists. After which they change their state to Coloc.
In state Coloc, an agent declares its current node as $home$, while also changes its state to Initial  [^3] The state Initial is defined in Algorithm 2 and then executes algorithm PerpExplore-Coloc-Pbl.
5.1.1 Correctness and Complexity
In this section, we discuss the correctness and complexity of our algorithm PerpExplore-Scat-Pbl.
Lemma 5.1.
PerpExplore-Scat-Pbl guarantees to solve PerpExploration-BBh problem if no agent is consumed by the byzantine black hole.
Proof 5.2.
Let us define a segment for an agent $a_{1}$ be $Seg(a_{1})$, where $Seg(a_{1})$ is defined to be the collection of consecutive nodes in the clockwise direction, starting from its own $home$ and its clockwise nearest $home$ (which is the $home$ of another agent say, $a_{2}$). Note that, since no agent is consumed by the black hole, any agent, say $a_{1}$, while executing the Algorithm 3, starts moving in clockwise direction in state Forward, reaches the nearest clockwise $home$, i.e., the last node of this segment $Seg(a_{1})$, then changes its state to Wait1, and after waiting for $n-size$ rounds, again changes its state to Fetch. In state Fetch, the agent basically moves counter-clockwise, until and unless it reaches its own $home$, which is nothing but the first node of this segment $Seg(a_{1})$. Also, note that, each agent, changes their state to Forward and Fetch together at the same round, while exploring their respective segments. Now, note that, the union of all these segments is basically the whole ring $R$ (i.e., more specifically $\cup^{4}_{i=1}Seg(a_{i})=R$) . Hence, this guarantees that the ring is perpetually explored if no agent is consumed by the black hole.
Lemma 5.3.
If an agent is consumed by the byzantine black hole, then within an additional finite rounds, there exists exactly one agent, which changes its state to Gather1.
Proof 5.4.
Let $a_{1}$ be an agent exploring $Seg(a_{1})$. Without loss of generality, let the black hole be in $Seg(a_{1})$. Also let $a_{2}$ be another agent which is exploring the segment $Seg(a_{2})$, where $home$ of $a_{2}$ is the common node between $Seg(a_{1})$ and $Seg(a_{2})$. Also, let $a_{3}$ be an agent which is exploring the segment $Seg(a_{3})$, where the $home$ of $a_{3}$ is the common node between $Seg(a_{2})$ and $Seg(a_{3})$. Note that, $a_{1}$ can be consumed by the black hole, either in state Forward or in state Fetch.
  • If $a_{1}$ is consumed while in state Forward, then it fails to bring back the pebble from the $home$ of $a_{2}$, while it is scheduled to change its state to Fetch, after ending the states Forward and Wait1. Note that, since we have assumed that the black hole is in $Seg(a_{1})$, and $Seg(a_{1})\cap Seg(a_{2})=\{home\text{ of }a_{2}$}. Thus, $a_{2}$ must bring a pebble from the $home$ of $a_{3}$ to it’s own $home$, after it changes its state to Fetch. So, now as $a_{2}$ reaches its own $home$ along with the pebble it carried from the $home$ of $a_{3}$, it finds that there already exists another pebble, which the agent $a_{1}$ is unable to carry back to its own $home$, as it is already consumed by the black hole. This causes an anomaly for $a_{2}$ at its own $home$ (i.e., encountering more than one pebble). After detecting the anomaly $a_{2}$ immediately changes its state to Gather1. Note that, after $a_{1}$ is consumed by the black hole, it has taken an additional at most $size-1+(n-size)+1+size+1=n+size+1$ 333calculates the number of rounds it took $a_{2}$ to change to state Gather1, where at most $size-1$ rounds required in state Forward after $a_{1}$ gets consumed by the black hole, $(n-size)$ rounds for changing to state Wait1 and waiting, 1 round for changing from Wait1 to Fetch, then $size$ rounds for moving in state Fetch and returning back to its $home$, and lastly 1 round for changing to state Gather1 from Fetch. rounds to change the state of $a_{2}$ to Gather1. Further in this case we show that no other agents except $a_{2}$ changes to state Gather1. Let $t^{\prime}$ be the round when $a_{2}$ changes its state to Gather 1. And $t<t^{\prime}$ be the last round before $t^{\prime}$ when all the agents moved into state Forward. Then $t^{\prime}=t+n+1+size+1$. Next $a_{2}$ waits for $n-size-1$ rounds that is until round $t+2n+1$ and starts moving counter clockwise being in state Gather1. Note that by observation 1, the other agents stay at their home until round $t+3n+1$. So from $t+2n+2$ until $t+3n+1$, these $n-1$ rounds are enough for $a_{2}$ to meet the other two while they are still being in the state Wait2. So they change their state to Gather2 and eventually to Coloc. As an agent can not change state to Gather1 from state Coloc or Gather2, no other agent changes to state Gather1.
  • If $a_{1}$ is consumed while in state Fetch. In this case, $a_{2}$ after returning back to its $home$ carrying the pebble from the nearest clockwise $home$, i.e., the $home$ of $a_{3}$, finds no anomaly. It is because, $a_{1}$ has been consumed by the black hole while in state Fetch, so that means, the pebble at $home$ of $a_{2}$ has been carried by $a_{1}$. Since, $a_{2}$ does not encounter any anomaly, hence it drops the pebble at its $home$, changes to state Wait2, waits at $home$ for $12n-size-$ 1 rounds and then changes to state Forward again. So, while executing the same execution, after reaching the clockwise nearest $home$, i.e., $home$ of $a_{3}$, it again changes its state to Wait1, waits for $n-size-1$ rounds, and further changes its state to Fetch. Note that, here at the $home$ of $a_{3}$, there exists a pebble which was earlier dropped by $a_{3}$, while it changed its state from Fetch to Wait2. So, $a_{2}$ carries this pebble and reaches its own $home$, where it finds the anomaly. It is because, as $a_{1}$ has been consumed by the black hole, it is unable to carry the pebble dropped by $a_{2}$ during its last state change from Fetch to Wait2. This phenomenon triggers $a_{2}$ to change its state to Gather1. Note that, since $a_{1}$ has been consumed by the black hole, it takes an additional $(size-1)+1+(2n-size)+size+1+(n-size)+size+1=3n+size+2$ 222calculates the number of rounds it took $a_{2}$ to change to state Gather1, where $size-1$ rounds required in state Fetch after $a_{1}$ gets consumed by the black hole, 1 round for changing its state from Fetch to Wait2, after this the agent waits for $12n-size-$ 1 rounds and in one additional round changes to state Forward, in this state it takes $size$ rounds, 1 round required to change to state Wait1, then $n-size-1$ rounds the agent waits, whereas 1 more round required to change to state Fetch, finally $size$ rounds in state Fetch, and lastly 1 round for changing its state to Gather1rounds to change $a_{2}$’s state to Gather1. And as shown for the first case it can be similarly showed that except $a_{2}$ no other agent changes state to Gather1
Lemma 5.5.
After the first agent changes its state to Gather1, it takes an additional at most $22n-$ 2 rounds, for the remaining agents to gather at a single node (i.e., the nearest counter-clockwise $home$ of the black hole node) along with their pebbles.
Proof 5.6.
Let $a_{2}$ be the agent which changes to state Gather1 at $t$-th round (where $t>0$), due to the fact that it encounters an anomaly at its $home$ because $a_{1}$ (w.l.o.g) has been consumed by the black hole, and the length of the segment it explores be denoted by $size$ (i.e., length of $Seg(a_{2})=size$). Also, let $t_{0}$ ($0<t_{0}<t$) be the last round when each agent changed their states to Forward. Note that $t$ should be exactly equal to $t_{0}+size+1+(n-size-1)+1+size+1=t_{0}+n+size+2$ 333movement in state Forward for $size$ rounds, then one round for change to state Wait1, then stay in state Wait1 for $n-size-1$ rounds, after which change to state Fetch for one round, and move in state Fetch for $size$ rounds, and finally change to state Gather1 for one round. Note that $a_{2}$ after changing its state to Gather1 (current round is $t_{0}+n+size+2$, which is also the $t-$th round according to our definition of $t$), further waits for $n-size-1(\leq n-2$, as $size\geq 1$ ) rounds (current round $t_{0}+2n+1$), and starts moving clockwise from $(t_{0}+2n+2)-$th round. On the other hand, an alive agent which was in state Fetch at $t-$th round, must have changed its own state to Fetch at $(t_{0}+n+1)-$th round (refer Observation 5.1). From Fetch, it takes further at most $n$ (since length of a segment can be at most $n-1$) rounds to change to state Wait2. Note that any other alive agent except $a_{2}$ must change to state Wait2 after completion of state Fetch as by Lemma 5.3, exactly one agent (here $a_{2}$) already changed its state to Gather1 from Fetch at round $t=t_{0}+n+size+2$). So, any agent except $a_{2}$ at the round $(t_{0}+2n+1)-$th round must be in state Wait2, and should remain in this state until $(t_{0}+3n+2)-$th round, if no anomaly is detected (refer Observation 1) or meets with other agents. So, any alive agent except $a_{2}$ remains at their $home$ within $t_{0}+2n+1$ to $t_{0}+3n+2$ rounds. This is by Lemma 5.3, only $a_{2}$ being the first agent to detect anomaly can change its state to Gather1, whereas the remaining agent can only change its state to Gather2 or Coloc, when they meet with $a_{2}$, while in state Wait2. Further, note that an alive agent except $a_{2}$, waits at their respective $home$ for $n+1$ rounds since, $a_{2}$ starts moving clockwise being in state Gather1 (after $(t_{0}+2n+1)-$th round). Now, when $a_{2}$ reaches its clockwise neighbor (say) $a_{3}$, at its $home$. $a_{3}$ must be at state Wait2, in this case, $a_{3}$ changes its state to Gather2, and starts moving clockwise along with $a_{2}$ carrying their respective pebbles, until they meet another agent, (say) $a_{4}$, at its $home$, while in state Wait2. After which they all change to state Coloc, which also guarantees that they have at least 3 pebbles at the current node (since $a_{2}$ carried at least 2 pebbles from its own $home$ where it detects the anomaly, $a_{3}$ also carried its one pebble present at its own $home$). Now since no agent can change to state Gather1 from Gather2 or Coloc, $a_{2}$ is the only agent that changes to state Gather1. Note that, before gathering $a_{2}$, $a_{3}$ and $a_{4}$ can never be destroyed, as they traverse only the nodes in $\cup^{4}_{i=2}Seg(a_{i})$, whereas the black hole is in $Seg(a_{1})\backslash\{home$ of $a_{1}$, $home$ of $a_{2}\}$.
Remark 5.7.
In the discussion of our Algorithm 3, we have considered that 4 agents are scattered along 4 distinct nodes (i.e., each node with multiplicity 1). Here we describe how our algorithm (Algorithm 3) works for the remaining cases, (i.e., when 4 agents are scattered among 3 or 2 nodes initially) by slight modification. Observe that, in these remaining cases there exists at least one node, where multiplicity is greater than 1 initially. Let there exists a node with multiplicity $3$, in this case, these agents directly start executing PerpExplore-Coloc-Pbl, while execution if they encounter the fourth agent somewhere along $R$, they just ignore this agent (note that the IDs, of all the 3 initially co-located agents are already collected by each other) while executing their current algorithm. On the other hand, if initially there exists a node with multiplicity greater than 1 and less than 3, then in that case, only the lowest ID agent (say $a_{1}$) at the current node, changes its state Forward, whereas the other agent at the current node (say $a_{2}$) changes its state Backup-Wait at some round say $t_{0}$ $(>0)$. The agent in state Forward, continue executing the algorithm PerpExplore-Scat-Pbl, whereas the agent in state Backup-Wait, waits at the current node for $12n+$ 1 rounds (i.e., until $t_{0}+2n+1$-th round) and then checks for anomaly in the next round (i.e., if the current node has more pebble than current number of agents). If such anomaly exists after $t_{0}+2n+1$ round, then that implies that $a_{1}$ is already in state Gather1. It is because, $a_{1}$ while it returned back to its $home$ carrying a pebble in state Fetch at round $t_{0}+n+size+2$ round (where $size$ is the length of $Seg(a_{1})$), encounters the first anomaly, i.e., the number of pebble is more than number agents. In that case, it directly changes its state to Gather1, and further waits at $home$ for $n-size-1$ rounds, i.e., till $t_{0}+2n+1$ rounds. After which it starts to move clockwise, and on the other hand $a_{2}$ also starts moving clockwise, while changing its state to Gather2 from Backup-Wait. This guarantees that if the anomaly is detected at a multiplicity then both $a_{1}$ and $a_{2}$ moves together to gather, along with the third agent which is in state Wait2 (as described in Lemma 5.5) and changes its state to Gather2 whenever it finds #$\#Agent$ at its $home$ is more than one. Now if the anomaly is not detected at the multiplicity, that is $a_{2}$ at round $t_{0}+2n+1$ finds no anomaly then it moves into state Backup whereas $a_{1}$ is in state Wait2 at the same node. Both of them waits until round $13n+$ 1 and at round $23n+$ 2, $a_{1}$ changes its state to Forward again (refer to Observation 1) and $a_{2}$ changes its state to Backup-Wait again. If some other agent detects anomaly it must meet $a_{1}$ and $a_{2}$ at their $home$ at some round $t$ where $t_{0}+2n+2\leq t\leq t_{0}+3n+1$. In that case they find that there are 3 agents at their $home$ and all of them change to state Coloc and executes algorithm PerpExplore-Coloc-Pbl.
Next we can have the following theorem
Theorem 5.8.
A set of 4 synchronous scattered agents with one pebble each is necessary and sufficient to solve the PerpExploration-BBH problem on a ring $R$, when the initial starting locations of the agents are different along $R$.
5.2 Communication: Whiteboard
In this section, we discuss the algorithm PerpExplore-Scat-WhitBrd, which achieves PerpExploration-BBH, with the help of three agents on a ring $R$ with $n$ nodes, where each node has a whiteboard that can store $O(\log n)$ bits of data. Note that, in this model as well we have considered that the agents are placed arbitrarily placed along the nodes of the ring $R$ (each such node is a ‘safe node’). So, there can be two cases, first, all the three agents are initially placed at three different positions, second, two agents are together whereas the third agent is in a different position. We will provide algorithm for the first case only, as an algorithm for the second case can be easily designed by modifying and merging the algorithms for the first case and algorithm PerpExplore-Coloc-Pbl.
The idea of our algorithm PerpExplore-Scat-WhitBrd is as follows: the three agents separately explore three segments of the ring $R$, such that, their union is the whole ring $R$ and intersection of any two segment contains only a safe node (which is a $home$ for one agent). Thus, while exploring, only one agent can be consumed by the black hole, and in an additional finite number of rounds another agent can learn,the consumption of the earlier agent by the black hole (this agent is precisely the agent whose segment has a safe node common with the segment of the consumed agent. Also it learns about this consumption from the whiteboard of the common node). In which case, this agent gathers with the other alive agent, and locates the exact segment in which the earlier agent was exploring before its consumption by the black hole. After locating this segment, both these agents starts performing Cautious-Walk either in clockwise or counter-clockwise direction, based on the information they gained about the earlier agents movement direction before getting consumed by the black hole. The agents before getting consumed, marks each node of their corresponding segment with different markings (i.e., more precisely writes either left or right, after erasing the earlier direction marking at the current node) based on the direction of their movement. So, the two agents performing Cautious-Walk, search for the first node, which does not have the direction marking w.r.to the direction the earlier consumed agent was moving. This node which is devoid of this direction marking is indeed the black hole node, and at least one agent, among the agents performing Cautious-Walk detects it, and remains alive, which further perpetually explores the ring $R$, avoiding this node.
Let us consider the three agents (say, $a_{1}$, $a_{2}$ and $a_{3}$) are initially placed at three nodes of the ring $R$, which are not only safe nodes but are also recognised as the $home$ of these agents. Initially, the agents starting from their respective $home$ are assigned the task to explore a set of nodes, which we term as a segment of the corresponding agents. More precisely, a segment for an agent $a_{i}$ is defined as the set of consecutive nodes, starting from its $home$ and ending at the nearest clockwise $home$ (i.e., the $home$ of first clockwise placed agent), which is also termed as $Seg(a_{i})$. Note that $\cup^{3}_{i=1}Seg(a_{i})=R$. So if none of the agents are ever destroyed the ring will still be perpetually explored.
Let us first discuss the case, in which we describe the possible movements of the agent and the respective state changes they perform, until one agent gets consumed by the black hole, and another agent gets to know that an agent is already consumed by the black hole. After which, we will describe all the possible movements and state changes performed by the remaining two alive agent, between getting to know that already one has been consumed by the black hole, and finally detecting the black hole position.
Initially all agents start from state Initial at their respective $home$. In state Initial, an agent ( without loss of generality, say $a_{1}$) first clears the already present data (if at all) at the whiteboards of their respective $home$, then initializes $T_{time}=0$ ($T_{time}$ stores the number of rounds elapsed since the start of state Initial), and writes a message of the form (home, ID) at its $home$, where ID is the ID of the agent. This type of message is termed as “home" type message, which consists of two components, first component stores the message home and second component stores the ID of the agent writing, i.e., ID of the agent whose $home$ is the $Current-Node$. Further, it changes its state to Forward, and moves in clockwise direction.
In state Forward, the agent moves in clockwise direction while erasing the earlier direction marking (if exists), i.e., left and then writes the new direction marking, i.e., right in each node. The agent also increases the $T_{time}$ variable by one in each round. This process continues until the agent encounters a node which has a “home" type message. This “home" type message signifies that the agent has reached the end of segment $Seg(a_{1})$, i.e., in other words it has reached the nearest clockwise $home$, say $v_{c}$. Note that the length of a segment can be at most $n-2$, hence within $T_{time}=n-1$, an agent $a_{1}$ is bound to reach the last node of its own segment i.e., $v_{c}$. In any case irrespective of the current $T_{time}$, the agent waits at $v_{c}$ until $T_{time}=n-1$, after which in the next round, it checks for the following information at $v_{c}$. If the agent finds a message of type “visited" at $v_{c}$, the agent considers this as an anomaly and learns that an agent of which $v_{c}$ is the home (from the ID component of home type message at $v_{c}$) must have entered black hole while returning back from its clockwise nearest $home$ (refer Lemma 5.9). Then in this case, the agent stores the message of type dir in its local memory, where dir=(Counter-clockwise, NULL). In a dir type message, the first component is called a direction component which indicates the direction of the agent which gets consumed by the black hole, along which it was moving just before it gets consumed. On the other hand, the second component either stores ID of some agent or it stores $NULL$ message. Both these components are useful for certain state transitions. Next, the agent changes its state to Gather. Otherwise, if no anomaly is detected, then the agent simply writes the message (Visited, self ID) at $v_{c}$ and changes its state to Back-Wait.
Next, in state Back-Wait, an agent (here $a_{1}$) waits at $v_{c}$ until $T_{time}=2n-1$. In this waiting time $a_{1}$ waits for other agent to meet $a_{1}$ if the other agent detects any anomaly during its state Forward. While waiting, if it finds that the #$\#Agent$ at $v_{c}$ is more than 1, and the whiteboard at $v_{c}$ has a message of type dir, then the agent performs the following task. After noticing this, the agent moves along the counter clockwise direction, while changing its state to Gather2. Note that, our algorithm ensures that in state Back-Wait, if a dir type message is seen by an agent, then the direction component of this message must be in a counter-clockwise direction (refer Lemma 5.15). Otherwise, if no such dir type message is seen by $a_{1}$, and also $T_{time}=2n$, then in this round, the agent changes its state to Backtrack.
In state Backtrack, an agent (here $a_{1}$) starts to move in a counter-clockwise direction from $v_{c}$, and after each move erases the earlier direction, i.e., right, and writes the new direction, i.e., left, and also increments $T_{time}$ by 1. This process continues until it reaches its own $home$, i.e., reads a home type message with ID same to its own ID. Again note that, since a segment can be of length at most $n-2$, hence within $T_{time}=3n-1$, the agent reaches its own $home$. After which it does not move until $T_{time}=3n$. At $T_{time}=3n$, it checks whether its $home$ has visited type message, if so then directly changes its state to Initial-Wait. Otherwise, absence of such visited type message, creates an anomaly for the agent. This only happens when another agent, say $a_{j}$, ($j\neq 1$), for which $Seg(a_{1})\cap Seg(a_{j})=$ $home$ of $a_{1}$, does not arrive at $home$ of $a_{1}$ during its Forward state because of getting consumed at the black hole (refer Lemma 5.11). This instigates the agent $a_{1}$ to change its state to Gather and store the dir type message (clockwise,NULL) in its local memory..
In state Initial-Wait, the agent waits at its $home$ until $T_{time}=4n-1$. This waiting period is enough for an anomaly finding agent in state Backtrack to meet with it. If the agent finds that there are more than one agent at its $home$, and also there is a dir type message as well, then our algorithm ensures that the direction component of this dir type message must be in clockwise direction (refer Lemma 5.17). In which case, the agent starts moving clockwise and changes its state to Gather1. On the contrary, if $T_{time}=4n$ and no anomaly is detected, then the agent again moves back to state Initial.
An agent can change its state to Gather, either from state Forward, or from state Backtrack. Note that in either case, it carries the stored message of type dir (which has a direction component and an ID component which is either $NULL$ or stores the ID of an agent). So, in this state if the direction component of dir is along clockwise (resp. counter-clockwise), then the agent moves along clockwise (resp. counter-clockwise) direction until it finds a node with #$\#Agent$ more than one. Next, in case of counter-clockwise direction component, the agent updates the dir message to (Counter-clockwise, ID’) where ID’ is the ID of the other agent at the same node. Then the agent writes the updated dir message at the current node and changes state to Gather2. If the agent in state Gather met the other agent when the dir type message having clockwise direction component the, the agent simply just writes the message at the current node and moves to state Gather1.
Also note that exactly one agent can move into state Gather. This is because if an agent is consumed during Forward state then the anomaly is first found by a single agent whose $home$ is the clockwise nearest $home$ of the consumed agent. This agent changes its state to Gather. Now, before the other agent finds further anomaly the agent in state Gather meets the other agent and forces it to change into state Gather1. Similarly If an agent $a_{1}$ is consumed by the black hole during its Backtrack state, then the first anomaly is detected by a single agent (more precisely the agent for which the clockwise nearest $home$ is the $home$ of $a_{1}$). In this case it changes its state to Gather and forces the other agent to move into state Gather2 by meeting it before it finds further anomaly. Thus we have the following observation. {observation} There is exactly one agent that changes its state to Gather throughout the execution of Algorithm PerpExplore-Scat-Whitbrd. Also, The agent which changes its state to Gather, is the first agent to understand that, an agent has already been consumed by the black hole.
An agent can change its state to Gather1, only from the states Initial-Wait or Gather. An agent changes its state to Gather1 from Gather, only when while moving along clockwise direction finds another agent present, in which case it writes the corresponding message of type dir and correspondingly changes its state to Gather1. on the other hand, an agent (here $a_{1}$) can only change its state to Gather1 from Initial-Wait, if it is waiting at its $home$, and within which it encounters another agent at its $home$ along with it a message of type dir is written (in which the direction component of dir is along clockwise). Note that if $a_{1}$ in state Initial-Wait finds more than one agent at current node and dir type message at round $t$ then, the other agent must be in state Gather while written there at round $t-1$ and moved into state Gather1 in the same round. Then, during round $t$, $a_{1}$ first stores the dir type message before it moves to state Gather1 and moves according to the direction component of the dir type message (i.e., clockwise). Also, the other agent that changed its state to Gather1 from Gather in round $t-1$ also moves according to the dir type message during round $t$. Thus we have another observation {observation} If two agents move to state Gather1 , one moves from state Gather and the other moves from state Initial-Wait. Furthermore, they change to state Gather1 at the same node and leaves the node together at the same round and stays together.
In state Gather1, the agent which changes its state to Gather1 from Initial-Wait first stores the dir type message, as the other agent must already have stored this message in state Gather. Irrespective of which, the agent starts moving in a clockwise direction, until it finds a node which has a home (i.e., of the form (home, ID)) type message, where the ID component this message does not match with the IDs of the agents present at the current node. Note that our algorithm ensures that the current node is the $home$ of $a_{j}$, if $a_{j}$ is the agent to be consumed by the black hole (refer Lemma 5.19). Further, as there are two agents at the current node both in state Gather1 (Observation 5.2), so now the agent with the lowest ID, sets $Move=0$, also initializes the $Marking$ variable to $right$, since the direction component of dir is clockwise, and further changes its state to Cautious-Leader. On the other hand the other alive agent at the current node changes its state to Cautious-Follower, while initializing $Move=0$.
An agent can change its state to Gather2, only from the states Back-Wait or Gather.
An agent changes its state to Gather2 from Gather, only when while moving along counter-clockwise direction finds another agent present, in which case it updates the dir message and writes the corresponding message of type dir and correspondingly changes its state to Gather2.
On the other hand, an agent can only change its state to Gather2 from Back-Wait, if it is waiting at its corresponding segments nearest clockwise $home$, and within which it encounters another agent at its current node along with it a message of type dir is written (in which the direction component of dir is along counter-clockwise). After storing the dir type message in its local memory and changing its state to Gather2 from Back-Wait, an agent immediately moves counter-clockwise in the same round. This generates another observation similar to the Observation 5.2 as follows.
{observation}
If two agents move to state Gather2 , one moves from state Gather and the other moves from state Back-Wait. Furthermore, they change to state Gather2 at the same node and leaves the node together at the same round and stays together.
In state Gather2, the agent moves in a counter-clockwise direction, until it finds a node which has a home (i.e., of the form (home, ID’)) type message, where ID’ matches with ID component of dir type message in its local memory. Note that our algorithm ensures that the current node is the nearest clockwise $home$ of $a_{1}$, if $a_{1}$ is the agent to be consumed by the black hole (refer Lemma 5.19). Further, as there are two agents at the current node, both in state Gather2 (refer Observation 5.2), so now the agent with the lowest ID, sets $Move=0$, also initializes the $Marking$ variable to $left$, since the direction component of dir is counter-clockwise, and further changes its state to Cautious-Leader. On the other hand the other alive agent at the current node, changes its state to Cautious-Follower, while initializing $Move=0$.
In state Cautious-Leader, an agent updates $Move=1$ and moves along the direction component of dir. After moving when it is alone on a node, it checks if the current node is marked with $Marking$ (i.e., either $left$ or $right$). If the node is marked, then it moves in the opposite direction (i.e., the opposite of direction component of dir message) to meet with the agent in state Cautious-Follower. Otherwise, if there is no $Marking$ at the current node then it identifies the current node as the black hole node (only if it remains alive) and continues perpetual exploration avoiding this node. If an agent in state Cautious-Leader with $Move=1$ meets with another agent (i.e., the agent in state Cautious-Follower) it updates variable $Move$ to 0, and again moves according to the direction component.
In state Cautious-Follower, the agent after finding $Move=0$, initializes the variable $Wait=0$ and updates $Move=1$. If $Move=1$, then the agent checks whether $Wait<1$, if so then update $Wait=Wait+1$. When $Wait=1$, it checks if the current node has more than one agent (i.e., whether the agent in state Cautious-Leader has returned or not), if so then further update $Move=0$ and move along the direction component of dir, along with the agent in state Cautious-Leader. Otherwise, if the agent in state Cautious-Leader does not return, that means the number of agent at the current node is 1 when $Wait=1$, this symbolizes that the agent in state Cautious-Leader has been either consumed by the black hole or the agent in state Cautious-Leader finds that its current node (i.e., the next node along the direction component of dir for the agent in state Cautious-Follower) is the black hole and continues to perpetually move in the same direction, further avoiding this node. In any case, the agent executing Cautious-Leader has not returned implies the next node along direction component of dir for the agent executing Cautious-Follower is the black hole node. So, in this case, the agent declares the next node along the direction component of dir is the black hole node and continues perpetual exploration avoiding that node.
// Algorithm is written for an agent $r$
1 Input: $n$
[…]
4 In State Initial:
5 Clear Whiteboard at the $Current-Node$
6 $T_{time}=0$
7 Write (home, ID($r$)) at the $Current-Node$.
8 Change state to Forward and move clockwise
9
10In State Forward:
11 if $T_{time}<n$ then
12       $T_{time}=T_{time}+1$
13       if $Current-Node$ does not have any “home” type messages then
14
15            Write at the $Current-Node$ right and erase left (if exists) and move clockwise.
16
17
18else
19       $T_{time}=T_{time}+1$
20       if $Current-Node$ already have a “visited” type message then
21             Store the message of type dir, where dir=(Counter-clockwise,NULL)
22             and then change to state Gather.
23
24      else
25             Write (Visited, ID($r$)) at the $Current-Node$ and change to state Back-Wait
26
27
28In State Back-Wait:
29 if $T_{time}<2n$ then
30       $T_{time}=T_{time}+1$
31       if #$\#Agent$ at the $Current-Node>1$wedge have a message of type ”dir" then
32             if direction component of dir message is of type “counter-clockwise" then
33                   Store the dir type message in local memory
34                   Move along direction counter-clockwise, and change to state Gather2.
35
36
37
38else
39       Move in counter-clockwise direction and change to state Backtrack.
40
41In State Backtrack:
42 if $T_{time}<3n$ then
43       $T_{time}=T_{time}+1$.
44       if $Current-Node$ does not have any “home" type message then
45             Write at the $Current-Node$ left and erase right and move counter-clockwise.
46
47
48else
49       $T_{time}=T_{time}+1$
50       if $Current-Node$ already have a “visited" type message then
51             Change to state Initial-Wait.
52
53      else
54             Store the message of type dir, where dir=(Clockwise,NULL) and then change to state Gather.
55
56
57In state Initial-Wait:
58 if $T_{time}<4n$ then
59       $T_{time}=T_{time}+1$
60       if #$\#Agent$ at the $Current-Node>1$wedge have a message of type “dir" then
61             if direction component of dir message is of type “clockwise" then
62                   Store the dir type message in local memory
63                   Change to state Gather1 and move clockwise.
64
65
66
67else
68       Change to state Initial.
69
70In State Gather:
71 if direction component of dir is clockwise then
72       if #$\#Agents$ at $Current-Node$ is 1 then
73             Move in a clockwise direction.
74      else
75             Write at $Current-Node$ dir and move to state Gather1.
76
77
78else
79       if #$\#Agents$ at $Current-Node$ is 1 then
80             Move in a counter-clockwise direction.
81      else
             Update dir= (counter-clockwise,ID’) // ID’ is the ID of the other agent
82             Write at $Current-Node$ dir and move to state Gather2.
83
84
85
Algorithm 4 PerpExplore-Scat-WhitBrd
61
62In State Gather1:
63 if $Current-Node$ has “home" type message with the ID component of the message does not match with IDs of the agent present at the $Current-Node$  [^4] IDs present at the $Current-Node$ implies the IDs of the agents present at the $Current-Node$ then
64       if ID is the lowest among the set of IDs at the $Current-Node$ then
65             Set $Move=0$.
66             if direction component of dir is clockwise then
67                  Set $Marking=right$.
68            else
69                   $Marking=left$
70            Change its state to Cautious-Leader.
71
72      else
73             Set $Move=0$
74             Change to state Cautious-Follower.
75
76
77else
78       Move in a clockwise direction.
79
80In State Gather2:
81 if $Current-Node$ has a “home" type message with ID component same as ID component of dir then
82       if ID is the lowest among the set of IDs at the $Current-Node$ then
83             Set $Move=0$.
84             if direction component of dir is clockwise then
85                  Set $Marking=right$.
86            else
87                   $Marking=left$
88            Change its state to Cautious-Leader.
89
90      else
91             Set $Move=0$
92             Change to state Cautious-Follower.
93
94
95else
96       Move in a counter-clockwise direction.
97
98In State Cautious-Leader:
99 if $Move=0$ then
100       Update $Move=1$ and move according to the direction component of dir.
101
102else
103       if $Current-Node$ has exactly one agent then
104             if $Current-Node$ is already marked with $Marking$ then
105                   $Move=0$ and move opposite to the direction component of dir
106            else
107                   Declare $Current-Node$ as the black hole node and continue perpetual exploration avoiding this node.
108
109
110      else
111             Update $Move=0$ and moves according to the dirction component of dir type message in its local memory
112
113In State Cautious-Follower:
114 if $Move=0$ then
115       Set $Wait=0$ and update $Move=1$.
116
117else
118       if $Wait<1$ then
119             $Wait=Wait+1$.
120
121       if #$\#Agent$ at $Current-Node>1$ then
122             Update $Move=0$ and move according to the direction component of dir.
123
124      else
125             Declare the next node along the direction component of dir as the black hole node and continue perpetual exploration avoiding this node.
126
127
128
5.2.1 Correctness and Complexity
In the next two lemmas (Lemma 5.9 and Lemma 5.11) we first ensure that when we say an agent encounters an anomaly, is actually an anomaly. And what an agents interpret from these anomalies are true. Next in Lemma 5.13 we ensure that the agent that changes to state Gather can always identify the state of the consumed agent at the time it gets destroyed at the black hole.
Lemma 5.9.
Let $a_{i}$ and $a_{j}$ be two agents exploring the segments $Seg(a_{i})$ and $Seg(a_{j})$, such that $Seg(a_{i})\cap Seg(a_{j})=v_{c}$, where $v_{c}$ is also the $home$ of $a_{j}$. If during execution of PerpExplore-Scat-Whitbrd, there exists a round $t>0$, in which $a_{i}$ in state Forward finds a visited type message, then there exists a round $0<t^{\prime}<t$, in which $a_{j}$ has been consumed by the black hole, while exploring the segment $Seg(a_{j})$ in state Backtrack.
Proof 5.10.
Let us suppose, there does not exist any round $t^{\prime}<t$ in which $a_{j}$ has been consumed by the black hole in state Backtrack. This means that, either $a_{j}$ is yet to be consumed by the black hole at round $t^{\prime}$, for all $t^{\prime}<t$ or $a_{j}$ has been consumed by the black hole at round $t^{\prime}$, while it is in state Forward.
Case I: Let $a_{j}$ is yet to be consumed by the black hole at round $t^{\prime}$, for all $t^{\prime}<t$. Note that since $a_{i}$ is at $v_{c}$ and checks for the visited type message, this implies $T_{time}=n$. This means, at $(t-n-1)$-th round $a_{i}$ changed its state to Forward from Initial. This also implies $a_{j}$ was also in state Initial at $(t-n-1)-$th round at $v_{c}$. Hence, at this round $a_{j}$ must have cleared any whiteboard data at $v_{c}$ (as $v_{c}$ is the $home$ of $a_{j}$) and also has moved clockwise after changing the state to Forward. Observe that, the node $v_{c}$ can only be explored by $a_{i}$ and $a_{j}$. So, these two agents has the possibility to write the visited type message at $v_{c}$ after round $t-n-1$ and before round $t$. Further, note that in the next $n+1$ rounds after $(t-n-1)-$th round (i.e., until round $t$), $a_{j}$ cannot return back to $v_{c}$. So, $a_{j}$ can not write anything on $v_{c}$ after round $t-n-1$ and before round $t$. Also, $a_{i}$ can only write the visited type message at $v_{c}$ at the $t-$th round (i.e., when $T_{time}=n$ according to Algorithm 4). But, in this case, $a_{i}$ at $t-$th round before writing a message, finds that there already exists a visited type message, and this a contradiction, to the fact that $a_{j}$ has not been consumed by the black hole at round $t^{\prime}$, for all $t^{\prime}<t$.
Case II: Let $a_{j}$ has been consumed by the black hole at some round $t^{\prime}<t$, while in state Forward. This implies there exists a round $0<t^{\prime\prime}<t^{\prime}$ in which $a_{j}$ was in state Initial (this is the last time $a_{j}$ was in state Initial before it gets consumed by the black hole). Thus this means, $t^{\prime}\leq t^{\prime\prime}+n-2$ as a segment can be of length at most $n-2$. Then at round $t^{\prime\prime}$, $a_{i}$ was also in state Initial at its own $home$. Now, note that after $t^{\prime\prime}$, $a_{i}$ can visit $v_{c}$ within round $t^{\prime\prime}+1$ and $t^{\prime\prime}+n+1$ while in state Forward. Note that from round $t^{\prime\prime}+1$ and before round $t^{\prime\prime}+n+1$ no agent can write a visited type message at $v_{c}$ (by similar argument as in Case I). So during round $t^{\prime\prime}+n+1$, $a_{i}$ can not see any visited type message at $v_{c}$. So, $t>t^{\prime\prime}+n+1$. Also note that the next time after round $t^{\prime\prime}+n+1$, $a_{i}$ visits $v_{c}$ in state Forward, is earliest at round $t^{\prime\prime}+4n+1$. This implies $t\geq t^{\prime\prime}+4n+1$. Now, let us consider the agent $a_{k}$ where $Seg(a_{k})\cap Seg(a_{j})=home$ of $a_{k}$ and $Seg(a_{j})\cap Seg(a_{i})=home$ of $a_{i}$. Note that in round $t^{\prime\prime}+n+1$ both $a_{k}$ and $a_{i}$ moves into state Back-Wait (as none of them sees any visited type message). Next after waiting there upto round $t^{\prime\prime}+2n$, both of them changes their state to Backtrack at round $t^{\prime\prime}+2n+1$. Next they reach their corresponding $home$ and checks for a visited type message at round $t^{\prime\prime}+3n+1$. Note that $a_{i}$ finds the visited type message left by $a_{k}$ and changes to state Initial-Wait whereas, $a_{k}$ does not find any visited type message left by $a_{j}$ (as $a_{j}$ is consumed at the black hole at state Forward at round $t^{\prime}\leq t^{\prime\prime}+n-2<t^{\prime\prime}+3n+1$). So, $a_{k}$ changes to state Gather1 at round $t^{\prime\prime}+3n+1$ and moves clockwise with the message dir$=$ (Clockwise, NULL) until it finds $a_{i}$ at it’s $home$ in state Initial-Wait before round $t^{\prime\prime}+4n$ (as length of a segment can be of length at most $n-2$). Thus before $t^{\prime\prime}+4n+1$-th round $a_{i}$ changes its state to Gather1. Since no agent moves to state Forward from state Gather1, $a_{i}$ can not be at $v_{c}$ in state Forward during round $t$. So we again arrive at a contradiction. This implies $a_{j}$ can not be in state Forward while it is consumed at the black hole at round $t^{\prime}<t$.
Lemma 5.11.
Let $a_{i}$ and $a_{j}$ be two agents such that $v_{c}=Seg(a_{i})\cap Seg(a_{j})=home$ of $a_{i}$. If there exists a round $t>0$ such that $a_{i}$ checks and finds no visited type message on $v_{c}$ while in state Backtrack, then there must be a round $t^{\prime}<t$ in which $a_{j}$ was consumed by the black hole while in state Forward.
Proof 5.12.
Let us consider that there does not exists any round $t^{\prime}<t$ where $a_{j}$ was consumed by the black hole in state Forward. This implies that either $a_{j}$ is not consumed by the black hole for all rounds $t^{\prime}<t$ or it is consumed by the black hole at round $t^{\prime}<t$ while it was in state Backtrack.
Case-I: Let us consider that $a_{j}$ is yet to be consumed by the black hole at round $t^{\prime},\leavevmode\nobreak\ \forall t>t^{\prime}>0$. Note that, at round $t$, $a_{i}$ checks for a visited type message while in state Backtrack, and that means the current $T_{time}$ for any alive agent must be equal $3n$ (at $T_{time}=n$ every alive agent ends its state Forward, then until $T_{time}=2n$ every such agent is in state Back-Wait, and at $T_{time}=3n$ every alive agent checks for the message in state Backtrack). Note that, this means at time $t-3n-1$, both $a_{i}$ and $a_{j}$ was in state Initial at their respective $home$. Then at round $t-2n-1$, both of them must be in state Forward with $T_{time}=2n$, while $a_{j}$ is at $v_{c}$ and $a_{i}$ is at $home$ of $a_{k}$ ( $a_{k}$ is the third agent exploring $Seg(a_{k})$, such that $Seg(a_{i})\cap Seg(a_{k})=home$ of $a_{k}$ and $Seg(a_{k})\cap Seg(a_{j})=home$ of $a_{j}$). So, at this point, both $a_{i}$ and $a_{j}$ must have written a visited type message at their respective current nodes (i.e., $v_{c}$ and $home$ of $a_{k}$). Further observe that, within $t-2n$ round to $t-n-1$ round $a_{j}$ remains at $v_{c}$ while in state Back-Wait, and according to our algorithm no agent in state Back-Wait can alter the already stored information at their current node, i.e., within these rounds $a_{j}$ cannot erase the visited type message at $v_{c}$. Next, within $t-n$ to $t-1$ round, the only agent which can visit the node $v_{c}$ is $a_{i}$, while it is in state Backtrack. Note that, $a_{i}$ cannot alter any visited type message during state Backtrack. Thus, $a_{i}$ at round $t$ checks and finds a visited type message at $v_{c}$. This is a contradiction to the assumption that $a_{i}$ does not find any visited type message at round $t$. Thus, $a_{j}$ must have been consumed by the black hole at time $t^{\prime}$ for some $t^{\prime}<t$.
Case-II: Suppose $a_{j}$ has been consumed by the black hole at some round $t^{\prime}<t$, while in state Backtrack. Let $t^{\prime\prime}$ be the round when $a_{j}$ was in state Initial for the last time where $0<t^{\prime\prime}<t^{\prime}<t$. Note that $t^{\prime}>t^{\prime\prime}+2n+1$ (because an agent changes its state to Backtrack at round $t^{\prime\prime}+2n+1$). So at round $t^{\prime\prime}+n+1$ it was alive and at $v_{c}$ in state Forward. During this round it also writes a visited type message at $v_{c}$. Now as we discussed earlier in Case-I, from round $t^{\prime\prime}+n+1$ till round $t-1$ no agent can erase or alter the visited type message at $v_{c}$. Thus at round $t$, $a_{i}$ must find the visited type message at $v_{c}$ upon checking while in state Backtrack. This is a contradiction to the fact that at round $t$, $a_{i}$ checks and finds no visited type message at $v_{c}$ while it is in state Backtrack. Thus if $a_{j}$ is consumed by the black hole at round $t^{\prime}$, then it must be in state Forward.
Lemma 5.13.
The agent which changes its state to Gather, correctly identifies the state in which an agent was in, just before it gets consumed by the black hole.
Proof 5.14.
Let $a_{i},a_{j}$ and $a_{k}$ be three agents initially at three different nodes (i.e., three different $home$ for three agents) of $R$. Let $Seg(a_{i})\cap Seg(a_{j})=home$ of $a_{i}$, $Seg(a_{i})\cap Seg(a_{k})=home$ of $a_{k}$ and $Seg(a_{k})\cap Seg(a_{j})=home$ of $a_{j}$. Without loss of generality let $a_{j}$ be the agent that is consumed by the black hole at a round $t>0$. Now we have two cases according th the state of $a_{j}$ at the time of getting consumed.
Case-I: $a_{j}$ gets consumed by the black hole while in state Forward at round $t$. Let $t^{\prime}<t$ be the round when $a_{j}$ was in state Initial for the last time. Note that at round $t^{\prime}$, all agents are in state Initial. We claim that $a_{i}$ is the agent to change its state to Gather. If possible let $a_{k}$ changes its state to Gather. Then first we show that it must change its state to Gather before round $t^{\prime}+3n+1$. Otherwise, since $a_{j}$, fails to reach $home$ of $a_{i}$ at round $t^{\prime}+n+1$, it does not write any visited type message there but $a_{i}$ reaches $home$ of $a_{k}$ at round $t^{\prime}+n+1$ and writes a visited type message at $home$ of $a_{k}$. Also, since, $a_{i}$ and $a_{j}$ are the only two agent which visits $home$ of $a_{i}$., this implies, no message can be written at $home$ of $a_{i}$ on and between rounds $t^{\prime}$ and $t^{\prime}+3n$. So, when $a_{i}$ returns $home$ and finds no visited type message round $t^{\prime}+3n+1$ while in state Backtrack, it changes its state to Gather but $a_{k}$ doesn’t do so as it sees the visited type message at its $home$, left by $a_{i}$. So, if $a_{k}$ is the agent to change state to Gather it must be before round $t^{\prime}+3n+1$. Thus $a_{k}$ must move into state Gather from state Forward at round $t^{\prime}+n+1$. This can only occur if $a_{k}$ finds a visited type message at $home$ of $a_{j}$ at round $t^{\prime}+n+1$. But this is not possible as $a_{j}$ have already erased all previous data on its $home$ at round $t^{\prime}$ while in state Initial, and there is no other agent which can visit $home$ of $a_{j}$ and alter the data between $t^{\prime}$ and $t^{\prime}+n+1$ rounds. Hence, at round $t^{\prime}+n+1$ when $a_{k}$ visits the $home$ of $a_{j}$ it does not find any visited type message. Thus $a_{i}$ must be the agent to change its state to Gather. Now to prove that $a_{i}$ stores the dir type message with direction component clockwise (which means $a_{j}$ was in state Forward while it was consumed), we have to show that $a_{i}$ changes its state to Gather from Backtrack. If possible let $a_{i}$ changes its state to Gather from state Forward. Then it must be at round $t^{\prime}+n+1$ when $a_{i}$ checks and finds a visited type message at the $home$ of $a_{k}$. Now at round $t^{\prime}$, $a_{k}$ which was in state Initial, cleared all previous data on its $home$. This implies the visited type message that $a_{i}$ finds at round $t^{\prime}+n+1$ must be written there after round $t^{\prime}$. But as only $a_{i}$ can be there after $t^{\prime}$ and before $t^{\prime}+n+1$, and since it can not alter any data on $home$ of $a_{k}$ before round $t^{\prime}+n+1$, it finds no visited type message at $home$ of $a_{k}$ at round $t^{\prime}+n+1$. Thus $a_{i}$ can not change its state to Gather from state Forward. Hence, it must change its state to Gather from state Backtrack at round $t^{\prime}+3n+1$ and according to the algorithm PerpExplore-Scat-WhitBrd, the dir type message that $a_{i}$ stores must have direction component clockwise.
Case-II: Let $a_{j}$ gets consumed by the black hole at round $t$ while in state Backtrack. We have to show that the agent that changes the state to Gather must store the dir type message with direction component counter-clockwise (as any agent can only move in counter clockwise direction, in state Backtrack). It will be enough to show that the agent that changes its state to Gather must have changed it from state Forward (as only in this case the state changing agent stores the dir type message having a counter-clockwise direction component in its local memory). Let $t^{\prime}$ be the round when $a_{j}$ was in state Initial for the last time. We claim that $a_{i}$ and $a_{k}$ can not change to state Gather before round $t^{\prime}+3n+1$. Note that at round $t^{\prime}$ both $a_{i}$ and $a_{k}$ are at their corresponding $home$ in state Initial. Now if any one of them changes to state Gather, the earliest it can happen is at round $t^{\prime}+n+1$ when both of them are in state Forward. In this case the agent that changes to state Gather must have found a visited type message at the current node (i.e., for $a_{i}$ it is $home$ of $a_{k}$ and for $a_{k}$ it is $home$ of $a_{j}$) at round $t^{\prime}+n+1$. Note that during round $t^{\prime}$ any previous messages are erased from both $home$ of $a_{k}$ and $home$ of $a_{j}$ by $a_{k}$ and $a_{j}$, respectively and no other agent can visit and alter data at these nodes before $t^{\prime}+n+1$ round. Hence none of $a_{i}$ and $a_{k}$ finds any visited type message at their current nodes at round $t^{\prime}+n+1$. Hence, none of these agents changes their state to Gather at round $t^{\prime}+n+1$. Next they can only change their state to Gather at the round $t^{\prime}+3n+1$ when both of them (i.e., $a_{i}$ and $a_{k}$) are at their respective $home$. An agent among them changes its state to Gather at round $t^{\prime}+3n+1$ if it finds no visited type message at their current node (i.e., corresponding $home$). Note that $t^{\prime}+3n+1>t>t^{\prime}+n+1$ (as $a_{j}$ was in state Backtrack at round $t$). So, at round $t^{\prime}+n+1$ all of $a_{i},a_{j}$ and $a_{k}$ are at nodes $home$ of $a_{k},home$ of $a_{i}$ and $home$ of $a_{j}$ in state Forward and writes a visited type message in the nodes, respectively. These messages can not be altered by any agent until round $t^{\prime}+3n+1$ (by similar argument as in Case-I). Note that at round $t^{\prime}+3n+1$ both $a_{i}$ and $a_{k}$ are at their corresponding $home$ and both of them finds a visited type message at these nodes left by $a_{j}$ and $a_{i}$ respectively (at round $t^{\prime}+n+1$). So, none of them changes to state Gather even at round $t^{\prime}+3n+1$. Next they move into state Initial-Wait and wait at their $home$ until $t^{\prime}+4n$ round. Now, since no agent in state Initial-Wait can change its state to Gather2 until it meets with an agent in state Gather and all alive agent at round $t^{\prime}+3n+2$ are at state Initial-Wait all of them (i.e., $a_{i}$ and $a_{k}$) waits and changes state to Initial again at round $t^{\prime}+4n+1$. Next at round $t^{\prime}+4n+2$ both $a_{i}$ and $a_{k}$ erase all previous data at their corresponding $home$. and changes status to Forward again. Note that the next time $a_{i}$ and $a_{k}$ can change the state to gather must be at the round $t^{\prime}+5n+2$ when both of them are in state Forward, and are currently at the $home$ of $a_{k}$ and $home$ of $a_{j}$, respectively. Further, note that since $home$ of $a_{k}$ does not have any visited type message at round $t^{\prime}+5n+2$ (as $a_{k}$ erased any data at round $t^{\prime}+4n+2$ and no other agent can alter data there after $t^{\prime}+4n+2$ and before $t^{\prime}+5n+2$), so $a_{i}$ can not change to state Gather at round $t^{\prime}+5n+2$. Also observe that the visited type message at the $home$ of $a_{j}$ written by $a_{k}$ during round $t^{\prime}+3n+1$ is still there at round $t^{\prime}+5n+2$ (after $t^{\prime}+3n+1$ before $t^{\prime}+4n+2$ since $a_{j}$ is already consumed before $t^{\prime}+3n+1$, no agent is on $home$ of $a_{j}$ to erase the data, also after $t^{\prime}+4n+1$ till $t^{\prime}+5n+1$ only $a_{k}$ can visit $home$ of $a_{j}$ in state Forward and Back-Wait, but in these states it does not alter any data at the $home$ of $a_{j}$) $a_{k}$ finds it during the round $t^{\prime}+5n+2$ in state Forward and changes its state to Gather.
Lemma 5.15.
If an agent finds a dir type message while it is in state Back-Wait, the direction component of this message must be counter-clockwise.
Proof 5.16.
Suppose an agent $a_{i}$ in state Back-Wait gets a dir type message with clockwise direction component at some round $t$. This implies there exists another agent $a_{k}$ that has changed its state to Gather after storing a dir type message having direction component clockwise at some round $t^{\prime}<t$. This can only happen if $a_{k}$ was in state Backtrack at the beginning of round $t^{\prime}$. So at the beginning of round $t^{\prime}$, $a_{i}$ was also in state Backtrack and thus at round $t^{\prime}$, $a_{i}$ changes its state to Initial-Wait. Note that $a_{k}$ meets and shares dir type message with $a_{i}$ while $a_{i}$ is still at state Initial-Wait. This contradicts our assumption that $a_{i}$ gets dir message at state Back-Wait. Thus, if $a_{i}$ finds a dir type message in state Back-Wait then it must have the direction component counter-clockwise.
With similar argument we can also proof the following lemma.
Lemma 5.17.
If an agent finds a dir type message while it is in state Initial-Wait, the direction component of this message must be clockwise.
Definition 5.18 (Cautious start node).
Let $a_{i}$, $a_{j}$ and $a_{k}$ be three agents executing the algorithm PerpExplore-Scat-Whitbrd, and suppose $a_{j}$ be the first agent to enter the black hole, while exploring the segment $Seg(a_{j})$. Let $v_{1}$ be the $home$ of $a_{j}$ and $v_{2}$ be the furthest node from $v_{1}$ along clockwise direction which is inside $Seg(a_{j})$. We define the Cautious start node to be $v_{1}$ if $a_{j}$ is consumed by the black hole during state Forward. Otherwise, if $a_{j}$ is consumed by the black hole in state Backtrack then, $v_{2}$ is defined to be the Cautious start node.
Lemma 5.19.
Let the first agent is consumed by the black hole at some round $t>0$, then there exists a round $t^{\prime}>t$, at which the remaining alive agents reach the cautious start node and changes their state to Cautious-Leader and Cautious-Follower.
Proof 5.20.
Let $a_{i}$, $a_{j}$ and $a_{k}$ be three agents, exploring $Seg(a_{i})$, $Seg(a_{j})$ and $Seg(a_{k})$, respectively, where $Seg(a_{i})\cap Seg(a_{j})=home$ of $a_{i}$, $Seg(a_{i})\cap Seg(a_{k})=home$ of $a_{k}$ and $Seg(a_{k})\cap Seg(a_{j})=home$ of $a_{j}$. Let $a_{j}$ be the agent which gets consumed by the black hole, while exploring $Seg(a_{j})$, then we have the following cases:
Case-I: $a_{j}$ is consumed by the black hole while it is moving in a clockwise direction, i.e., in state Forward at round $t>0$. Note that in this case the cautious start node is the $home$ of $a_{j}$. This implies there exists a round $0<t_{0}<t<t_{0}+n$, where $a_{j}$ was in state Initial. Note that, $a_{j}$ can not write any visited type message at $home$ of $a_{i}$, as it gets consumed before reaching that node. So, at round $t_{0}+3n+1$, when $a_{i}$ is at its own $home$ in state Backtrack and checks any visited type message, it finds none exists. This triggers $a_{i}$ to change its state to Gather from Backtrack with the corresponding dir = (clockwise, NULL) (which is indeed the correct direction, refer Lemma 5.13). Further, the agent finds $a_{k}$, which is currently waiting at its own $home$ in state Initial-Wait (as it does not find any anomaly, so from Backtrack it changed to state Initial-Wait at round $t_{0}+3n+1$). So, the moment $a_{i}$ reaches the $home$ of $a_{k}$, it takes one additional round to store the dir type message at the current node and then changes its state to Gather1. On the other hand, whenever $a_{k}$ finds a dir type message is written at its current node, it also changes its state to Gather1, i.e., at round $t_{0}+3n+1<t^{\prime\prime}<t_{0}+4n$ both agents change their state to Gather1. After which, they together start moving in a clockwise direction (5.2), until they reach a node which has a home type message with the ID component, different from the IDs of $a_{i}$ and $a_{k}$. Note that this must be the $home$ of $a_{j}$, as the home type message $a_{j}$ has written at round $t_{0}$, cannot be erased by any other agent except $a_{j}$. Also $a_{j}$ can only erase this in state Initial at round $t_{0}+4n+2$ if it would have returned back, but since it is already consumed between round $t_{0}+1$ and $t_{0}+n-1$, hence this possibility never arises, so the home type message remains, when $a_{i}$ and $a_{k}$ together reaches this node. After which they change their state to Cautious-Leader and Cautious-Follower depending on their IDs.
Case-II: $a_{j}$ is consumed by the black hole while it is moving in a counter-clockwise direction, i.e., in state Backtrack at round $t>0$. Note that in this case the cautious start node is the $home$ of $a_{i}$. Let $t_{0}$ be the round when $a_{j}$ was in state Initial the last time. This implies $t_{0}+2n<t<t_{0}+3n$, now by similar argument as explained in Case-II of Lemma 5.13, $a_{k}$ changes its state to Gather while storing the dir = (counter-clockwise,NULL) message, at the $home$ of $a_{j}$ from state Forward, and starts moving in a counter-clockwise direction. Note that at round $t_{0}+5n+2$ as $a_{i}$ did not find any anomaly, so it changes its state to Back-Wait from Forward. Hence, at round $t_{1}$, where $t_{1}<t_{0}+6n$, $a_{k}$ finds $a_{i}$, while $a_{i}$ is still in state Back-Wait. This triggers $a_{k}$ to change its state to Gather2 at round $t_{1}$, while updating the dir type message to (Counter-clockwise, ID’), where ID’ is the ID of $a_{i}$. Then at the same round $a_{k}$ writes the updated message at the current node (i.e., $home$ of $a_{k}$). Whenever $a_{i}$ sees this dir type message (at round $t_{1}+1$) it also changes its state to Gather2. Next in state Gather2, both start to move counter-clockwise (from round $t_{1}+2$) and continues to move until they find a home type message with ID matching the ID of the dir type message. Note that, this node is nothing but the $home$ of $a_{i}$ (as the ID component of dir type message stores the ID of $a_{i}$). Note that at round $t_{0}+4n+2$, $a_{i}$ written a home type message at its own $home$. This message can be erased only again at round $t_{0}+8n+2$ (i.e., when $a_{i}$ reaches its $home$ again in state Initial). But since $a_{i}$ changed its state to Gather2 before $t_{0}+6n+1$ it cannot move back to state Initial again. So,when $a_{k}$ and $a_{i}$ reaches $home$ of $a_{i}$, they finds the home type message. Hence, both $a_{k}$ and $a_{i}$ reach the cautious start node within $t_{0}+7n$ and further changes their states to Cautious-Leader and Cautious-Follower, depending on their IDs.
Lemma 5.21.
Let $a_{i}$ and $a_{k}$ be the two agents which start the states Cautious-Leader and Cautious-Follower from the cautious start node, then within finite rounds of executing algorithm PerpExplore-Scat-WhitBrd, at least one agent detects the location of the black hole.
Proof 5.22.
Let $a_{j}$ be the agent that has been consumed by the black hole at round $t>0$. Now there are two cases based on the state of $a_{j}$ at round $t$.
Case-I: $a_{j}$ was in state Forward at round $t$. In that case the node $v_{1}=home$ of $a_{j}$ is the cautious start node. Let $v_{2}$ be the farthest node along clockwise direction in $Seg(a_{j})$ (i.e., the clockwise nearest $home$ of $a_{j}$). Now if there exists any round $t^{\prime}<t$ when $a_{j}$ was in state Backtrack then it must have started its state Backtrack from $v_{2}$ and moved counter-clockwise until $v_{1}$ while erasing all right markings from each node in $Seg(a_{j})$. So when $a_{j}$ started the state Forward at its $home$ before being consumed, all nodes in $Seg(a_{j})$ are without any right marking. This case will happen also when there are no round $t^{\prime}<t$ where $a_{j}$ was in state Backtrack, i.e., there does not exist any round before $t^{\prime}$, when $a_{j}$ was in state Forward. So before $a_{j}$ is consumed at the black hole , say $v_{b}$, at round $t$, it has marked right at all nodes starting from the next node of $v_{1}$ in clockwise direction, upto the node just before $v_{b}$ ($v_{b}$ can not be marked as before marking it the agent $a_{j}$ is consumed there). Let without loss of generality, $a_{i}$ is in state Cautious-Leader and $a_{j}$ is in state Cautious-Follower at $v_{1}$ at some round $t_{0}>t$. Then $a_{i}$ always moves ahead alone in clockwise direction to a new node $v$.Then it moves back only if sees the right marking and brings $a_{k}$ to $v$ along with it. Note that $a_{i}$ always sees a right marking until $v_{b}$. Now when it moves to $v_{b}$, if it is not consumed it must see no such right marking as $a_{j}$ failed to mark it at round $t$. In this case $a_{i}$ leaves the node in clockwise direction to a new node. So here $a_{i}$ is able to detect the black hole. On the other hand if $a_{i}$ is consumed while it visits $v_{b}$, then it can not return back to $a_{k}$. When $a_{k}$ sees that $a_{i}$ has not returned it interprets that $a_{i}$ must have been consumed at the black hole which is the next node along clockwise direction. Thus in this case also at least one agent can detect the black hole.
Case-II: $a_{j}$ was in state Backtrack at round $t$. In this situation, let without loss of generality, the node $v_{2}$ be the $home$ of $a_{i}$ and, it is also the cautious start node. Let $v_{1}$ be the farthest node along counter-clockwise direction in $Seg(a_{j})$ (i.e., the $home$ of $a_{j}$). Now, if there exists any round $t^{\prime}<t$ when $a_{j}$ was in state Forward, then it must have started this state from $v_{1}$ and moved in a clockwise direction until it reaches the node $v_{2}$ while erasing all the left markings from each node it traverses in $Seg(a_{j})$. This means that, when $a_{j}$ started the state Backtrack from $v_{2}$, there does not exist any node in $Seg(a_{j})$ with left markings. So before $a_{j}$ is consumed by the black hole node $v_{b}$ (say) at round $t$, it must have marked all the nodes with left, starting from the next counter-clockwise node of $v_{2}$ to the adjacent clockwise neighbor of $v_{b}$ (as before writing this message at $v_{b}$, the agent gets consumed by the black hole). Let without loss of generality, $a_{i}$ be the lowest ID agent among $a_{i}$ and $a_{k}$, hence it starts in state Cautious-Leader, whereas $a_{k}$ starts in state Cautious-Follower, at some round $t_{0}>t$. This means $a_{i}$ is the first agent to move alone in the next counter-clockwise neighbour say, $v$. After which, only if it sees a left message then only it moves back in clockwise direction at the node of $a_{k}$, and in the next round both these agents reach the node $v$. Observe, $a_{i}$ always finds a left message until the node $v_{b}$. Whenever it reaches $v_{b}$, either it gets consumed by the black hole, or it finds that no left marking is present at the current node. This triggers $a_{i}$ to detect the current node to be the black hole and moves in counter-clockwise direction to a new node. Otherwise, if it also gets consumed by the black hole, then in the next round it is unable to return to $a_{k}$, which triggers $a_{k}$ to conclude that $a_{i}$ must have been consumed by the black hole as well, and it correctly detects the black hole to be the next node in the counter-clockwise direction. Thus for each scenario there exists at least one agent which is able to correctly detect the black hole location.
Note that within at most $3n$ number of rounds after both alive agent starts cautious walk from cautious start node, the black hole will be detected by at least one agent. So from Lemma 5.19 and Lemma 5.21 within atmost $10n$ rounds after the first agent is consumed by the black hole there exists atleast one agent that knows the exact location of the black hole which can now explore the ring $R$ perpetually avoiding the black hole. So if all three agents strats from three different nodes the PerpExploration-BBH will be solved if each node has a whiteborad of memory $O(\log n)$. Now for the case where three agents starts from two different nodes we have the following remark.
Remark 5.23.
Let $a_{i},a_{j}$ and $a_{k}$ be three agents that starts from two initial nodes, say $home_{1}$ and $home_{2}$. By Pigeon hole principle, exactly one of $home_{1}$ and $home_{2}$ initially must have two agents. Without loss of generality let $home_{1}$ has two agents, say $a_{i}$ and $a_{k}$, initially. In this case, the agents having multiplicity greater than one at the current node does not move. On the other hand the singleton agent, i.e., $a_{j}$ starting from $home_{2}$ moves clockwise marking each node with message right. If $a_{j}$ reaches $home_{1}$ before being consumed by the black hole, $home_{1}$ now has three agents co-located. Thus from here the agents execute the whiteboard version of PerpExplore-Coloc-Pbl (refer Remark 4.24). On the other hand, let us consider the case when $a_{j}$ gets consumed before reaching $home_{1}$. Note that irrespective of the location of $home_{2}$ it takes at most $n-1$ rounds for $a_{j}$ to reach $home_{1}$ from the beginning. So $a_{i}$ and $a_{j}$ waits for $n$ rounds and finds that no one has arrived yet. In this case, both of $a_{i}$ and $a_{k}$ moves to $home_{2}$ along clockwise direction and starts to perform the cautious walk , where lowest ID agent among $a_{i}$ and $a_{k}$ changes its state to Cautious-Leader, whereas the other agent changes its state to Cautious-Follower. Next, the agent executing Cautious-Leader searches for the marking right. As argued earlier, within at most $4n$ rounds after an agent is consumed, at least one of the remaining alive agents can detect the black hole and continue to explore the ring avoiding that node.
So from Remark 4.24 and from the results in this section we have the following theorem
Theorem 5.24.
Three synchronous agents are necessary and sufficient to solve the problem PerpExploration-BBH on a ring $R$, when each node of $R$ has a whiteboard of $O(\log n)$ bits of memory irrespective of their starting location.
6 Conclusion

### v2

@@
-Pritam Goswami    Adri Bhattacharya    Raja Das    Partha Sarathi Mandal
+Perpetual Exploration of a Ring in Presence of Byzantine Black Hole
 Abstract
-Perpetual exploration is a fundamental problem in the domain of mobile agents, where an agent needs to visit each node infinitely often.
-This issue has received lot of attention, mainly for ring topologies, presence of black holes adds more complexity.
-A black hole can destroy any incoming agent without any observable trace.
-In [2, 18], the authors considered this problem in the context of Periodic data retrieval.
-They introduced a variant of black hole called gray hole (where the adversary chooses whether to destroy an agent or let it pass) among others and showed that 4 asynchronous and co-located agents are essential to solve this problem (hence perpetual exploration) in presence of such a gray hole if each node of the ring has a whiteboard.
-This paper investigates the exploration of a ring in presence of a “byzantine black hole”.
-In addition to the capabilities of a gray hole, in this variant, the adversary chooses whether to erase any previously stored information on that node.
-Previously, one particular initial scenario (i.e., agents are co-located) and one particular communication model (i.e., whiteboard) are investigated.
-Now, there can be other initial scenarios where all agents may not be co-located.
-Also, there are many weaker models of communications (i.e., Face-to-Face, Pebble) where this problem is yet to be investigated.
-The agents are synchronous.
-The main results focus on minimizing the agent number while ensuring that perpetual exploration is achieved even in presence of such a node under various communication models and starting positions.
-Further, we achieved a better upper and lower bound result (i.e., 3 agents) for this problem (where the malicious node is a generalized version of a gray hole), by trading-off scheduler capability, for co-located and in presence of a whiteboard.
+Perpetual exploration stands as a fundamental problem in the domain of distributed mobile agent algorithms, where the objective is to ensure that each node within a graph is visited by at least one agent infinitely often.
+While this issue has received significant attention, particularly concerning ring topologies, the presence of malicious nodes, referred to as black holes, adds more complexity.
+A black hole can destroy any incoming agent without leaving any trace of its existence.
+In [2, 18], the authors have considered this problem in the context of periodic data retrieval.
+They introduced a variant of black hole called gray hole (where the adversary chooses whether to destroy an agent or let it pass) among others, and showed that 4 asynchronous and co-located agents are necessary and sufficient to solve the periodic data retrieval problem (hence perpetual exploration) in the presence of such a gray hole if each of the nodes of the ring has a whiteboard.
+This paper investigates the exploration of a ring by introducing a realistic variant of gray hole, termed as a “Byzantine black hole”.
+In addition to the capabilities of a gray hole, the adversary can also choose whether to erase any previously stored information on that node.
+Note that in [2, 18], the authors only considered one particular initial scenario (i.e., agents are initially co-located) and one specific communication model (i.e., whiteboard).
+Now, there can be many other initial scenarios where all agents might not be co-located (i.e., they may be scattered).
+Also, there are many weaker communications models such as Face-to-Face, Pebble where this perpetual exploration problem is yet to be investigated, in presence of a Byzantine black hole.
+The main results of our paper emphasize minimizing the number of agents while guaranteeing that they perform the perpetual exploration on a ring even in the presence of a Byzantine black hole under different communication models and for different starting scenarios.
+On the positive side, as a byproduct of our work, we achieved a better upper and lower bound result (i.e., 3 agents) for perpetual exploration in the presence of a Byzantine black hole (which is a more generalized version of a gray hole), by trading-off the scheduler capability, when the agents are initially co-located, and each node contains a whiteboard.
 keywords:
-Mobile Agents, Exploration, Ring, Black Hole, Byzantine Fault
+Mobile Agents, Exploration, Ring, Black Hole, Malicious host, Byzantine Fault
+category:
+[…]
 1 Introduction
 Exploring a set of nodes in a network is one of the fundamental tasks in the domain of distributed computing by mobile agents, formulated in the year of 1951 by Shannon [20].
-Now the of security of these mobile agents while exploring these networks is one of the fundamental issues to address.
-Among all the possible security threats that is addressed yet in literature, two among them are the most prominent, i.e., the threats from a malicious agent [17] and the threats from a malicious host [13].
+Now, the security of these mobile agents while exploring these networks is one of the important issues that need to be addressed.
+Among all the possible security threats that are addressed yet in literature, two among them are the most prominent, i.e., the threats from a malicious agent [17] and the threats from a malicious host [13].
 In this paper, we are interested in the latter case, where the threats are from a malicious host.
 This host is a stationary node in the network, which has the ability to destroy any incoming agents without leaving any trace of its existence.
-So, first task of the mobile agents operating in the network, must be to locate this malicious node.
-Note that, the most trivial optimisation parameter to ensure while locating the black hole, is that minimum of number of agents gets consumed or destroyed by this node.
+So, the first task of the mobile agents operating in the network must be to locate this malicious node.
+Note that the most trivial optimisation parameter to ensure while locating this malicious host (also termed as a black hole), is that a minimum number of agents gets destroyed by this node.
 This problem of locating the black hole by mobile agents is termed as black hole search (also termed as BHS problem) problem.
-BHS problem is studied from the year of 2006, when Dobrev et al.
+BHS problem is studied from the year 2006, when Dobrev et al.
 [12] first introduced it.
-After which, till date there has been many variations to this problem, some of them are [7, 8, 11, 13, 15].
-This problem has various real life implications, such as the black hole can be a virus in the network or it can be some kind of crash failure, such that this node resembles the characteristic of a black hole, after the failure.
-Observe that, in order to detect the black hole, there needs to be some agent which has to visit that particular node.
-Further, since any agent visiting the node gets destroyed, so there must be some communication tool, which can render this information to other alive agents, such that at least one agent remains alive, knowing the location of the black hole.
-Three such communication tool has been predominantly used in literature: a) whiteboard model [14], in which there is a storage capacity at each node, which an agent can use to leave a message by reading its contents and writing some new information, b) pebble model [16], an agent can carry a movable token from one node to another, c) face-to-face model [10], in this case an agent can share and communicate with another agent when they are at the same node and at the same time.
-In addition to the communication tools, the initial locations of the agents (i.e., whether the agents are initially scattered [15] or they are co-located [10]) is also one of the important parameters, generally studied in literature.
-Further, the most studied version of black hole has fairly a basic nature, i.e., only destroying any incoming agent.
-Note that, in reality black hole’s may not be such simple, it may have many ways to disrupt the movement or harm an agent.
-Considering this phenomenon, we in this paper have tried to consider a black hole which has more capabilities other than just destroying any incoming agent.
-In our case, black hole may or may not kill any incoming agent, it may do so based on an adversarial scheduler which decides when to destroy an incoming agent and when not to.
-Whenever it decides not to destroy an agent, it simply behaves as any other node in the network, disguising it from the rest of the nodes, creating no anomaly for the visiting agent.
-In addition to this, we have also considered that the black hole has further capabilities, it can also choose whether to destroy the message (i.e., stored data in case of whiteboard, and placed token in case of pebble) at that node along with the incoming agent or not.
-This choice is also maintained by an adversarial scheduler.
-We call this kind of black hole as a byzantine black hole.
-Our aim in this paper is to solve the problem of perpetual exploration in a network, i.e., visiting every node in the network infinitely often by the mobile agents, while detecting the presence of this byzantine black hole.
-After detection, the task of the agents must be visit the remaining nodes of the network infinitely often, whereas avoiding this byzantine black hole node.
-Previously, in [2, 18] the authors introduced this model of black hole, which has more capabilities other than just destroying a black hole.
-They considered the following black hole characteristics: a black hole can fake agents, change the whiteboard contents, or change the ports different from the requested ones.
-In those papers, their aim was to achieve perpetual exploration in a ring (which they term as periodic data retrieval problem) by a team of asynchronous mobile agents, which are initially co-located, under these various black hole characteristics.
-In our paper, we chose one such black hole (which has more capability than a normal black hole) and we investigated this perpetual exploration problem in a ring, by a team of synchronous agents, under all possible initial configurations (i.e., the agents can be co-located or the agents can be scattered), as well as under all possible communication tools (i.e., the agents can either communicate among themselves via whiteboard, or using a pebble, or while they meet in a face-to-face manner).
-In each case, we are able to design algorithms, which serve our purpose of detection of this malicious node, as well as execute perpetual exploration.
+After which, till date, there have been many variations to this problem, some of them are [7, 8, 11, 13, 15].
+This problem has various real life implications, such as the black hole can be a virus in the network or it can be some crash failure, such that this node resembles the characteristic of a black hole, after the failure.
+Observe that, to detect the black hole, there needs to be some agent that has to visit that particular node.
+Further, since any agent visiting the node gets destroyed, there must be some communication tool, that can render this information to other alive agents, such that at least one agent remains alive, knowing the location of the black hole.
+Three such communication tools have been predominantly used in literature: a) whiteboard model [14], in which there is a storage capacity at each node, which an agent can use to leave a message by reading its contents and writing some new information, b) pebble model [16], an agent can carry a movable token from one node to another, c) face-to-face model [10], in this case, an agent can share and communicate with another agent when they are at the same node and at the same time.
+In addition to the communication tools, the initial locations of the agents (i.e., whether the agents are initially scattered [15] or they are co-located [10]) is also one of the important parameters, generally studied in the literature.
+Further, the most studied version of a black hole has a fairly basic nature, i.e., only destroying any incoming agent.
+Note that, in reality black holes may not be so simple; they may have many ways to disrupt the movement or harm an agent.
+Considering this phenomenon, we in this paper have tried to consider a black hole that has more capabilities other than just destroying any incoming agent.
+In our case a black hole may or may not kill any incoming agent; it may do so based on an adversary, which decides when to destroy an incoming agent and when not to.
+Whenever it decides not to destroy an agent, it simply behaves like any other node in the network, disguising it from the rest of the nodes, and creating no anomaly for the visiting agent.
+In addition to this, we have also considered that the black hole has further capabilities; it can also choose whether to destroy the message (i.e., stored data in case of a whiteboard, and place the token in case of pebble) at that node along with the incoming agent.
+This choice is also maintained by an adversary as well.
+We call this kind of black hole a Byzantine black hole.
+Our aim in this paper is to solve the problem of perpetual exploration in a network, i.e., visiting every node in the network infinitely often other than the Byzantine black hole, by the mobile agents.
+Previously, in [2, 18] the authors introduced a set of models for a black hole, which has more capabilities other than just destroying an agent, they term these malicious nodes as gray hole, gray+ hole and red hole, respectively based on their capabilities.
+They considered the following characteristics: they can fake agents, change the whiteboard contents, change the ports different from the requested ones, or can change the FIFO ordering as well.
+In this context, they solved perpetual exploration in a ring (which they term as periodic data retrieval problem) by a team of asynchronous mobile agents, under these aforementioned various black hole characteristics, in which the agents are initially co-located and each node in a network has a whiteboard.
+On the other hand, the results of above-mentioned papers only holds for the case, when initially the agents are co-located and each node of the ring has a whiteboard.
+Note that there can be many other initial positions for the agent to start with, and also, whiteboard is a very powerful communication tool in this domain of study of mobile agents.
+So, in this paper, we investigate these gaps in the context of perpetual exploration, considering the presence of one Byzantine black hole using synchronous agents.
+Also note that the position of a Byzantine black hole can be arbitrary (except the starting locations of the agent) but fixed.
+The Byzantine black hole we considered is a generalized version of the gray hole, as it also can choose whether to erase any previous data stored at that node, the moment it acts as a black hole.
 1.1 Related Works
-The black hole search (i.e., BHS) problem is a prominent sub-problem of exploration problem, studied in literature.
+The black hole search (i.e., BHS) problem is a prominent variation of exploration problem studied in the literature.
 A survey of which can be found in [19].
 This problem is investigated under various topologies (such as trees [8], rings [13], tori [5] and in arbitrary and unknown networks [7, 11]).
-All these discussed networks yet are static in nature.
-Recently, there has been a lot of interest on dynamic networks.
+All these discussed networks are static in nature.
+Recently, there has been a lot of interest in dynamic networks.
 The following papers [3, 4, 10], studied the BHS problem on dynamic ring, dynamic torus and dynamic cactus graph, where the underlying condition is that, irrespective of how many edges are dynamic in nature, the network must remain connected at any time interval (which is also termed as 1-interval connected).
-In rings, the BHS problem has been studied for different variants, the most predominant among them are choice of schedulers (i.e., synchronous [6] and asynchronous [1]), communication tools (i.e., face-to-face [7], pebble [16] and whiteboard [1]) and initial position of the agents (i.e., colocated [1] and scattered [6]).
-The most relevant papers, related to our work are the papers by Královič et al.
+In rings, the BHS problem has been studied for different variants; the most predominant among them are choice of schedulers (i.e., synchronous [6] and asynchronous [1]), communication tools (i.e., face-to-face [7], pebble [16] and whiteboard [1]) and initial position of the agents (i.e., co-located [1] and scattered [6]).
+The most relevant papers related to our work are the papers by Královič et al.
 [18] and by Bampas et al.
@@
 [18] is the first to introduce a variant of this black hole, where the black hole has the ability to either choose to destroy an agent or let it pass (which they term as gray hole).
-Further they extended the notion of gray hole, where the gray hole has these following additional capabilities: it has the ability to alter the run time environment (i.e., changing the whiteboard information), or it has the ability to not to maintain communication protocol (i.e., do not maintain FIFO order).
-They solved this problem under asynchronous scheduler on a ring, only when the agents are initially co-located and each node in the network has a whiteboard.
-The following results are obtained by them, they gave an upper bound of 9 agents for performing periodic data retrieval (i.e., which is equivalent to perpetual exploration) in presence of a gray hole, further, in addition to gray hole, when the whiteboard is unreliable as well, they proposed an upper bound of 27 agents.
+Further they extended the notion of gray hole, where the gray hole has the following additional capabilities: it has the ability to alter the run time environment (i.e., changing the whiteboard information), or it has the ability to not to maintain communication protocol (i.e., do not maintain FIFO order).
+They solved this problem under an asynchronous scheduler on a ring, only when the agents are initially co-located and each node in the network has a whiteboard.
+The following results are obtained by them, they gave an upper bound of 9 agents for performing periodic data retrieval (i.e., which is equivalent to perpetual exploration) in the presence of a gray hole; further, in addition to gray hole, when the whiteboard is unreliable as well, they proposed an upper bound of 27 agents.
 Next, Bampas et al.
@@
 As far as we are aware, we are the first to investigate the perpetual exploration problem of a ring under different communication tools (i.e., face-to-face, pebble and whiteboard) as well as for different initial positions (i.e., co-located and scattered), for a variant of gray hole, where it can erase any previously stored information but can not alter it.
-We term this type of gray hole as a byzantine black hole.
+We term this type of gray hole a Byzantine black hole.
 In the following part, we discuss the results we have obtained.
-Our Contribution: We, in this paper, investigate the perpetual exploration problem, by a team of synchronous mobile agents, of a ring $R$ of size $n$, in presence of a byzantine black hole.
+Our Contribution: In this paper, we investigate the perpetual exploration problem, by a team of synchronous mobile agents, of a ring $R$ of size $n$, in the presence of a Byzantine black hole.
 First, we consider the case when the agents are initially co-located.
 We obtain the following results.
-A: For Face-to-Face model of communication we obtain that $\lceil\log(n-1)\rceil+3$ agents are sufficient to perpetually explore $R$.
-B: For Pebble model of communication, we obtain that 3 agents are necessary and sufficient to perpetually explore $R$.
-C: For Whiteboard model as well, we achieve the same lower and upper bounds as mentioned in B.
-This result shows that, by weakening the scheduler from asynchronous (as assumed in [2]) to synchronous, the tight bound on number of agents in order to perpetually explore $R$, reduces from 4 to 3.
+A: For Pebble model of communication, we obtain that 3 agents are necessary and sufficient to perpetually explore $R$.
+B: For Face-to-Face model of communication, we obtain that 5 agents are sufficient to perpetually explore $R$.
+C: For Whiteboard model as well, we achieve the same lower and upper bounds as mentioned in A.
+This result shows that, by considering the scheduler to be synchronous instead of asynchronous (as assumed in [2]), the tight bound on the number of agents to perpetually explore $R$, reduces from 4 to 3.
 Next, we consider the case, when the agents are initially scattered, and in this context, we obtain the following results:
-D: For Pebble model of communication, we show that 4 agents are necessary and sufficient to perpetually explore $R$.
-E: For Whiteboard model of communication, we obtain an improved bound of 3 agents (in comparison to D), which is necessary and sufficient to perpetually explore the ring $R$.
-In the following table, we have summarized the results.
+D: For Pebble model of communication, we show that 4 agents are necessary and sufficient to explore $R$ perpetually .
+E: For Whiteboard model of communication, we obtain an improved bound of 3 agents (in comparison to D), which is necessary and sufficient to explore the ring $R$ perpetually.
+In the following Table 1, we have summarized the results.
 Whiteboard Pebble Face-to-Face
-Co-located Upper Bound 3 3 $\lceil\log(n-1)\rceil+3$
-Lower Bound 3 3
+Co-located Upper Bound 3 3 5
+Lower Bound 3 3 3
 Scattered Upper Bound 3 4
 Lower Bound 3 4 Non-Constant [9]
-Table 1: Table of results
+Table 1: Summary of our results
+Organisation: Rest of the paper is organised as follows.
+Section 2 presents model and preliminaries.
+In Section 3, we discuss some impossibility results.
+In Section 4 and 5, we propose perpetual exploration algorithms for the agents, when the agents are initially co-located and scattered, respectively.
+Lastly, we conclude in Section 6.
+Note that due to page restriction, detailed description of the algorithms and proofs of theorems and lemmas are referred to the Appendix.
 2 Model and Preliminaries
-In this paper , the considered underlying topology of the network is an oriented ring $R=\{v_{0},v_{1},\dots v_{n-1}\}$.
-Each node $v_{i}$ (where $i\in\{0,1,\dots,n-1$}) is unlabeled and has two ports connecting $v_{(i-1)\bmod{n}}$ and $v_{(i+1)\bmod{n}}$ labeled left and right consistently.
-A set $A=\{a_{0},a_{1},\dots a_{k-1}\}$ of $k$ agents operate in $R$.
-We consider two types initial position for the set $A$ of agents.
-In the first type we consider, each agent in $A$ is co-located at a node, which we term as $home$.
-In the second type we consider, the agents can start from several distinct nodes, which we term as scattered initial position.
-Each agent has the knowledge of the underlying topology $R$ and has some computational capabilities, thus it requires $\mathcal{O}(\log n)$ bits of internal memory.
-The agents have unique IDs of size $\lceil\log k\rceil$ bits taken from the set $[1,k]$, which are perceived by other agents when they are co-located.
-The agents are autonomous and they execute same set of rules (i.e., they execute the same algorithm).
-We consider three types of communication that the agents have, in order to communicate with other agents. These models of communication are as follows:
-  • Face-to-Face: In this model of communication, an agent can communicate with another agent when they are co-located.
-  • Pebble: In this model, the agents are equipped with a movable token (also termed as pebble), which signifies a single bit of information, the agents can use this pebble in order to communicate with other agents.
-  • Whiteboard: In this case, each node of $R$ contains $\mathcal{O}(\log k)$ bits of memory, which can be used to store and maintain information.
-Any agent can read the existing information or write any new information at the whiteboard of its current node.
-Note that, fair mutual exclusion is maintained, i.e., concurrent access of the whiteboard data is not permitted.
-The agents operate in synchronous rounds and in each round every agent becomes active and takes a local snapshot of its surroundings.
-For an agent at a node $v$ in some round $r$, the snapshot contains two ports incident to $v$, already stored data’s on the memory of $v$ (if any, only in case of whiteboard model of communication), number of pebbles located at $v$ (if any, only in case of pebble model of communication), contents from its own local memory and IDs of other agents on $v$.
-Based on this snapshot an agent executes some action.
-This actions includes a communication step and a move step.
-In a communication step an agent can communicate implicitly or explicitly with other agents according to the communication models discussed above.
-In the move step the agent can move to a neighbour node by following a port incident to $v$.
-Thus if an agent at node $v$ in round $r$, decides to move during the move step, at round $r+1$ it resides on a neighbour node of $v$.
-All these actions are atomic, so an agent cannot distinguish another agent concurrently passing through the same edge, instead it can only interact with another agent (based on its communication model) only when it reaches another node.
-A black hole is a stationary malicious node in an underlying graph, which has the ability to destroy any visiting agent, without leaving any trace of its existence.
-In this paper we have considered one unique node of $R$ to be a black hole.
-Further, the black hole nature of the node is controlled by an adversary.
-We term this kind of node as Byzantine Black Hole, which is defined as follows.
-Definition 2.1 (Byzantine Black Hole).
-Let $G$ be an underlying graph of a network with agents.
-A node $u$ of $G$ is called a Byzantine black hole if sometimes it behaves as a black hole, whereas, in rest of the time it behaves as a normal node.
-Moreover, this transition between the black hole nature and normal nature of the node is controlled by the adversary.
-We, in this paper, have assumed that the underlying graph contains a single byzantine black hole, whereas the other nodes are normal nodes, and they are termed as safe nodes.
-Note that, the starting positions of each agent must be a safe node.
-Here, we assume that, if the adversary decides to activate the black hole nature of the byzantine black hole node, then it does so at the beginning of its corresponding round, and the node retains that nature until the end of this current round.
-Furthermore, we have considered that our byzantine black hole has the ability to always destroy any incoming agent, during its black hole nature.
-In addition it also has the ability to choose whether to destroy any information present on that node during the black hole nature of that node.
-So, our aim in this paper, is to perpetually explore the ring $R$, with minimum number of agents. Next, we formally define our problem.
+In this paper, we consider the underlying topology of the network as an oriented ring $R=\{v_{0},v_{1},\dots,v_{n-1}\}$.
+Each node $v_{i}$ (where $i\in\{0,1,\dots,n-1\}$) is unlabeled and has two ports connecting $v_{(i-1)\bmod{n}}$ and $v_{(i+1)\bmod{n}}$, labeled consistently as left and right.
+A set $A=\{a_{0},a_{1},\dots,a_{k-1}\}$ of $k$ agents operates in $R$.
+We consider two types of initial positions for the set $A$ of agents.
+In the first type, each agent in $A$ is co-located at a node, which we term as their home.
+In the second type, the agents can start from several distinct nodes, which we term as scattered initial positions.
+Each agent has knowledge of the underlying topology $R$ and possesses some computational capabilities, thus requiring $\mathcal{O}(\log n)$ bits of internal memory.
+The agents have unique IDs of size $\mathcal{O}(\log k)$ bits taken from the set $[0,k^{c}]$ ($c$ is a constant), which are perceived by other agents when they are co-located.
+The agents are autonomous and execute the same set of rules (i.e., they execute the same algorithm).
+We consider three types of communication that the agents have in order to communicate with other agents.
+Face-to-Face (F2F): In this model, an agent can communicate with another agent when they are co-located.
+Pebble: In this model, the agents are equipped with a movable token (also termed as “pebble”), which signifies a single bit of information.
+The agents can carry a pebble from one node in a fairly mutually exclusive way and also can drop it on any other node.
+Agents use this pebble to mark some special nodes, for other agents to distinguish it.
+Whiteboard: In this case, each node of $R$ contains $\mathcal{O}(\log n)$ bits of memory, which can be used to store and maintain information.
+Any agent can read the existing information or write any new information on the whiteboard of its current node.
+Note that fair mutual exclusion is maintained, i.e., concurrent access to the whiteboard data is not permitted.
+The agents operate in synchronous rounds, and in each round, every agent becomes active and takes a local snapshot of its surroundings.
+For an agent at a node $v$ in some round $r$, the snapshot contains two ports incident to $v$, already stored data on the memory of $v$ (if any, only in the case of the whiteboard model of communication), the number of pebbles located at $v$ (if any, only in the case of the pebble model of communication), contents from its local memory, and IDs of other agents on $v$.
+Based on this snapshot, an agent executes some action.
+This action includes a communication step and a move step.
+In a communication step, an agent can communicate implicitly or explicitly with other agents according to the communication models discussed above.
+In the move step, the agent can move to a neighbouring node by following a port incident to $v$.
+Thus, if an agent at node $v$ in round $r$ decides to move during the move step, in round $r+1$, it resides on a neighbour node of $v$.
+All these actions are atomic, so an agent cannot distinguish another agent concurrently passing through the same edge; instead, it can only interact with another agent (based on its communication model) when it reaches another node.
+A black hole is a stationary malicious node in an underlying graph, which has the ability to destroy any visiting agent without leaving any trace of its existence.
+Furthermore, the black hole nature of the node is controlled by an adversary.
+In addition to this, whenever the black hole nature is activated, the adversary can also choose to destroy any information stored at that node.
+We term this kind of node a Byzantine Black Hole.
+In this paper, we assume that the underlying graph contains a single Byzantine black hole, while the other nodes are normal nodes, termed as safe nodes.
+It is assumed that the starting positions of each agent must be a safe node.
+This Byzantine black hole node is unknown to the agent.
+Here, we assume that if the adversary decides to activate the black hole nature of the Byzantine black hole node, it does so at the beginning of its corresponding round, and the node retains that nature until the end of this current round.
+Furthermore, we have considered that our Byzantine black hole has the ability always to destroy any incoming agent during its black hole nature and also choose to destroy any information present on that node.
+This paper aims to perpetually explore the ring $R$ with the minimum number of agents.
+Next, we formally define our problem:
 Definition 2.1 (PerpExploration-BBH).
-Given a ring network $R$ with $n$ nodes, where one node ($v_{b}$) is a Byzantine black hole, and with a set of agents $A$ positioned on R, the PerpExploration-BBH asks the agents in $A$ to move such a way that, each node of $R$ except $v_{b}$ is visited by at least one agent infinitely often.
+Given a ring network $R$ with $n$ nodes, where one node ($v_{b}$) is a Byzantine black hole, and with a set of agents $A$ positioned on $R$, the PerpExploration-BBH, asks the agents in $A$ to move in such a way that each node of $R$, except $v_{b}$, is visited by at least one agent infinitely often.
 3 Impossibility Results
-In the first result we prove that two agents can not solve PerpExploration-BBH in the strongest communication model of whiteboard.
+Here, at beginning, we state the first impossibility result which gives us a lower bound on minimum number of agents required to solve PerpExploration-BBH.
 Theorem 3.1.
-A set of two synchronous agents in a ring $R$ of size $n$ cannot solve PerpExploration-BBH, even in presence of whiteboard if number of possible consecutive black hole positions is greater or equals to 3.
-Proof 3.2.
-Let $v_{1},v_{2}$ and $v_{3}$ be three possible consecutive black hole positions in a ring $R$.
-Let two agents $a_{1}$ and $a_{2}$ are sufficient to solve the PerpExploration-BBH problem on $R$.
+A set of two synchronous agents in a ring $R$ of size $n$ cannot solve PerpExploration-BBH, even in the presence of a whiteboard if number of possible consecutive black hole positions is at least 3.
+The main idea of the proof of the above Theorem 3.1 is that we create a scenario using adversarial techniques, where after one agent is destroyed by the Byzantine black hole $v_{b}$ (say), the other and only alive agent ends up with at least two possible choices for the position of the black hole, which in turn creates confusion for the agent.
+Hence it is unable to correctly detect the black hole node among these two sets of nodes.
+In this scenario it is impossible for the only alive agent to successfully explore the whole ring except $v_{b}$ perpetually, either without correctly detecting the Byzantine black hole position, or without getting destroyed as well.
+As a direct implication of Theorem 3.1, we can have the following corollaries.
+(Please see Appendix in Section 7 for the detailed proof)
+Corollary 3.2.
+A set of three synchronous agents are required to solve the PerpExploration-BBH on a ring $R$ with $n$ nodes, where each node of $R$ has a whiteboard.
+Note that this lower bound of 3 agents is also true for agents with the pebble model of communication, as the pebble model of communication can be easily simulated in the whiteboard communication model.
+Corollary 3.3.
+A set of two agents, each equipped with $O(\log n)$ pebbles, can not solve the PerpExploration-BBH problem on a ring $R$ with $n$ nodes.
+Our next result further improves the lower bound for number of agents, when agents are scattered and have pebble model of communication.
+Theorem 3.4.
+A set of 3 scattered agents, each equipped with a pebble, can not solve the PerpExploration-BBH problem on a ring $R$ with $n$ nodes.
+We have proved Theorem 3.4 using an adversarial technique.
+We created an instance of three copies of the same ring $R$ (i.e., $R_{1}$, $R_{2}$ and $R_{3}$) in such a way that the position of Byzantine black hole ($v_{b})$ are different in each of them but they are consecutive in $R$.
+Also, in each case, the agents are initially placed at the nodes of $R$, which are at a distance of equal length between each other.
+Now, when an agent is destroyed by $v_{b}$, the other two agents can not distinguish between these three copies (due to same information on nodes for each of them).
+So, the problem reduces to two agents with 3 pebbles solving PerpExploration-BBH where a number of possible consecutive black hole positions is at least three.
+Then by Corollary 3.3, we can have the desired result (For detailed proof, see Appendix in Section 8).
+Next, corollary is a direct consequence of Theorem 3.4.
+Corollary 3.5.
+A set of four scattered agents, each with a pebble, are required to solve PerpExploration-BBH on a ring $R$ with $n$ nodes.
+4 Perpetual Exploration with Co-located Agents
+In this section we assume that initially agents are co-located at a node which is termed as home.
+On the basis of this assumption here we investigate the sufficiency for the number of agents to solve PerpExploration-BBH problem under different model of communication.
+4.1 Pebble Model of Communication
+In this section, we consider the communication model, where the starting node has $k-1$ identical and movable tokens (termed as pebbles), where $k$ is the total number of agents deployed.
+A pebble can be carried by an agent from one node to another.
+This pebble acts as a mode of communication for the agents, as the agents can perceive the presence of a pebble at the current node.
+Moreover, an agent can also perceive the presence of other agents which are co-located at the current round (i.e., gather the IDs of the other co-located agents).
+Our main aim is to prove the following theorem in this section.
+Theorem 4.1.
+A team of 3 co-located, synchronous agents are necessary and sufficient to solve PerpExploration-BBH on a ring $R$, with $n$ nodes under the pebble model of communication,with the presence of two pebbles initially co-located with the agents.
+The necessary part follows from Corollary 3.3.
+For the sufficiency part we present an algorithm, PerpExplore-Coloc-Pbl.
+We have provided a very brief description of the algorithm.
+For detailed description and pseudo code see Appendix (Section 9).
+Brief Description of the Algorithm: Given a ring $R$ with a Byzantine black hole node $v_{b}$, a set of agents, $A=\{a_{1},a_{2},a_{3}\}$ (where we assume, ID of $a_{1}\leavevmode\nobreak\ <$ ID of $a_{2}\leavevmode\nobreak\ <$ ID of $a_{3}$), are initially co-located at $home$.
+The agents have no knowledge about the position of the node $v_{b}$, the only set of knowledge the agents have are: the total number of nodes in the underlying topology $R$ (i.e., $n$), and also the knowledge that the initial node, i.e., $home$ is safe.
+So, the remaining arc of $n-1$ nodes in $R$ is a suspicious region (which we term as $S$, where the cardinality of $S$ i.e., $|S|$ denotes the length or size of the suspicious region) for each of the three agents.
+The main idea of this algorithm is as follows: initially two agents $a_{1}$ and $a_{2}$ explores the ring perpetually, while $a_{3}$ waits at the $home$.
+$a_{1}$ and $a_{2}$ explores the ring $R$, executing the rule described as follows.
+If all the three agents are at $home$ along with two pebbles, then in the $r-$th round ($r\geq 0$), $a_{1}$ moves clockwise with one pebble and it waits at round $r+1$ for $a_{2}$ to arrive.
+In round $(r+1)$, $a_{2}$, leaving the remaining pebble at $home$ moves to the next node along clockwise direction to meet $a_{1}$.
+In the subsequent round, i.e., $(r+2)-$th round, $a_{1}$ after meeting $a_{2}$ moves again to the next node in the clockwise direction and waits for 3 rounds (i.e., till $(r+5)-$th round) for $a_{2}$ if and only if $a_{1}$ is not at $home$.
+In the mean time at $(r+3)-$th round, $a_{2}$ leaves its current node, moves one step in counter-clockwise direction, collects the left behind pebble.
+Then in $(r+4)-$th round $a_{2}$ moves again to the earlier node in clockwise direction.
+Subsequently, in $(r+5)-$th round, $a_{2}$ again leaves the pebble at its current node, and moves in a clockwise direction, to meet $a_{1}$ (which is waiting at the current node for $a_{2}$ to arrive), for better reference see Fig.
+1.
+Note that, when $a_{2}$ meets $a_{1}$ outside $home$, the pebble it was carrying is left by $a_{2}$ at the nearest counter-clockwise node of the meeting node (i.e., where $a_{2}$ meets with $a_{1}$).
+Now from round $(r+6)$ onwards to $(r+9)-$th round, the same execution repeats, as it has occurred between round $(r+2)$ to round $(r+5)$.
+The only difference is that, if both agents (i.e., $a_{1}$ and $a_{2}$) at round $r+2$ were at a node $v_{1}$, at round $r+6$, they are on the nearest clockwise node of $v_{1}$ (i.e., at $v_{2}$, say).
+This process continues until $a_{1}$ reaches $home$.
+In this case $a_{1}$ does not move clockwise, until $a_{2}$ meets it with the pebble it was carrying, in that case, it waits further until $a_{2}$ brings that pebble back at $home$.
+For this to happen after reaching $home$, 5 rounds are sufficient for waiting.
+When $a_{2}$ reaches $home$ with the pebble, then all 3 agents are again at the $home$ with 2 pebbles.
+This way the 3 agents explore the ring.
+The exploration can hamper if either $a_{1}$ or, $a_{2}$, or a pebble is destroyed by the Byzantine black hole $v_{b}$ while the above mentioned procedure is executed by them.
+Figure 1: An execution of PerpExplore-Coloc-Pbl, starting from the configuration where $a_{0}$ and $a_{1}$ are together on a vertex, to the configuration where $a_{0}$ and $a_{1}$ are on the same vertex again, which is the clockwise neighbour of the earlier vertex.
+The red node is the byzantine black hole, the green node is the $home$ and red boxes are pebbles.
+Firstly, let only $a_{1}$ is destroyed at some round $t$.
+Then there must exists a round $t^{\prime}<t$, when $a_{2}$ was with $a_{1}$ at some node $v_{0}$ (say) for the last time.
+This means at round $t^{\prime}$, $a_{1}$ moved clockwise to the next node $v_{1}$ along with the pebble it was carrying, and it must have been destroyed before or at the round when $a_{2}$ reaches $v_{1}$ again (in this case $v_{1}$ is $v_{b}$).
+Now when $a_{2}$ reaches $v_{1}$ there can be two cases.
+Either $a_{2}$ remains alive or (as the adversary may choose not to destroy $a_{2}$), it can get destroyed as well.
+For the initial case, $a_{2}$ identifies the Byzantine black hole by not seeing $a_{1}$ there and then it can leave that node and can explore the ring perpetually avoiding the node.
+The latter case is equivalent to both $a_{1}$ and $a_{2}$ are destroyed by the Byzantine black hole $v_{b}=v_{1}$.
+Note that since outside $home$ whenever $a_{2}$ reaches the same node of $a_{1}$, it leaves behind a pebble at the nearest counter-clockwise node (here $v_{0}$).
+For this case, another agent (i.e., $a_{3}$) which was waiting at $home$, finds none of the two exploring agents returns at $home$, even after waiting for a sufficient number of rounds.
+This incident triggers $a_{3}$ to move clockwise until it finds a pebble (one that is left behind by $a_{2}$ at $v_{0}$).
+Whenever the pebble is found, $a_{3}$ knows that the next clockwise node is the Byzantine black hole and it starts exploring $R$ avoiding that node.
+Now, there can be two other cases that can hamper the above mentioned exploring procedure.
+For the first case, let only $a_{2}$ is destroyed at some round $t>0$.
+Let $t^{\prime}<t$ be the last round before $t$ when both $a_{1}$ and $a_{2}$ were together at a node $v_{0}$.
+At round $t^{\prime}$, $a_{1}$, moves clockwise to the next node $v_{1}$.
+Now if $a_{2}$ is not destroyed in the next 3 rounds, then it meets $a_{1}$ at $v_{1}$ again, contrary to our assumption.
+So, if only $a_{2}$ is destroyed then it must be at any of the rounds $t^{\prime}+1,t^{\prime}+2$ or $t^{\prime}+3$.
+In this 3 round $a_{2}$ can be either on $v_{0}$ or $v_{-1}$ where, $v_{-1}$ is the counter-clockwise nearest node of $v_{0}$.
+In this case, when $a_{1}$ finds $a_{2}$ has not arrived at $v_{1}$ even after waiting for 3 rounds, in which case, it understands that either $v_{0}$ or $v_{-1}$ is the Byzantine black hole.
+This knowledge triggers $a_{1}$ to move clockwise until it meets $a_{3}$ at $home$, leaving the pebble along with it behind at $v_{1}$.
+When $a_{3}$ sees that even after certain round of waiting only $a_{1}$ is able to arrive at $home$, and that too without the pebble it was supposed to be carrying, $a_{3}$ understands that $a_{1}$ must have detected an anomaly.
+In this case both $a_{1}$ and $a_{3}$ moves counter-clockwise together to find the pebble left by $a_{1}$.
+When they find it they declare $v_{1}$ as $home$.
+After this they start exploring the ring in different directions in the following way.
+$a_{1}$ moves clockwise until $v_{-1}$ and $a_{3}$ starts moving counter-clockwise until $v_{0}$.
+After that, they both move in opposite direction until they again reach $home$ (which is the new $home$) and waits for sufficient number of rounds to meet each other.
+This way the exploration keeps going on until one among $a_{1}$ or $a_{3}$ is destroyed.
+Let $a_{3}$ fails to reach $home$.
+In that case, $a_{1}$ will detect that $v_{0}$ must be the Byzantine black hole and it can start exploring the ring $R$ avoiding $v_{0}$.
+On the other hand if $a_{1}$ fails to reach then it must be destroyed by the Byzantine black hole which is nothing but the node $v_{-1}$.
+Knowing this $a_{3}$ then explores the ring avoiding that node.
+Now there can be another case when no agents but a pebble is destroyed at $v_{b}$.
+This pebble must be the pebble that is left behind at $v_{-1}$ by the agent $a_{2}$ when it reaches $v_{0}$ to meet $a_{1}$.
+In this case, only the pebble can be destroyed before $a_{2}$ reaches to collect the pebble back.
+So, when $a_{2}$ reaches $v_{-1}$, it finds out that there is no pebble.
+This leads to $a_{2}$, knowing that $v_{-1}$ is the Byzantine black hole and in which case it starts exploring the ring $R$ avoiding that node.
+If $a_{2}$ is also destroyed while it reaches $v_{-1}$ to collect the pebble, this case is similar to the case where we described the algorithm when only $a_{2}$ is destroyed.
+Sketch of Correctness: To prove the correctness of algorithm PerpExplore-Coloc-Pbl we prove the following Theorem. Here we just present a sketch of the proof, for details see Appendix (Section 9.7).
+Theorem 4.2.
+Algorithm PerpExplore-Coloc-Pbl solves PerpExploration-BBH on a ring $R$ with 3 co-located and synchronous agents under the pebble model of communication with the presence of two pebbles initially co-located with the agents.
+We first prove that if no agents are destroyed then either the agents continue with the exploration of ring $R$ without knowing the exact location of the Byzantine black hole, or there exists at least one agent that knows the exact location of the Byzantine black hole which can then explore the ring indefinitely avoiding the Byzantine black hole (See Lemma 9.1 in Appendix).
+Next we proved that, if one agent is destroyed while exploring the ring, then either there exists one agent that identifies the Byzantine black hole uniquely and continues to indefinitely explore all nodes of the ring, except the Byzantine black hole or the exploration continues while the length of the suspicious region (i.e., $|S|$) decreases to 2 from $n-1$ (Lemma 9.3 in Appendix).
+For the later case we proved that the exploration continues until another agent is destroyed by the Byzantine black hole (Remark 9.5 in Appendix).
+In this case when two agents are destroyed by the Byzantine black hole, the only alive agent can uniquely identify the Byzantine black hole without moving on to it (Lemma 9.6 in Appendix).
+Thus it can then perform the exploration of $R$ avoiding the Byzantine black hole.
+This proves Theorem 4.2.
+4.2 Face-to-Face Model of Communication
+In this section, we consider the Face-to-Face (also termed as F2F) model and prove the following theorem.
+Theorem 4.3.
+A team of 5 co-located and synchronous agents are sufficient to solve PerpExploration-BBH on a ring $R$ with $n$ nodes under the F2F model of communication.
+In order to prove Theorem 4.3, we discuss an algorithm with 5 co-located agents, where each agent can communicate among themselves at the same node at the same round (i.e., each agent have F2F model of communication).
+Initially 3 lowest ID agents, say, $a_{1}$, $a_{2}$ and $a_{3}$ are chosen, where fourth lowest ID agent say, $a_{4}$, is associated with $a_{1}$ and the largest ID agent, say $a_{5}$, is associated with $a_{2}$.
+The idea of the algorithm resembles to that of the algorithm PerpExplore-Coloc-Pbl.
+Note that as in this case there is no existence of pebbles, hence, we configure the behaviour of the agents $a_{4}$ and $a_{5}$ in such a way that they act as pebbles, associated with $a_{1}$ and $a_{2}$ in PerpExplore-Coloc-Pbl, respectively.
+In PerpExplore-Coloc-Pbl when we say that an agent $a_{i}$ carries a pebble $p$, in this case we mean that the agent $a_{i}$ communicates a message carry with its associated agent $a_{i^{\prime}}$ such that both these agents simultaneously move together until further new instruction is communicated.
+On the contrary as per PerpExplore-Coloc-Pbl when we say that an agent $a_{i}$ drops a pebble $p$ at a node $v$, then in this case we mean that $a_{i}$ communicates a message drop with $a_{i^{\prime}}$, which in turn instructs $a_{i^{\prime}}$ not to move further and remain stationary at the node $v$ until further instruction is communicated.
+Hence, with this terminology, a team of 5 agents can execute the algorithm PerpExplore-Coloc-Pbl, and the correctness also follows similarly.
+This shows that a team of 5 co-located and synchronous agents are sufficient to solve PerpExploration-BBH under F2F model of communication.
+This proves Theorem 4.3.
+4.3 Whiteboard Model of Communication
+The algorithm PerpExplore-Coloc-Pbl can be simulated in whiteboard model of communication.
+A pebble on a node can be simulated by a bit of information, which is marked on the whiteboard of that node.
+Note that, collecting a pebble from the node is simulated by erasing the same bit of information, on that node and dropping the pebble can be simulated by marking the node with a bit of information as well on the whiteboard of that node.
+From this we get the following result for whiteboard model of communication.
+Theorem 4.4.
+A team of 3 synchronous co-located agents are necessary and sufficient to solve PerpExploration-BBH on a ring $R$ if each node are equipped with a whiteboard having constant memory.
+Previously in [2], for asynchronous scheduler the tight bound for number of agents to solve this problem was 4.
+Now from Theorem 4.4, we get a trade-off between reducing the optimal number of agents required vs the scheduler, in presence of a more generalized version of gray hole as well (i.e., Byzantine black hole).
+5 Perpetual Exploration with Scattered Agents
+Here in this section, we discuss the problem of PerpExploration-BBH, when the agents are initially scattered on more than one nodes, where each of these starting nodes are assumed to be safe.
+We investigate this problem, under different communication models.
+First, we discuss the pebble model of communication and subsequently we discuss the whiteboard model of communication.
+5.1 Pebble Model of Communication
+Our goal in this section is to prove the following theorem.
+Theorem 5.1.
+A team of 4 synchronous scattered agents with one pebble each is necessary and sufficient to solve the PerpExploration-BBH problem on a ring $R$ with $n$ nodes, when the agents are initially scattered on $R$.
+The necessary part follows from Corollary 3.5.
+For the sufficient part, we present an algorithm PerpExplore-Scat-Pbl, that can solve the PerpExploration-BBH problem in the above context.
+In this algorithm we assumed that 4 agents are initially scattered at 4 different nodes of $R$.
+To include the remaining cases where the agents are initially scattered at less than or equal to 3 nodes, a slight modification of PerpExplore-Scat-Pbl is enough which we describe in Remark 10.7 in Appendix.
+Here we describe only the main idea of the algorithm.
+The detailed description and pseudocodes of the algorithm can also be found at the Appendix (Section 10.1).
+Brief Idea of the Algorithm: Let us consider, the starting position of $a_{0}$ to be the first, after which the starting position of $a_{1},a_{2}$ and $a_{3}$, respectively follows in clockwise order.
+Let $h_{i}$ be the starting node of an agent $a_{i}$ (i.e., $home$ of $a_{i}$), where $i\in\{0,1,2,3\}$.
+By $Seg(a_{i})$ we define the clockwise arc starting from the node $h_{i}$ and ending at $h_{(i+1)\pmod{4}}$ (called segment of $a_{i}$).
+If no agents are destroyed by the Byzantine black hole $v_{b}$ then the exploration of $R$ goes as follows.
+Agent $a_{i}$ moves clockwise along $Seg(a_{i})$ leaving its pebble at $h_{i}$ until it reaches $h_{(i+1)\pmod{4}}$ (i.e., the end of $Seg(a_{i})$).
+An agent can distinguish the node $h_{(i+1)\pmod{4}}$ by seeing the pebble left there by the agent $a_{(i+1)\pmod{4}}$.
+When $a_{i}$ reaches $h_{(i+1)\pmod{4}}$ traversing $Seg(a_{i})$ for the first time, it knows the length of the segment $Seg(a_{i})$ and stores the length in its own memory.
+For further traversals it does not depend on seeing pebbles at the end nodes of the segment.
+It can simply use the length of the segment.
+After $a_{i}$ reaches $h_{(i+1)\pmod{4}}$, it waits there for a certain number of rounds so that the other agents can in the meantime reach the endpoints of there corresponding segments, while moving along a clockwise direction.
+After this waiting, $a_{i}$ collects the pebble (if exists) from $h_{(i+1)\pmod{4}}$ (the pebble left by $a_{(i+1)\pmod{4}}$) and starts moving counter-clockwise along with the collected pebble until it reaches its own $home$, i.e., $h_{i}$ along with the pebble it was carrying.
+The waiting time at $h_{(i+1)\pmod{4}}$ also ensures that all agents starts moving counter-clockwise at the same round.
+Also note that, if $a_{i}$ does not find any pebble to collect at $h_{(i+1)\pmod{4}}$ it starts moving counter-clockwise without any pebble (this case can only happen if $a_{(i+1)\pmod{4}}$ is destroyed at the Byzantine black hole node $v_{b}$, while it was returning back after collecting the pebble from $h_{(i+2)\pmod{4}}$ in some previous round).
+After $a_{i}$ reaches $h_{i}$ moving counter-clockwise, it again waits for another set of rounds before it repeats the whole process again.
+Again, the waiting time at $h_{i}$ is configured in such a way that, all agents start repeating the process at the same round (for details on the precise value of these waiting times, refer to the detailed description of the algorithm in Appendix in Section 10.1).
+Note that, since $\cup_{i=0}^{3}Seg(a_{i})=R$, if no agents are destroyed, the perpetual exploration of $R$ happens.
+Figure 2: (a) $h_{i}$’s are $home$ marked as green.
+Agent $a_{i}$ is on $h_{i}$ initially with a pebble.
+We name the pebble initially at $h_{i}$ as $p_{i}$, but in reality they are anonymous.
+(a-b) Each agent moves clockwise until the next $home$ (i.e., $h_{(i+1)\pmod{4}}$) without carrying any pebble.
+The agents already reached ( here $a_{1}$ and $a_{3}$) waits for others.
+(c-e) All agents are on their clockwise nearest $home$.
+They start moving counter clockwise together with the pebble present at their current location towards their initial $home$.
+The agents which already reaches their $home$, wait for others to reach their $home$.
+The exploration can be hampered only if, an agent is destroyed at the Byzantine black hole node $v_{b}$.
+Note that, since $Seg(a_{i})\cap Seg(a_{j})$ is either empty or consists of a safe node and in addition to that, as a segment is explored by only one agent.
+Hence, at most one agent can be destroyed while exploring their respective segments, as described above.
+Without loss of generality, let $v_{b}\in Seg(a_{j})$, for some $j\in\{0,1,2,3\}$.
+Now there are two cases.
+Case-I: Let $a_{j}$ is destroyed while moving clockwise.
+For this case, $a_{j}$ fails to collect the pebble at $h_{(j+1)\pmod{4}}$ left by $a_{(j+1)\pmod{4}}$ and return $h_{j}$.
+So, when $a_{(j+1)\pmod{4}}$, returns $h_{(j+1)\pmod{4}}$ after moving counter-clockwise along with the pebble it has collected from $h_{(j+2)\pmod{4}}$, it finds two pebble at $h_{(j+1)\pmod{4}}$.
+This is considered by $a_{(j+1)\pmod{4}}$ as an anomaly and it learns that $v_{b}$ must be in the arc, which is in the counter-clockwise direction starting from $h_{(j+1)\pmod{4}}$.
+In this scenario, $a_{(j+1)\pmod{4}}$ waits a certain number of rounds (if needed) to ensure that every other alive agents have reached their corresponding $home$.
+Then $a_{(j+1)\pmod{4}}$ starts moving clockwise with both the pebbles.
+The aim of this move is to meet and gather with the other alive agents.
+Note that the other alive agents wait at their corresponding $home$, after moving along the counter-clockwise direction.
+It is because, at their respective $home$ they do not detect any anomaly.
+The waiting time is provided in such a way that it is enough for $a_{(j+1)\pmod{4}}$ to detect anomaly and meet both the remaining alive agents after moving clockwise, while the other agents are waiting at their $home$.
+The agent $a_{(j+1)\pmod{4}}$ first meets $a_{(j+2)\pmod{4}}$ at $h_{(j+2)\pmod{4}}$ while it moves clockwise after detecting anomaly.
+Then both of them moves together, while $a_{(j+2)\pmod{4}}$ carries the pebble, which it was earlier carrying back to $home$.
+They move until they meet with $a_{(j+3)\pmod{4}}$ at $h_{(j+3)\pmod{4}}$.
+Note that at this moment 3 agents are at a node (which is $h_{(j+3)\pmod{4}}$), with at least 3 pebbles (one carried by $a_{(j+3)\pmod{4}}$ and two carried by $a_{(j+2)\pmod{4}}$).
+In this case they execute the algorithm PerpExplore-Coloc-Pbl and achieve PerpExploration-BBH of the ring $R$.
+Case-II: Let $a_{j}$ be destroyed at $v_{b}$ while it was moving counter-clockwise along with the pebble it collected from $h_{(j+1)\pmod{4}}$.
+Then in this case, when all alive agents return to their corresponding $home$, none of them finds any anomaly as all agent sees exactly one pebble at their $home$.
+So, they wait and start the exploration again at the same round.
+Note that in this scenario, all agents except $a_{j}$, move clockwise to the end points of their corresponding segments, waits there and collects pebble (if any) and moves back to their corresponding $home$ again.
+Now, since $a_{j}$ was destroyed earlier, the pebble left by $a_{(j+1)\pmod{4}}$ was picked by no agents and thus, while $a_{(j+1)\pmod{4}}$ returns back to $h_{(j+1)\pmod{4}}$ it finds two pebbles and does the same execution as explained in Case-I.
+The basic idea of this algorithm is to gather three agents with at least 3 pebbles (refer Remark 10.1) in the expense of one agent and then execute PerpExplore-Coloc-Pbl to solve the PerpExplroation-BBH problem.
+Sketch of Correctness: To proof the correctness of the algorithm PerpExplore-Scat-Pbl, we have to prove the following theorem.
+Here we only provide the proof idea.
+For details see Appendix (Section 10.2).
+Theorem 5.2.
+Algorithm PerpExplore-Scat-Pbl solves PerpExploration-BBH problem of a ring $R$ with 4 synchronous and scattered agents under the pebble model of communication where each agents are equipped with a pebble.
+First in Corollary 10.4, we proved that algorithm PerpExplore-Scat-Pbl guarantees exploration if no agents are destroyed.
+This corollary is a direct consequence of the Lemma 10.2.
+Further we showed that if an agent is destroyed then exactly one agent detects anomaly and that agent gathers with the remaining two alive agents within finite rounds (Lemma 10.5 in Appendix).
+The rest of the proof follows from Theorem 4.2.
+5.2 Whiteboard Model of Communication
+The main aim of this section is to prove the following theorem.
+Theorem 5.3.
+A team of 3 synchronous agents are necessary and sufficient to solve the problem PerpExploration-BBH on a ring $R$ with $n$ nodes, when each node of $R$ has a whiteboard of $O(\log n)$ bits of memory, irrespective of their starting location.
+The necessary part follows from Corollary 3.2.
+The sufficiency part for co-located initial situation follows from Theorem 4.4.
+For the other cases where the agents are initially scattered at more than one starting nodes, we propose an algorithm PerpExplore-Scat-Whitbrd.
+This algorithm is designed for the case when all the agents are starting at different nodes.
+Note that, this algorithm can be easily modified a bit to include the case where initially three agents are scattered at two distinct nodes (refer the modification in Appendix, Section 11.3).
+For detailed description and the pseudo code, see Appendix (Section 11.1).
+Brief Description of the Algorithm: Let $a_{0},a_{1}$ and $a_{2}$ be three agents starting from the nodes $h_{0},h_{1}$ and $h_{2}$, respectively, where these nodes are in a clockwise order.
+Let $Seg(a_{i})$ (also called ‘Segment of $a_{i}$’) be the clockwise arc starting from $h_{i}$ and ending at $h_{(i+1)\pmod{3}}$.
+The algorithm first ensures that the ring is explored perpetually if no agents are destroyed.
+In order to do this, the algorithm goes as follows.
+An agent, say $a_{i}$, first erases all previously stored data (if any) at $h_{i}$.
+Then it writes the message (home, ID($a_{i}$)) at $h_{i}$ and starts moving clockwise.
+This type of message is called a home type message that indicates it is $home$ of $a_{i}$.
+$a_{i}$ moves clockwise until it reaches $h_{(i+1)\pmod{3}}$.
+It distinguishes $h_{(i+1)\pmod{3}}$ by seeing the home type message left by $a_{(i+1)\pmod{3}}$.
+When $a_{i}$ moves clockwise, it also marks each node of $Seg(a_{i})$, except $h_{i}$ and $h_{(i+1)\pmod{3}}$, by writing right, after erasing any previous such markings (if at all exists) at each such nodes.
+After $a_{i}$ reaches $h_{(i+1)\pmod{3}}$ it waits for a certain number of rounds (if needed), so that the other agents (say $a_{j}$) gets enough time to reach endpoints of their corresponding segment (i.e.,$h_{(j+1)\pmod{3}}$).
+After this waiting, each alive agent $a_{i}$ write the message (visited, ID($a_{i}$)) at $h_{(i+1)\pmod{3}}$ at the same round.
+This type of messages is termed as a visited type message, which indicates that an agent with its corresponding ID, also mentioned in the visited type message, has visited the respective node.
+Then each $a_{i}$ waits for $n$ rounds at $h_{(i+1)\pmod{3}}$ and then starts moving counter-clockwise at the same round.
+$a_{i}$ moves counter-clockwise until it reaches $h_{i}$, i.e., its own $home$.
+While moving counter-clockwise, $a_{i}$ erases previously written right marking (which it marked while moving along clockwise direction) from each nodes of $Seg(a_{i})$ and writes left there upon arriving (except at $h_{i}$ and at $h_{(i+1)\pmod{3}}$).
+When $a_{i}$ reaches its own $home$ (i.e., $h_{i}$), it waits again for a certain number of rounds, upon seeing the visited type message, left there by $a_{(i-1)\pmod{3}}$.
+This waiting period is enough for each of the other alive agents to reach their corresponding $home$.
+After this waiting period is over, all of the agents, starts repeating the same procedure again together, from the same round.
+This procedure ensures that, if no agent is destroyed at the Byzantine black hole node $v_{b}$, then the perpetual exploration of $R$ continues.
+This can only be hampered, only if an agent gets destroyed at the node $v_{b}$.
+Without loss of generality let $v_{b}\in Seg(a_{j})$ for some, $j\in\{0,1,2\}$.
+So only $a_{j}$ can be destroyed at $v_{b}$, while performing this exploration.
+It is because $a_{j}$ is the only agent to visit each node $u$ of $Seg(a_{j})$, where $u\in Seg(a_{j})\backslash\{h_{j},h_{(j+1)\pmod{3}}\}$.
+There can be two cases.
+Case-I: Let $a_{j}$ be destroyed while it is moving in clockwise direction along $Seg(a_{j})$.
+This implies it fails to reach $h_{(j+1)\pmod{3}}$ and also fails to write a visited type message there.
+So, when $a_{(j+1)\pmod{3}}$ returns to its $home$ (i.e., $h_{(j+1)\pmod{3}}$), it finds no visited type message (as it should’ve, if $a_{i}$ was not destroyed).
+$a_{(j+1)\pmod{3}}$ interprets from this that, the segment which is adjacent to its own segment in counter-clockwise direction has the Byzantine black hole and an agent must have been destroyed there while moving clockwise.
+This information triggers it to move clockwise with the aim of gathering with the other agent (i.e., $a_{(j+2)\pmod{3}}$).
+Note that, when $a_{(j+1)\pmod{3}}$ starts moving, the other agent is waiting and the waiting time is sufficient for the moving agent to meet it.
+After they meet, $a_{(j+1)\pmod{3}}$ shares the information about its direction of movement in the whiteboard of $h_{(j+2)\pmod{3}}$.
+With this, they again start moving clockwise together until they reach $h_{j}$.
+The agents can distinguish $h_{j}$ by the home type message written there by $a_{j}$ before it was destroyed.
+Note that in the clockwise direction each node after $h_{j}$ up to the previous node of $v_{b}$ were marked right by $a_{j}$ before it was destroyed.
+So from $h_{j}$, $a_{(j+1)\pmod{3}}$ and $a_{(j+2)\pmod{3}}$ starts moving cautiously clockwise.
+That is, while both agents ($a_{L}$ and $a_{H}$, where $a_{L}$ and $a_{H}$ are the agent with lowest ID and highest ID among $a_{(j+1)\pmod{3}}$ and $a_{(j+2)\pmod{3}}$ respectively) are at a node $v_{0}$, $a_{L}$ moves clockwise to the next node, say $v_{1}$ while the other agent waits at $v_{0}$.
+If at $v_{1}$, $a_{L}$ sees the right marking, it interprets $v_{1}$ is safe.
+So, it comes back to $v_{0}$.
+At $v_{0}$, seeing that $a_{L}$ has returned, $a_{H}$ also interprets $v_{1}$ to be safe.
+So, in the next round both $a_{L}$ and $a_{H}$ moves to $v_{1}$ together and repeats the process from $v_{1}$ again.
+When $a_{L}$ reaches $v_{b}$, it sees no right marking there.
+$a_{L}$ if alive, interprets this by determining the current node to be Byzantine black hole.
+So, it starts perpetual exploration of $R$, except $v_{b}$.
+Otherwise, if $a_{L}$ gets destroyed at $v_{b}$, it does not return to the previous node where $a_{H}$ is waiting.
+Even after waiting, when $a_{H}$ sees $a_{L}$ has not returned, it interprets this incident as the next clockwise node is the Byzantine black hole and starts exploring the ring $R$ avoiding that node.
+Thus for this case the perpetual exploration continues.
+Case-II: Let $a_{j}$ be destroyed at $v_{b}$ while it is moving counter-clockwise.
+In this case both the alive agents $a_{(j+1)\pmod{3}}$ and $a_{(j+2)\pmod{3}}$ find visited type message after returning to their corresponding $home$.
+This is because, $a_{j}$ is destroyed at $v_{b}$ after writing visited type message at $h_{(j+1)\pmod{3}}$.
+So, all alive agents after waiting, starts repeating the exploration procedure again.
+The agents first erase the whiteboard content at their corresponding $home$ and then writes the corresponding home type message before it starts moving clockwise until the end of their corresponding segment.
+Note that since $a_{j}$ was destroyed earlier (even before the repetition starts), $a_{(j-1)\pmod{3}}$ finds the visited type message which was written there earlier by it, is still present (which was supposed to be erased if $a_{j}$ was alive).
+From this information, $a_{(j-1)\pmod{3}}$ interprets that, $v_{b}$ must be in the segment adjacent to its own segment in the clockwise direction.
+It also interprets that, an agent must have been destroyed at $v_{b}$ while it was moving counter-clockwise.
+This is because, if the agent was destroyed while it was moving clockwise then it would have been detected by the other agents when they are at their $home$ (as described in Case-I), which is before the start of the next clockwise move (note that the agents have already started next clockwise move).
+So, this information triggers $a_{(j-1)\pmod{3}}$ to move counter-clockwise with the aim to gather with $a_{(j-2)\pmod{3}}$.
+Since $a_{(j-1)\pmod{3}}$ starts its move while $a_{(j-2)\pmod{3}}$ waits at $home$ and also, since the waiting time is sufficient, $a_{(j-1)\pmod{3}}$ meets $a_{(j-2)\pmod{3}}$ during the waiting period at $h_{(j-1)\pmod{3}}$.
+So, after they meet, $a_{(j-1)\pmod{3}}$, communicates the direction of its move to $a_{(j-2)\pmod{3}}$ on the whiteboard of $h_{(j-1)\pmod{3}}$ and they move together until they reach $h_{(j-2)\pmod{3}}(=h_{(j+1)\pmod{3}})$ together.
+They distinguish the node by the home type message left there by $a_{(j-2)\pmod{3}}$ before the start of moving clockwise.
+Note that, in the counter-clockwise direction, the nodes in $Seg(a_{j})$, starting from the next node of $h_{(j+1)\pmod{3}}$ up to the previous node of $v_{b}$ are the only nodes those were marked left by $a_{j}$ before it was destroyed.
+So, from $h_{(j+1)\pmod{3}}$, the alive agents $a_{(j+1)\pmod{3}}$ and $a_{(j+2)\pmod{3}}$ start moving cautiously in the counter-clockwise direction.
+Among $a_{(j+1)\pmod{3}}$ and $a_{(j+2)\pmod{3}}$, an agent with lowest ID is denoted by $a_{L}$ and the agent with highest ID is denoted by $a_{H}$.
+The cautious walk is same as described in Case-I, except for the fact that here after moving one node in the counter-clockwise direction, the agent $a_{L}$ searches for the left marking.
+If it finds such marking it returns to $a_{H}$ and then both moves counter-clockwise and repeats the process.
+On the other hand if $a_{L}$ does not find any left marking (it must be $v_{b}$) and it stays alive, $a_{L}$ moves out of that node and starts exploring $R$ avoiding that node.
+On the contrary, if $a_{L}$ gets destroyed then $a_{H}$ sees that $a_{L}$ has not returned while it should have.
+From this incident it interprets next counter-clockwise node is $v_{b}$ and it starts exploring $R$ avoiding that node.
+Sketch of Correctness: Here we give a brief glimpse on how we proved the following theorem. For details see Appendix (Section 11.2)
+Theorem 5.4.
+Algorithm PerpExplore-Scat-Whitbrd solves PerpExploration-BBH problem of a ring $R$ with $n$ nodes and with 3 synchronous agents initially scattered under the whiteboard model of communication
+In order to prove this theorem, we first define a cautious start node.
+Let $a_{i}$ be the agent destroyed at the Byzantine black hole first.
+We define cautious start node to be $h_{i}$, if $a_{i}$ was moving clockwise when destroyed, otherwise it is $h_{(i+1)\pmod{3}}$.
+We first prove that after one agent is destroyed, the remaining two agents gather at the cautious start node.
+This result can be found in details in Lemma 11.11 in Appendix.
+Note that $a_{i}$ marks each node on the arc between the cautious start node and the Byzantine black hole with either left or right markings depending on the direction it was moving before it was destroyed.
+We proved that the alive agent can know the direction of $a_{i}$ before it was destroyed.
+This can be found in details in Lemma 11.1 and Lemma 11.3 in Appendix.
+After the alive agents reach the cautious start node, they start moving cautiously looking for the marking based on the direction of $a_{i}$ before it was destroyed.
+We proved that at least an agent among the two moving cautiously will stay alive knowing the exact location of the Byzantine black hole (Lemma 11.13 in Appendix).
+Hence this agent can explore $R$ perpetually avoiding the Byzantine black hole.
+This proves Theorem 5.4.
+6 Conclusion
+The paper addresses perpetual exploration of a ring network in the presence of a malicious node, which we call as a Byzantine black hole.
+This problem, termed as PerpExploration-BBH, is explored under three communication models (Face-To-Face, Pebble, and Whiteboard) considering various initial scenarios (co-located or scattered agents) with the aim of minimizing number of agents.
+We proposed optimal results (in terms of number of agents), for Pebble and Whiteboard communication models, under both initial scenarios.
+Further, an upper bound of 5 agents and a lower bound of 3 agents is provided for Face-to-Face model, in case of co-located agents.
+Future research could focus on proposing an optimal bound for Face-To-Face co-located scenario, whereas proposing constructive lower and upper bounds for Face-To-Face scattered scenario.
+Additionally, investigating this problem in different scheduler models can be another way of future direction.
+Appendix
+7 Proof of Theorem 3.1
+Statement: A set of two synchronous agents in a ring $R$ of size $n$ cannot solve
+PerpExploration-BBH, even in presence of whiteboard if number of possible
+consecutive Byzantine black hole positions is at least 3.
+Proof: Let $v_{1},v_{2}$ and $v_{3}$ be three possible consecutive Byzantine black hole positions in a ring $R$.
+Let two agents $a_{1}$ and $a_{2}$ be sufficient to solve the PerpExploration-BBH problem on $R$.
 Thus there exists an algorithm $\mathcal{A}$ such that $a_{1}$ and $a_{2}$ can solve PerpExploration-BBH by executing $\mathcal{A}$.
@@
 Let us take two copies $R_{1}$ and $R_{2}$ of the same ring $R$.
-In $R_{1}$, $v_{1}$ is the black hole and in $R_{2}$, $v_{2}$ is the black hole.
-Let the adversary in $R_{1}$ destroys $a_{1}$ at round $t-1$ and in $R_{2}$ destroys $a_{1}$ in round $t$.
+In $R_{1}$, $v_{1}$ is the Byzantine black hole and in $R_{2}$, $v_{2}$ is the Byzantine black hole.
+Let the adversary in $R_{1}$ destroy $a_{1}$ at round $t-1$ and in $R_{2}$ destroy $a_{1}$ at round $t$.
 We claim that at round $t$ and $t-1$, $a_{2}$ can not be on either of $v_{1}$ or $v_{2}$.
-Note that at $t$-th round $a_{2}$ can not be on $v_{2}$ otherwise, both $a_{1}$ and $a_{2}$ gets destroyed at the same round and no other agent is left to explore further.
-Similarly at round $t-1$ $a_{2}$ can not be on $v_{1}$.
+Note that at $t$-th round $a_{2}$ can not be on $v_{2}$ otherwise, both $a_{1}$ and $a_{2}$ get destroyed at the same round and no other agent is left to explore further.
+Similarly, at round $t-1$, $a_{2}$ can not be on $v_{1}$.
 Now consider the case when at $t-$ th round $a_{2}$ is at $v_{1}$.
@@
 We first argue that at round $t-1$, $a_{2}$ can not be at $u_{0}$.
-Otherwise, since the whiteboard content at nodes except at $v_{1}$ and $v_{2}$ are same in $R_{1}$ and $R_{2}$ at round $t$ (due to same execution by $a_{1}$ and $a_{2}$ upto round $t-1$ and at round $t$ for $a_{2}$ in both $R_{1}$ and $R_{2}$).
-Now since at round $t$ all previous datas on whiteboard except $v_{1}$ and $v_{2}$ does not help $a_{2}$ to distinguish between $R_{1}$ and $R_{2}$, the problem now can be thought of solving PerpExploration-BBH with one agent where number of possible consecutive black hole position is atleast two.
+Otherwise, since the whiteboard content at nodes except at $v_{1}$ and $v_{2}$ are the same in $R_{1}$ and $R_{2}$ at round $t$ (due to the same execution by $a_{1}$ and $a_{2}$ up to round $t-1$ and at round $t$ for $a_{2}$ in both $R_{1}$ and $R_{2}$).
+Now since at round $t$, all previous data on the whiteboard except $v_{1}$ and $v_{2}$ does not help $a_{2}$ to distinguish between $R_{1}$ and $R_{2}$, the problem now can be thought of solving PerpExploration-BBH with one agent where the number of possible consecutive Byzantine black hole position is atleast two.
 Now this is impossible to solve.
-So, at $t-1$-th round $a_{2}$ can not be on $u_{0}$ We now only have to prove that $a_{2}$ can not be at $v_{2}$ during round $t-1$.
-Let in $R_{2}$, adversary activates the black hole at $v_{2}$ at round $t-1$ which destroys $a_{2}$ at round $t-1$.
-Now since the effect of this destruction of agent $a_{2}$ does not effect the inputs of $a_{1}$ at round $t-1$ on $v_{1}$, it move to $v_{2}$ at round $t$ where the black hole is activated again by adversary destroying $a_{1}$ too.
+So, at $t-1$-th round $a_{2}$ can not be on $u_{0}$.
+We now only have to prove that $a_{2}$ can not be at $v_{2}$ during round $t-1$.
+Let in $R_{2}$, adversary activates the Byzantine black hole at $v_{2}$ at round $t-1$ which destroys $a_{2}$ at round $t-1$.
+Now since the effect of this destruction of agent $a_{2}$ does not affect the inputs of $a_{1}$ at round $t-1$ on $v_{1}$, it moves to $v_{2}$ at round $t$ where the Byzantine black hole is activated again by adversary destroying $a_{1}$ too.
 Thus $a_{2}$ must remain outside of $v_{1}$ and $v_{2}$ during round $t$ and $t-1$ while executing $\mathcal{A}$.
 So, at round $t$ since all whiteboard content are same for $R_{1}$ and $R_{2}$ (except for nodes $v_{1}$ and $v_{2}$ whose content only differs after round $t-2$ for $R_{1}$ and $R_{2}$ which can not be known by $a_{2}$ as it was not there during that rounds), they can not help $a_{2}$ to distinguish between $R_{1}$ and $R_{2}$.
-So the problem now can be thought of as solving PerpExploration-BBH with one agent where number of possible consecutive black hole position is at least 2.
+So the problem now can be thought of as solving PerpExploration-BBH with one agent where the number of possible consecutive Byzantine black hole positions is at least 2.
 And this is impossible.
-So there doesn’t exist any algorithm that solves PerpExploration-BBH with two agents on a Ring $R$ where number of possible consecutive positions of black hole is at least 3.
-Note that since the $n$ can be sufficiently large so that initially number of possible consecutive black hole positions is atleast three we have the following corollary.
-Corollary 3.3.
-It is impossible to solve PerpExploration-BBH on a ring $R$ with $n$ nodes using two synchronous agents even if the nodes are equipped with whiteboards.
-Corollary 3.4.
-A set of three synchronous agents are required to solve the PerpExploration-BBH on a ring $R$ with $n$ nodes, where each node has a whiteboard.
-Corollary 3.5.
-A set of two agents, each equipped with $O(\log n)$ pebbles can not solve the PerpExploration-BBH problem on a ring $R$ with $n$ nodes.
-Now the following theorem improves the lower bound for the case when agents are scattered and each agent is equipped with a pebble, whereas there is no whiteboard present at the nodes of $R$.
-The theorem is as follows.
-Theorem 3.6.
-A set of 3 scattered agents, each equipped with a pebble can not solve the PerpExploration-BBH problem on a ring $R$ with $n$ nodes.
-Proof 3.7.
-Let $a_{1},a_{2}$ and $a_{3}$ be three agents equipped with a pebble each.
-The agents are placed on three nodes $h_{1},h_{2}$ and $h_{3}$ initially in such a way that distance between $h_{i}$ and $h_{j}$ is same for all $i,j\in\{1,2,3\}$, where we consider $h_{j}$ to be the nearest node of $h_{i}$ in the clockwise direction.
-Without loss of generality let this distances be sufficiently large.
-Further, let there exists an algorithm $\mathcal{A}$ that solves the PerpExploration-BBH problem in this setting.
-Let without loss of generality $a_{1}$ be the first agent to explore the third node from its corresponding starting position (i.e., $h_{1}$) in any of the clockwise or counter-clockwise direction, when each agent starts executing the algorithm $\mathcal{A}$.
-Suppose by following $\mathcal{A}$, $a_{1}$ can visit the third node in clockwise direction (without loss of generality) first at a round say $t>0$.
-Let $v_{1},v_{2}$ and $v_{3}$ be those set of three nodes from $h_{1}$ in the clockwise direction.
-Let $C_{1}$,$C_{2}$ and $C_{3}$ be three scenarios where in $C_{i}$, $v_{i}$ is the black hole.
-We claim that, $a_{1}$ can not carry its pebble during any execution of $\mathcal{A}$.
-Otherwise in scenario $C_{1}$, it would be consumed along with its pebble, and since the distances between two consecutive $h_{i}$ are sufficiently large, hence other agents would have no idea that an agent is already consumed.
+So there doesn’t exist any algorithm that solves PerpExploration-BBH with two agents on a Ring $R$ where the number of possible consecutive positions of the Byzantine black hole is at least 3.
+8 Proof of Theorem 3.4
+Statement: A set of 3 scattered agents, each equipped with a pebble can not solve the PerpExploration-BBH problem on a ring $R$ with $n$ nodes.
+Proof: Let $a_{1},a_{2}$, and $a_{3}$ be three agents equipped with a pebble each.
+The agents are placed on three nodes $h_{1},h_{2}$ and $h_{3}$ initially in such a way that the distance between $h_{i}$ and $h_{j}$ is the same for all $i,j\in\{1,2,3\}$, where we consider $h_{j}$ to be the nearest node of $h_{i}$ in the clockwise direction.
+Without loss of generality let this distance be sufficiently large.
+Further, let there exist an algorithm $\mathcal{A}$ that solves the PerpExploration-BBH problem in this setting.
+Let without loss of generality $a_{1}$ be the first agent to explore the third node from its corresponding starting position (i.e., $h_{1}$) in any of the clockwise or counter-clockwise directions, when each agent starts executing the algorithm $\mathcal{A}$.
+Suppose by following $\mathcal{A}$, $a_{1}$ can visit the third node in the clockwise direction (without loss of generality) first at a round say $t>0$.
+Let $v_{1},v_{2}$ and $v_{3}$ be those sets of three nodes from $h_{1}$ in the clockwise direction.
+Let $C_{1}$,$C_{2}$ and $C_{3}$ be three scenarios where in $C_{i}$, $v_{i}$ is the Byzantine black hole.
+We claim that $a_{1}$ can not carry its pebble during any execution of $\mathcal{A}$.
+Otherwise, in scenario $C_{1}$, it would be destroyed along with its pebble, and since the distances between two consecutive $h_{i}$ are sufficiently large, hence other agents would have no idea that an agent is already destroyed.
 This is equivalent to solving the PerpExploration-BBH with two agents having a pebble each and as $n$ is sufficiently large this is impossible due to Corollary 3.3.
-Now suppose adversary chooses to activate the black hole whenever $a_{1}$ reaches there for the first time.
-In this case for all $C_{1}$, $C_{2}$ and $C_{3}$, at round $t$, agents $a_{2}$ and $a_{3}$ has no idea about where $a_{1}$ is consumed even if it knows that $a_{1}$ is consumed, as the distances between two consecutive $h_{i}$ are sufficiently large so their exploration region does not intersects till round $t$ and timeout (or,waiting for other agent) strategy do not work as for all $C_{i}$ an agent can get same timeout output.
-Also for all $C_{i}$s the position of the pebbles and alive agents will be same at round $t$.
-Now for the alive agents, the number of nodes for possible position of the black hole must be greater than or equal to 3 at round $t$.
-So the situation at round $t$ is similar to the problem of solving PerpExploration-BBH on a ring $R$ with two agents having a total of 3 pebbles where the number of possible consecutive positions of black hole is greater or equal to 3.
-Since we have assumed $\mathcal{A}$ solves the problem thus, $\mathcal{A}$ can also solve the problem of PerpExploration-BBH with two agents where number of possible positions of black hole is greater or equal to 3.
+Now suppose the adversary chooses to activate the Byzantine black hole whenever $a_{1}$ reaches there for the first time.
+In this case for all $C_{1}$, $C_{2}$ and $C_{3}$, at round $t$, agents $a_{2}$ and $a_{3}$ have no idea about where $a_{1}$ is destroyed even if they know that $a_{1}$ is destroyed, as the distances between two consecutive $h_{i}$ are sufficiently large so their exploration region does not intersect till round $t$ and timeout (or, waiting for other agents) strategy do not work as for all $C_{i}$ an agent can get same timeout output.
+Also for all $C_{i}$s the position of the pebbles and alive agents will be the same at round $t$.
+Now for the alive agents, the number of nodes for the possible position of the Byzantine black hole must be greater than or equal to 3 at round $t$.
+So the situation at round $t$ is similar to the problem of solving PerpExploration-BBH on a ring $R$ with two agents having a total of 3 pebbles where the number of possible consecutive positions of the Byzantine black hole is greater or equal to 3.
+Since we have assumed $\mathcal{A}$ solves the problem thus, $\mathcal{A}$ can also solve the problem of PerpExploration-BBH with two agents where the number of possible positions of the Byzantine black hole is greater or equal to 3.
 But due to Theorem 3.1 it is impossible.
-Hence there can never exists any algorithm that solves PerpExploration-BBH with three scattered agents each of which are equipped with a pebble.
-Corollary 3.8.
-A set of four scattered agents each with a pebble, are required to solve PerpExploration-BBH on a ring $R$ with $n$ nodes.
-4 Co-located Agents
-4.1 Communication: Face-to-Face
-In this section, we discuss the PerpExploration-BBH problem, under the conditions, where the agents are initially located at a safe node $home$, and they can only communicate with another agent whenever they are at the same node at the same round.
-Here considering this scenario we provide an algorithm PerpExplore-F2F (Algorithm 1) to solve the PerpExploration-BBH problem with $\lceil\log(n-1)\rceil+3$ agents.
-Description of PerpExplore-F2F: Given a ring $R$ with a byzantine black hole node $v_{b}$, a set of agents $A=\{a_{1},a_{2},\dots,a_{k}\}$, where $k=\lceil\log(n-1)\rceil+3$, are initially co-located at a safe node, which we term as $home$.
-Note that, the agents have no knowledge about the position of $v_{b}$.
-The agents have the knowledge of $n$ (which is the total number of nodes in $R$), whereas they know that $home$ is safe, as it is the initial position of the set of agents, $A$.
-Moreover, they have no idea about the remaining $n-1$ nodes of $R$, as $v_{b}$ can be any node among these $n-1$ nodes.
-We call this arc containing these $n-1$ nodes as the suspicious region, and it is denoted by $S$ with cardinality $|S|$.
-The main idea of the PerpExplore-F2F algorithm is from the $home$ only two agents, say $a_{1}$ and $a_{2}$ starts exploring the ring perpetually.
-The algorithm goes like this, one agent (preferably the agent with lowest Id, among the set of agents which are yet to be consumed by $v_{b}$) moves clockwise until it reaches the ’middle node’ of the suspicious region $S$, while the other agent (preferably the agent with second lowest Id, among the set of agents which have not been consumed by $v_{b}$) also performs the same movement, but along counter-clockwise direction, after which they return to the $home$ and this process perpetually continues.
-Now while executing this movement by these two agents, there can be two possible scenarios, the moment at least one agent gets destroyed by the byzantine black hole, $v_{b}$.
-Firstly, both the agents can get destroyed (this happens only when $v_{b}$ is the middle node of $S$), whereas secondly, only one agent is destroyed while the other agent following its usual course of algorithm, returns to $home$.
-For the initial case, the remaining agents at home can uniquely determines the byzantine black hole, $v_{b}$, as the middle node of $S$ and consecutively starts exploring the ring while avoiding $v_{b}$.
-For the latter case, the agents at the $home$ along with the agent that has returned to $home$, eventually gets to know that, the other agent has not returned and which in turn concludes that the arc, along which the missing agent was exploring contains the node $v_{b}$.
-This way our algorithm guarantees that, in the worst case, the size of the suspicious region decreases by a factor of at least half of the previous size, in the expense of one agent.
-So this implies, when $\lceil\log n\rceil$ agents are destroyed in $v_{b}$, the size of the suspicious region becomes exactly 1 (where this node in the suspicious region is nothing but $v_{b}$).
-In this case the remaining alive agent can determine the exact position of $v_{b}$ and hence, can perpetually explore the ring $R$ avoiding $v_{b}$.
-Next we will give a brief description of our algorithm PerpExplore-F2F.
-The algorithm works as follows, initially each agent at $home$ starts from state Initial0.
-In this state, an agent initializes the following variables: $S_{lft}$ to 1, $S_{rgt}$ to $n-1$,where $S_{lft}$ and $S_{rgt}$ denotes the distance to the left and the right end nodes of the suspicious region $S$ from the $home$, in the clockwise direction.
-Also in this state an agent declares its current node as the $home$.
-Note that, the agents stay at this state only for one round at the beginning of the execution of our algorithm, after which they enter the state Initial1.
-Now, after an agent enters the state Initial1, it calculates $|S|$ (where $|S|=|S_{rgt}-S_{lft}+1|$), initializes $Tnodes$ to 0 (where $Tnodes$ is the distance an agent traverses in order to reach its destination node).
-Next, it checks the variables $|S|$ and $k$ (where $k$ is the number of agents present at the $Current-Node$, i.e., $home$).
-If it finds $|S|=2$ and $k=2$, i.e., the length of the suspicious region is 2 and the total number of agents present at $home$ is also 2, then both the agents execute the sequence $SEQ0$ of states, where $SEQ0$ is defined as follows:
-[…]
-According to this sequence, the agent moves in to state Detection (after satisfying both the conditions $|S|=2$ and $k=2$ in state Initial1).
-In this state, the lowest Id agent is instructed to move clockwise and reach its destination position $S_{lft}$, whereas, the other agent reaches its destination $S_{rgt}$ along a counter-clockwise direction, while after each movement updating $Tnodes=Tnodes+1$.
-Now, after reaching their respective positions, the agents return back to $home$ along the same path in which they reached their destination nodes, again incrementing $Tnodes$ by 1 after each move.
-So, the total number of rounds required by the agent, starting from Initial1 in order to reach $home$ is $Tnodes+1$.
-After reaching $home$, they are further instructed to wait for $22n-Tnodes+$ 2 rounds.
-In the following round it again checks whether, #$\#Agent$ at $home$ is still $k$ i.e., 2, if so then again it changes to state Initial1, and the sequence $SEQ0$ is executed repeatedly.
-Otherwise, if the current agent is of lowest Id then declares the byzantine black hole node is at the location of $S_{rgt}$, otherwise if the current agent is not of lowest Id, then it declares that this byzantine black hole node is at $S_{lft}$.
-Irrespective of which, they continue performing perpetual exploration of $R$, by avoiding this black hole node.
-Hence, the total number of rounds required to complete the execution of this sequence is $42n+$ 4.
-On the contrary, if either $|S|\neq 2$ or $k\neq 2$.
-Note that, if $|S|=1$, then this node is nothing but the byzantine black hole node, and the remaining agents (such agent always exists, refer Lemma 4.11) detect this, and further continue perpetual exploration, avoiding this node.
-Otherwise, for $|S|\geq 2$ and $k\geq 3$ in state Initial1, the agents update the following variables: $k$ to $k-2$ (where $k-2$ is the number of agent which enters the next state Stagnant), locates the next destination node, where it calculates the variable $Dest=S_{lft}+\lceil\frac{|S|}{2}\rceil-1$, which is the clockwise distance of the destination node from $home$, $W_{time}$ to 00 (where $W_{time}$ signifies the waiting time at destination node).
-Agents in Initial1 state can change their state in to one of the following three states: Forward-Left, Forward-Right and Stagnant.
-An agent moves in to Forward-Left, if it is the agent with lowest Id among all the agents present at $home$, whereas an agent moves in to Forward-Right, if the agent is having second lowest Id among all the agents present at $home$.
-Lastly, an agent enters the state Stagnant, if it is neither with lowest Id nor with second lowest Id.
-According to the state an agent enters after Initial1, we define three sequence of states $SEQ1$, $SEQ2$ and $SEQ3$, respectively, which an agent can follow in the exact order after Initial1 state, based on whether the agent enters Forward-Left or Forward-Right or Stagnant.
-If the agent enters the state Forward-Left then it starts following $SEQ1$, or if it enters the state Forward-Right then it follows $SEQ2$ or if it enters Stagnant state then it follows $SEQ3$, respectively.
-Note that, in each sequence of states the agent eventually reaches back to Initial1, if not it is destroyed at the node $v_{b}$, while performing any of the states: Forward-Left (resp, Forward-Right) or Wait-Left (resp, Wait-Right) or Backtrack-Left (resp, Backtrack-Right).
-Also note that if one agent stated executing sequence $SEQ1$, then there must be another agent that executes sequence $SEQ2$ and the other agents execute the sequence $SEQ3$ starting from the same round (refer Line 10-17 of Algorithm  1).
-We define $SEQ1$, $SEQ2$ and $SEQ3$ as follows:
-[…]
-[…]
-[…]
-Our algorithm guarantees that, each agent starting from Initial1 again enters the state Initial1 at the same round, after following any of these three sequences.
-Definition 4.1 (Iteration).
-We define an iteration of our algorithm as the number of rounds required to execute any one of these sequences (i.e., $SEQ0$, $SEQ1$, $SEQ2$ and $SEQ3$) by an agent.
-Note that, each iteration has exactly $42n+$ 4 rounds.
-Each agent performs these iterations until this byzantine black hole is detected, after which, they explore the ring $R$ perpetually just by avoiding this black hole node.
-Next, we describe how the agents execute the sequences in order to perform the PerpExplore-F2F algorithm in one iteration.
-1 Input: $n,k$;
-2 States: {Initial0, Initial1, Stagnant, Forward-Left, Forward-Right, Backtrack-Left, Backtrack-Right, Wait-Left, Wait-Right, Wait1-Left, Wait1-Right}
-3
-4In State Initial0:
-5 $S_{lft}=$ 1, $S_{rgt}=$ $n-1$;
-6
-7Declare $Current-Node$ as $home$ and update to state Initial1.
-8
-9In State Initial1:
-10 $|S|=|S_{rgt}-S_{lft}+1|$, $Tnodes=0$;
-11 if $S_{lft}\neq S_{rgt}$ then
-12       if $|S|=2\wedge k=2$ then
-13             Move to state Detection.
-14
-15      else
-16             $k=k-2$, $Dest=S_{lft}+\lceil\frac{|S|}{2}\rceil-1$, $W_{time}=0$;
-17             if $LowestID$ then
-18                   Update state to Forward-Left.;
-19
-20             else if $SecondLowestID$ then
-21                   Update to state Forward-Right.;
-22
-23             else
-24                   Move to state Stagnant.;
-25
-26
-27
-28else
-29       Declare $S_{lft}$ as the byzantine black hole location and continue exploration avoiding that node.
-30
-31
-32In State Stagnant:
-33 Wait at the $Current-Node$ for $22n+$ 2 rounds.
-34 if #$\#Agent=k$ then
-35       Declare $Dest$ is the byzantine black hole location and continue exploration avoiding that node.
-36
-37 else if #$\#Agent=k+1$ then
-38       Update $k=k+1$ and collect the updated $S_{lft}$, $S_{rgt}$ and update accordingly.
-39       Move to state Initial1.
-40
-41 else
-42       Update $k=k+2$ and move to state Initial1.
-43
-44
-[…]
-[…]
-47
-48In State Wait-Left:
-49
-50if $W_{time}<n-Tnodes$  then
-51       $W_{time}++$
-52else
-53       Enter state Backtrack-Left and move $right$
-54
-55
-56In State Wait-Right:
-57
-58if  $W_{time}<n-Tnodes$  then
-59       $W_{time}++$
-60else
-61       Enter state Backtrack-Right and move $left$
-62
-63
-Algorithm 1 PerpExplore-F2F
-Execution of the sequence $SEQ1$: As discussed earlier, if the agent is of lowest Id in the in the state Initial1, then it is instructed to change their state to Forward-Left.
-In state Forward-Left, an agent reaches the destination node, which is $Dest$ distance away from $home$, along the clockwise direction, while after each move incrementing the variable $Tnodes$ by 1.
-Now, the moment it reaches the destination node, it is further instructed to change its state to Wait-Left.
-Note that, an agent stays in Forward-Left state for exactly $Tnodes+1$ rounds while following the sequence $SEQ1$ in an iteration.
-After changing its state to Wait-Left, the agent waits at the destination node for $n-Tnodes-1$ rounds, after which it is again instructed to move in to another state, which is Backtrack-Left while already starting to move one hop along counter-clockwise direction.
-Observe that an agent spends exactly $n-Tnodes$ round in state Wait-Left while executing $SEQ1$ in an iteration.
-Next in the state Backtrack-Left, an agent immediately starts moving along counter-clockwise direction, until the agent reaches the $home$ node, which is $Tnodes-1$ distance apart from its current node.
-While it reaches $home$, it again changes to state Wait1-Left.
-An agent stays at the Backtrack-Left state for exactly $Tnodes$ rounds.
-State Wait1-Left, instructs the agent to wait at the current node, i.e., $home$ for $n-Tnodes$ many rounds.
-This waiting time is enough for another agent that started executing sequence $SEQ2$ to return back to the $home$ if not destroyed by the byzantine black hole.
-After waiting, the agent checks the number of agent present at the current node.
-If #$\#Agent$ increases to $k+1$ (where $k$ agents are in state Stagnant), then in that case, it updates $k$ to $k+1$ and understands that the second lowest Id agent has not returned to $home$.
-In this situation, the only possibility of the other agent not returning back is that it has entered the black hole, along its traversal.
-In this situation it updates $S_{lft}$ to the destination node, which is $Dest$ distance apart from $home$ (i.e., basically the suspicious region becomes half, because initially $S_{lft}$ was 1, whereas now it is updated to $Dest$), and also communicates this information to the remaining $k-1$ agents at $home$, and directly moves into state Initial1.
-42
-[…]
-[…]
-45 In State Wait1-Left:
-46 Update $d=0$ and wait at the $Current-Node$ for $n-Tnodes$ rounds.
-47 if #$\#Agent=k+1$ then
-48       Update $k=k+1$ and $S_{lft}=Dest$ and communicate $S_{lft}$ to remaining agents at $Current-Node$.
-49       In the next round move to state Initial1.
-50
-51else
-52       Update $k=k+2$.
-53       In the next round move to state Initial1.
-54
-55
-56In State Wait1-Right:
-57 Update $d=0$ and wait at the $Current-Node$ for $n-Tnodes$ rounds.
-58 if #$\#Agent=k+1$ then
-59       Update $k=k+1$ and $S_{rgt}=Dest$ and communicate $S_{rgt}$ to remaining agents at $Current-Node$.
-60       In the next round move to state Initial1.
-61
-62else
-63       Update $k=k+2$.
-64       In the next round move to state Initial1.
-65
-66
-67In State Detection:
-68 If lowest Id then move to $S_{lft}$ clockwise, otherwise move to $S_{rgt}$ counter-clockwise, and after each move, update $Tnodes=Tnodes+1$.
-69 Return back to $home$ and again update $Tnodes=Tnodes+1$ after each move.
-70 if $Current-Node$ is $home$ then
-71       Wait at $home$ for $22n-Tnodes+$ 2 rounds.
-72       if #$\#Agent=k$ then
-73             Move to state Initial1.
-74
-75      else
-76             If lowest Id then declare byzantine black hole is $S_{rgt}$, otherwise declare byzantine black hole is $S_{lft}$, and then continue exploration avoiding this node.
-77
-78
-79
-On the other hand, if #$\#Agent$ at $home$ increases from $k$ to $k+2$ (i.e., both lowest Id and second lowest Id has returned back to $home$) then the agent simply waits for one round, after which it directly moves to state Initial1.
-Here note that an agent stays in the state Wait1-Left for $n-Tnodes+2$ rounds.
-Hence for an agent executing sequence $SEQ1$, the length of an iteration is $42n+$ 4.
-Execution of the sequence $SEQ2$.
-This execution if followed by the agent, which in state Initial1 has the second lowest Id.
-In this sequence, the agent from Initial1 moves to state Forward-Right.
-The state Forward-Right is exactly the same as the state Forward-Left which has been discussed above, the only difference is that the agent moves along a counter-clockwise direction, instead of clockwise direction.
-After this state, the agent moves to state Wait-Right, in this state again similar to the earlier state Wait-Left, the agent waits for $n-Tnodes-1$ many rounds at the destination node.
-After which it directly moves along clockwise direction for one hop and moves to state Backtrack-Right.
-This state resembles Backtrack-Left, instead the fact that in this state the agent moves along clockwise direction for $Tnodes-1$ distance, in which case it reaches $home$.
-After reaching $home$, it again waits for $n-Tnodes$ rounds while in state Wait1-Right, and after which performs similar checks as discussed earlier in Wait1-Left, but here, if after the waiting period, the total number of agents at $home$ is $k+1$ (where $k$ many agents are in state Stagnant), it similarly modifies the suspicious region by updating $S_{rgt}$ to $Dest$ (while earlier $S_{lft}$ was updated to $Dest$ in state Wait1-Left).
-Lastly, it ends this sequence $SEQ2$ by moving in to state Initial1 again.
-Similar to the length of an iteration for an agent executing $SEQ1$, the length of an iteration for any agent executing $SEQ2$ is again $42n+$ 4.
-Execution of the sequence $SEQ3$.
-This execution is followed by the agent, which in state Initial1 finds that it is neither of lowest Id nor of second lowest Id.
-In this state, the agent is instructed to move to state Stagnant.
-While in state Stagnant, it waits for $22n+$ 2 rounds at $home$, after which in $(2n+3)$-th round it checks the number of agent present at $home$.
-If ##\##Agent is $k$ (i.e., only the agents which moved to state Stagnant), then the agent understands that the lowest Id agent (which started sequence $SEQ1$) and second lowest Id agent (which started sequence $SEQ2$), both has entered the black hole, and since $R$ has only one black hole node, so this scenario of both agents entering the black hole can only occur when the node is common to both of these agents.
-This common node is nothing but the destination node, hence if this scenario happens, then the agent in state Stagnant, concludes that the black hole is none other than the destination node, and continues perpetual exploration avoiding this node.
-On the contrary, if #$\#Agent$ is $k+1$ (i.e., one among the agents executing $SEQ1$ and $SEQ2$ has returned, whereas the other agent has entered the black hole).
-In this scenario, updates either $S_{lft}$ or $S_{rgt}$ (based on the fact, whether the agent executing $SEQ1$ or the agent executing $SEQ2$ has returned), and then moves to state Initial1 while updating $k$ to $k+1$, ending sequence $SEQ3$.
-On the other hand, if both the agents executing $SEQ1$ and $SEQ2$ return to $home$ (i.e., ##\##Agent is $k+2$), then the agents in state Stagnant simply moves to state Initial1 while updating $k$ to $k+2$, and ending the sequence $SEQ3$.
-Note that here also, the iteration is of length $42n+$ 4 rounds, which leads to the following observation.
-{observation}
-For any $i$ ($i>0$), the $i$-th iteration for each of the $k$ alive agents, starts at the same round and at the same node, i.e., $home$.
-{observation}
-For any $i$ ($i>0$), each agent moves to the state Initial1 at the end of this $i$-th iteration.
-{observation}
-If $|S|=2$ and $k=2$, then both these agents follow the sequence $SEQ0$, whereas, if $|S|\geq 2$ and $k\geq 3$, then one agent follows $SEQ1$, another agent follows $SEQ2$, whereas the remaining agents follow the sequence $SEQ3$.
-That is,during an iteration if an agent follows $SEQ0$ then no agent follows $SEQ1$, $SEQ2$ or $SEQ3$ in that iteration, and vice-versa.
-4.1.1 Correctness and Complexity
-In this section we give the correctness and complexity analysis of our algorithm PerpExplore-F2F.
-Lemma 4.2.
-During any iteration, if exactly one agent gets destroyed by the byzantine black hole, then at the end of this iteration, the length of the suspicious region becomes at most $\lceil\frac{|S|}{2}\rceil+1$, where $|S|\geq 2$ is the length of the suspicious region at the start of this iteration.
-Proof 4.3.
-Let us consider at the $i$-th iteration of our Algorithm 1, the suspicious region be $S$ (where, $|S|>1$), so the agent starts from the state Initial1, and can encounter one of the following scenarios: first, $|S|=2$ (i.e., $S=\{S_{lft},S_{rgt}\}$) and $k=2$, second, $|S|\geq 2$ and $k>2$ (refer line 8 of Algorithm 1).
-If the first scenario is satisfied, then both the agents execute the sequence $SEQ0$ (refer description of PerpExplore-F2F), in which they perpetually visit the nodes $S_{lft}$ and $S_{rgt}$, respectively and return back to $home$, until one of these agent fails to return $home$, i.e., enters the byzantine black hole (refer state Detection in Algorithm 1).
-So, while the other agent returns $home$ (which was visiting $S_{rgt}$, without loss of generality), it is instructed to wait for $22n-Tnodes+$ 2 rounds (refer line 63 of Algorithm 1).
-After waiting, it finds that the agent visiting $S_{lft}$ is yet to return $home$, then this phenomenon helps the alive agent conclude that the agent visiting $S_{lft}$ must have entered the black hole (because $22n-Tnodes+$ 2 rounds is sufficient number of rounds for any agent visiting the node $S_{lft}$ to return back to $home$, if not consumed by the black hole), which is none other than the node $S_{lft}$ (as it is the only node in the suspicious region that this agent visits).
-Hence, in this case the black hole node is detected and also $|S|$ becomes 1 at the end of $i$-th iteration, which satisfies our claim.
-On the otherhand, if $|S|\geq 2$ and $k>3$, the agents are either following $SEQ1$ or $SEQ2$ or $SEQ3$, among them exactly one agent follows $SEQ1$ (say, $a_{1}$), exactly one agent follows $SEQ2$ (say, $a_{2}$), whereas the rest of the agent follows $SEQ3$ (where these agents stay at $home$ during this iteration, refer the description of execution of $SEQ3$).
-This means that the agent which can be destroyed by the black hole can be either $a_{1}$ or $a_{2}$ (as $home$ is always a safe node).
-If $a_{1}$ is destroyed, then $a_{2}$ during this iteration, after returning back to $home$ (while in state Backtrack-Right), waits at $home$ for $n-Tnodes$ many rounds (while in state Wait1-Right) for $a_{1}$ to return.
-Since, $a_{1}$ has already been consumed by the black hole, so it fails to return to $home$ within the waiting time of $a_{2}$.
-This observation of $a_{2}$ at $home$ instigates it to modify the suspicious region by updating $S_{rgt}$ to $Dest$, and $a_{2}$ communicates this updated information to the rest of the agents executing $SEQ3$ at $home$.
-It is because, $a_{1}$ only explores the nodes of the suspicious region that lies between $S_{lft}$ and $Dest$.
-Let $S_{new}$ be the modified suspicious region, then $|S_{new}|=|Dest-S_{lft}+1|=|S_{lft}+\lceil\frac{|S|}{2}\rceil-1-S_{lft}+1|=\lceil\frac{|S|}{2}\rceil<\lceil\frac{|S|}{2}\rceil+1$.
-Now consider the other case, where instead of $a_{1}$, $a_{2}$ be the agent which enters the black hole during an iteration.
-In this situation as well, similar to the argument explained above in case of $a_{2}$, $a_{1}$ will get to know $a_{2}$ has entered the black hole during the iteration, while $a_{1}$ waits for $n-Tnodes$ rounds at $home$.
-Observe, $a_{2}$ traverses the nodes of the suspicious region that lies between $Dest$ and $S_{rgt}$.
-So, $a_{1}$ updates $S_{lft}$ to $Dest$ in order to get the updated suspicious region $S_{new}$ and communicate with the rest of the agents executing $SEQ3$ at $home$.
-This implies $|S_{new}|=|S_{rgt}-Dest+1|=|S_{rgt}-S_{lft}-\lceil\frac{|S|}{2}\rceil+2|=||S|-\lceil\frac{|S|}{2}\rceil+1|\leq\lceil\frac{|S|}{2}\rceil+1$.
-So, in each scenario, our claim holds.
-{observation}
-Note that, if $|S|$ is odd then $|S_{new}|=\lceil\frac{|S|}{2}\rceil$, otherwise, $\lceil\frac{|S|}{2}\rceil\leq|S_{new}|\leq\lceil\frac{|S|}{2}\rceil+1$.
-Lemma 4.4.
-If during an iteration, exactly two agents gets destroyed by the byzantine black hole, then in that iteration, the position of the black hole is detected by the remaining alive agents.
-Proof 4.5.
-This situation of both agents being consumed by the black hole cannot arise when $|S|=2$ and $k=2$, because in this case both agents explore disjoint nodes of the suspicious region, in which only one node can be the black hole.
-On the other hand when $|S|\geq 2$ and $k>3$, this situation of two agents entering the black hole, can only arise for the agents $a_{1}$ and $a_{2}$, where $a_{1}$ and $a_{2}$ are following $SEQ1$ and $SEQ2$, respectively.
-Now, since the node (say $v_{b}$) at distance $Dest$ in the clockwise direction from $home$ is the only common node that both $a_{1}$ and $a_{2}$ visits while executing an iteration.
-So, if both these agents gets destroyed, then $v_{b}$ must be the node, in which the black hole is located.
-Now, as per our algorithm, the agents executing $SEQ3$ waits at $home$ for $22n+$ 2 rounds, after which they find that both $a_{1}$ and $a_{2}$ are yet to return $home$, whereas $22n+$ 2 rounds are sufficient for these two agents to return back to $home$ while executing their respective sequences, if they have not entered the black hole.
-This scenario helps the remaining agent conclude that both $a_{1}$ and $a_{2}$ has entered the black hole, and the node $v_{b}$ is the exact position of the byzantine black hole.
-Lemma 4.6.
-Let $|S_{0}|$ be the initial length of the suspicious region where $|S_{0}|=2^{p}+l$, $p\geq 1$ and $0\leq l<2^{p}$.
-Let $S_{i}$ be the suspicious region after it is modified for the $i-$th time.
-Then, $|S_{i}|\leq\lceil\frac{|S_{0}|}{2^{i}}\rceil+2$.
-Proof 4.7.
-Let $|S_{0}|$ be the initial length of the suspicious region where $|S_{0}|=2^{p}+l$, $p\geq 1$ and $0\leq l<2^{p}$.
-Let $S_{i}$ be the suspicious region after $i$ modifications.
-We now prove this lemma by induction.
-For the base case let $i=1$.
-Then from Lemma 4.2, $|S_{1}|\leq\lceil\frac{|S_{0}|}{2}\rceil+1<\lceil\frac{|S_{0}|}{2}\rceil+2$.
-So the lemma is true for $i=1$.
-Let the lemma be true for $i=j-1$.
-That is, $|S_{j-1}|\leq\lceil\frac{|S_{0}|}{2^{j-1}}\rceil+2$.
-Now to prove the lemma we have to show that $|S_{j}|\leq\lceil\frac{|S_{0}|}{2^{j}}\rceil+2$.
-Now by Lemma 4.2, $|S_{j}|\leq\lceil\frac{|S_{j-1}|}{2}\rceil+1\leq\lceil\frac{\lceil\frac{|S_{0}|}{2^{j-1}}\rceil+2}{2}\rceil+1\leq\lceil\frac{|S_{0}|}{2^{j}}\rceil+2$.
-Hence, by mathematical induction we can conclude that $|S_{i}|\leq\lceil\frac{|S_{0}|}{2^{i}}\rceil+2$, for any $i\geq 1$
-Lemma 4.8.
-If $|S_{t}|>2$ , then there exists at least one agent which is following the sequence $SEQ3$ in our Algorithm 1 in any iteration until the next modification of $S_{t}$, where $S_{t}$ is the suspicious region after $t$ modifications .
-Proof 4.9.
-Let $S_{0}$ be the initial suspicious region.
-Then, $|S_{0}|=2^{p}+l$ where $p\geq 1$ and $0\leq l<2^{p}$.
-Let $S_{i}$ be the suspicious region after $i$ modifications of it.
-Let for some $t>0$, $|S_{t}|>2$.
-This implies that during all the previous $t$ iterations where the suspicious region is modified, only one agent is consumed by the byzantine black hole.
-As otherwise $|S_{j}|=1$ for some $0<j<t$, and already the byzantine black hole is detected.
-This implies after the $t$-th modification, number of agents consumed is exactly $t$.
-So the remaining number of agent after $t$-th modification of suspicious region is exactly $\lceil\log|S_{0}|\rceil+3-t$ (since $\lceil\log|S_{0}|\rceil+3$, are the total number of agents executing PerpExplore-F2F along $R$).
-Now in any iteration until the next modification of $S_{t}$, there is an agent that follows $SEQ3$ if $\lceil\log|S_{0}|\rceil+3-t\geq 3\implies t\leq\lceil\log|S_{0}|\rceil$.
-Now we suppose, let $t>\lceil\log|S_{0}|\rceil$. Now there are two cases.
-Case-I : $|S_{0}|=2^{p}$ (i.e., $l=0$).
-Then, $t>p$.
-Now by Lemma 4.6, $|S_{p}|\leq 3$.
-Also, number of alive agent is exactly 3 after the $p-$th modification.
-Now if $|S_{p}|$ is not less than or equal to 2, then the agents will again follow the sequences $SEQ1,SEQ2$ and $SEQ3$ until the next modification of $S_{p}$.
-Now , $|S_{p+1}|\leq 2$ ( Observation 4.1.1).
-Now this is a contradiction as $2<|S_{t}|\leq|S_{p+1}|\leq 2$.
-Case-II : Now let us consider the case where $|S_{0}|=2^{p}+l$ where $02^{p}>l>$ 0.
-Then $t>p+1$.
-Now by Lemma 4.6, $|S_{p}|\leq 4$.
-Also, number of alive agent is exactly 4 after the $p-$th modification.
-Now if $|S_{p}|$ is not less than or equal to 2, then the agents will again follow the sequences $SEQ1,SEQ2$ and $SEQ3$ until the next modification of $S_{p}$.
-Now, $|S_{p+1}|\leq 3$ (by Observation 4.1.1).
-Again if $|S_{P+1}|$ is $\geq 2$ then the agents will again follow the sequences $SEQ1,SEQ2$ and $SEQ3$ until the next modification of $S_{p+1}$.
-Now again by Observation 4.1.1, $|S_{P+2}|\leq 2$.
-Now this again is a contradiction as $2<|S_{t}|\leq|S_{p+2}|\leq 2$.
-Hence $t\leq\lceil\log|S_{0}|\rceil$. Thus, there must exist an agent following $SEQ3$ in any iteration after the $t-$th modification of suspicious region until the next one where $|S_{t}|>2$
-From the above proof we can have the following corollary
-Corollary 4.10.
-If $|S_{t}|>2$ where $S_{t}$ is the suspicious region after $t$ modifications, then there is atleast 3 alive agents for all iterations after the $t$-th modification of the suspicious region and until the next modification.
-Lemma 4.11.
-Let $S_{t}$ be the suspicious region after $t$-th modification of $S_{0}$, where $S_{0}$ is the initial suspicious region.
-If $|S_{t}|=2$, then Algorithm 1, ensures that the number of agent which has not been consumed by the byzantine black hole is at least 2, for all iterations after the $t$-th modification of the suspicious region and until the next modification.
-Proof 4.12.
-As per our claim, $|S_{t}|=2$ after the $t$-th modification of $S_{0}$.
-This implies that after the $(t-1)$-th modification, we have either $|S_{t-1}|=3$ or $|S_{t-1}|=4$.
-Now we have the following cases.
-Case-I: If $|S_{0}|=2^{p}$ and $|S_{t-1}|=4$, then by Lemma 4.6, we have $|S_{t-1}|\leq\lceil\frac{|S_{0}|}{2^{t-1}}\rceil+2\implies 2\leq 2^{p-t+1}\implies 1\leq p-t+1$, i.e., $t-1\leq p-1$.
-Now since in each iteration where the suspicious region is modified exactly one agent gets destroyed, this means after $t-1$ modifications of suspicious region, at most $p-1$ agents gets destroyed by the black hole.
-This implies out of $p+3$ agents executing Algorithm 1 on $S_{0}$, now at least $4$ agents are alive to explore the suspicious region $S_{t-1}$.
-Hence, in the worst case, exactly one agent among them can enter the black hole at the $t$-th (note that 2 agents cannot enter the black hole, because in that case $|S_{t}|$ will be 1, by Lemma 4.4) modification whereas the updated length of suspicious region $S_{t}$, i.e., $|S_{t}|=2$.
-Now for the case when $|S_{t-1}|=3$, by similar argument $t-1\leq p$, so at least 3 agents among $p+3$ are now alive to explore $S_{t-1}$.
-Again, by consumption of one agent in the worst case, the updated suspicious region length $|S_{t}|$ will be 2.
-So, in both the scenarios, we show that when $|S_{t}|=2$ at least 2 agents are alive.
-Case-II: If $|S_{0}|=2^{p}+l<2^{p+1}$, where $0<l<2^{p}$ and $|S_{t-1}|=4$, in this case we have $p+4$ agents operating along $S_{0}$ (since, $\lceil\log|S_{0}|\rceil+3$ are the total number of agents) then again by Lemma 4.6, we have $|S_{t-1}|\leq\lceil\frac{|S_{0}|}{2^{t-1}}\rceil+2\implies 2<2^{p-t+2}\implies t-1<p$, so by earlier argument discussed in case-I, at least 5 (=$p+4-(p-1)$) agents are alive.
-So, in the worst case, exactly one agent gets consumed among 5 agents for the modification of $S_{t-1}$, which leads to the $t$-th modification, where $|S_{t}|=2$.
-On the contrary, if $|S_{t-1}|=3$, then similarly we have $11<2^{p-t+2}\implies t-1<p+$ 1, i.e., at least 4 agents are alive ($=p+4-p$ agents), so again in the worst case exactly one agent gets consumed among 4 agents for the modification of $S_{t-1}$, which leads to the $t$-th modification, where $|S_{t}|=2$.
-So, in this case as well, in both the scenarios when $|S_{t}|=2$, at least 2 agents are alive.
-This holds our claim.
-Theorem 4.13.
-PerpExplore-F2F solves the PerpExplore-BBH problem with $\lceil\log(n-1)\rceil+3$ agents having face to face communication and starting from the same node on a ring $R$ with $n$ nodes.
-Proof 4.14.
-We prove this theorem, by first ensuring that in any iteration of our Algorithm 1, every node is visited by at least one agent.
-Second, we show that, in any iteration there exists at least one alive agent.
-Let $|S|$ be the length of the suspicious region and $k$ be the number of alive agents at the beginning of iteration $i$.
-Let $|S|=2$ and $k=2$, in this case we have, $S=\{S_{lft},S_{rgt}\}$.
-So, the two agents $a_{1}$ and $a_{2}$ (say) follows the sequence $SEQ0$, in which $a_{1}$ is instructed to visit $S_{lft}$ along clockwise direction, whereas $a_{2}$ is instructed to visit $S_{rgt}$ along counter-clockwise direction, while performing an iteration.
-This implies that, if $S_{lft}$ is at distance $l$ ($l>0$) in a clockwise direction, $a_{1}$ after traversing a distance of $l$ reaches $S_{lft}$, whereas $a_{2}$ traverses a distance of $n-l-1$ in counter-clockwise direction to visit $S_{rgt}$, in an iteration.
-This in turn shows that, the agents starting from $home$, visits each node of $R$ while performing an iteration.
-Note that, $|S|>2$ and $k=2$ never occurs by Corollary 4.10.
-So, the only possibility remains is that $|S|>2$ and $k>2$, in this situation, one among them say, $a_{1}$ executes the sequence $SEQ1$, another say, $a_{2}$ executes the sequence $SEQ2$, whereas the remaining agents does not leave $home$ while executing $SEQ3$, in an iteration.
-Observe that, $a_{1}$ traverses in a clockwise direction for $Dest$ distance, and reaches the destination node.
-Similarly, $a_{2}$ also reaches this same node, while traversing along a counter-clockwise direction for $n-Dest$ distance.
-This shows that, in any iteration, there exists two such agents, which explore the whole ring, hence exploration is guaranteed in each such iteration.
-Next, we show that in every iteration, number of alive agents is at least 1. Let us consider during the $i$-th iteration, all the agents gets consumed by the byzantine black hole.
-Note that, at the beginning of the $i-$th iteration, number of alive agent can not be 1.
-By contradiction, let us suppose there is only one agent alive at the beginning of $i-$th iteration which gets consumed by the byzantine black hole during this iteration.
-This implies $|S|\geq 2$ at the beginning of this iteration.
-Now by Lemma 4.11 and Corollary 4.10, there must exists at least two alive agents at the beginning of the $i-$th iteration which contradicts our assumption of one alive agent.
-Also observe that number of alive agent at the start of iteration $i$, can not be strictly greater than 2.
-Otherwise, let number of alive agents $k>2$ at the beginning of iteration $i$.
-Since during iteration $i$ all the $k$ agents are destroyed, this implies $|S|\geq 2$ Now since $|S|\geq 2$ and $k>2$, our algorithm ensures that, in the worst case during this iteration at most two agents can be consumed by the byzantine black hole.
-So after the iteration number of alive agents is at least $k-2\geq 1$.
-This contradicts our assumption that all agents are destroyed at the end of this iteration.
-Now the only remaining case is that exactly two agents are alive at the beginning of $i-$th iteration.
-Similarly as both of the agents are destroyed during the $i-$th iteration $|S|\geq 2$ at the beginning of this iteration.
-Note that $|S|$ will be exactly 2, for $k=2$, as by Corollary 4.10, we have if $|S|>2$ then $k$ must be at least 3.
-Now according to the algorithm, both the agents follows $SEQ0$, which instructs them to follow disjoint paths.
-Hence, this concludes that during this iteration, only can be consumed by the black hole, while there exists another agent which is alive even after the iteration, contradicting our assumption.
-Hence for all iterations number of alive agents must be greater or equals to 1.
-Thus, algorithm PerpExplore-F2F solves the PerpExplore-BBH problem with $\lceil\log(n-1)\rceil+3$ agents having face to face communication and starting from the same node on a ring $R$ with $n$ nodes.
-4.2 Communication: Pebble
-In this section, we consider the model, where each agent has an access to a movable token (termed as pebble) which can be carried by an agent from one node to another.
-This pebble acts as a mode of communication for the agents, as the agents can perceive the presence of a pebble at the current node.
-Moreover, an agent can also perceive the presence of other agents which are co-located at the current round (i.e., gather the Ids of the other co-located agents).
-Next, we discuss the idea of our algorithm PerpExplore-Coloc-Pbl (Algorithm 2) which solves the PerpExploration-BBH problem with 3 agents, each accompanying a pebble.
-Description of PerpExplore-Coloc-Pbl: Given a ring $R$ with a byzantine black hole node $v_{b}$, a set of agents $A=\{a_{1},a_{2},a_{3}\}$ (where w.l.o.g ID of $a_{1}\leavevmode\nobreak\ <$ ID of $a_{2}\leavevmode\nobreak\ <$ ID of $a_{3}$), are initially co-located at a safe node, which we term as $home$.
-In this case as well, the agents have no knowledge about the position of the node $v_{b}$, the only set of knowledge the agents have are: the total number of nodes in the underlying topology $R$ (i.e., $n$), and also the knowledge that the initial node, i.e., $home$ is safe.
-So, the remaining arc of $n-1$ nodes in $R$ is a suspicious region (which we again term as $S$, where the cardinality of this arc is defined by $|S|$) for each of the three agents.
-The idea of this algorithm is as follows: if all the three agents are at $home$ along with their pebble, then in the $r-$th round ($r>0$) $a_{1}$ moves clockwise with its pebble.
-In the next round, i.e., $(r+1)-$th round, $a_{2}$, leaving its pebble at $home$ moves to the next node along clockwise direction to meet $a_{1}$.
-In the subsequent round, i.e., $(r+2)-$th round, $a_{1}$ after meeting $a_{2}$ moves again to the next node in the clockwise direction.
-Note that $a_{1}$ moves to next node even if $a_{2}$ meets $a_{1}$ without pebble if and only if $a_{1}$ is not at $home$, and waits for 3 rounds (i.e., till $(r+5)-$th round).
-In $(r+3)-$th round, $a_{2}$ leaves its current node, moves one step in counter-clockwise direction, collects the left behind pebble, and in $(r+4)-$th round moves again to the earlier node in clockwise direction.
-Subsequently, in $(r+5)-$th round, $a_{2}$ again leaves the pebble at its current node, and moves in a clockwise direction, to meet $a_{1}$.
-This process continues, until each of $a_{1}$ and $a_{2}$ along with the two pebbles they carried, reaches $home$.
-This movement of $a_{1}$ and $a_{2}$ guarantees exploration unless at least one agent is destroyed by the black hole, while they are exploring the ring by following the above mentioned rules.
-In the meantime the third agent $a_{3}$ along with a pebble waits at $home$ for these remaining agents to return.
-Note that, if all the agents along with the pebbles carried by them, return back to $home$, then again they start the above mentioned execution from $home$, this guarantees that until an agent enters the black hole, perpetual exploration is performed.
-Now, if any agent ($a_{1}$ or $a_{2}$ or both $a_{1}$ and $a_{2}$), enters the black hole, then our algorithm guarantees that either perpetual exploration is performed by the remaining alive agents or, the byzantine black hole will be detected within finite rounds by at least one alive agent, and it will continue performing perpetual exploration, avoiding the byzantine black hole, and independent of the position of other alive agents or pebbles.
-In the following part we give a detailed description of our algorithm PerpExplore-Coloc-Pbl.
-Initially all the agents are in state Initial.
+Hence there can never exist any algorithm that solves PerpExploration-BBH with three scattered agents each of which is equipped with a pebble.
+9 Algorithm PerpExplore-Coloc-Pbl
+9.1 Detailed Description and Pseudocode
+In the following part, we give a detailed description of our algorithm PerpExplore-Coloc-Pbl.
+Initially, all the agents are in state Initial.
 In this state, an agent first declares its $Current-Node$ as $home$, after which initializes the variable $T_{time}=0$ (where, $T_{time}$ is the number of rounds passed since the agent has moved from state Initial) and gathers the ID of the remaining agents currently at $home$.
-Next, the agent with lowest ID, i.e., $a_{1}$ moves to state Leader, the agent with second lowest ID, i.e., $a_{2}$ moves to state Follower-Find, whereas the remaining agent, i.e., $a_{3}$ moves to state Backup.
+Next, the agent with the lowest ID, i.e., $a_{1}$ moves to state Leader, the agent with the second lowest ID, i.e., $a_{2}$ moves to state Follower-Find, whereas the remaining agent, i.e., $a_{3}$ moves to state Backup.
 We now define an iteration for this algorithm.
-An iteration is defined to be a collection of $14n+$ 1 consecutive rounds starting from the latest round where all 3 agents along with 3 pebbles are at $home$ in state Initial.
-Note that, a new iteration fails to execute, if either at least one pebble or an agent gets destroyed by the black hole in the current iteration.
-More precisely, the meaning of failing an iteration implies that, an agent after ending its current iteration does not again start a new iteration by moving in to state Initial.
-Suppose, all the agents along with the pebbles successfully execute the $i$-th iteration and reaches $home$, i.e., neither any pebble nor any agent gets destroyed by the black hole.
-So, according to our algorithm, $a_{1}$ being the lowest ID, follows the following sequence $SEQ1$, whereas $a_{2}$ being the second lowest ID follows $SEQ2$, and the $a_{3}$ follows the sequence $SEQ3$.
+An iteration is defined to be a collection of $14n+$ 1 consecutive rounds starting from the latest round where all 3 agents along with 2 pebbles are at $home$ in state Initial.
+Note that, a new iteration fails to execute if either at least one pebble or an agent gets destroyed by the Byzantine black hole in the current iteration.
+More precisely, the meaning of failing an iteration implies that an agent after ending its current iteration does not again start a new iteration by moving into state Initial.
+Suppose, all the 3 agents along with 2 pebbles successfully execute the $i$-th iteration and reach $home$, i.e., neither any pebble nor any agent gets destroyed by the Byzantine black hole.
+In this situation, according to our algorithm, $a_{1}$ being the lowest ID, follows the following sequence $SEQ1$, whereas $a_{2}$ being the second lowest ID follows $SEQ2$, and the $a_{3}$ follows the sequence $SEQ3$.
 The above sequences are as follows:
@@
 Execution of sequence $SEQ1$: $a_{1}$ performs this sequence, in which it first changes to state Leader, after finding it as the lowest ID agent at $home$ in state Initial.
-In state Leader, the agent performs the following checks: if the current node is $home$, then it checks that if the current node has the agent with second lowest ID, whereas the number of pebble at $home$ is 3 and $T_{time}\geq 2$.
+In state Leader, the agent performs the following checks: if the current node is $home$, then it checks that if the current node has the agent with the second lowest ID, whereas the number of pebbles at $home$ is 2 and $T_{time}\geq 2$.
 If all these conditions are satisfied, then the agent waits till $T_{time}=4n$ after which it moves to state Initial during $T_{time}=4n+1$.
-Otherwise, if it finds that it is with the second lowest ID agent and number of pebble at $home$ is also 3, but $T_{time}<2$, in this case the agent initializes $W_{time}=0$ (where $W_{time}$ is the waiting time of the agent) and then moves one hop along with the pebble to the next clockwise node.
-Otherwise, if the agent finds that neither it is with the second lowest ID agent nor the number of pebbles at $home$ is 3, then it waits for 5 rounds (i.e., waits when $W_{time}<6$), after which if still the above condition persists, then concludes $S_{lft}$ as the node which is at a clockwise distance of $n-2$ from $home$ and $S_{rgt}$ the node at a clockwise distance of $n-1$.
+Otherwise, if it finds that it is with the second lowest ID agent and the number of pebbles at $home$ is also 2, but $T_{time}<2$, in this case, the agent initializes $W_{time}=0$ (where $W_{time}$ is the waiting time of the agent) and then moves one hop along with the pebble to the next clockwise node.
+Otherwise, if the agent finds that neither it is with the second lowest ID agent nor the number of pebbles at $home$ is 2, then it waits for 5 rounds (i.e., waits when $W_{time}<6$), after which if still the above condition persists, then concludes $S_{lft}$ as the node which is at a clockwise distance of $n-2$ from $home$ and $S_{rgt}$ the node at a clockwise distance of $n-1$.
 Further, it updates $W_{time}=0$ and moves to the next state Detection.
-On the contrary, if the current node of $a_{1}$ is not $home$, then also it performs the following checks: first, it checks whether it is with the second lowest ID agent (i.e., $a_{2}$ in this case), if so, then initializes $W_{time}=0$ and then moves one hop clockwise to the next node along with the pebble it is accompanying.
-Otherwise, if it is not with the second lowest ID agent, then waits for 3 rounds (i.e., waits when $W_{time}<4$), after which if still it is without the second lowest ID agent, then the agent moves to state Report-Leader, while leaving the pebble at the current node and moving one hop to the clockwise node.
-Note that until and unless there is no anomalies detected by the agent $a_{1}$, it ends an iteration, by changing to state Initial from Leader.
+On the contrary, if the current node of $a_{1}$ is not $home$, then also it performs the following checks: first, it checks whether it is with the second lowest ID agent (i.e., $a_{2}$ in this case), and if so, then initializes $W_{time}=0$ and then moves one hop clockwise to the next node along with the pebble it is accompanying.
+Otherwise, if it is not with the second lowest ID agent, then wait for 3 rounds (i.e., waits when $W_{time}<4$), after which if still it is without the second lowest ID agent, then the agent moves to state Report-Leader while leaving the pebble (it is accompanying) at the current node and moving one hop to the clockwise node.
+Note that until and unless there are no anomalies detected by the agent $a_{1}$, it ends an iteration, by changing to state Initial from Leader.
 Otherwise, if any anomaly is detected, then $a_{1}$ ends the current iteration by either changing to state Detection or Report-Leader from the state Leader, in which case, it never changes to state either Leader or Initial.
-Execution of sequence $SEQ2$: $a_{2}$ being the second lowest Id agent in state Initial, performs this sequence in an iteration.
+Execution of sequence $SEQ2$: $a_{2}$ being the second lowest ID agent in state Initial, executes this sequence in an iteration.
 On successful completion of this iteration, an agent starting from Initial, performs the sub-sequence ${Follower-SEQ}$, $n$ times and then again changes to state Initial.
 The sub-sequence ${Follower-SEQ}$ symbolizes the sequence Follower-Find $\rightarrow$ Follower-Collect $\rightarrow$ Follower-Find, which an agent performs $n$ times until it again changes to state Initial (i.e., ${Follower-SEQ}$ is denoted by ${Follower-SEQ}^{n}$).
-While executing ${Follower-SEQ}$, the agent in state Follower-Find checks, whether the current node does not contain the lowest ID agent and the $Move$ parameter is also set to 0 (the parameter $Move$ is either 0 or 1, when 1 it symbolizes that the current node must contain the lowest ID agent, when 0, it means the current node is the adjacent counter clockwise node with respect to the node which contains the lowest ID agent), if so then it moves one hop clockwise leaving the pebble at the current node, while updating $Move$ to 1.
-Otherwise, if it finds $Move$ to be 1 whereas the current node does not contain the lowest ID agent, then it stops the current iteration, detects that current node is black hole and starts performing perpetual exploration, avoiding the black hole node.
+While executing ${Follower-SEQ}$, the agent in state Follower-Find checks, whether the current node does not contain the lowest ID agent, and the $Move$ parameter is also set to 0 (the parameter $Move$ is either 0 or 1, when 1 it symbolizes that the current node must contain the lowest ID agent, when 0, it means the current node is the adjacent counter-clockwise node with respect to the node which contains the lowest ID agent), if so then it moves one hop clockwise leaving the pebble at the current node, while updating $Move$ to 1.
+Otherwise, if it finds $Move$ to be 1 whereas the current node does not contain the lowest ID agent, then it stops the current iteration, detects that the current node is a Byzantine black hole, and starts performing perpetual exploration, avoiding the Byzantine black hole node.
 On the other hand, if the current node contains the lowest ID agent, then $a_{2}$ directly changes to state Follower-Collect.
@@
 If it finds a pebble, then moves one hop clockwise along with the pebble.
-Now, if $T_{time}=4n+1$ and current node is $home$, then stops performing $Follower-SEQ$ and moves to state Initial.
+Now, if $T_{time}=4n+1$ and the current node is $home$, then stops performing $Follower-SEQ$ and moves to state Initial.
 Otherwise, updates $Move$ to 0 and changes to state Follower-Find.
-If the pebble is not found, then it concludes current node to be the black hole and starts performing perpetual exploration avoiding this node.
+If the pebble is not found, then it concludes the current node to be the Byzantine black hole and starts performing perpetual exploration avoiding this node.
 Note that, the agent only changes to state Initial only when it reaches $home$ again, which is after it performs ${Follower-SEQ}$ for $n$ times starting from $SEQ2$.
-If at any point, if $a_{2}$ finds any irregularities, based on whether a pebble is present or not, then only it directly concludes the black hole position.
+If at any point, $a_{2}$ finds any irregularities, based on whether a pebble is present or not, then only it directly concludes the Byzantine black hole position.
 In all other cases, it follows the ${Follower-SEQ}$ sequence and changes to state Follower-Collect.
@@
 In this sequence, after being in state Initial, it changes to state Backup, in which the agent first waits at $home$ till $T_{time}=4n$, after which it checks the following details, and accordingly either moves to state Initial if no irregularities are detected, otherwise moves to state Find-Pebble or Find-BH, based on the irregularity it detects.
-  • The current node, i.e., $home$ has only the lowest and highest ID agents, i.e., $a_{1}$ and $a_{3}$, whereas the number of pebble at $home$ is 1.
+  • The current node, i.e., $home$ has only the lowest and highest ID agents, i.e., $a_{1}$ and $a_{3}$, whereas the number of pebble at $home$ is 0.
 If this condition is satisfied, then the agent changes its state to Find-Pebble.
-Note that, this condition is satisfied only when $a_{2}$ fails to reach $home$ while performing this iteration, i.e., enters the black hole, while $a_{1}$ has detected some anomalies, which instigates it to leave the pebble at the current position at which it detects the anomaly, and henceforth moves back to $home$ while in state Report-Leader.
-  • $home$ has only the highest and lowest ID agents, i.e., $a_{1}$ and $a_{3}$, but number of pebble at $home$ is 2, which implies that $a_{2}$ has failed to return to $home$ even when $T_{time}=4n$.
+Note that, this condition is satisfied only when $a_{2}$ fails to reach $home$ while performing this iteration, i.e., enters the Byzantine black hole, while $a_{1}$ has detected some anomalies, which instigates it to leave the pebble at the current position at which it detects the anomaly, and henceforth moves back to $home$ while in the state Report-Leader.
+  • $home$ has only the highest and lowest ID agents, i.e., $a_{1}$ and $a_{3}$, but the number of pebbles at $home$ is 1, which implies that $a_{2}$ has failed to return to $home$ even when $T_{time}=4n$.
 In this case, $a_{3}$ concludes $S_{lft}$ to be the node which is at a clockwise distance of $n-2$ from $home$, whereas $S_{rgt}$ to be the node which is at a clockwise distance of $n-1$ from $home$.
 Finally, it updates $W_{time}=0$ and then moves to state Detection.
-  • $home$ has all the three agents and moreover, number of pebbles present at $home$ is also 3.
-In this case, $a_{3}$ changes to state Initial, completing the iteration without detecting any anomalies.
+  • $home$ has all three agents and the number of pebbles present at $home$ is 2. In this case, $a_{3}$ changes to state Initial, completing the iteration without detecting any anomalies.
   • After $4n$ rounds, there is no other agent except $a_{3}$ present at $home$.
-In this case, the agent move one hop, clockwise without the pebble present at $home$, and then changes to state Find-BH.
-This case arises only when both the agents $a_{1}$ and $a_{2}$ fail to reach $home$, while performing this iteration.
-Now, next we define the states Report-Leader, Find-Pebble, Find-BH, Detection. An agent moves to either of these states only when it detects some anomalies.
+In this case, the agent moves one hop clockwise, and then changes to state Find-BH.
+This case arises only when both the agents $a_{1}$ and $a_{2}$ fail to reach $home$ while performing this iteration.
+Next we define the states Report-Leader, Find-Pebble, Find-BH, Detection. An agent moves to either of these states only when it detects some anomalies.
 The state Report-Leader, is executed by only the lowest ID agent, i.e., $a_{1}$, while it is in state Leader.
-The anomaly detected is as follows, $a_{1}$, after reaching a new node that is not $home$, along with the pebble it is carrying, waits for at most 3 rounds, for second lowest ID agent, i.e., $a_{2}$ to arrive (which is in state Follower-Find).
-If $a_{2}$ does not arrive even after the waiting, $a_{1}$ concludes that the agent (i.e., $a_{2}$) has entered the black hole, this triggers $a_{1}$ to move to state Report-Leader.
+The anomaly detected is as follows, $a_{1}$, after reaching a new node that is not $home$, along with the pebble it is carrying, waits for at most 3 rounds, for the second lowest ID agent, i.e., $a_{2}$ to arrive (which is in state Follower-Find).
+If $a_{2}$ does not arrive even after the waiting, $a_{1}$ concludes that the agent (i.e., $a_{2}$) has entered the Byzantine black hole, this triggers $a_{1}$ to move to state Report-Leader.
 In this state, it moves clockwise until it finds the highest ID agent, i.e., $a_{3}$ at $home$.
-Then waits until $T_{time}=4n$, after which changes state to Find-Pebble.
+Then waits until $T_{time}=4n$, after which it changes its state to Find-Pebble.
+The state Find-Pebble, is performed together by only $a_{1}$ and $a_{3}$, i.e., the lowest and highest ID agent.
+This state can be reached by $a_{1}$ only via Report-Leader state, whereas by $a_{3}$ from the state Backup.
+The main idea for the agents in this state is to move counter-clockwise till they find the pebble left by $a_{1}$ when it detects some anomaly (i.e., $a_{2}$ has entered the Byzantine black hole) and move into state Report-Leader.
+This pebble acts as a marker which indicates that the Byzantine black hole is either the counter-clockwise neighbor of this node or, it is at a counter-clockwise distance of 2 hops from it.
+This node at which it finds the pebble is declared to be the new $home$, from which they conclude $S_{lft}$ and $S_{rgt}$ to be the nodes which are at a clockwise distance of $n-2$ and $n-1$, respectively from this new $home$.
+Finally, they update $W_{time}=0$ and then change to state Detection.
+Note that, this state is performed by an agent during the last round of the current iteration, (i.e., when $T_{times}=4n+1$).
+The state Find-BH, is only executed by the highest ID agent, i.e., $a_{3}$.
+This state is executed by $a_{3}$, only when it finds no other agent at $home$, at the last round of the current iteration.
+This situation can only occur, when both the agents $a_{1}$ and $a_{2}$ have entered the Byzantine black hole.
+So, this triggers $a_{3}$ to change to state Find-BH, in which it moves along a clockwise direction until a pebble is found.
+Whenever a pebble is found, it concludes the next node to be the Byzantine black hole, and starts performing perpetual exploration avoiding this node.
+Finally, the state Detection, is performed by an agent (i.e., $a_{1}$ or $a_{3}$) only if they know the position of $S_{lft}$ and $S_{rgt}$, which are not only two consecutive nodes in a suspicious region $S$.
+Moreover, these two nodes are exactly at a distance of clockwise $n-1$ and $n-2$ distance from $home$ (the $home$ can be the initial starting node, or it can be the updated $home$ as well).
+In this state, the agent with the lowest ID, i.e., $a_{1}$, moves clockwise till it reaches $S_{rgt}$, and then again returns back to $home$.
+After returning back, if it finds $a_{3}$, then again it performs the same procedure.
 1 Input: $n,\leavevmode\nobreak\ k=3$;
@@
 17 if $Current-Node=home$ then
-18       if $With\leavevmode\nobreak\ Second\leavevmode\nobreak\ Lowest\leavevmode\nobreak\ Id\wedge\#Pebble=3\wedge T_{time}\geq 2$  then
+18       if $with\leavevmode\nobreak\ Second\leavevmode\nobreak\ Lowest\leavevmode\nobreak\ ID\wedge\#Pebble=2\wedge T_{time}\geq 2$  then
 19             Wait at the current node till $T_{time}=4n$.
@@
 21
-22       else if $With\leavevmode\nobreak\ Second\leavevmode\nobreak\ Lowest\leavevmode\nobreak\ Id\wedge\#Pebble=3\wedge T_{time}<2$ then
+22       else if $With\leavevmode\nobreak\ Second\leavevmode\nobreak\ Lowest\leavevmode\nobreak\ ID\wedge\#Pebble=2\wedge T_{time}<2$ then
 23             $W_{time}=0$.
@@
 53 else if not with lowest ID $\wedge$ $Move=1$ then
-54       Detect the $Current-Node$ as black hole and continue exploring the ring perpetually avoiding this node.
+54       Detect the $Current-Node$ as Byzantine black hole and continue exploring the ring perpetually avoiding this node.
 55
@@
 68else
-69       Detect the $Current-Node$ as black hole and continue perpetually exploring the ring avoiding this node.
+69       Detect the $Current-Node$ as Byzantine black hole and continue perpetually exploring the ring avoiding this node.
 Algorithm 1 PerpExplore-Coloc-Pbl
@@
 62 Wait until $T_{time}=4n$.
-63 if $Current-Node$ has only the lowest and highest ID agents and #$\#Pebble=1$ then
+63 if $Current-Node$ has only the lowest and highest ID agents and #$\#Pebble=0$ then
 64       Move to state Find-Pebble.
 65
-66 else if $Current-Node$ has only the lowest and highest ID agents and #$\#Pebble=2$ then
+66 else if $Current-Node$ has only the lowest and highest ID agents and #$\#Pebble=1$ then
 67       $S_{lft}=$ node at a clockwise distance of $n-2$ from $home$
@@
 70
-71 else if $Current-Node$ has both lowest and second lowest Id agent and #$\#Pebble=3$ then
+71 else if $Current-Node$ has both lowest and second lowest ID agent and #$\#Pebble=2$ then
 72       Change to state Initial.
@@
 74 else if $Current-Node$ has no other agent then
-75       Move in a clockwise direction without pebble and change to state Find-BH.
+75       Move in a clockwise direction and change to state Find-BH.
 76
@@
 80if a pebble is found then
-81       Conclude the next node along clockwise direction is the black hole, and continue perpetual exploration avoiding the black hole node.
+81       Conclude the next node along clockwise direction is the Byzantine black hole, and continue perpetual exploration avoiding the Byzantine black hole node.
 82
@@
 87In State Find-Pebble:
-88 Move counter-clockwise without the pebble.
+88 Move in a counter-clockwise direction.
 89 if a pebble is found then
@@
 101       Moves clockwise till the node at a distance $n-2$ from $home$ then returns back to $home$ along the same path and checks for the other agent.
-102       If other agent is there it repeats the same procedure as stated in line 86 of Algorithm 2. Otherwise detects $S_{rgt}$ as the black hole and start exploring avoiding that node.
+102       If other agent is there it repeats the same procedure as stated in line 86 of Algorithm 1. Otherwise detects $S_{rgt}$ as the Byzantine black hole and start exploring avoiding that node.
 103else
 104       Moves one hop counter-clockwise from $home$ then returns back to $home$ along the same path then waits for $62n-$ 6 rounds and checks for the other agent.
-105       If other agent is there it repeats the same procedure as stated in line 90 of Algorithm 2. Otherwise detects $S_{lft}$ as the black hole and start exploring avoiding that node.
-The state Find-Pebble, is performed together by only $a_{1}$ and $a_{3}$, i.e., the lowest and highest ID agent.
-This state can be reached by $a_{1}$ only via Report-Leader state, whereas by $a_{3}$ from the state Backup.
-The main idea for the agents in this state is to move counter-clockwise till they find the pebble left by $a_{1}$ when it detects some anomaly (i.e., $a_{2}$ has entered the byzantine black hole) and moved into state Report-Leader.
-This pebble acts as a marker which indicates that black hole is either the counter-clockwise neighbour of this node or, it is at a counter clockwise distance of 2 hop from it.
-This node at which it finds the pebble is declared to be the new $home$, from which they conclude $S_{lft}$ and $S_{rgt}$ to be the nodes which are at a clockwise distance of $n-2$ and $n-1$, respectively from this new $home$.
-Finally, they updates $W_{time}=0$ and then changes to state Detection.
-Note that, this state is performed by an agent during the last round of the current iteration, (i.e., when $T_{times}=4n+1$).
-The state Find-BH, is only executed by the highest ID agent, i.e., $a_{3}$.
-This state is executed by $a_{3}$, only when it finds no other agent at $home$, at the last round of the current iteration.
-This situation can only occur, when both the agents $a_{1}$ and $a_{2}$ has entered the black hole.
-So, this triggers $a_{3}$ to change to state Find-BH, in which it moves along a clockwise direction, until a pebble is found.
-Whenever, a pebble is found, it concludes the next node to be the black hole, and starts performing perpetual exploration avoiding this node.
-Finally, the state Detection, is performed by an agent (i.e., $a_{1}$ or $a_{3}$) only if they know the position of $S_{lft}$ and $S_{rgt}$, which are not only two consecutive nodes in a suspicious region $S$.
-Moreover, these two nodes are exactly at a distance of clockwise $n-1$ and $n-2$ distance from $home$ (the $home$ can be the initial starting node, or it can be the updated $home$ as well).
-In this state, the agent with lowest ID, i.e., $a_{1}$, moves clockwise till it reaches $S_{rgt}$, and then again returns back to $home$.
-After returning back, if it finds $a_{3}$, then again it performs the same procedure.
-On the contrary, if it does not find $a_{3}$, then it understands that $S_{rgt}$ is the black hole node, and hence continues perpetual exploration avoiding this node.
+105       If other agent is there it repeats the same procedure as stated in line 90 of Algorithm 1. Otherwise detects $S_{lft}$ as the Byzantine black hole and start exploring avoiding that node.
+On the contrary, if it does not find $a_{3}$, then it understands that $S_{rgt}$ is the Byzantine black hole node, and hence continues perpetual exploration avoiding this node.
 On the other hand, $a_{3}$ in this state moves counter-clockwise direction from $home$, i.e., reaches $S_{lft}$, after which it returns back to $home$ again following the same path and then waits for $62n-$ 6 rounds at $home$ and then checks for the other agent, i.e., $a_{1}$.
-If it finds $a_{1}$, then again continues to perform the same movement, otherwise it concludes that $S_{lft}$ is the black hole node, and hence starts the perpetual exploration of the ring just by avoiding this black hole node.
-4.2.1 Correctness and Complexity
-In this section we discuss the correctness and complexity of our algorithm PerpExplore-Coloc-Pbl.
+If it finds $a_{1}$, then again continues to perform the same movement, otherwise, it concludes that $S_{lft}$ is the Byzantine black hole node, and hence starts the perpetual exploration of the ring just by avoiding this Byzantine black hole node.
+9.2 Correctness and Complexity
+In this section, we discuss the correctness and complexity of our algorithm PerpExplore-Coloc-Pbl.
 Lemma 9.1.
-If no agent gets consumed by the byzantine black hole, while performing an iteration of Algorithm 2, then our algorithm ensures that in that iteration, either the ring $R$ is explored or there is one agent that knows exactly the location of the black hole.
-Proof 4.16.
-Suppose till the completion of $i$-th iteration $(i>0)$ no agent is consumed by the black hole, then in any iteration $0<j\leq i$, the agent $a_{1}$ follows the sequence $SEQ1$, whereas $a_{2}$ follows the sequence $SEQ2$, as instructed starting from the initial $home$ node.
+If no agent gets destroyed by the Byzantine black hole, while performing an iteration of Algorithm 1, then our algorithm ensures that in that iteration, either the ring $R$ is explored when no agent knows the exact Byzantine black hole location or, there is one agent that knows exactly the location of the Byzantine black hole.
+Proof 9.2.
+Suppose till the completion of $i$-th iteration $(i>0)$ no agent is destroyed by the Byzantine black hole, then in any iteration $0<j\leq i$, the agent $a_{1}$ follows the sequence $SEQ1$, whereas $a_{2}$ follows the sequence $SEQ2$, as instructed starting from the initial $home$ node.
 Considering this scenario we have two cases:
-  • No agents and no pebbles are destroyed: In this case, while executing the sequence $SEQ1$, the agent after changing its state to Leader from Initial, visits a new node after every 4 rounds in clockwise direction.
-Since, no agent has been consumed in this iteration as well, so $a_{1}$ un-obstructively visits each node in a clockwise direction, and reaches $home$ to finally end this sequence $SEQ1$, while again changing its state to Initial.
-This shows that, at least one agent visits each node of $R$ in an iteration completing the exploration of $R$.
-  • No agent gets destroyed but some pebbles are consumed by the black hole: Note that, this situation can only occur when the black hole consumes the pebble carried and released by $a_{2}$, while $a_{2}$ is not at the node containing the pebble, executing the sequence $SEQ2$.
+  • No agents and no pebbles are destroyed: In this case, while executing the sequence $SEQ1$, the agent after changing its state to Leader from Initial, visits a new node after every 4 rounds in the clockwise direction.
+Since no agent has been destroyed in this iteration as well, so $a_{1}$ un-obstructively visits each node in a clockwise direction, and reaches $home$ to finally end this sequence $SEQ1$, while again changing its state to Initial.
+This shows that at least one agent visits each node of $R$ in an iteration completing exploration of $R$.
+  • No agent gets destroyed but some pebbles are destroyed by the Byzantine black hole: Note that, this situation can only occur when the Byzantine black hole consumes the pebble carried and released by $a_{2}$, while $a_{2}$ is not at the node containing the pebble, executing the sequence $SEQ2$.
 It is because, in an iteration, only $a_{1}$ and $a_{2}$ moves away from $home$ while executing their respective sequences $SEQ1$ and $SEQ2$.
-And the pebble carried by $a_{1}$ ( i.e., the agent with smallest ID) cannot be destroyed without destroying $a_{1}$, as $a_{1}$ always carries the pebble at each node while executing its respective sequence.
-So, in this situation, while $a_{2}$ ( i.e., the agent with second lowest ID) in state Follower-Collect moves back to the adjacent node in counter-clockwise direction, in order to collect the already released pebble, absence of which triggers $a_{2}$ to conclude the current node to be the black hole, and which leads the agent to immediately leave the black hole node.
-In this case, the consumption of a pebble leads to the black hole detection.
+And the pebble carried by $a_{1}$ ( i.e., the agent with the smallest ID) cannot be destroyed without destroying $a_{1}$, as $a_{1}$ always carries the pebble at each node while executing its respective sequence.
+So, in this situation, while $a_{2}$ ( i.e., the agent with the second lowest ID) in the state Follower-Collect moves back to the adjacent node in the counter-clockwise direction, in order to collect the already released pebble, absence of which triggers $a_{2}$ to conclude the current node to be the Byzantine black hole, and which leads the agent to immediately leave the Byzantine black hole node.
+In this case, the consumption of a pebble leads to the Byzantine black hole detection.
 {observation}
-If no agent gets consumed by the byzantine black hole while performing an iteration of Algorithm 2 but some pebbles are consumed by the black hole, then by Lemma 4.15, one agent (more precisely, the agent with second lowest ID, i.e., $a_{2}$) detects the exact location of the black hole.
-Further, this agent continues to perpetually explore the ring $R$ by avoiding the black hole node.
+If no agent gets destroyed by the Byzantine black hole while performing an iteration of Algorithm 1 but some pebbles are destroyed by the Byzantine black hole, then by Lemma 9.1, one agent (more precisely, the agent with second lowest ID, i.e., $a_{2}$) detects the exact location of the Byzantine black hole.
+Further, this agent continues to perpetually explore the ring $R$ by avoiding the Byzantine black hole node.
 Lemma 9.3.
@@
   1. 1.
-    The exact location of the black hole is detected by at least one agent.
+    The exact location of the Byzantine black hole is detected by at least one agent.
   2. 2.
-    All alive agents become colocated and they agree about two consecutive nodes $S_{lft}$ and $S_{rgt}$, one among which is the black hole.
-Proof 4.18.
-Note that the agent which has entered the black hole is either $a_{1}$ or $a_{2}$, based on which we have the following conditions:
-  • $a_{1}$ falls in to the black hole: This situation can only occur when $a_{1}$ following sequence $SEQ1$ is in state Leader, visits the black hole node along with its pebble.
-Now, based on the adversarial choice of byzantine black hole, the pebble may or may not be destroyed by the black hole.
-Irrespective of which, $a_{2}$ within 3 additional rounds reaches the black hole node while in state Follower-Find, and absence of $a_{1}$ triggers the agent $a_{2}$ to determine the current node as black hole, and immediately leave the current node.
-  • $a_{2}$ falls in to the black hole: This can only happen when $a_{2}$ is not with $a_{1}$ at the black hole node.
+    All alive agents become co-located and they agree about two consecutive nodes $S_{lft}$ and $S_{rgt}$, one among which is the Byzantine black hole.
+Proof 9.4.
+Note that the agent that has entered the Byzantine black hole is either $a_{1}$ or $a_{2}$, based on which we have the following conditions:
+  • $a_{1}$ falls in to the Byzantine black hole: This situation can only occur when $a_{1}$ following sequence $SEQ1$ is in state Leader, visits the Byzantine black hole node along with its pebble.
+Now, based on the adversarial choice of the Byzantine black hole, the pebble may or may not be destroyed by the Byzantine black hole.
+Irrespective of which, $a_{2}$ within 3 additional rounds reaches the Byzantine black hole node while in the state Follower-Find, and the absence of $a_{1}$ triggers the agent $a_{2}$ to determine the current node as the Byzantine black hole, and immediately leave the current node.
+  • $a_{2}$ falls in to the Byzantine black hole: This can only happen when $a_{2}$ is not with $a_{1}$ at the Byzantine black hole node.
 Otherwise, both agents will be destroyed, contradicting our claim.
-Note that, $a_{2}$ can either be with the pebble or alone, while it is consumed by the black hole.
+Note that, $a_{2}$ can either be with the pebble or alone, while it is destroyed by the Byzantine black hole.
 Also note that, since $a_{2}$ is executing the sequence $SEQ2$, hence the distance between $a_{1}$ and $a_{2}$ can be at most 2.
-So during an iteration, if $a_{2}$ is destroyed by a black hole, then the black hole must be any one of the two consecutive counter-clockwise nodes, from the current position of $a_{1}$.
+So during an iteration, if $a_{2}$ is destroyed by a Byzantine black hole, then the Byzantine black hole must be any one of the two consecutive counter-clockwise nodes, from the current position of $a_{1}$.
 If the current node is not $home$, in this case, $a_{1}$ gets to know about this fact, only when it finds the absence of $a_{2}$ even after waiting for 3 rounds, at the current node.
-Now, according to our algorithm, it leaves the pebble accompanied by it, and moves from the current node in clockwise direction to $home$ with state Report-Leader.
+Now, according to our algorithm, it leaves the pebble accompanied by it and moves from the current node in the clockwise direction to $home$ with state Report-Leader.
 After, reaching $home$, it waits until the end of this iteration, i.e., until $4n$ rounds from the start of this iteration and moves to state Find-Pebble.
-During the $14n+$ 1-th round from the start of the current iteration, $a_{3}$ finds that it is only with $a_{1}$, whereas only one pebble exits, this leads the agent to change its state to Find-Pebble.
+During the $14n+$ 1-th round from the start of the current iteration, $a_{3}$ finds that it is only with $a_{1}$, whereas no pebble exits, this leads the agent to change its state to Find-Pebble.
 In this state, both the agents move counter-clockwise from $home$ until they encounter a pebble.
-Note that, this pebble is the one left by $a_{1}$ after determining the fact that $a_{2}$ has entered the black hole.
-So, now both these agents $a_{1}$ and $a_{3}$, declare the node with pebble as $home$, whereas also denote the adjacent counter-clockwise node from the current node as $S_{lft}$, whereas the other node adjacent to $S_{lft}$ in counter-clockwise direction as $S_{rgt}$.
-On the other hand, if the current node of $a_{1}$ is $home$, while $a_{2}$ is destroyed by the black hole, then $a_{2}$ can never reach $home$ along with the pebble it was carrying by following the sequence $SEQ2$.
-So, at the end of this iteration, both $a_{1}$ and $a_{3}$ finds that there are two agents (i.e., $a_{1}$ and $a_{3}$) and two pebbles, respectively at $home$.
-This leads both of them to conclude that, the adjacent counter-clockwise node to be $S_{rgt}$, whereas the adjacent counter-clockwise node of $S_{rgt}$ as $S_{lft}$.
-So, in each case we showed that either of our two conditions hold.
+Note that, this pebble is the one left by $a_{1}$ after determining the fact that $a_{2}$ has entered the Byzantine black hole.
+So, now both these agents $a_{1}$ and $a_{3}$, declare the node with the pebble as $home$, whereas also denote the adjacent counter-clockwise node from the current node as $S_{lft}$, whereas the other node adjacent to $S_{lft}$ in the counter-clockwise direction as $S_{rgt}$.
+On the other hand, if the current node of $a_{1}$ is $home$, while $a_{2}$ is destroyed by the Byzantine black hole, then $a_{2}$ can never reach $home$ along with the pebble it was carrying by following the sequence $SEQ2$.
+So, at the end of this iteration, both $a_{1}$ and $a_{3}$ find that there are two agents (i.e., $a_{1}$ and $a_{3}$) and one pebble, respectively at $home$.
+This leads both of them to conclude that, the adjacent counter-clockwise node is $S_{rgt}$, whereas the adjacent counter-clockwise node of $S_{rgt}$ is $S_{lft}$.
+So, in each case, we showed that either of our two conditions holds.
 Remark 9.5.
-Following from Lemma 4.17, if exactly one agent gets destroyed by the byzantine black hole and the exact location of the black hole is detected by at least one agent, then our algorithm ensures that the ring is perpetually explored by at least one alive agent while avoiding the black hole node.
-On the contrary, all alive agents become co-located and they agree about two consecutive nodes $S_{lft}$ and $S_{rgt}$, among which one is a black hole, then until and unless another agent falls in to the black hole, our algorithm ensures that the ring is perpetually explored, by the remaining alive agents.
-Note that, by Lemma 4.20 while any one among them gets destroyed, then the other agent detects the exact black hole node, and then further perpetually explores the ring $R$, avoiding the black hole node.
+Following from Lemma 9.3, if exactly one agent gets destroyed by the Byzantine black hole and the exact location of the Byzantine black hole is detected by at least one agent, then our algorithm ensures that the ring is perpetually explored by at least one alive agent while avoiding the Byzantine black hole node.
+On the contrary, all alive agents become co-located and they agree about two consecutive nodes $S_{lft}$ and $S_{rgt}$, among which one is a Byzantine black hole, then until and unless another agent falls into the Byzantine black hole, our algorithm ensures that the ring is perpetually explored, by the remaining alive agents.
+Note that, by Lemma 9.6 while anyone among them gets destroyed, then the other agent detects the exact Byzantine black hole node, and then further perpetually explores the ring $R$, avoiding the Byzantine black hole node.
 Lemma 9.6.
-Algorithm PerpExplore-Coloc-Pbl ensures that if two among three agents gets consumed by the byzantine black hole, then the remaining agent knows the exact location of the black hole within some additional finite rounds without being destroyed.
-Proof 4.21.
-Let us suppose, by contradiction, there exists a round $r$ ($>0$), within which two agents gets destroyed by the black hole, whereas the alive agent is unable to detect the exact location.
+Algorithm PerpExplore-Coloc-Pbl ensures that if two among three agents get destroyed by the Byzantine black hole, then the remaining agent knows the exact location of the Byzantine black hole within some additional finite rounds without being destroyed.
+Proof 9.7.
+Let us suppose, by contradiction, there exists a round $r$ ($>0$), within which two agents get destroyed by the Byzantine black hole, whereas the alive agent is unable to detect the exact location.
 Now we have the following cases:
-  • Both agents gets consumed by the black hole in the same iteration: Note that in this case, the consumed agents are $a_{1}$ and $a_{2}$, respectively, which while executing the sequences $SEQ1$ and $SEQ2$ enters the black hole.
-Also note that, $a_{1}$ and $a_{2}$ can only be together at a node, when $a_{2}$ has left the pebble it is carrying in the counter-clockwise node.
+  • Both agents get destroyed by the Byzantine black hole in the same iteration: Note that in this case, the destroyed agents are $a_{1}$ and $a_{2}$, respectively, which while executing the sequences $SEQ1$ and $SEQ2$ enters the Byzantine black hole.
+Also, note that $a_{1}$ and $a_{2}$ can only be together at a node when $a_{2}$ has left the pebble it is carrying in the counter-clockwise node.
 So, now $a_{3}$ after the end of this iteration, finds that no agent among $a_{1}$ and $a_{3}$ has reached $home$, which triggers $a_{3}$ to change to state Find-BH.
-In this state, it moves clockwise, until it finds a pebble (this pebble is the one left by $a_{2}$ in the counter-clockwise node of the black hole).
-Whenever a pebble is found, $a_{3}$ concludes that the next node is the black hole, which contradicts our claim.
-  • There exists an iteration where exactly one agent is destroyed and the other one is destroyed during round $r$: Now in this case let $r^{\prime}<r$ be a round when exactly an agent is destroyed during an iteration, say $i>0$.
-Then by Lemma 4.17, within finitely many additional rounds $r_{0}$, where $r_{0}+r^{\prime}<r$, both the alive agents becomes colocated and agree about two consecutive nodes $S_{lft}$ and $S_{rgt}$, one among which must be the black hole.
-Observe that according to our algorithm, the node which is clockwise adjacent to $S_{rgt}$ is determined to be the $home$ (if not already so).
-Note that the other case according to Lemma  4.17 where after exactly one agent is destroyed by the black hole during an iteration at a round $r^{\prime}<r$, there is an agent that knows the exact location of the black hole within a finite additional rounds and the agent starts exploring the ring avoiding the black hole and so even after round $r$ it stays alive knowing the exact location of the black hole contradicting our assumption.
-Now, when both the agents are colocated during a round say $r_{0}+r^{\prime}<r$ and knows that the black hole must be any of the two consecutive nodes $S_{lft}$ or, $S_{rgt}$.
+In this state, it moves clockwise, until it finds a pebble (this pebble is the one left by $a_{2}$ in the counter-clockwise node of the Byzantine black hole).
+Whenever a pebble is found, $a_{3}$ concludes that the next node is the Byzantine black hole, which contradicts our claim.
+  • There exists an iteration where exactly one agent is destroyed and the other one is destroyed during round $r$: Now, in this case, let $r^{\prime}<r$ be a round when exactly an agent is destroyed during an iteration, say $i>0$.
+Then by Lemma 9.3, within finitely many additional rounds $r_{0}$, where $r_{0}+r^{\prime}<r$, both the alive agents become co-located and agree about two consecutive nodes $S_{lft}$ and $S_{rgt}$, one among which must be the Byzantine black hole.
+Observe that according to our algorithm, the node that is clockwise adjacent to $S_{rgt}$ is determined to be the $home$ (if not already so).
+Note that in the other case according to Lemma  9.3 where after exactly one agent is destroyed by the Byzantine black hole during an iteration at a round $r^{\prime}<r$, there is an agent that knows the exact location of the Byzantine black hole within a finite additional round and the agent starts exploring the ring avoiding the Byzantine black hole and so even after round $r$ it stays alive knowing the exact location of the Byzantine black hole contradicting our assumption.
+Now, when both the agents are co-located during a round say $r_{0}+r^{\prime}<r$ and know that the Byzantine black hole must be any of the two consecutive nodes $S_{lft}$ or $S_{rgt}$.
 In this case, they move into state Detection.
-After which one agent moves clockwise until it reaches $S_{lft}$ and the other one moves counter clockwise until it reaches $S_{right}$ and moves back to $home$.
+After which one agent moves clockwise until it reaches $S_{lft}$ and the other one moves counter-clockwise until it reaches $S_{rgt}$ and moves back to $home$.
 The agent that visits $S_{rgt}$ waits for the other agent at $home$ for $62n-$ 6 rounds which is sufficient for the other agent to return back at $home$.
-Now when during round $r$, another agent among these two gets destroyed, within at most $(2n-4)$ rounds they fails to meet at $home$.
-Which trigers the remaining alive agent to determine exactly which one among $S_{lft}$ or, $S_{rgt}$ is the black hole.
+Now when during round $r$, another agent among these two gets destroyed, within at most $(2n-4)$ rounds (among the total $62n-$ 6 rounds, within which both agents are to meet) they fail to meet at $home$.
+Which triggers the remaining alive agent to determine exactly which one among $S_{lft}$ or $S_{rgt}$ is the Byzantine black hole.
 This also contradicts the claim we made earlier.
-Thus, The algorithm PerpExplore-Coloc-Pbl ensures that, if two agent among three agents gets consumed by the byzantine black hole, the remaining agent must know the exact location of the black hole in some finite additional rounds without getting destroyed.
-Corollary 4.22.
-Number of alive agents must be at least 2, when the exact position of the byzantine black hole is not known to them.
-From Lemma 4.20 and algorithm PerpExplore-Coloc-Pbl we have the following theorem
-Theorem 4.23.
-A set of 3 synchronous colocated agents equipped with one pebble each are necessary and sufficient to solve PerpExploration-BBH on a ring $R$.
-Remark 4.24 (Algorithm on whiteboard model of communication).
-The PerpExplore-Coloc-Pbl algorithm (i.e., Algorithm 2 ) also works for colocated agents with whiteboard model of communication.
-The agents can easily simulate a pebble on a node by a bit marked on that nodes memory.
-Also, carrying a pebble from one node to another can be simulated by removing the marking from the current node’s memory and move to a new node where the agent writes a new bit.
-Thus 3 colocated agents are also sufficient to solve Perpexplore-BBH when the nodes are equipped with whiteboards.
-5 Scattered Agents
-5.1 Communication: Pebble
+Thus, the algorithm PerpExplore-Coloc-Pbl ensures that, if two among three agents get destroyed by the Byzantine black hole, the remaining agent must know the exact location of the Byzantine black hole in some finite additional rounds without getting destroyed.
+10 Algorithm PerpExplore-Scat-Pbl
+10.1 Detailed description and Pseudocode of the Algorithm
 In this section, we discuss the model where the agents are placed arbitrarily along the nodes of the ring $R$ (note that each such node must be a safe node), where each agent has a movable token, which it can carry along with it, and acts as a mode of inter agent communication.
@@
 We here describe and present the pseudocode considering the first case only.
-The case with 3 agents in one node can be dealt by instructing the colocated agents to execute PerpExplore-Coloc-Pbl when there are three agents on the current node (i.e., from the beginning).
+The case with 3 agents in one node can be dealt by instructing the co-located agents to execute PerpExplore-Coloc-Pbl when there are three agents on the current node (i.e., from the beginning).
 The algorithm we discuss here can also be used for the remaining cases with slight modifications.
 These modifications are described in the Remark 10.7
-The main idea of the PerpExplore-Scat-Pbl algorithm is that 4 agents perpetually explore the ring, when no agents are consumed by the black hole.
+The main idea of the PerpExplore-Scat-Pbl algorithm is that 4 agents perpetually explore the ring, when no agents are destroyed by the Byzantine black hole.
 Our algorithm ensures that during this exploration at most one agent can be destroyed.
 In this scenario, the remaining agents within further finite time, gather at a single node (which is the starting node of either of these 4 agents, hence it is safe), and after which they together start executing the algorithm PerpExplore-Coloc-Pbl.
+Remark 10.1.
+As mentioned in PerpExplore-Coloc-Pbl requires, 3 agents and 2 pebbles to perform PerpExploration-BBH, but in this case we instruct the agents to perform PerpExplore-Coloc-Pbl, whenever three agents and at least three pebbles are able to gather at a node.
+These agents performs PerpExplore-Coloc-Pbl with a slight modification that the excess pebble (or pebbles) remain at their gathered node, which they term as $home$.
+So, all the conditions that are mentioned, remains same in this case as well, only the condition where an agent changes to certain state $s_{i}$ by observing $t_{i}$ (say) number of pebbles at $home$, transforms to the fact that the agent changes to state $s_{i}$ by observing $t_{i}+x$ (where $x=\text{total \# pebbles}-2$) number of pebbles at $home$.
 At the beginning of the Algorithm 2 all agents are in state Initial1.
@@
 Now since during each move at state Forward, an agent increases the variable $size$ by 1, the $size$ variable must store the length between an agents own $home$ and the clockwise nearest $home$ when the agent finds another pebble.
-Beyond this point (i.e., when $s=1$) the agents will use the $size$ variable as reference to calculate how much it should move, as after this, as there is a possibility that the agent may not find a pebble at the clockwise nearest $home$.
+Beyond this point (i.e., when $s=1$) the agents will use the $size$ variable as reference to calculate how much it should move, as after this, there is a possibility that the agent may not find a pebble at the clockwise nearest $home$.
 On the contrary, if $s=1$, this implies the agent already knows how much distance it should travel from its own $home$ in order to reach the clockwise nearest $home$.
@@
 If these two conditions are satisfied then the agent just increments $W_{time}$ by 1.
-On the other hand, if it finds that $W_{time}<2n-size$ but #$\#Agent$ at the current node is 2, then it changes its state to Gather2 and updates $W_{time}$ to 0.
+On the other hand, if it finds that $W_{time}<2n-size$ but #$\#Agent$ at the current node is 2, then it changes its state to Gather2 and moves clockwise and updates $W_{time}$ to 0.
 Otherwise, if it finds $W_{time}<2n-size$ but #$\#Agent$ at the current node is 3 then it updates $W_{time}$ to 0, whereas changes its state to Coloc.
 Lastly, whenever it finds $W_{time}=2n-size$, it changes its state to Forward.
-Note that if no agent gets destroyed by the black hole, then our algorithm guarantees that an agent can neither changes its state to Gather1 or in state Gather2.
-It is because, an agent, say $a_{1}$ can only find more than one pebble at its $home$ node, only when another agent say $a_{2}$ (exploring the counter-clockwise nodes of $home$ of $a_{1}$), which was supposed to fetch the pebble left by $a_{1}$ at its $home$, is destroyed.
-In this case, whenever $a_{1}$ returns while in state fetch with another pebble, it finds more than one pebble.
-So the first agent, which first finds this anomaly changes its state to Gather1 from the state Fetch.
-Now this agent must initiate gathering with the remaining two agents (as it is the one which has first detected the anomaly of more than one pebble at its $home$).
-Note that, if while $a_{1}$ starts its gathering phase, moving in a clockwise direction, there is a possibility that other agents are still moving either in clockwise or counter-clockwise direction.
-So, to address this challenge, the algorithm instructs $a_{1}$ (i.e., the agent which is in state Gather1) to wait at the current node (i.e., its $home$) for precisely $n-size-1$ rounds, and then start moving clockwise direction.
-This waiting time of $a_{1}$ guarantees all the other alive agents to reach their respective $home$ while in state Fetch.
-Now, in order to avoid further movement of remaining agents (i.e., again change to state Forward, since these agents are yet to detect any anomaly, which makes gathering a bit challenging), our algorithm instructs the remaining agents to wait at the current node for a certain number of rounds (i.e., more precisely, $12n-size-$ 1).
-This waiting period is sufficient, for the agent $a_{1}$ to meet with the other alive agents while in state Gather1.
-Note that, if no agent detects any anomaly, then they directly change their state to Wait2 from the state Fetch.
-If no such agent exists which encounters any anomaly and changes its state to Gather1 or Gather2 then this waiting time of $12n-size-$ 1 rounds, also ensures that all of them must move to state Forward at the exact same round (i.e, ($23n+$ 2)-th round from the previous time it moved to state Forward) [^1] At $(n+1)$-th round an agent moves to state Fetch after that it moves for $size$ rounds to reach its $home$ in counter-clockwise direction and changes to state Wait2.
-So at round $n+3+size$ an agent changes to state Wait2.
-After that waits for further $12n-size-$ 1 rounds and changes to state forward, taking a total of $23n+$ 2 rounds .
-Now we have the following observation {observation} Let $t$ be a round when all alive agents moved into state Forward.
-If no agent detects any anomaly (i.e., no agent finds more than one pebble at home) then, at round $t+3n+2$ all the agents together move into state Forward again.
-An agent in state Gather1, is the first agent to detect the anomaly, hence it is first instructed to wait at the current node, i.e., its $home$ for the first $n-size-1$ rounds, and then start moving in a clockwise direction, while accompanying all the pebbles at the current node  [^2] Note that we have considered that in this case an agent, while in state Gather1 or Gather2 can carry more than one pebble, which can be easily restricted to one agent can carry only one pebble, but in that case the agent needs to return and carry one pebble each time and reach back to its current node.
-This movement in the clockwise direction continues until the number of agent at the current node is exactly 3, after which it changes to state Coloc.
-An agent in state Gather2 (i.e., the alive agents which are not the first to detect the anomaly), first checks whether the current node has exactly 3 agents on it, if not then it moves in a counter-clockwise direction, while accompanying the pebbles present at its $home$,
+[…]
 1 Input: $n$
@@
 65 else if $W_{time}<2n-size\wedge\#Agent$ at $Current-Node$ is 2 then
-66       Initialize $W_{time}=0$ and change to state Gather2.
+66       Initialize $W_{time}=0$ and change to state Gather2 and moves clockwise.
 67
@@
 Algorithm 2 PerpExplore-Scat-Pbl
-until it satisfies the condition, where it finds a node on which a total of 3 agents exists. After which they change their state to Coloc.
+An agent in state Gather1, is the first agent to detect the anomaly, hence it is first instructed to wait at the current node, i.e., its $home$ for the first $n-size-1$ rounds, and then start moving in a clockwise direction, while accompanying all the pebbles at the current node  [^2] Note that we have considered that in this case an agent, while in state Gather1 or Gather2 can carry more than one pebble, which can be easily restricted to one agent can carry only one pebble, but in that case the agent needs to return and carry one pebble each time and reach back to its current node.
+This movement in the clockwise direction continues until the number of agent at the current node is exactly 3, after which it changes to state Coloc.
+An agent in state Gather2 (i.e., the alive agents which are not the first to detect the anomaly), first checks whether the current node has exactly 3 agents on it, if not then it moves in a counter-clockwise direction, while accompanying the pebbles present at its $home$, until it satisfies the condition, where it finds a node on which a total of 3 agents exists.
+After which they change their state to Coloc.
 In state Coloc, an agent declares its current node as $home$, while also changes its state to Initial  [^3] The state Initial is defined in Algorithm 1 and then executes algorithm PerpExplore-Coloc-Pbl.
-5.1.1 Correctness and Complexity
+10.2 Correctness and Complexity
 In this section, we discuss the correctness and complexity of our algorithm PerpExplore-Scat-Pbl.
+Before that, let us define an iteration of the algorithm PerpExplore-Scat-Pbl.
+An iteration is a set of $13n+$ 1 consecutive rounds starting from the round in which all alive agents change to state Forward together.
+Next we prove the following lemma.
 Lemma 10.2.
-PerpExplore-Scat-Pbl guarantees to solve PerpExploration-BBh problem if no agent is consumed by the byzantine black hole.
-Proof 5.2.
-Let us define a segment for an agent $a_{1}$ be $Seg(a_{1})$, where $Seg(a_{1})$ is defined to be the collection of consecutive nodes in the clockwise direction, starting from its own $home$ and its clockwise nearest $home$ (which is the $home$ of another agent say, $a_{2}$).
-Note that, since no agent is consumed by the black hole, any agent, say $a_{1}$, while executing the Algorithm 3, starts moving in clockwise direction in state Forward, reaches the nearest clockwise $home$, i.e., the last node of this segment $Seg(a_{1})$, then changes its state to Wait1, and after waiting for $n-size$ rounds, again changes its state to Fetch.
-In state Fetch, the agent basically moves counter-clockwise, until and unless it reaches its own $home$, which is nothing but the first node of this segment $Seg(a_{1})$.
-Also, note that, each agent, changes their state to Forward and Fetch together at the same round, while exploring their respective segments.
-Now, note that, the union of all these segments is basically the whole ring $R$ (i.e., more specifically $\cup^{4}_{i=1}Seg(a_{i})=R$) .
-Hence, this guarantees that the ring is perpetually explored if no agent is consumed by the black hole.
+If no agents are destroyed by the Byzantine black hole until the completion of $p$-th iteration, then for each $j-$th iteration, every node of the ring $R$ is visited by at least one agent, where $j\in\mathbb{N}^{+},\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ j\leq p$.
+Proof 10.3.
+Let $a_{0},a_{1},a_{2}$ and $a_{3}$ be 4 agents initially located at $h_{0},h_{1},h_{2}$ and $h_{3}$ respectively where, $h_{0},h_{1},h_{2},h_{3}$ are in clockwise order starting from $h_{0}$.
+By $Seg(a_{i})$ we denote the clockwise arc starting from $h_{i}$ and ending at $h_{(i+1)\pmod{4}}$.
+Note that in any of the $j-$th iteration ($j\leq p$), an agent $a_{i}$ (for all $i\in\{0,1,2,3\}$) moves clockwise along $Seg(a_{i})$ starting from $h_{i}$ while in state Forward and reaches $h_{(i+1)\pmod{4}}$.
+Then $a_{i}$ in the same iteration moves back to $h_{i}$ again while being in state Fetch.
+Thus each node of $Seg(a_{i})$ is explored twice in $j$-th iteration by $a_{i}$, where $j\leq p$.
+Since for any $v\in R$, $v\in Seg(a_{j}),$ for some $j\in\{0,1,2,3\}$, and in each iteration $a_{j}$ explores each node of $Seg(a_{j})$ until $p-$th iteration, the result follows.
+Corollary 10.4.
+If no agents are destroyed, the agents can perpetually explore a ring execuing the algorithm PerpExplore-Scat-Pbl.
+Next we prove, after the consumption of one agent by the Byzantine black hole, during a subsequent iteration or at the current iteration, at most one agent can change its state to Gather1 and gathers with two more agents before that iteration ends.
 Lemma 10.5.
-If an agent is consumed by the byzantine black hole, then within an additional finite rounds, there exists exactly one agent, which changes its state to Gather1.
-Proof 5.4.
-Let $a_{1}$ be an agent exploring $Seg(a_{1})$.
-Without loss of generality, let the black hole be in $Seg(a_{1})$.
-Also let $a_{2}$ be another agent which is exploring the segment $Seg(a_{2})$, where $home$ of $a_{2}$ is the common node between $Seg(a_{1})$ and $Seg(a_{2})$.
-Also, let $a_{3}$ be an agent which is exploring the segment $Seg(a_{3})$, where the $home$ of $a_{3}$ is the common node between $Seg(a_{2})$ and $Seg(a_{3})$.
-Note that, $a_{1}$ can be consumed by the black hole, either in state Forward or in state Fetch.
-  • If $a_{1}$ is consumed while in state Forward, then it fails to bring back the pebble from the $home$ of $a_{2}$, while it is scheduled to change its state to Fetch, after ending the states Forward and Wait1.
-Note that, since we have assumed that the black hole is in $Seg(a_{1})$, and $Seg(a_{1})\cap Seg(a_{2})=\{home\text{ of }a_{2}$}.
-Thus, $a_{2}$ must bring a pebble from the $home$ of $a_{3}$ to it’s own $home$, after it changes its state to Fetch.
-So, now as $a_{2}$ reaches its own $home$ along with the pebble it carried from the $home$ of $a_{3}$, it finds that there already exists another pebble, which the agent $a_{1}$ is unable to carry back to its own $home$, as it is already consumed by the black hole.
-This causes an anomaly for $a_{2}$ at its own $home$ (i.e., encountering more than one pebble).
-After detecting the anomaly $a_{2}$ immediately changes its state to Gather1.
-Note that, after $a_{1}$ is consumed by the black hole, it has taken an additional at most $size-1+(n-size)+1+size+1=n+size+1$ 333calculates the number of rounds it took $a_{2}$ to change to state Gather1, where at most $size-1$ rounds required in state Forward after $a_{1}$ gets consumed by the black hole, $(n-size)$ rounds for changing to state Wait1 and waiting, 1 round for changing from Wait1 to Fetch, then $size$ rounds for moving in state Fetch and returning back to its $home$, and lastly 1 round for changing to state Gather1 from Fetch.
-rounds to change the state of $a_{2}$ to Gather1.
-Further in this case we show that no other agents except $a_{2}$ changes to state Gather1.
-Let $t^{\prime}$ be the round when $a_{2}$ changes its state to Gather 1.
-And $t<t^{\prime}$ be the last round before $t^{\prime}$ when all the agents moved into state Forward.
-Then $t^{\prime}=t+n+1+size+1$.
-Next $a_{2}$ waits for $n-size-1$ rounds that is until round $t+2n+1$ and starts moving counter clockwise being in state Gather1.
-Note that by observation 1, the other agents stay at their home until round $t+3n+1$.
-So from $t+2n+2$ until $t+3n+1$, these $n-1$ rounds are enough for $a_{2}$ to meet the other two while they are still being in the state Wait2.
-So they change their state to Gather2 and eventually to Coloc.
-As an agent can not change state to Gather1 from state Coloc or Gather2, no other agent changes to state Gather1.
-  • If $a_{1}$ is consumed while in state Fetch.
-In this case, $a_{2}$ after returning back to its $home$ carrying the pebble from the nearest clockwise $home$, i.e., the $home$ of $a_{3}$, finds no anomaly.
-It is because, $a_{1}$ has been consumed by the black hole while in state Fetch, so that means, the pebble at $home$ of $a_{2}$ has been carried by $a_{1}$.
-Since, $a_{2}$ does not encounter any anomaly, hence it drops the pebble at its $home$, changes to state Wait2, waits at $home$ for $12n-size-$ 1 rounds and then changes to state Forward again.
-So, while executing the same execution, after reaching the clockwise nearest $home$, i.e., $home$ of $a_{3}$, it again changes its state to Wait1, waits for $n-size-1$ rounds, and further changes its state to Fetch.
-Note that, here at the $home$ of $a_{3}$, there exists a pebble which was earlier dropped by $a_{3}$, while it changed its state from Fetch to Wait2.
-So, $a_{2}$ carries this pebble and reaches its own $home$, where it finds the anomaly.
-It is because, as $a_{1}$ has been consumed by the black hole, it is unable to carry the pebble dropped by $a_{2}$ during its last state change from Fetch to Wait2.
-This phenomenon triggers $a_{2}$ to change its state to Gather1.
-Note that, since $a_{1}$ has been consumed by the black hole, it takes an additional $(size-1)+1+(2n-size)+size+1+(n-size)+size+1=3n+size+2$ 222calculates the number of rounds it took $a_{2}$ to change to state Gather1, where $size-1$ rounds required in state Fetch after $a_{1}$ gets consumed by the black hole, 1 round for changing its state from Fetch to Wait2, after this the agent waits for $12n-size-$ 1 rounds and in one additional round changes to state Forward, in this state it takes $size$ rounds, 1 round required to change to state Wait1, then $n-size-1$ rounds the agent waits, whereas 1 more round required to change to state Fetch, finally $size$ rounds in state Fetch, and lastly 1 round for changing its state to Gather1rounds to change $a_{2}$’s state to Gather1.
-And as shown for the first case it can be similarly showed that except $a_{2}$ no other agent changes state to Gather1
-Lemma 5.5.
-After the first agent changes its state to Gather1, it takes an additional at most $22n-$ 2 rounds, for the remaining agents to gather at a single node (i.e., the nearest counter-clockwise $home$ of the black hole node) along with their pebbles.
-Proof 5.6.
-Let $a_{2}$ be the agent which changes to state Gather1 at $t$-th round (where $t>0$), due to the fact that it encounters an anomaly at its $home$ because $a_{1}$ (w.l.o.g) has been consumed by the black hole, and the length of the segment it explores be denoted by $size$ (i.e., length of $Seg(a_{2})=size$).
-Also, let $t_{0}$ ($0<t_{0}<t$) be the last round when each agent changed their states to Forward.
-Note that $t$ should be exactly equal to $t_{0}+size+1+(n-size-1)+1+size+1=t_{0}+n+size+2$ 333movement in state Forward for $size$ rounds, then one round for change to state Wait1, then stay in state Wait1 for $n-size-1$ rounds, after which change to state Fetch for one round, and move in state Fetch for $size$ rounds, and finally change to state Gather1 for one round.
-Note that $a_{2}$ after changing its state to Gather1 (current round is $t_{0}+n+size+2$, which is also the $t-$th round according to our definition of $t$), further waits for $n-size-1(\leq n-2$, as $size\geq 1$ ) rounds (current round $t_{0}+2n+1$), and starts moving clockwise from $(t_{0}+2n+2)-$th round.
-On the other hand, an alive agent which was in state Fetch at $t-$th round, must have changed its own state to Fetch at $(t_{0}+n+1)-$th round (refer Observation 5.1).
-From Fetch, it takes further at most $n$ (since length of a segment can be at most $n-1$) rounds to change to state Wait2.
-Note that any other alive agent except $a_{2}$ must change to state Wait2 after completion of state Fetch as by Lemma 5.3, exactly one agent (here $a_{2}$) already changed its state to Gather1 from Fetch at round $t=t_{0}+n+size+2$).
-So, any agent except $a_{2}$ at the round $(t_{0}+2n+1)-$th round must be in state Wait2, and should remain in this state until $(t_{0}+3n+2)-$th round, if no anomaly is detected (refer Observation 1) or meets with other agents.
-So, any alive agent except $a_{2}$ remains at their $home$ within $t_{0}+2n+1$ to $t_{0}+3n+2$ rounds.
-This is by Lemma 5.3, only $a_{2}$ being the first agent to detect anomaly can change its state to Gather1, whereas the remaining agent can only change its state to Gather2 or Coloc, when they meet with $a_{2}$, while in state Wait2.
-Further, note that an alive agent except $a_{2}$, waits at their respective $home$ for $n+1$ rounds since, $a_{2}$ starts moving clockwise being in state Gather1 (after $(t_{0}+2n+1)-$th round).
-Now, when $a_{2}$ reaches its clockwise neighbor (say) $a_{3}$, at its $home$.
-$a_{3}$ must be at state Wait2, in this case, $a_{3}$ changes its state to Gather2, and starts moving clockwise along with $a_{2}$ carrying their respective pebbles, until they meet another agent, (say) $a_{4}$, at its $home$, while in state Wait2.
-After which they all change to state Coloc, which also guarantees that they have at least 3 pebbles at the current node (since $a_{2}$ carried at least 2 pebbles from its own $home$ where it detects the anomaly, $a_{3}$ also carried its one pebble present at its own $home$).
-Now since no agent can change to state Gather1 from Gather2 or Coloc, $a_{2}$ is the only agent that changes to state Gather1.
-Note that, before gathering $a_{2}$, $a_{3}$ and $a_{4}$ can never be destroyed, as they traverse only the nodes in $\cup^{4}_{i=2}Seg(a_{i})$, whereas the black hole is in $Seg(a_{1})\backslash\{home$ of $a_{1}$, $home$ of $a_{2}\}$.
+In an iteration, at most one agent can change its state to Gather1 executing PerpExplore-Scat-Pbl.
+Furthermore, before the end of that iteration the agent gathers with two more agents at a safe node with at least 3 pebbles.
+Proof 10.6.
+Let $a_{0},a_{1},a_{2}$ and $a_{3}$ be 4 agents initially located at $h_{0},h_{1},h_{2}$ and $h_{3}$ respectively where, $h_{0},h_{1},h_{2},h_{3}$ are in clockwise order starting from $h_{0}$.
+By $Seg(a_{i})$ we denote the clockwise arc starting from $h_{i}$ and ending at $h_{(i+1)\pmod{4}}$.
+Without loss of generality, let the first agent that is destroyed by the Byzantine black hole be $a_{0}$ during the $p-$th iteration.
+So, the Byzantine black hole $v_{b}$, must be in $Seg(a_{0})$.
+Since, no agents changes to state Gather1, before an agent is destroyed.
+Hence, until $(p-1)-$th iteration, there does not exist any agent, which changes its state to Gather1.
+Now, we have following two cases, which are based on the state $a_{0}$ was in, before it gets destroyed by the Byzantine black hole.
+Case-I: $a_{0}$ was in state Forward when it was destroyed.
+So, $a_{0}$ fails to reach $h_{1}$ to bring the pebble, left by $a_{1},$ back to $h_{0}$ during $p-$th iteration.
+So, when $a_{1}$ reaches $h_{1}$ at the $p-$th iteration, along with the pebble it collected from $h_{2}$, it finds two pebble at the node $h_{1}$.
+This happens at $(n+2+|Seg(a_{1})|)-$th round of $p-$th iteration.
+In the same round $a_{1}$ changes to state Gather1.
+Next it waits an additional $n-|Seg(a_{1})|-1$ rounds and then in the $(2n+2)-$ th round of the $p-$ th iteration, it starts moving clockwise along with both the pebbles.
+Note that, all other alive agents (i.e., $a_{2}$ and $a_{3}$) reaches their home by the $(2n+2)-$th round of the $p-$th iteration.
+Also upon reaching their respective $home$, $a_{2}$ and $a_{3}$ both sees exactly one pebble (i.e., the pebble they carried back to their $home$) so they change their state to Wait2.
+By Observation 1, from $(2n+2)$-th round of the iteration upto the end of the iteration (i.e., $(3n+1)-$th round of the iteration), both $a_{2}$ and $a_{3}$ stays at $h_{2}$ and $h_{3}$, respectively in state Wait2.
+So, after $a_{1}$ starts moving clockwise in the $(2n+2)-$th round, it first meets $a_{2}$ in round $2n+1+|Seg(a_{1})|$ (which is $\leq 3n-1$ as $|Seg(a_{1})|\leq n-2$) of the $p-$th iteration.
+And then meets $a_{3}$ at round $2n+1+|Seg(a_{1})|+|Seg(a_{2})|\leq 3n$ (as $|Seg(a_{1})|+|Seg(a_{2})|\leq n-1)$ which is also in $p-$th iteration.
+Now note that, the moment $a_{1}$ meets $a_{2}$, it changes to state Gather2 and starts moving together with $a_{1}$ along with its own pebble until they meet $a_{3}$.
+When all three of them gathers together at $h_{3}$ there is atleast 3 pebbles at $h_{3}$ (two pebbles carried by $a_{1}$, one pebble carried by $a_{2}$, and one pebble remains for $a_{3}$).
+In this scenario, they change their state to Coloc.
+Note that in $p-$th iteration only $a_{1}$ changes to state Gather1 as other agents can not change to state Gather1 from either from Wait2, or, Gather2, or, Coloc.
+So, for this case the claimed result holds.
+Case-II: Let $a_{0}$ was in state Fetch while it was destroyed by the Byzantine black hole node $v_{b}$.
+In this case, none of the alive agents see more than one pebble at their $home$ after they return there in the $p-$th iteration.
+This is because, here the the pebbles left by agents $a_{0},a_{1},a_{2}$ and $a_{3}$ has been collected by agents $a_{3},a_{0},a_{1}$ and $a_{2}$, respectively and when they reach their corresponding $home$, the pebble they carried are the only one there.
+Thus all of them (except $a_{0}$ as it was destroyed during $p-$th iteration in state Fetch) moves to state Wait2 after reaching $home$ at the $p-$th iteration and after the end of the iteration moved to state Forward and starts the $(p+1)-$th iteration.
+Since $a_{0}$ was destroyed in the previous iteration it remains unavailable to collect the pebble from $h_{1}$, left by $a_{1}$ in the $(p+1)-th$ iteration.
+Note that here in this iteration both $a_{1}$ and $a_{2}$ collects pebble from $h_{2}$ and $h_{3}$ ( left by $a_{2}$ and $a_{3}$ respectively) and returns back to $h_{1}$ and $h_{2}$ respectively.
+But $a_{3}$ upon reaching $h_{0}$ does not find any pebble to collect.
+In this case it returns to $h_{3}$ without any pebble.
+Note that here only $a_{1}$ sees more than one pebble upon returning back to $h_{1}$ in the $(p+1)-$th iteration.
+So only $a_{1}$ changes to state Gather1 at round $n+2+|Seg(a_{1})|$.
+Now by similar argument as in Case-I, the rest follows.
+After 3 agents gather at a safe node with at least 3 pebbles they execute the algorithm PerpExplore-Coloc-Pbl which solves the problem of PerpExploration-BBH on a ring $R$ for 3 co-located agents with at least one pebble each agent (refer, Theorem 4.2).
+Thus this proves Theorem 5.2.
+which is stated again below.
+Statement of Theorem 5.2: Algorithm PerpExplore-Scat-Pbl solves PerpExploration-BBH problem of a ring $R$ with 4 synchronous and scattered agents under the pebble model of communication where each agents are equipped with a pebble.
+10.3 Modification of PerpExplore-Scat-Pbl to include the cases where starting positions has multiplicity
 Remark 10.7.
@@
 Observe that, in these remaining cases there exists at least one node, where multiplicity is greater than 1 initially.
-Let there exists a node with multiplicity $3$, in this case, these agents directly start executing PerpExplore-Coloc-Pbl, while execution if they encounter the fourth agent somewhere along $R$, they just ignore this agent (note that the IDs, of all the 3 initially co-located agents are already collected by each other) while executing their current algorithm.
+Let there exists a node with multiplicity $3$, in this case, these agents directly start executing PerpExplore-Coloc-Pbl.
+While execution of the algorithm, if they encounter the fourth agent somewhere along $R$, they just ignore this agent (note that the IDs, of all the 3 initially co-located agents are already collected by each other) while executing their current algorithm.
 On the other hand, if initially there exists a node with multiplicity greater than 1 and less than 3, then in that case, only the lowest ID agent (say $a_{1}$) at the current node, changes its state Forward, whereas the other agent at the current node (say $a_{2}$) changes its state Backup-Wait at some round say $t_{0}$ $(>0)$.
@@
 If such anomaly exists after $t_{0}+2n+1$ round, then that implies that $a_{1}$ is already in state Gather1.
-It is because, $a_{1}$ while it returned back to its $home$ carrying a pebble in state Fetch at round $t_{0}+n+size+2$ round (where $size$ is the length of $Seg(a_{1})$), encounters the first anomaly, i.e., the number of pebble is more than number agents.
-In that case, it directly changes its state to Gather1, and further waits at $home$ for $n-size-1$ rounds, i.e., till $t_{0}+2n+1$ rounds.
+It is because, $a_{1}$ while it returned back to its $home$ carrying a pebble in state Fetch at round $t_{0}+n+size+2$ round (where $size$ is the length of $Seg(a_{1})$), it encounters the first anomaly, i.e., the number of pebble is more than number agents.
+In that case, $a_{1}$ directly changes its state to Gather1, and further waits at $home$ for $n-size-1$ rounds, i.e., till $t_{0}+2n+1$ rounds.
 After which it starts to move clockwise, and on the other hand $a_{2}$ also starts moving clockwise, while changing its state to Gather2 from Backup-Wait.
-This guarantees that if the anomaly is detected at a multiplicity then both $a_{1}$ and $a_{2}$ moves together to gather, along with the third agent which is in state Wait2 (as described in Lemma 5.5) and changes its state to Gather2 whenever it finds #$\#Agent$ at its $home$ is more than one.
+This guarantees that if the anomaly is detected at a multiplicity then both $a_{1}$ and $a_{2}$ moves together to gather with the third agent, which is in state Wait2.
 Now if the anomaly is not detected at the multiplicity, that is $a_{2}$ at round $t_{0}+2n+1$ finds no anomaly then it moves into state Backup whereas $a_{1}$ is in state Wait2 at the same node.
@@
 In that case they find that there are 3 agents at their $home$ and all of them change to state Coloc and executes algorithm PerpExplore-Coloc-Pbl.
-Next we can have the following theorem
-Theorem 5.8.
-A set of 4 synchronous scattered agents with one pebble each is necessary and sufficient to solve the PerpExploration-BBH problem on a ring $R$, when the initial starting locations of the agents are different along $R$.
-5.2 Communication: Whiteboard
+11 Algorithm PerpExplore-Scat-Whitbrd
 In this section, we discuss the algorithm PerpExplore-Scat-WhitBrd, which achieves PerpExploration-BBH, with the help of three agents on a ring $R$ with $n$ nodes, where each node has a whiteboard that can store $O(\log n)$ bits of data.
@@
 So, there can be two cases, first, all the three agents are initially placed at three different positions, second, two agents are together whereas the third agent is in a different position.
-We will provide algorithm for the first case only, as an algorithm for the second case can be easily designed by modifying and merging the algorithms for the first case and algorithm PerpExplore-Coloc-Pbl.
-The idea of our algorithm PerpExplore-Scat-WhitBrd is as follows: the three agents separately explore three segments of the ring $R$, such that, their union is the whole ring $R$ and intersection of any two segment contains only a safe node (which is a $home$ for one agent).
-Thus, while exploring, only one agent can be consumed by the black hole, and in an additional finite number of rounds another agent can learn,the consumption of the earlier agent by the black hole (this agent is precisely the agent whose segment has a safe node common with the segment of the consumed agent.
-Also it learns about this consumption from the whiteboard of the common node).
-In which case, this agent gathers with the other alive agent, and locates the exact segment in which the earlier agent was exploring before its consumption by the black hole.
-After locating this segment, both these agents starts performing Cautious-Walk either in clockwise or counter-clockwise direction, based on the information they gained about the earlier agents movement direction before getting consumed by the black hole.
-The agents before getting consumed, marks each node of their corresponding segment with different markings (i.e., more precisely writes either left or right, after erasing the earlier direction marking at the current node) based on the direction of their movement.
-So, the two agents performing Cautious-Walk, search for the first node, which does not have the direction marking w.r.to the direction the earlier consumed agent was moving.
-This node which is devoid of this direction marking is indeed the black hole node, and at least one agent, among the agents performing Cautious-Walk detects it, and remains alive, which further perpetually explores the ring $R$, avoiding this node.
-Let us consider the three agents (say, $a_{1}$, $a_{2}$ and $a_{3}$) are initially placed at three nodes of the ring $R$, which are not only safe nodes but are also recognised as the $home$ of these agents.
+We will provide algorithm for the first case only, as an algorithm for the second case can be easily designed by modifying and merging the algorithms for the first case and algorithm PerpExplore-Coloc-Pbl (as discussed in Remark 11.15 in Appendix, Section 11.3).
+11.1 Detailed description and Pseudocode of the algorithm
+Let us consider the three agents (say, $a_{1}$, $a_{2}$ and $a_{3}$) are initially placed at three nodes of the ring $R$, which are not only safe nodes but are also recognized as the $home$ of these agents.
 Initially, the agents starting from their respective $home$ are assigned the task to explore a set of nodes, which we term as a segment of the corresponding agents.
@@
 So if none of the agents are ever destroyed the ring will still be perpetually explored.
-Let us first discuss the case, in which we describe the possible movements of the agent and the respective state changes they perform, until one agent gets consumed by the black hole, and another agent gets to know that an agent is already consumed by the black hole.
-After which, we will describe all the possible movements and state changes performed by the remaining two alive agent, between getting to know that already one has been consumed by the black hole, and finally detecting the black hole position.
-Initially all agents start from state Initial at their respective $home$.
-In state Initial, an agent ( without loss of generality, say $a_{1}$) first clears the already present data (if at all) at the whiteboards of their respective $home$, then initializes $T_{time}=0$ ($T_{time}$ stores the number of rounds elapsed since the start of state Initial), and writes a message of the form (home, ID) at its $home$, where ID is the ID of the agent.
-This type of message is termed as “home" type message, which consists of two components, first component stores the message home and second component stores the ID of the agent writing, i.e., ID of the agent whose $home$ is the $Current-Node$.
-Further, it changes its state to Forward, and moves in clockwise direction.
-In state Forward, the agent moves in clockwise direction while erasing the earlier direction marking (if exists), i.e., left and then writes the new direction marking, i.e., right in each node.
+Let us first discuss the case, in which we describe the possible movements of the agent and the respective state changes they perform, until one agent gets destroyed by the Byzantine black hole, and another agent gets to know that an agent is already destroyed by the Byzantine black hole.
+After which, we will describe all the possible movements and state changes performed by the remaining two alive agents, between getting to know that already one has been destroyed by the Byzantine black hole, and finally detecting the Byzantine black hole’s position.
+Initially, all agents start from state Initial at their respective $home$.
+In state Initial, an agent ( without loss of generality, say $a_{1}$) first clear the already present data (if at all) at the whiteboards of their respective $home$, then initializes $T_{time}=0$ ($T_{time}$ stores the number of rounds elapsed since the start of state Initial), and writes a message of the form (home, ID) at its $home$, where ID is the ID of the agent.
+This type of message is termed as “home" type message, which consists of two components, the first component stores the message home and the second component stores the ID of the agent writing, i.e., ID of the agent whose $home$ is the $Current-Node$.
+Further, it changes its state to Forward and moves in the clockwise direction.
+In state Forward, the agent moves in the clockwise direction while erasing the earlier direction marking (if exists), i.e., left and then writes the new direction marking, i.e., right in each node.
 The agent also increases the $T_{time}$ variable by one in each round.
-This process continues until the agent encounters a node which has a “home" type message.
-This “home" type message signifies that the agent has reached the end of segment $Seg(a_{1})$, i.e., in other words it has reached the nearest clockwise $home$, say $v_{c}$.
+This process continues until the agent encounters a node that has a “home" type message.
+This “home" type message signifies that the agent has reached the end of segment $Seg(a_{1})$, i.e., in other words, it has reached the nearest clockwise $home$, say $v_{c}$.
 Note that the length of a segment can be at most $n-2$, hence within $T_{time}=n-1$, an agent $a_{1}$ is bound to reach the last node of its own segment i.e., $v_{c}$.
 In any case irrespective of the current $T_{time}$, the agent waits at $v_{c}$ until $T_{time}=n-1$, after which in the next round, it checks for the following information at $v_{c}$.
-If the agent finds a message of type “visited" at $v_{c}$, the agent considers this as an anomaly and learns that an agent of which $v_{c}$ is the home (from the ID component of home type message at $v_{c}$) must have entered black hole while returning back from its clockwise nearest $home$ (refer Lemma 5.9).
+If the agent finds a message of type “visited" at $v_{c}$, the agent considers this as an anomaly and learns that an agent of which $v_{c}$ is the home (from the ID component of home type message at $v_{c}$) must have entered Byzantine black hole while returning back from its clockwise nearest $home$ (refer Lemma 11.1).
 Then in this case, the agent stores the message of type dir in its local memory, where dir=(Counter-clockwise, NULL).
-In a dir type message, the first component is called a direction component which indicates the direction of the agent which gets consumed by the black hole, along which it was moving just before it gets consumed.
-On the other hand, the second component either stores ID of some agent or it stores $NULL$ message.
+In a dir type message, the first component is called a direction component which indicates the direction of the agent that gets destroyed by the Byzantine black hole, along which it was moving just before it gets destroyed.
+On the other hand, the second component either stores ID of some agent or stores $NULL$ message.
 Both these components are useful for certain state transitions.
@@
 Next, in state Back-Wait, an agent (here $a_{1}$) waits at $v_{c}$ until $T_{time}=2n-1$.
-In this waiting time $a_{1}$ waits for other agent to meet $a_{1}$ if the other agent detects any anomaly during its state Forward.
+In this waiting time, $a_{1}$ waits for the other agent to meet $a_{1}$ if the other agent detects any anomaly during its state Forward.
 While waiting, if it finds that the #$\#Agent$ at $v_{c}$ is more than 1, and the whiteboard at $v_{c}$ has a message of type dir, then the agent performs the following task.
-After noticing this, the agent moves along the counter clockwise direction, while changing its state to Gather2.
-Note that, our algorithm ensures that in state Back-Wait, if a dir type message is seen by an agent, then the direction component of this message must be in a counter-clockwise direction (refer Lemma 5.15).
+After noticing this, the agent moves along the counter-clockwise direction, while changing its state to Gather2.
+Note that, our algorithm ensures that while in state Back-Wait, if a dir type message is seen by an agent, then the direction component of this message must be in a counter-clockwise direction (refer Lemma 11.7).
 Otherwise, if no such dir type message is seen by $a_{1}$, and also $T_{time}=2n$, then in this round, the agent changes its state to Backtrack.
 In state Backtrack, an agent (here $a_{1}$) starts to move in a counter-clockwise direction from $v_{c}$, and after each move erases the earlier direction, i.e., right, and writes the new direction, i.e., left, and also increments $T_{time}$ by 1.
-This process continues until it reaches its own $home$, i.e., reads a home type message with ID same to its own ID.
+This process continues until it reaches its own $home$, i.e., reads a home type message with ID same as its own ID.
 Again note that, since a segment can be of length at most $n-2$, hence within $T_{time}=3n-1$, the agent reaches its own $home$.
@@
 At $T_{time}=3n$, it checks whether its $home$ has visited type message, if so then directly changes its state to Initial-Wait.
-Otherwise, absence of such visited type message, creates an anomaly for the agent.
-This only happens when another agent, say $a_{j}$, ($j\neq 1$), for which $Seg(a_{1})\cap Seg(a_{j})=$ $home$ of $a_{1}$, does not arrive at $home$ of $a_{1}$ during its Forward state because of getting consumed at the black hole (refer Lemma 5.11).
-This instigates the agent $a_{1}$ to change its state to Gather and store the dir type message (clockwise,NULL) in its local memory..
+Otherwise, the absence of such visited type message, creates an anomaly for the agent.
+This only happens when another agent, say $a_{j}$, ($j\neq 1$), for which $Seg(a_{1})\cap Seg(a_{j})=$ $home$ of $a_{1}$, does not arrive at $home$ of $a_{1}$ during its Forward state because of getting destroyed at the Byzantine black hole (refer Lemma 11.3).
+This instigates the agent $a_{1}$ to change its state to Gather and store the dir type message (clockwise, NULL) in its local memory.
 In state Initial-Wait, the agent waits at its $home$ until $T_{time}=4n-1$.
-This waiting period is enough for an anomaly finding agent in state Backtrack to meet with it.
-If the agent finds that there are more than one agent at its $home$, and also there is a dir type message as well, then our algorithm ensures that the direction component of this dir type message must be in clockwise direction (refer Lemma 5.17).
-In which case, the agent starts moving clockwise and changes its state to Gather1.
+This waiting period is enough for an anomaly finding agent in state Backtrack, to meet with it.
+If the agent finds that there is more than one agent at its $home$, and also there is a dir type message as well, then our algorithm ensures that the direction component of this dir type message must be in clockwise direction (refer Lemma 11.9).
+In this case, the agent starts moving clockwise and changes its state to Gather1.
 On the contrary, if $T_{time}=4n$ and no anomaly is detected, then the agent again moves back to state Initial.
-An agent can change its state to Gather, either from state Forward, or from state Backtrack.
-Note that in either case, it carries the stored message of type dir (which has a direction component and an ID component which is either $NULL$ or stores the ID of an agent).
+An agent can change its state to Gather, either from state Forward or from state Backtrack.
+Note that in either case, it carries the stored message of type dir (which has a direction component and an ID component that is either $NULL$ or stores the ID of an agent).
 So, in this state if the direction component of dir is along clockwise (resp.
@@
 counter-clockwise) direction until it finds a node with #$\#Agent$ more than one.
-Next, in case of counter-clockwise direction component, the agent updates the dir message to (Counter-clockwise, ID’) where ID’ is the ID of the other agent at the same node.
+Next, in case of a counter-clockwise direction component, the agent updates the dir message to (Counter-clockwise, ID’) where ID’ is the ID of the other agent at the same node.
 Then the agent writes the updated dir message at the current node and changes state to Gather2.
-If the agent in state Gather met the other agent when the dir type message having clockwise direction component the, the agent simply just writes the message at the current node and moves to state Gather1.
-Also note that exactly one agent can move into state Gather.
-This is because if an agent is consumed during Forward state then the anomaly is first found by a single agent whose $home$ is the clockwise nearest $home$ of the consumed agent.
+If the agent in the state Gather met the other agent when the dir type message has a clockwise direction component, the agent simply just writes the message at the current node and moves to state Gather1.
+Also, note that exactly one agent can move into state Gather.
+This is because if an agent is destroyed during Forward state then the anomaly is first found by a single agent whose $home$ is the clockwise nearest $home$ of the destroyed agent.
 This agent changes its state to Gather.
 Now, before the other agent finds further anomaly the agent in state Gather meets the other agent and forces it to change into state Gather1.
-Similarly If an agent $a_{1}$ is consumed by the black hole during its Backtrack state, then the first anomaly is detected by a single agent (more precisely the agent for which the clockwise nearest $home$ is the $home$ of $a_{1}$).
-In this case it changes its state to Gather and forces the other agent to move into state Gather2 by meeting it before it finds further anomaly.
+Similarly, If an agent $a_{1}$ is destroyed by the Byzantine black hole during its Backtrack state, then the first anomaly is detected by a single agent (more precisely the agent for which the clockwise nearest $home$ is the $home$ of $a_{1}$).
+In this case, it changes its state to Gather and forces the other agent to move into state Gather2 by meeting it before it finds further anomaly.
 Thus we have the following observation.
 {observation} There is exactly one agent that changes its state to Gather throughout the execution of Algorithm PerpExplore-Scat-Whitbrd.
-Also, The agent which changes its state to Gather, is the first agent to understand that, an agent has already been consumed by the black hole.
+Also, The agent which changes its state to Gather, is the first agent to understand that, an agent has already been destroyed by the Byzantine black hole.
 An agent can change its state to Gather1, only from the states Initial-Wait or Gather.
@@
 on the other hand, an agent (here $a_{1}$) can only change its state to Gather1 from Initial-Wait, if it is waiting at its $home$, and within which it encounters another agent at its $home$ along with it a message of type dir is written (in which the direction component of dir is along clockwise).
-Note that if $a_{1}$ in state Initial-Wait finds more than one agent at current node and dir type message at round $t$ then, the other agent must be in state Gather while written there at round $t-1$ and moved into state Gather1 in the same round.
+Note that if $a_{1}$ in state Initial-Wait finds more than one agent at the current node and dir type message at round $t$ then, the other agent must be in state Gather while written there at round $t-1$ and moved into state Gather1 in the same round.
 Then, during round $t$, $a_{1}$ first stores the dir type message before it moves to state Gather1 and moves according to the direction component of the dir type message (i.e., clockwise).
 Also, the other agent that changed its state to Gather1 from Gather in round $t-1$ also moves according to the dir type message during round $t$.
-Thus we have another observation {observation} If two agents move to state Gather1 , one moves from state Gather and the other moves from state Initial-Wait.
-Furthermore, they change to state Gather1 at the same node and leaves the node together at the same round and stays together.
+Thus we have another observation {observation} If two agents move to state Gather1, one moves from state Gather, and the other moves from state Initial-Wait.
+Furthermore, they change to state Gather1 at the same node and leave the node together at the same round and stay together.
 In state Gather1, the agent which changes its state to Gather1 from Initial-Wait first stores the dir type message, as the other agent must already have stored this message in state Gather.
-Irrespective of which, the agent starts moving in a clockwise direction, until it finds a node which has a home (i.e., of the form (home, ID)) type message, where the ID component this message does not match with the IDs of the agents present at the current node.
-Note that our algorithm ensures that the current node is the $home$ of $a_{j}$, if $a_{j}$ is the agent to be consumed by the black hole (refer Lemma 5.19).
-Further, as there are two agents at the current node both in state Gather1 (Observation 5.2), so now the agent with the lowest ID, sets $Move=0$, also initializes the $Marking$ variable to $right$, since the direction component of dir is clockwise, and further changes its state to Cautious-Leader.
-On the other hand the other alive agent at the current node changes its state to Cautious-Follower, while initializing $Move=0$.
-An agent can change its state to Gather2, only from the states Back-Wait or Gather.
-An agent changes its state to Gather2 from Gather, only when while moving along counter-clockwise direction finds another agent present, in which case it updates the dir message and writes the corresponding message of type dir and correspondingly changes its state to Gather2.
-On the other hand, an agent can only change its state to Gather2 from Back-Wait, if it is waiting at its corresponding segments nearest clockwise $home$, and within which it encounters another agent at its current node along with it a message of type dir is written (in which the direction component of dir is along counter-clockwise).
-After storing the dir type message in its local memory and changing its state to Gather2 from Back-Wait, an agent immediately moves counter-clockwise in the same round.
-This generates another observation similar to the Observation 5.2 as follows.
-{observation}
-If two agents move to state Gather2 , one moves from state Gather and the other moves from state Back-Wait.
-Furthermore, they change to state Gather2 at the same node and leaves the node together at the same round and stays together.
-In state Gather2, the agent moves in a counter-clockwise direction, until it finds a node which has a home (i.e., of the form (home, ID’)) type message, where ID’ matches with ID component of dir type message in its local memory.
-Note that our algorithm ensures that the current node is the nearest clockwise $home$ of $a_{1}$, if $a_{1}$ is the agent to be consumed by the black hole (refer Lemma 5.19).
-Further, as there are two agents at the current node, both in state Gather2 (refer Observation 5.2), so now the agent with the lowest ID, sets $Move=0$, also initializes the $Marking$ variable to $left$, since the direction component of dir is counter-clockwise, and further changes its state to Cautious-Leader.
-On the other hand the other alive agent at the current node, changes its state to Cautious-Follower, while initializing $Move=0$.
-In state Cautious-Leader, an agent updates $Move=1$ and moves along the direction component of dir.
-After moving when it is alone on a node, it checks if the current node is marked with $Marking$ (i.e., either $left$ or $right$).
-If the node is marked, then it moves in the opposite direction (i.e., the opposite of direction component of dir message) to meet with the agent in state Cautious-Follower.
-Otherwise, if there is no $Marking$ at the current node then it identifies the current node as the black hole node (only if it remains alive) and continues perpetual exploration avoiding this node.
-If an agent in state Cautious-Leader with $Move=1$ meets with another agent (i.e., the agent in state Cautious-Follower) it updates variable $Move$ to 0, and again moves according to the direction component.
-In state Cautious-Follower, the agent after finding $Move=0$, initializes the variable $Wait=0$ and updates $Move=1$.
-If $Move=1$, then the agent checks whether $Wait<1$, if so then update $Wait=Wait+1$.
-When $Wait=1$, it checks if the current node has more than one agent (i.e., whether the agent in state Cautious-Leader has returned or not), if so then further update $Move=0$ and move along the direction component of dir, along with the agent in state Cautious-Leader.
-Otherwise, if the agent in state Cautious-Leader does not return, that means the number of agent at the current node is 1 when $Wait=1$, this symbolizes that the agent in state Cautious-Leader has been either consumed by the black hole or the agent in state Cautious-Leader finds that its current node (i.e., the next node along the direction component of dir for the agent in state Cautious-Follower) is the black hole and continues to perpetually move in the same direction, further avoiding this node.
-In any case, the agent executing Cautious-Leader has not returned implies the next node along direction component of dir for the agent executing Cautious-Follower is the black hole node.
-So, in this case, the agent declares the next node along the direction component of dir is the black hole node and continues perpetual exploration avoiding that node.
+Irrespective of which, the agent starts moving in a clockwise direction until it finds a node that has a home (i.e., of the form (home, ID)) type message, where the ID component this message does not match with the IDs of the agents present at the current node.
+Note that our algorithm ensures that the current node is the $home$ of $a_{j}$ if $a_{j}$ is the agent to be destroyed by the Byzantine black hole (refer Lemma 11.11).
+Further, as there are two agents at the current node both in state Gather1 (Observation 11.1), so now the agent with the lowest ID, sets $Move=0$, also initializes the $Marking$ variable to $right$, since the direction component of dir is clockwise, and further changes its state to Cautious-Leader.
+On the other hand, the other alive agent at the current node changes its state to Cautious-Follower, while initializing $Move=0$.
 // Algorithm is written for an agent $r$
@@
 106            else
-107                   Declare $Current-Node$ as the black hole node and continue perpetual exploration avoiding this node.
+107                   Declare $Current-Node$ as the Byzantine black hole node and continue perpetual exploration avoiding this node.
 108
@@
 124      else
-125             Declare the next node along the direction component of dir as the black hole node and continue perpetual exploration avoiding this node.
+125             Declare the next node along the direction component of dir as the Byzantine black hole node and continue perpetual exploration avoiding this node.
 126
@@
 128
-5.2.1 Correctness and Complexity
+An agent can change its state to Gather2, only from the states Back-Wait or Gather.
+An agent changes its state to Gather2 from Gather, only when while moving along counter-clockwise direction finds another agent present, in which case it updates the dir message and writes the corresponding message of type dir and correspondingly changes its state to Gather2.
+On the other hand, an agent can only change its state to Gather2 from Back-Wait, if it is waiting at its corresponding segments nearest clockwise $home$, and within which it encounters another agent at its current node along with it a message of type dir is written (in which the direction component of dir is along counter-clockwise).
+After storing the dir type message in its local memory and changing its state to Gather2 from Back-Wait, an agent immediately moves counter-clockwise in the same round.
+This generates another observation similar to the Observation 11.1 as follows.
+{observation}
+If two agents move to state Gather2, one moves from state Gather, and the other moves from state Back-Wait.
+Furthermore, they change to state Gather2 at the same node and leave the node together at the same round and stay together.
+In state Gather2, the agent moves in a counter-clockwise direction, until it finds a node which has a home (i.e., of the form (home, ID’)) type message, where ID’ matches with ID component of dir type message in its local memory.
+Note that our algorithm ensures that the current node is the nearest clockwise $home$ of $a_{1}$ if $a_{1}$ is the agent to be destroyed by the Byzantine black hole (refer Lemma 11.11).
+Further, as there are two agents at the current node, both in state Gather2 (refer Observation 11.1), so now the agent with the lowest ID, sets $Move=0$, also initializes the $Marking$ variable to $left$, since the direction component of dir is counter-clockwise, and further changes its state to Cautious-Leader.
+On the other hand, the other alive agent at the current node, changes its state to Cautious-Follower, while initializing $Move=0$.
+In state Cautious-Leader, an agent updates $Move=1$ and moves along the direction component of dir.
+After moving when it is alone on a node, it checks if the current node is marked with $Marking$ (i.e., either $left$ or $right$).
+If the node is marked, then it moves in the opposite direction (i.e., the opposite of direction component of dir message) to meet with the agent in state Cautious-Follower.
+Otherwise, if there is no $Marking$ at the current node then it identifies the current node as the Byzantine black hole node (only if it remains alive) and continues perpetual exploration avoiding this node.
+If an agent in the state Cautious-Leader with $Move=1$ meets with another agent (i.e., the agent in the state Cautious-Follower) it updates variable $Move$ to 0 and again moves according to the direction component.
+In state Cautious-Follower, the agent after finding $Move=0$, initializes the variable $Wait=0$ and updates $Move=1$.
+If $Move=1$, then the agent checks whether $Wait<1$, if so then update $Wait=Wait+1$.
+When $Wait=1$, it checks if the current node has more than one agent (i.e., whether the agent in state Cautious-Leader has returned or not), if so then further update $Move=0$ and moves along the direction component of dir, along with the agent in state Cautious-Leader.
+Otherwise, if the agent in state Cautious-Leader does not return, that means the number of agents at the current node is 1 when $Wait=1$, this symbolizes that the agent in state Cautious-Leader has been either destroyed by the Byzantine black hole or the agent in state Cautious-Leader finds that its current node (i.e., the next node along the direction component of dir for the agent in state Cautious-Follower) is the Byzantine black hole and continues to perpetually move in the same direction, further avoiding this node.
+In any case, the agent executing Cautious-Leader has not returned implies the next node along the direction component of dir for the agent executing Cautious-Follower is the Byzantine black hole node.
+So, in this case, the agent declares the next node along the direction component of dir is the Byzantine black hole node and continues perpetual exploration avoiding that node.
+11.2 Correctness and Complexity
 In the next two lemmas (Lemma 11.1 and Lemma 11.3) we first ensure that when we say an agent encounters an anomaly, is actually an anomaly.
 And what an agents interpret from these anomalies are true.
-Next in Lemma 5.13 we ensure that the agent that changes to state Gather can always identify the state of the consumed agent at the time it gets destroyed at the black hole.
+Next in Lemma 11.5 we ensure that the agent that changes to state Gather can always identify the state of the destroyed agent at the time it gets destroyed at the Byzantine black hole.
 Lemma 11.1.
 Let $a_{i}$ and $a_{j}$ be two agents exploring the segments $Seg(a_{i})$ and $Seg(a_{j})$, such that $Seg(a_{i})\cap Seg(a_{j})=v_{c}$, where $v_{c}$ is also the $home$ of $a_{j}$.
-If during execution of PerpExplore-Scat-Whitbrd, there exists a round $t>0$, in which $a_{i}$ in state Forward finds a visited type message, then there exists a round $0<t^{\prime}<t$, in which $a_{j}$ has been consumed by the black hole, while exploring the segment $Seg(a_{j})$ in state Backtrack.
-Proof 5.10.
-Let us suppose, there does not exist any round $t^{\prime}<t$ in which $a_{j}$ has been consumed by the black hole in state Backtrack.
-This means that, either $a_{j}$ is yet to be consumed by the black hole at round $t^{\prime}$, for all $t^{\prime}<t$ or $a_{j}$ has been consumed by the black hole at round $t^{\prime}$, while it is in state Forward.
-Case I: Let $a_{j}$ is yet to be consumed by the black hole at round $t^{\prime}$, for all $t^{\prime}<t$.
+If during the execution of PerpExplore-Scat-Whitbrd, there exists a round $t>0$, in which $a_{i}$ in state Forward finds a visited type message, then there exists a round $0<t^{\prime}<t$, in which $a_{j}$ has been destroyed by the Byzantine black hole while exploring the segment $Seg(a_{j})$ in state Backtrack.
+Proof 11.2.
+Let us suppose, there does not exist any round $t^{\prime}<t$ in which $a_{j}$ has been destroyed by the Byzantine black hole in state Backtrack.
+This means that either $a_{j}$ is yet to be destroyed by the Byzantine black hole at round $t^{\prime}$, for all $t^{\prime}<t$ or $a_{j}$ has been destroyed by the Byzantine black hole at round $t^{\prime}$, while it is in state Forward.
+Case I: Let $a_{j}$ is yet to be destroyed by the Byzantine black hole at round $t^{\prime}$, for all $t^{\prime}<t$.
 Note that since $a_{i}$ is at $v_{c}$ and checks for the visited type message, this implies $T_{time}=n$.
-This means, at $(t-n-1)$-th round $a_{i}$ changed its state to Forward from Initial.
+This means at $(t-n-1)$-th, round $a_{i}$ changed its state to Forward from Initial.
 This also implies $a_{j}$ was also in state Initial at $(t-n-1)-$th round at $v_{c}$.
-Hence, at this round $a_{j}$ must have cleared any whiteboard data at $v_{c}$ (as $v_{c}$ is the $home$ of $a_{j}$) and also has moved clockwise after changing the state to Forward.
+Hence, at this round, $a_{j}$ must have cleared any whiteboard data at $v_{c}$ (as $v_{c}$ is the $home$ of $a_{j}$) and also has moved clockwise after changing the state to Forward.
 Observe that, the node $v_{c}$ can only be explored by $a_{i}$ and $a_{j}$.
-So, these two agents has the possibility to write the visited type message at $v_{c}$ after round $t-n-1$ and before round $t$.
+So, these two agents have the possibility to write the visited type message at $v_{c}$ after round $t-n-1$ and before round $t$.
 Further, note that in the next $n+1$ rounds after $(t-n-1)-$th round (i.e., until round $t$), $a_{j}$ cannot return back to $v_{c}$.
 So, $a_{j}$ can not write anything on $v_{c}$ after round $t-n-1$ and before round $t$.
-Also, $a_{i}$ can only write the visited type message at $v_{c}$ at the $t-$th round (i.e., when $T_{time}=n$ according to Algorithm 4).
-But, in this case, $a_{i}$ at $t-$th round before writing a message, finds that there already exists a visited type message, and this a contradiction, to the fact that $a_{j}$ has not been consumed by the black hole at round $t^{\prime}$, for all $t^{\prime}<t$.
-Case II: Let $a_{j}$ has been consumed by the black hole at some round $t^{\prime}<t$, while in state Forward.
-This implies there exists a round $0<t^{\prime\prime}<t^{\prime}$ in which $a_{j}$ was in state Initial (this is the last time $a_{j}$ was in state Initial before it gets consumed by the black hole).
-Thus this means, $t^{\prime}\leq t^{\prime\prime}+n-2$ as a segment can be of length at most $n-2$.
+Also, $a_{i}$ can only write the visited type message at $v_{c}$ at the $t-$th round (i.e., when $T_{time}=n$).
+But, in this case, $a_{i}$ at $t-$th round before writing a message, finds that there already exists a visited type message, and this a contradiction, to the fact that $a_{j}$ has not been destroyed by the Byzantine black hole at round $t^{\prime}$, for all $t^{\prime}<t$.
+Case II: Let $a_{j}$ has been destroyed by the Byzantine black hole at some round $t^{\prime}<t$, while in state Forward.
+This implies there exists a round $0<t^{\prime\prime}<t^{\prime}$ in which $a_{j}$ was in state Initial (this is the last time $a_{j}$ was in state Initial before it gets destroyed by the Byzantine black hole).
+Thus this means, that $t^{\prime}\leq t^{\prime\prime}+n-2$, as a segment can be of length at most $n-2$.
 Then at round $t^{\prime\prime}$, $a_{i}$ was also in state Initial at its own $home$.
 Now, note that after $t^{\prime\prime}$, $a_{i}$ can visit $v_{c}$ within round $t^{\prime\prime}+1$ and $t^{\prime\prime}+n+1$ while in state Forward.
-Note that from round $t^{\prime\prime}+1$ and before round $t^{\prime\prime}+n+1$ no agent can write a visited type message at $v_{c}$ (by similar argument as in Case I).
+Note that from round $t^{\prime\prime}+1$ and before round $t^{\prime\prime}+n+1$ no agent can write a visited type message at $v_{c}$ (by a similar argument as in Case I).
 So during round $t^{\prime\prime}+n+1$, $a_{i}$ can not see any visited type message at $v_{c}$.
@@
 Now, let us consider the agent $a_{k}$ where $Seg(a_{k})\cap Seg(a_{j})=home$ of $a_{k}$ and $Seg(a_{j})\cap Seg(a_{i})=home$ of $a_{i}$.
-Note that in round $t^{\prime\prime}+n+1$ both $a_{k}$ and $a_{i}$ moves into state Back-Wait (as none of them sees any visited type message).
-Next after waiting there upto round $t^{\prime\prime}+2n$, both of them changes their state to Backtrack at round $t^{\prime\prime}+2n+1$.
-Next they reach their corresponding $home$ and checks for a visited type message at round $t^{\prime\prime}+3n+1$.
-Note that $a_{i}$ finds the visited type message left by $a_{k}$ and changes to state Initial-Wait whereas, $a_{k}$ does not find any visited type message left by $a_{j}$ (as $a_{j}$ is consumed at the black hole at state Forward at round $t^{\prime}\leq t^{\prime\prime}+n-2<t^{\prime\prime}+3n+1$).
+Note that in round $t^{\prime\prime}+n+1$ both $a_{k}$ and $a_{i}$ move into state Back-Wait (as none of them sees any visited type message).
+Next after waiting there up to round $t^{\prime\prime}+2n$, both of them change their state to Backtrack at round $t^{\prime\prime}+2n+1$.
+Next, they reach their corresponding $home$ and check for a visited type message at round $t^{\prime\prime}+3n+1$.
+Note that $a_{i}$ finds the visited type message left by $a_{k}$ and changes to state Initial-Wait whereas, $a_{k}$ does not find any visited type message left by $a_{j}$ (as $a_{j}$ is destroyed at the Byzantine black hole at state Forward at round $t^{\prime}\leq t^{\prime\prime}+n-2<t^{\prime\prime}+3n+1$).
 So, $a_{k}$ changes to state Gather1 at round $t^{\prime\prime}+3n+1$ and moves clockwise with the message dir$=$ (Clockwise, NULL) until it finds $a_{i}$ at it’s $home$ in state Initial-Wait before round $t^{\prime\prime}+4n$ (as length of a segment can be of length at most $n-2$).
@@
 So we again arrive at a contradiction.
-This implies $a_{j}$ can not be in state Forward while it is consumed at the black hole at round $t^{\prime}<t$.
+This implies $a_{j}$ can not be in state Forward while it is destroyed at the Byzantine black hole at round $t^{\prime}<t$.
 Lemma 11.3.
 Let $a_{i}$ and $a_{j}$ be two agents such that $v_{c}=Seg(a_{i})\cap Seg(a_{j})=home$ of $a_{i}$.
-If there exists a round $t>0$ such that $a_{i}$ checks and finds no visited type message on $v_{c}$ while in state Backtrack, then there must be a round $t^{\prime}<t$ in which $a_{j}$ was consumed by the black hole while in state Forward.
-Proof 5.12.
-Let us consider that there does not exists any round $t^{\prime}<t$ where $a_{j}$ was consumed by the black hole in state Forward.
-This implies that either $a_{j}$ is not consumed by the black hole for all rounds $t^{\prime}<t$ or it is consumed by the black hole at round $t^{\prime}<t$ while it was in state Backtrack.
-Case-I: Let us consider that $a_{j}$ is yet to be consumed by the black hole at round $t^{\prime},\leavevmode\nobreak\ \forall t>t^{\prime}>0$.
+If there exists a round $t>0$ such that $a_{i}$ checks and finds no visited type message on $v_{c}$ while in state Backtrack, then there must be a round $t^{\prime}<t$ in which $a_{j}$ was destroyed by the Byzantine black hole while in state Forward.
+Proof 11.4.
+Let us consider that there does not exist any round $t^{\prime}<t$ where $a_{j}$ was destroyed by the Byzantine black hole in the state Forward.
+This implies that either $a_{j}$ is not destroyed by the Byzantine black hole for all rounds $t^{\prime}<t$ or it is destroyed by the Byzantine black hole at round $t^{\prime}<t$ while it was in state Backtrack.
+Case-I: Let us consider that $a_{j}$ is yet to be destroyed by the Byzantine black hole at round $t^{\prime},\leavevmode\nobreak\ \forall t>t^{\prime}>0$.
 Note that, at round $t$, $a_{i}$ checks for a visited type message while in state Backtrack, and that means the current $T_{time}$ for any alive agent must be equal $3n$ (at $T_{time}=n$ every alive agent ends its state Forward, then until $T_{time}=2n$ every such agent is in state Back-Wait, and at $T_{time}=3n$ every alive agent checks for the message in state Backtrack).
-Note that, this means at time $t-3n-1$, both $a_{i}$ and $a_{j}$ was in state Initial at their respective $home$.
+Note that, this means at time $t-3n-1$, both $a_{i}$ and $a_{j}$ were in state Initial at their respective $home$.
 Then at round $t-2n-1$, both of them must be in state Forward with $T_{time}=2n$, while $a_{j}$ is at $v_{c}$ and $a_{i}$ is at $home$ of $a_{k}$ ( $a_{k}$ is the third agent exploring $Seg(a_{k})$, such that $Seg(a_{i})\cap Seg(a_{k})=home$ of $a_{k}$ and $Seg(a_{k})\cap Seg(a_{j})=home$ of $a_{j}$).
 So, at this point, both $a_{i}$ and $a_{j}$ must have written a visited type message at their respective current nodes (i.e., $v_{c}$ and $home$ of $a_{k}$).
-Further observe that, within $t-2n$ round to $t-n-1$ round $a_{j}$ remains at $v_{c}$ while in state Back-Wait, and according to our algorithm no agent in state Back-Wait can alter the already stored information at their current node, i.e., within these rounds $a_{j}$ cannot erase the visited type message at $v_{c}$.
-Next, within $t-n$ to $t-1$ round, the only agent which can visit the node $v_{c}$ is $a_{i}$, while it is in state Backtrack.
+Further observe that, within $t-2n$ round to $t-n-1$ round $a_{j}$ remains at $v_{c}$ while in state Back-Wait, and according to our algorithm no agent in state Back-Wait can alter the already stored information at their current node, i.e., within these rounds, $a_{j}$ cannot erase the visited type message at $v_{c}$.
+Next, within $t-n$ to $t-1$ round, the only agent that can visit the node $v_{c}$ is $a_{i}$, while it is in state Backtrack.
 Note that, $a_{i}$ cannot alter any visited type message during state Backtrack.
@@
 This is a contradiction to the assumption that $a_{i}$ does not find any visited type message at round $t$.
-Thus, $a_{j}$ must have been consumed by the black hole at time $t^{\prime}$ for some $t^{\prime}<t$.
-Case-II: Suppose $a_{j}$ has been consumed by the black hole at some round $t^{\prime}<t$, while in state Backtrack.
-Let $t^{\prime\prime}$ be the round when $a_{j}$ was in state Initial for the last time where $0<t^{\prime\prime}<t^{\prime}<t$.
+Thus, $a_{j}$ must have been destroyed by the Byzantine black hole at time $t^{\prime}$ for some $t^{\prime}<t$.
+Case-II: Suppose $a_{j}$ has been destroyed by the Byzantine black hole at some round $t^{\prime}<t$, while in state Backtrack.
+Let $t^{\prime\prime}$ be the round when $a_{j}$ was in state Initial for the last time where $0<t^{\prime\prime}<t^{\prime}$.
 Note that $t^{\prime}>t^{\prime\prime}+2n+1$ (because an agent changes its state to Backtrack at round $t^{\prime\prime}+2n+1$).
-So at round $t^{\prime\prime}+n+1$ it was alive and at $v_{c}$ in state Forward.
-During this round it also writes a visited type message at $v_{c}$.
+So at the round, $t^{\prime\prime}+n+1$ it was alive and at $v_{c}$ in state Forward.
+During this round, it also writes a visited type message at $v_{c}$.
 Now as we discussed earlier in Case-I, from round $t^{\prime\prime}+n+1$ till round $t-1$ no agent can erase or alter the visited type message at $v_{c}$.
@@
 This is a contradiction to the fact that at round $t$, $a_{i}$ checks and finds no visited type message at $v_{c}$ while it is in state Backtrack.
-Thus if $a_{j}$ is consumed by the black hole at round $t^{\prime}$, then it must be in state Forward.
+Thus if $a_{j}$ is destroyed by the Byzantine black hole at round $t^{\prime}$, then it must be in state Forward.
 Lemma 11.5.
-The agent which changes its state to Gather, correctly identifies the state in which an agent was in, just before it gets consumed by the black hole.
-Proof 5.14.
-Let $a_{i},a_{j}$ and $a_{k}$ be three agents initially at three different nodes (i.e., three different $home$ for three agents) of $R$.
+The agent which changes its state to Gather, correctly identifies the state in which an agent was in, just before it gets destroyed by the Byzantine black hole.
+Proof 11.6.
+Let $a_{i},a_{j}$, and $a_{k}$ be three agents initially at three different nodes (i.e., three different $home$ for three agents) of $R$.
 Let $Seg(a_{i})\cap Seg(a_{j})=home$ of $a_{i}$, $Seg(a_{i})\cap Seg(a_{k})=home$ of $a_{k}$ and $Seg(a_{k})\cap Seg(a_{j})=home$ of $a_{j}$.
-Without loss of generality let $a_{j}$ be the agent that is consumed by the black hole at a round $t>0$.
-Now we have two cases according th the state of $a_{j}$ at the time of getting consumed.
-Case-I: $a_{j}$ gets consumed by the black hole while in state Forward at round $t$.
+Without loss of generality let $a_{j}$ be the agent that is destroyed by the Byzantine black hole at a round $t>0$.
+Now we have two cases according to the state of $a_{j}$ at the time of getting destroyed.
+Case-I: $a_{j}$ gets destroyed by the Byzantine black hole while in state Forward at round $t$.
 Let $t^{\prime}<t$ be the round when $a_{j}$ was in state Initial for the last time.
@@
 We claim that $a_{i}$ is the agent to change its state to Gather.
-If possible let $a_{k}$ changes its state to Gather.
+If possible let $a_{k}$ change its state to Gather.
 Then first we show that it must change its state to Gather before round $t^{\prime}+3n+1$.
 Otherwise, since $a_{j}$, fails to reach $home$ of $a_{i}$ at round $t^{\prime}+n+1$, it does not write any visited type message there but $a_{i}$ reaches $home$ of $a_{k}$ at round $t^{\prime}+n+1$ and writes a visited type message at $home$ of $a_{k}$.
-Also, since, $a_{i}$ and $a_{j}$ are the only two agent which visits $home$ of $a_{i}$., this implies, no message can be written at $home$ of $a_{i}$ on and between rounds $t^{\prime}$ and $t^{\prime}+3n$.
+Also, since, $a_{i}$ and $a_{j}$ are the only two agents which visit $home$ of $a_{i}$., this implies, no message can be written at $home$ of $a_{i}$ on and between rounds $t^{\prime}$ and $t^{\prime}+3n$.
 So, when $a_{i}$ returns $home$ and finds no visited type message round $t^{\prime}+3n+1$ while in state Backtrack, it changes its state to Gather but $a_{k}$ doesn’t do so as it sees the visited type message at its $home$, left by $a_{i}$.
@@
 Thus $a_{i}$ must be the agent to change its state to Gather.
-Now to prove that $a_{i}$ stores the dir type message with direction component clockwise (which means $a_{j}$ was in state Forward while it was consumed), we have to show that $a_{i}$ changes its state to Gather from Backtrack.
-If possible let $a_{i}$ changes its state to Gather from state Forward.
+Now to prove that $a_{i}$ stores the dir type message with direction component clockwise (which means $a_{j}$ was in state Forward while it was destroyed), we have to show that $a_{i}$ changes its state to Gather from Backtrack.
+If possible let $a_{i}$ change its state to Gather from state Forward.
 Then it must be at round $t^{\prime}+n+1$ when $a_{i}$ checks and finds a visited type message at the $home$ of $a_{k}$.
 Now at round $t^{\prime}$, $a_{k}$ which was in state Initial, cleared all previous data on its $home$.
-This implies the visited type message that $a_{i}$ finds at round $t^{\prime}+n+1$ must be written there after round $t^{\prime}$.
+This implies the visited type message that $a_{i}$ finds at round $t^{\prime}+n+1$ must be written there, after round $t^{\prime}$.
 But as only $a_{i}$ can be there after $t^{\prime}$ and before $t^{\prime}+n+1$, and since it can not alter any data on $home$ of $a_{k}$ before round $t^{\prime}+n+1$, it finds no visited type message at $home$ of $a_{k}$ at round $t^{\prime}+n+1$.
@@
 Hence, it must change its state to Gather from state Backtrack at round $t^{\prime}+3n+1$ and according to the algorithm PerpExplore-Scat-WhitBrd, the dir type message that $a_{i}$ stores must have direction component clockwise.
-Case-II: Let $a_{j}$ gets consumed by the black hole at round $t$ while in state Backtrack.
-We have to show that the agent that changes the state to Gather must store the dir type message with direction component counter-clockwise (as any agent can only move in counter clockwise direction, in state Backtrack).
+Case-II: Let $a_{j}$ gets destroyed by the Byzantine black hole at round $t$ while in state Backtrack.
+We have to show that the agent that changes the state to Gather must store the dir type message with direction component counter-clockwise (as any agent can only move in the counter-clockwise direction, in state Backtrack).
 It will be enough to show that the agent that changes its state to Gather must have changed it from state Forward (as only in this case the state changing agent stores the dir type message having a counter-clockwise direction component in its local memory).
@@
 We claim that $a_{i}$ and $a_{k}$ can not change to state Gather before round $t^{\prime}+3n+1$.
-Note that at round $t^{\prime}$ both $a_{i}$ and $a_{k}$ are at their corresponding $home$ in state Initial.
+Note that at round $t^{\prime}$ both $a_{i}$ and $a_{k}$ are at their corresponding $home$ in the state Initial.
 Now if any one of them changes to state Gather, the earliest it can happen is at round $t^{\prime}+n+1$ when both of them are in state Forward.
-In this case the agent that changes to state Gather must have found a visited type message at the current node (i.e., for $a_{i}$ it is $home$ of $a_{k}$ and for $a_{k}$ it is $home$ of $a_{j}$) at round $t^{\prime}+n+1$.
-Note that during round $t^{\prime}$ any previous messages are erased from both $home$ of $a_{k}$ and $home$ of $a_{j}$ by $a_{k}$ and $a_{j}$, respectively and no other agent can visit and alter data at these nodes before $t^{\prime}+n+1$ round.
+In this case, the agent that changes to state Gather must have found a visited type message at the current node (i.e., for $a_{i}$ it is $home$ of $a_{k}$ and for $a_{k}$ it is $home$ of $a_{j}$) at round $t^{\prime}+n+1$.
+Note that during round $t^{\prime}$ any previous messages are erased from both $home$ of $a_{k}$ and $home$ of $a_{j}$ by $a_{k}$ and $a_{j}$, respectively, and no other agent can visit and alter data at these nodes before $t^{\prime}+n+1$ round.
 Hence none of $a_{i}$ and $a_{k}$ finds any visited type message at their current nodes at round $t^{\prime}+n+1$.
 Hence, none of these agents changes their state to Gather at round $t^{\prime}+n+1$.
-Next they can only change their state to Gather at the round $t^{\prime}+3n+1$ when both of them (i.e., $a_{i}$ and $a_{k}$) are at their respective $home$.
+Next, they can only change their state to Gather at the round $t^{\prime}+3n+1$ when both of them (i.e., $a_{i}$ and $a_{k}$) are at their respective $home$.
 An agent among them changes its state to Gather at round $t^{\prime}+3n+1$ if it finds no visited type message at their current node (i.e., corresponding $home$).
 Note that $t^{\prime}+3n+1>t>t^{\prime}+n+1$ (as $a_{j}$ was in state Backtrack at round $t$).
-So, at round $t^{\prime}+n+1$ all of $a_{i},a_{j}$ and $a_{k}$ are at nodes $home$ of $a_{k},home$ of $a_{i}$ and $home$ of $a_{j}$ in state Forward and writes a visited type message in the nodes, respectively.
-These messages can not be altered by any agent until round $t^{\prime}+3n+1$ (by similar argument as in Case-I).
-Note that at round $t^{\prime}+3n+1$ both $a_{i}$ and $a_{k}$ are at their corresponding $home$ and both of them finds a visited type message at these nodes left by $a_{j}$ and $a_{i}$ respectively (at round $t^{\prime}+n+1$).
+So, at the round $t^{\prime}+n+1$ all of $a_{i},a_{j}$ and $a_{k}$ are at nodes $home$ of $a_{k},home$ of $a_{i}$ and $home$ of $a_{j}$ in state Forward and writes a visited type message in the nodes, respectively.
+These messages can not be altered by any agent until round $t^{\prime}+3n+1$ (by a similar argument as in Case-I).
+Note that at round $t^{\prime}+3n+1$ both $a_{i}$ and $a_{k}$ are at their corresponding $home$ and both of them find a visited type message at these nodes left by $a_{j}$ and $a_{i}$ respectively (at round $t^{\prime}+n+1$).
 So, none of them changes to state Gather even at round $t^{\prime}+3n+1$.
-Next they move into state Initial-Wait and wait at their $home$ until $t^{\prime}+4n$ round.
-Now, since no agent in state Initial-Wait can change its state to Gather2 until it meets with an agent in state Gather and all alive agent at round $t^{\prime}+3n+2$ are at state Initial-Wait all of them (i.e., $a_{i}$ and $a_{k}$) waits and changes state to Initial again at round $t^{\prime}+4n+1$.
+Next, they move into state Initial-Wait and wait at their $home$ until $t^{\prime}+4n$ round.
+Now, since no agent in state Initial-Wait can change its state to Gather2 until it meets with an agent in state Gather and all alive agents at round $t^{\prime}+3n+2$ are at state Initial-Wait all of them (i.e., $a_{i}$ and $a_{k}$) waits and changes state to Initial again at round $t^{\prime}+4n+1$.
 Next at round $t^{\prime}+4n+2$ both $a_{i}$ and $a_{k}$ erase all previous data at their corresponding $home$.
@@
 Note that the next time $a_{i}$ and $a_{k}$ can change the state to gather must be at the round $t^{\prime}+5n+2$ when both of them are in state Forward, and are currently at the $home$ of $a_{k}$ and $home$ of $a_{j}$, respectively.
-Further, note that since $home$ of $a_{k}$ does not have any visited type message at round $t^{\prime}+5n+2$ (as $a_{k}$ erased any data at round $t^{\prime}+4n+2$ and no other agent can alter data there after $t^{\prime}+4n+2$ and before $t^{\prime}+5n+2$), so $a_{i}$ can not change to state Gather at round $t^{\prime}+5n+2$.
-Also observe that the visited type message at the $home$ of $a_{j}$ written by $a_{k}$ during round $t^{\prime}+3n+1$ is still there at round $t^{\prime}+5n+2$ (after $t^{\prime}+3n+1$ before $t^{\prime}+4n+2$ since $a_{j}$ is already consumed before $t^{\prime}+3n+1$, no agent is on $home$ of $a_{j}$ to erase the data, also after $t^{\prime}+4n+1$ till $t^{\prime}+5n+1$ only $a_{k}$ can visit $home$ of $a_{j}$ in state Forward and Back-Wait, but in these states it does not alter any data at the $home$ of $a_{j}$) $a_{k}$ finds it during the round $t^{\prime}+5n+2$ in state Forward and changes its state to Gather.
+Further, note that since $home$ of $a_{k}$ does not have any visited type message at round $t^{\prime}+5n+2$ (as $a_{k}$ erased any data at round $t^{\prime}+4n+2$ and no other agent can alter data there, after $t^{\prime}+4n+2$ and before $t^{\prime}+5n+2$), so $a_{i}$ can not change to state Gather at round $t^{\prime}+5n+2$.
+Also observe that the visited type message at the $home$ of $a_{j}$ written by $a_{k}$ during round $t^{\prime}+3n+1$ is still there at round $t^{\prime}+5n+2$ (after $t^{\prime}+3n+1$ before $t^{\prime}+4n+2$ since $a_{j}$ is already destroyed before $t^{\prime}+3n+1$, no agent is on $home$ of $a_{j}$ to erase the data, also after $t^{\prime}+4n+1$ till $t^{\prime}+5n+1$ only $a_{k}$ can visit $home$ of $a_{j}$ in state Forward and Back-Wait, but in these states it does not alter any data at the $home$ of $a_{j}$) $a_{k}$ finds it during the round $t^{\prime}+5n+2$ in state Forward and changes its state to Gather.
 Lemma 11.7.
 If an agent finds a dir type message while it is in state Back-Wait, the direction component of this message must be counter-clockwise.
-Proof 5.16.
+Proof 11.8.
 Suppose an agent $a_{i}$ in state Back-Wait gets a dir type message with clockwise direction component at some round $t$.
-This implies there exists another agent $a_{k}$ that has changed its state to Gather after storing a dir type message having direction component clockwise at some round $t^{\prime}<t$.
+This implies there exists another agent $a_{k}$ that has changed its state to Gather after storing a dir type message having a direction component clockwise at some round $t^{\prime}<t$.
 This can only happen if $a_{k}$ was in state Backtrack at the beginning of round $t^{\prime}$.
-So at the beginning of round $t^{\prime}$, $a_{i}$ was also in state Backtrack and thus at round $t^{\prime}$, $a_{i}$ changes its state to Initial-Wait.
+So at the beginning of round $t^{\prime}$, $a_{i}$ was also in state Backtrack, and thus at round $t^{\prime}$, $a_{i}$ changes its state to Initial-Wait.
 Note that $a_{k}$ meets and shares dir type message with $a_{i}$ while $a_{i}$ is still at state Initial-Wait.
 This contradicts our assumption that $a_{i}$ gets dir message at state Back-Wait.
-Thus, if $a_{i}$ finds a dir type message in state Back-Wait then it must have the direction component counter-clockwise.
-With similar argument we can also proof the following lemma.
+Thus, if $a_{i}$ finds a dir type message in the state Back-Wait then it must have the direction component counter-clockwise.
+With a similar argument, we can also prove the following lemma.
 Lemma 11.9.
@@
 Definition 11.10 (Cautious start node).
-Let $a_{i}$, $a_{j}$ and $a_{k}$ be three agents executing the algorithm PerpExplore-Scat-Whitbrd, and suppose $a_{j}$ be the first agent to enter the black hole, while exploring the segment $Seg(a_{j})$.
-Let $v_{1}$ be the $home$ of $a_{j}$ and $v_{2}$ be the furthest node from $v_{1}$ along clockwise direction which is inside $Seg(a_{j})$.
-We define the Cautious start node to be $v_{1}$ if $a_{j}$ is consumed by the black hole during state Forward.
-Otherwise, if $a_{j}$ is consumed by the black hole in state Backtrack then, $v_{2}$ is defined to be the Cautious start node.
+Let $a_{i}$, $a_{j}$, and $a_{k}$ be three agents executing the algorithm PerpExplore-Scat-Whitbrd, and suppose $a_{j}$ be the first agent to enter the Byzantine black hole, while exploring the segment $Seg(a_{j})$.
+Let $v_{1}$ be the $home$ of $a_{j}$ and $v_{2}$ be the furthest node from $v_{1}$ along the clockwise direction which is inside $Seg(a_{j})$.
+We define the Cautious start node to be $v_{1}$ if $a_{j}$ is destroyed by the Byzantine black hole during state Forward.
+Otherwise, if $a_{j}$ is destroyed by the Byzantine black hole in state Backtrack then, $v_{2}$ is defined to be the Cautious start node.
 Lemma 11.11.
-Let the first agent is consumed by the black hole at some round $t>0$, then there exists a round $t^{\prime}>t$, at which the remaining alive agents reach the cautious start node and changes their state to Cautious-Leader and Cautious-Follower.
-Proof 5.20.
+Let the first agent be destroyed by the Byzantine black hole at some round $t>0$, then there exists a round $t^{\prime}>t$, at which the remaining alive agents reach the cautious start node and change their state to Cautious-Leader and Cautious-Follower.
+Proof 11.12.
 Let $a_{i}$, $a_{j}$ and $a_{k}$ be three agents, exploring $Seg(a_{i})$, $Seg(a_{j})$ and $Seg(a_{k})$, respectively, where $Seg(a_{i})\cap Seg(a_{j})=home$ of $a_{i}$, $Seg(a_{i})\cap Seg(a_{k})=home$ of $a_{k}$ and $Seg(a_{k})\cap Seg(a_{j})=home$ of $a_{j}$.
-Let $a_{j}$ be the agent which gets consumed by the black hole, while exploring $Seg(a_{j})$, then we have the following cases:
-Case-I: $a_{j}$ is consumed by the black hole while it is moving in a clockwise direction, i.e., in state Forward at round $t>0$.
-Note that in this case the cautious start node is the $home$ of $a_{j}$.
+Let $a_{j}$ be the agent that gets destroyed by the Byzantine black hole while exploring $Seg(a_{j})$, and then we have the following cases:
+Case-I: $a_{j}$ is destroyed by the Byzantine black hole while it is moving in a clockwise direction, i.e., in state Forward at round $t>0$.
+Note that in this case, the cautious start node is the $home$ of $a_{j}$.
 This implies there exists a round $0<t_{0}<t<t_{0}+n$, where $a_{j}$ was in state Initial.
-Note that, $a_{j}$ can not write any visited type message at $home$ of $a_{i}$, as it gets consumed before reaching that node.
-So, at round $t_{0}+3n+1$, when $a_{i}$ is at its own $home$ in state Backtrack and checks any visited type message, it finds none exists.
+Note that, $a_{j}$ can not write any visited type message at $home$ of $a_{i}$, as it gets destroyed before reaching that node.
+So, at round, $t_{0}+3n+1$, when $a_{i}$ is at its own $home$ in state Backtrack and checks any visited type message, it finds none exists.
 This triggers $a_{i}$ to change its state to Gather from Backtrack with the corresponding dir = (clockwise, NULL) (which is indeed the correct direction, refer Lemma 11.5).
@@
 Note that this must be the $home$ of $a_{j}$, as the home type message $a_{j}$ has written at round $t_{0}$, cannot be erased by any other agent except $a_{j}$.
-Also $a_{j}$ can only erase this in state Initial at round $t_{0}+4n+2$ if it would have returned back, but since it is already consumed between round $t_{0}+1$ and $t_{0}+n-1$, hence this possibility never arises, so the home type message remains, when $a_{i}$ and $a_{k}$ together reaches this node.
+Also $a_{j}$ can only erase this in state Initial at round $t_{0}+4n+2$ if it would have returned back, but since it is already destroyed between round $t_{0}+1$ and $t_{0}+n-1$, hence this possibility never arises, so the home type message remains, when $a_{i}$ and $a_{k}$ together reaches this node.
 After which they change their state to Cautious-Leader and Cautious-Follower depending on their IDs.
-Case-II: $a_{j}$ is consumed by the black hole while it is moving in a counter-clockwise direction, i.e., in state Backtrack at round $t>0$.
-Note that in this case the cautious start node is the $home$ of $a_{i}$.
+Case-II: $a_{j}$ is destroyed by the Byzantine black hole while it is moving in a counter-clockwise direction, i.e., in state Backtrack at round $t>0$.
+Note that in this case, the cautious start node is the $home$ of $a_{i}$.
 Let $t_{0}$ be the round when $a_{j}$ was in state Initial the last time.
@@
 Hence, at round $t_{1}$, where $t_{1}<t_{0}+6n$, $a_{k}$ finds $a_{i}$, while $a_{i}$ is still in state Back-Wait.
-This triggers $a_{k}$ to change its state to Gather2 at round $t_{1}$, while updating the dir type message to (Counter-clockwise, ID’), where ID’ is the ID of $a_{i}$.
-Then at the same round $a_{k}$ writes the updated message at the current node (i.e., $home$ of $a_{k}$).
+This triggers $a_{k}$ to change its state to Gather2 at round $t_{1}$ while updating the dir type message to (Counter-clockwise, ID’), where ID’ is the ID of $a_{i}$.
+Then at the same round, $a_{k}$ writes the updated message at the current node (i.e., $home$ of $a_{k}$).
 Whenever $a_{i}$ sees this dir type message (at round $t_{1}+1$) it also changes its state to Gather2.
-Next in state Gather2, both start to move counter-clockwise (from round $t_{1}+2$) and continues to move until they find a home type message with ID matching the ID of the dir type message.
+Next in state Gather2, both start to move counter-clockwise (from round $t_{1}+2$) and continue to move until they find a home type message with ID matching the ID of the dir type message.
 Note that, this node is nothing but the $home$ of $a_{i}$ (as the ID component of dir type message stores the ID of $a_{i}$).
@@
 So,when $a_{k}$ and $a_{i}$ reaches $home$ of $a_{i}$, they finds the home type message.
-Hence, both $a_{k}$ and $a_{i}$ reach the cautious start node within $t_{0}+7n$ and further changes their states to Cautious-Leader and Cautious-Follower, depending on their IDs.
+Hence, both $a_{k}$ and $a_{i}$ reach the cautious start node within $t_{0}+7n$ and further change their states to Cautious-Leader and Cautious-Follower, depending on their IDs.
 Lemma 11.13.
-Let $a_{i}$ and $a_{k}$ be the two agents which start the states Cautious-Leader and Cautious-Follower from the cautious start node, then within finite rounds of executing algorithm PerpExplore-Scat-WhitBrd, at least one agent detects the location of the black hole.
-Proof 5.22.
-Let $a_{j}$ be the agent that has been consumed by the black hole at round $t>0$. Now there are two cases based on the state of $a_{j}$ at round $t$.
+Let $a_{i}$ and $a_{k}$ be the two agents that start the states Cautious-Leader and Cautious-Follower from the cautious start node, then within finite rounds of executing algorithm PerpExplore-Scat-WhitBrd, at least one agent detects the location of the Byzantine black hole.
+Proof 11.14.
+Let $a_{j}$ be the agent that has been destroyed by the Byzantine black hole at round $t>0$. Now there are two cases based on the state of $a_{j}$ at round $t$.
 Case-I: $a_{j}$ was in state Forward at round $t$.
-In that case the node $v_{1}=home$ of $a_{j}$ is the cautious start node.
-Let $v_{2}$ be the farthest node along clockwise direction in $Seg(a_{j})$ (i.e., the clockwise nearest $home$ of $a_{j}$).
-Now if there exists any round $t^{\prime}<t$ when $a_{j}$ was in state Backtrack then it must have started its state Backtrack from $v_{2}$ and moved counter-clockwise until $v_{1}$ while erasing all right markings from each node in $Seg(a_{j})$.
-So when $a_{j}$ started the state Forward at its $home$ before being consumed, all nodes in $Seg(a_{j})$ are without any right marking.
-This case will happen also when there are no round $t^{\prime}<t$ where $a_{j}$ was in state Backtrack, i.e., there does not exist any round before $t^{\prime}$, when $a_{j}$ was in state Forward.
-So before $a_{j}$ is consumed at the black hole , say $v_{b}$, at round $t$, it has marked right at all nodes starting from the next node of $v_{1}$ in clockwise direction, upto the node just before $v_{b}$ ($v_{b}$ can not be marked as before marking it the agent $a_{j}$ is consumed there).
+In that case, the node $v_{1}=home$ of $a_{j}$ is the cautious start node.
+Let $v_{2}$ be the farthest node along the clockwise direction in $Seg(a_{j})$ (i.e., the clockwise nearest $home$ of $a_{j}$).
+Now if there exists any round $t^{\prime}<t$ when $a_{j}$ was in the state Backtrack then it must have started its state Backtrack from $v_{2}$ and moved counter-clockwise until $v_{1}$ while erasing all right markings from each node in $Seg(a_{j})$.
+So when $a_{j}$ started the state Forward at its $home$ before being destroyed, all nodes in $Seg(a_{j})$ are without any right marking.
+This case will happen also when there is no round $t^{\prime}<t$ where $a_{j}$ was in state Backtrack, i.e., there does not exist any round before $t^{\prime}$, when $a_{j}$ was in state Forward.
+So before $a_{j}$ is destroyed at the Byzantine black hole, say $v_{b}$, at round $t$, it has marked right at all nodes starting from the next node of $v_{1}$ in the clockwise direction, up to the node just before $v_{b}$ ($v_{b}$ can not be marked as before marking it the agent $a_{j}$ is destroyed there).
 Let without loss of generality, $a_{i}$ is in state Cautious-Leader and $a_{j}$ is in state Cautious-Follower at $v_{1}$ at some round $t_{0}>t$.
-Then $a_{i}$ always moves ahead alone in clockwise direction to a new node $v$.Then it moves back only if sees the right marking and brings $a_{k}$ to $v$ along with it.
+Then $a_{i}$ always moves ahead alone in the clockwise direction to a new node $v$.Then it moves back only if sees the right marking and brings $a_{k}$ to $v$ along with it.
 Note that $a_{i}$ always sees a right marking until $v_{b}$.
-Now when it moves to $v_{b}$, if it is not consumed it must see no such right marking as $a_{j}$ failed to mark it at round $t$.
-In this case $a_{i}$ leaves the node in clockwise direction to a new node.
-So here $a_{i}$ is able to detect the black hole.
-On the other hand if $a_{i}$ is consumed while it visits $v_{b}$, then it can not return back to $a_{k}$.
-When $a_{k}$ sees that $a_{i}$ has not returned it interprets that $a_{i}$ must have been consumed at the black hole which is the next node along clockwise direction.
-Thus in this case also at least one agent can detect the black hole.
+Now when it moves to $v_{b}$, if it is not destroyed it must see no such right marking as $a_{j}$ failed to mark it at round $t$.
+In this case, $a_{i}$ leaves the node in the clockwise direction to a new node.
+So here $a_{i}$ is able to detect the Byzantine black hole.
+On the other hand, if $a_{i}$ is destroyed while it visits $v_{b}$, then it can not return back to $a_{k}$.
+When $a_{k}$ sees that $a_{i}$ has not returned it interprets that $a_{i}$ must have been destroyed at the Byzantine black hole which is the next node along the clockwise direction.
+Thus in this case also at least one agent can detect the Byzantine black hole.
 Case-II: $a_{j}$ was in state Backtrack at round $t$.
@@
 This means that, when $a_{j}$ started the state Backtrack from $v_{2}$, there does not exist any node in $Seg(a_{j})$ with left markings.
-So before $a_{j}$ is consumed by the black hole node $v_{b}$ (say) at round $t$, it must have marked all the nodes with left, starting from the next counter-clockwise node of $v_{2}$ to the adjacent clockwise neighbor of $v_{b}$ (as before writing this message at $v_{b}$, the agent gets consumed by the black hole).
+So before $a_{j}$ is destroyed by the Byzantine black hole node $v_{b}$ (say) at round $t$, it must have marked all the nodes with left, starting from the next counter-clockwise node of $v_{2}$ to the adjacent clockwise neighbor of $v_{b}$ (as before writing this message at $v_{b}$, the agent gets destroyed by the Byzantine black hole).
 Let without loss of generality, $a_{i}$ be the lowest ID agent among $a_{i}$ and $a_{k}$, hence it starts in state Cautious-Leader, whereas $a_{k}$ starts in state Cautious-Follower, at some round $t_{0}>t$.
-This means $a_{i}$ is the first agent to move alone in the next counter-clockwise neighbour say, $v$.
-After which, only if it sees a left message then only it moves back in clockwise direction at the node of $a_{k}$, and in the next round both these agents reach the node $v$.
-Observe, $a_{i}$ always finds a left message until the node $v_{b}$.
-Whenever it reaches $v_{b}$, either it gets consumed by the black hole, or it finds that no left marking is present at the current node.
-This triggers $a_{i}$ to detect the current node to be the black hole and moves in counter-clockwise direction to a new node.
-Otherwise, if it also gets consumed by the black hole, then in the next round it is unable to return to $a_{k}$, which triggers $a_{k}$ to conclude that $a_{i}$ must have been consumed by the black hole as well, and it correctly detects the black hole to be the next node in the counter-clockwise direction.
-Thus for each scenario there exists at least one agent which is able to correctly detect the black hole location.
-Note that within at most $3n$ number of rounds after both alive agent starts cautious walk from cautious start node, the black hole will be detected by at least one agent.
-So from Lemma 5.19 and Lemma 5.21 within atmost $10n$ rounds after the first agent is consumed by the black hole there exists atleast one agent that knows the exact location of the black hole which can now explore the ring $R$ perpetually avoiding the black hole.
-So if all three agents strats from three different nodes the PerpExploration-BBH will be solved if each node has a whiteborad of memory $O(\log n)$.
-Now for the case where three agents starts from two different nodes we have the following remark.
+This means $a_{i}$ is the first agent to move alone in the next counter-clockwise neighbor say, $v$.
+After which, only if it sees a left message then only it moves back in the clockwise direction at the node of $a_{k}$, and in the next round both these agents reach the node $v$.
+Observe that, $a_{i}$ always finds a left message until the node $v_{b}$.
+Whenever it reaches $v_{b}$, either it gets destroyed by the Byzantine black hole, or it finds that no left marking is present at the current node.
+This triggers $a_{i}$ to detect the current node to be the Byzantine black hole and moves in the counter-clockwise direction to a new node.
+Otherwise, if it also gets destroyed by the Byzantine black hole, then in the next round it is unable to return to $a_{k}$, which triggers $a_{k}$ to conclude that $a_{i}$ must have been destroyed by the Byzantine black hole as well, and it correctly detects the Byzantine black hole to be the next node in the counter-clockwise direction.
+Thus for each scenario, there exists at least one agent that is able to correctly detect the Byzantine black hole location.
+Note that within at most $3n$ number of rounds after both alive agent starts cautious walk from the cautious start node, the Byzantine black hole will be detected by at least one agent.
+So from Lemma 11.11 and Lemma 11.13 within at most $10n$ rounds after the first agent is destroyed by the Byzantine black hole there exists at least one agent that knows the exact location of the Byzantine black hole which can now explore the ring $R$ perpetually avoiding the Byzantine black hole.
+So if all three agents start from three different nodes the PerpExploration-BBH will be solved if each node has a whiteboard of memory $O(\log n)$.
+This proves the Theorem 5.4.
+We state the theorem again here for convenience.
+Statement of theorem 5.4: Algorithm PerpExplore-Scat-Whitbrd solves
+PerpExploration-BBH problem of a ring $R$ with $n$ nodes and with 3 synchronous agents initially scattered under the whiteboard model of communication
+11.3 Modification to the algorithm to include the cases where the initial starting nodes can have multiplicity
 Remark 11.15.
-Let $a_{i},a_{j}$ and $a_{k}$ be three agents that starts from two initial nodes, say $home_{1}$ and $home_{2}$.
+Let $a_{i},a_{j}$, and $a_{k}$ be three agents that start from two initial nodes, say $home_{1}$ and $home_{2}$.
 By Pigeon hole principle, exactly one of $home_{1}$ and $home_{2}$ initially must have two agents.
-Without loss of generality let $home_{1}$ has two agents, say $a_{i}$ and $a_{k}$, initially.
-In this case, the agents having multiplicity greater than one at the current node does not move.
+Without loss of generality let $home_{1}$ have two agents, say $a_{i}$ and $a_{k}$, initially.
+In this case, the agents having multiplicity greater than one at the current node do not move.
 On the other hand the singleton agent, i.e., $a_{j}$ starting from $home_{2}$ moves clockwise marking each node with message right.
-If $a_{j}$ reaches $home_{1}$ before being consumed by the black hole, $home_{1}$ now has three agents co-located.
-Thus from here the agents execute the whiteboard version of PerpExplore-Coloc-Pbl (refer Remark 4.24).
-On the other hand, let us consider the case when $a_{j}$ gets consumed before reaching $home_{1}$.
+If $a_{j}$ reaches $home_{1}$ before being destroyed by the Byzantine black hole, $home_{1}$ now has three agents co-located.
+Thus from here the agents execute the whiteboard version of PerpExplore-Coloc-Pbl (refer Subsection 4.3).
+On the other hand, let us consider the case when $a_{j}$ gets destroyed before reaching $home_{1}$.
 Note that irrespective of the location of $home_{2}$ it takes at most $n-1$ rounds for $a_{j}$ to reach $home_{1}$ from the beginning.
 So $a_{i}$ and $a_{j}$ waits for $n$ rounds and finds that no one has arrived yet.
-In this case, both of $a_{i}$ and $a_{k}$ moves to $home_{2}$ along clockwise direction and starts to perform the cautious walk , where lowest ID agent among $a_{i}$ and $a_{k}$ changes its state to Cautious-Leader, whereas the other agent changes its state to Cautious-Follower.
+In this case, both $a_{i}$ and $a_{k}$ move to $home_{2}$ along the clockwise direction and start to perform the cautious walk, where the lowest ID agent among $a_{i}$ and $a_{k}$ changes its state to Cautious-Leader, whereas the other agent changes its state to Cautious-Follower.
 Next, the agent executing Cautious-Leader searches for the marking right.
-As argued earlier, within at most $4n$ rounds after an agent is consumed, at least one of the remaining alive agents can detect the black hole and continue to explore the ring avoiding that node.
-So from Remark 4.24 and from the results in this section we have the following theorem
-Theorem 5.24.
-Three synchronous agents are necessary and sufficient to solve the problem PerpExploration-BBH on a ring $R$, when each node of $R$ has a whiteboard of $O(\log n)$ bits of memory irrespective of their starting location.
-6 Conclusion
+As argued earlier, within at most $4n$ rounds after an agent is destroyed, at least one of the remaining alive agents, detect the Byzantine black hole and continue to explore the ring avoiding that node.